A Bibliography of Publications in Communications of the ACM : 1990–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

11 April 2024 Version 2.91

Title word cross-reference

$150 [1852]. 3 [1862]. N [24]. O_{2} [226]. p × 64 [164].

-Fold [24].

1 [1302]. 12 [488]. 1989 [276]. 1989-1990 [275]. 1990s [444, 135, 214].

2 [1187]. 20/20 [1581]. 2000 [1488, 1476, 1309]. 20th [71]. 21st [241, 1661, 1239, 1421, 468, 1116]. 22nd [401].

3.0 [1650]. 3D [1689, 331, 458].

4 [439]. 400 [918]. 47s [1833].

5 [582]. 50th [980].

’89 [26].

’90s [1203]. ’92 [39]. ’95 [779]. ’96 [965]. 9X [398, 399, 397].

= [1396, 186].

Abandoned [1385]. Above [834]. Abstract [1179]. abstracting [1442]. abstractions [1865]. abuses [1866]. academe [1313, 1808]. Academic [646, 262, 1459, 1846, 1549, 60]. Academic-Industry [262]. Acadia [1458]. acceleration [317]. acceptance [1490].

Access [736, 1070, 1454, 1326, 234, 1459, 11, 1520, 1721, 1513, 1591, 1518, 205, 475, 826, 1146]. Accessing [1512]. Accidental [222, 375]. Accountability [958, 606]. accounting [1482]. Accreditation [259]. Accredited [1055, 259, 418, 673, 790]. ACE [457]. achieve [1376]. achieved [1082]. achievements [231]. achieving [1620]. ACM [401, 71, 979, 323, 442, 962, 961, 983, 1056, 1094, 1088, 1207, 1294, 1366, 1418, 1413, 1205, 1296, 1292, 1369, 1412, 1373, 1209, 1295, 1367, 25, 1449, 1489, 1869, 322, 831, 1328, 321, 404, 1299, 388, 278, 490, 1211, 241, 488, 1136, 232, 189, 1768, 34, 1826, 779]. ACM/IEEE [189]. ACM/IEEE-CS [189]. ACM’s [980]. Acquisition [1078, 1437, 125, 574]. Across [951, 1456]. Act [1832, 1255, 1093, 1696]. Action [1710, 1892, 604]. Active [1705]. Activities [234]. Activity [884, 288]. ADA [321, 398, 392, 394, 395, 399, 396, 393, 397]. AdAM [1778]. Adaptability [1141]. Adapting [1687]. Adaptive [657, 1199, 1397, 1543]. added [984, 1697]. Addendum [1870]. addictive [1505]. Adding [1568, 616, 532]. Additional [289]. address [1605]. addressing [698, 1782]. adds [1587, 1418]. Adequacy [262]. adopting [1139]. adoption [253, 304, 1711]. Advanced [440, 1501, 461]. Advancement [1023, 772]. Advances [643]. Advantage [949, 1920]. Adventure [238]. advertisements [1861]. advertising [1500]. advice [497]. Africa [612, 951, 432]. After [1191, 1394]. Aftermath [1343]. again [1534, 1566]. Against [804, 1375, 264, 1414]. Age [1059, 1285, 1811, 279, 1606, 636]. Agency [898]. Agenda [27, 263, 72]. Agent [685, 842, 1753]. agent-based [1753]. Agents [1744, 688, 1736, 1741, 1745, 678, 945, 1746, 412, 689, 1737, 679, 680, 686, 690, 691, 1748, 677, 1503, 682, 676, 687, 1743, 681, 1749].

Aggravation [348]. Aglets [1751]. agora [1463]. agreed [1377]. agreed-upon [1377]. ahead [982, 1207, 1662]. AI [937, 593, 941, 625, 932, 49, 1344]. AI-based [49]. Aid [243]. Aids [992]. aim [88]. Air [622]. AK [1833]. AK-47s [1833]. Alan [55]. Alexandria [823]. Algol [519]. Algorithm [311, 1, 1180, 146, 1182, 85, 2, 357, 289, 1761]. algorithm-based [1761]. Algorithms [1187, 1012, 1950, 1946, 117, 76, 1129, 1306, 636]. All-optical [1821]. Allen [1412, 405]. alliance [663]. allocation [856]. Allowing [1377, 382]. along [650]. Alpha [435, 434, 433, 436]. Alphabets [45]. Already [1282]. Alt [979]. Altair [554]. Alternative [659, 62]. Alternatives [1001]. alters [1377]. always [1581, 851]. Amazing [1270]. Amazon [300]. Amdahl [1184]. Amendment [183]. America [1654]. American [71, 241, 474]. Amoeba [127]. Among [1891]. Amplifiers [431]. Amsterdam [616]. Analog [1779, 1907]. analogy [305]. Analysis [190, 413, 361, 111, 24, 1054, 362, 1701, 848, 1185, 1780, 333, 369, 286, 607, 383, 749, 886, 642, 510, 839, 1264, 364, 384, 1448, 357, 596, 363, 549, 1527, 643, 1761]. Analyst [1010]. Anchoring [1162]. Andrew [1188, 1189]. Animal [565]. Animation [1858, 1046, 1101, 458]. anniversary [980]. annotations [560]. Announcing [763]. Annual [274, 80, 71, 52, 401]. Anonymity [1177]. Anonymous [1719, 1721]. Answer [1112]. Answers [1183, 912, 467]. Anti [1103]. Anti-Mac [1103]. anticipate [1656, 1377]. anticircumvention [1871]. antiquities [1513]. Antivirus [1195]. Anytime [1341]. anyway [1508]. Anywhere [1341]. Apache [1772]. Apartness [612]. Application [498, 54, 1426, 880, 1897, 1395]. Applications [1160, 939, 937, 940, 943, 795, 938, 270, 1301, 926, 634, 1383, 627, 1595, 875, 890, 1567, 1402, 134, 628, 1777, 630, 163, 1708, 1619, 1841]. applicative [379]. Applied [110]. Applying [798, 1349, 1262, 957, 364, 1139, 613]. appraising [1839]. appreciation [819]. Apprenticing [838]. Approach [1031, 380, 73, 381, 880, 592, 286, 1801, 509, 1316, 252, 564, 508, 1922, 652, 656, 49, 668]. Approaches [959, 349, 1254, 704, 1553, 1529]. approved [1902]. Approximate [1183]. apps [1912]. Arcade [1345]. Architecting [409]. Architecture [424, 814, 1665, 434, 1201, 1629, 415, 1138, 1881, 664, 617, 203, 687, 1843]. architectures [1897, 1689, 1922, 1651, 1754, 783]. Archives [963]. Area [812, 249]. Argument [1764]. arm [1679]. Armchair [1764]. Arms [804]. array [357]. Arrays [1179, 1776, 871]. Art [193, 1132, 1407, 1548, 28]. Arthur [122]. articulator [318]. articulator-based [318]. Artificial [404, 945, 1196, 1855, 349]. Arts [1175]. Asia [984]. Asian [756, 492]. Ask [1293]. Asleep [1227]. Aspects [1141, 1475]. assess [391]. Assessing [1537, 1482]. Assessment [1432, 484, 865, 33, 302, 1686, 53, 843, 21, 121]. assignment [655, 1678]. assignments [1716]. assistant [684]. Assistive [315]. Assumptions [829]. Assurance [1198]. Asymmetries [1164]. Asynchronous [1035, 1378]. Atlantic [1061]. ATLAS [1931]. ATM [789, 787, 784, 788, 786, 785, 783, 782]. attained [1545]. Attention [1351]. attitudes [946]. attributes [596]. Auction [856]. auctions [1594]. Audio [421, 915, 332, 1544]. augmentative [316]. Augmented [1350, 520, 525, 526]. Augmenting [876, 532]. Australian [946]. authentication [19]. author [962]. authoring [893]. Authors [1203, 1013, 1036, 1564]. Automated [653, 713, 1747]. Automatic [726, 1554, 620, 332, 383]. Automatically [32]. Automating [301, 410]. Automation [1303]. Autonomous [945, 1556, 1822]. Availability [1031]. available [1082]. Average [155]. Aviation [610]. avoid [1502]. avoidance [860]. Avoiding [419]. AWACS [1896]. Award [1292, 1052, 724, 426, 1053, 741, 1412]. Awards [1292]. Aware [524]. Awareness [1358]. awesome [463]. AXP [435, 434].

B [178]. Babbage [1607, 290]. Back [520, 1435]. backbone [1818, 699]. Backpropagation [1182]. backtracks [1759]. Backup [1007]. backward [1545]. Bad [1266, 1732]. Badge [1095]. Balance [1764, 1297]. Balanced [62]. Bandwagon [1001]. bandwidth [1785, 75]. Bank [1654]. Barney [708]. Base [335]. Base/Database [335]. Based [1260, 599, 626, 1363, 996, 846, 1019, 627, 1552, 286, 1183, 176, 269, 683, 526, 1617, 1914, 1753, 1653, 815, 1782, 19, 370, 22, 1894, 657, 586, 7, 847, 240, 1499, 318, 656, 49, 619, 200, 1754, 1631, 1761, 1519, 1749, 1594, 617, 615]. basic [633]. battle [1744, 1611, 1698]. Bayesian [798, 796, 795]. Be [1873, 1115, 1368, 1082, 76, 1624]. beam [1862]. Bearing [757]. beasts [1315]. beat [1839]. Become [1233, 170]. Becomes [1052]. Becoming [116]. Been [223, 137, 1804]. beer [867]. Before [351, 554]. begin [1209]. begins [984]. Behavior [1388, 527, 1934, 363]. Behind [1132]. believable [689]. Bell [1061]. belong [583]. Beltway [1249]. Benchmarking [1572]. benefits [1546, 501, 1653, 92, 1875]. Berkeley [822]. Best [1224, 1002, 1546, 1436, 158, 1884]. Better [1225, 1662, 1546, 282]. Between [1150, 1220, 1222, 1099, 1551, 514]. Beware [757]. Beyond [1228, 1612, 1307, 1291]. bias [853]. bibliographic [1517]. Big [1312, 1074]. Bilingual [47, 1771]. Bill [445]. binaries [1446]. Binary [438, 2]. Binding [1652]. Biological [237, 517]. Biologically [1254]. Biology [237]. biometrics [1866]. Biosphere [219]. Birth [1245]. Bit [336, 1756]. Bit-rot [1756]. Bit-Tree [336]. BITNET [128]. BITS [738]. Blacksburg [1173]. Blinn [1947]. Bloom [1208]. blue [1919, 1086, 1343, 1344]. Board [18]. Body [1213, 1356]. Body-Image [1356]. Books [1274, 983, 1295, 1520]. Bookshelf [1038]. Boosting [876]. Bootstraps [1074]. Both [1325, 1884]. bottlenecks [1124]. Bottom [691, 1620]. bottom-line [1620]. Bottom-up [691]. Boundaries [913]. Bounding [6]. Bowl [80]. Brain [1282, 1343]. Brazilian [300]. Breach [387]. breakdown [1771]. Breaks [1284]. bridge [1563]. bridging [82]. Brief [494, 493]. Bringing [421, 1071]. British [861, 826]. broadband [422]. Broadening [430]. Brock [1375]. Brokerages [1065]. brokers [1651]. Brother [1074]. browser [1516]. Budget [66, 867, 304]. Bug [1279]. build [1708]. Building [1818, 465, 1122, 175, 1632, 1781, 1028, 457, 1563, 399, 824, 786, 263, 1743, 47, 602, 220, 1405, 472, 603]. builds [1320]. bureaucracies [255]. bus [7]. bus-based [7]. Business [1237, 1567, 1155, 1099, 998, 1764, 274, 1338, 1083, 634, 35, 198, 1586, 627, 1546, 52, 1376, 1714]. buy [1746, 1789]. Buying [1869, 339, 1934]. buzzwordism [1852].

C [563, 236, 844, 92, 1070, 1903]. C* [325]. C-ODA [1070]. C/C [1903]. Cache [81]. CAD [284]. Calculating [1476]. Calculation [1291, 153]. Calculator [777]. Calendar [1037]. California [816]. camel [1771]. CamMotion [1105]. campus [1456, 675, 1460]. Can [743, 1016, 1873, 1129, 1254, 489]. Candidates [1011]. CANDY [318]. capabilities [1568]. Capability [1317].

Capital [1336]. CAPS [243]. Care [1361, 1633, 1359, 1346]. career [802]. Careers [646]. carrying [1082]. Cartwheeling [1105]. Carver [1412]. Case [627, 1552, 985, 1386, 1724, 1936, 1466, 128, 1147, 1618, 1846, 178, 298, 1504, 1654, 304, 1618, 301, 306, 308, 773]. Case-based [627, 1552]. cash [1064]. causes [1899]. Cautionary [1298]. CAVE [332]. CD [816, 1089, 166]. CD-I [166]. CD-ROM [816, 1089]. CDA [1325, 1375]. ceiling [767]. Celebration [1368, 766, 948]. Cells [1181]. Celtic [1245]. Censorship [1068, 1146]. center [1522, 810]. Centered [1333, 817, 1028, 1330, 576, 1618, 1018]. centers [60]. centralization [195]. centric [1815, 1843]. Century [1239, 468, 1661, 1421, 1116]. CEO [1863]. CEOs [182]. ceremony [1292]. CERT [705]. Certification [1831]. Certifying [147]. Challenges [996, 1043, 505, 608, 722, 784, 640, 782]. challenging [1835]. Champagne [867]. Champion [235]. championship [241, 401, 71]. chance [797]. change [842, 1464, 859, 1492, 489, 888, 1110]. changed [1608]. Changes [1461, 1826]. Changing [130, 1449, 1543]. channel [1497]. Chaos [944, 573]. Chaotic [943, 776]. characters [1752]. Charade [522]. charges [1454]. Charles [290]. chasm [1313]. check [1761]. chemistry [818]. Chess [401, 71, 241]. CHI [159]. Chief [1415, 1373]. Children [1358, 318]. Chile [897]. China [1563, 1435, 1817]. Chinese [1550, 1517]. Chip [433]. chips [1777]. choices [1465, 563]. Choosing [534, 1321]. choppy [1663]. Chuck [1607]. CIDR [698]. CIM [1404]. Cinema [1218, 67]. CIO [1863]. CIOs [1594]. Circuits [1779, 1760]. Citedness [172]. Citizens [1474]. clandestine [842]. Class [95, 1876, 712]. classes [1321]. classification [726, 1449, 748, 180]. Classroom [1378].

Classrooms [930]. Clearing [1360]. Client [1546, 346, 969]. Client-Server [346]. Client/server [1546]. CLIM [215]. Clinical [1360]. Clipper [552]. Clipping [552, 344]. Clock [138]. clockwise [1895]. CLOS [211]. Clothing [1096]. Cluster [1031, 712]. Clustering [1183]. CM [1187, 582]. CM-2 [1187]. CM-5 [582]. CMM [1131]. CMU [1188]. COACH [685]. Cobol [1438, 1371, 1555]. Code [1430, 518, 323, 442, 1902, 653]. codes [43]. codevelopment [500]. Coding [164, 243, 46]. Coevolution [1195]. Cognition [850, 882, 1113]. Cognitive [853]. cognitively [1743]. coherence [1741, 883]. coin [408]. coincidences [100]. Cold [602]. Collaboration [249, 1423, 900, 903, 703, 1524, 258]. Collaborative [1260, 1128, 1259, 1262, 1503, 261, 258, 885, 1284, 884, 333, 679, 414, 1708, 1504, 479, 773]. Collaboratory [1363, 1021, 256]. Colleagues [151]. Collecting [1095, 1839]. Collection [1181, 1742, 1314]. Collective [876, 327]. college [1136]. collegiate [981]. Color [15]. Column [1851]. COM [1542]. combinational [1759]. Combinatorial [774, 1761]. Combining [1459, 1318, 1259]. COMDEX [65, 192]. Come [1285, 1491, 636]. Comes [915, 1473]. coming [149]. Commentary [993, 1816]. Commentators [447]. Comments [388, 151]. Commerce [978, 1069, 1060, 1679, 1876, 1835, 1753, 1745, 1889, 1064, 1655, 1590]. Commercial [1062, 940, 941, 625, 932, 1006]. Commercialization [731]. commission [40]. commitment [609]. Committee [1207, 391, 1294, 1366]. Common [77, 1630, 218, 1304, 215]. Communication [1035, 1031, 4, 1030, 1034, 919, 1033, 317, 316, 40, 1124, 423, 253, 1313, 1544, 1029]. Communications [1242, 546, 1093, 17, 461, 785, 962, 961, 1696].

Communism [390]. Communities [1283, 1023, 95, 1739]. Community [603, 38, 1515, 302, 36, 1942]. companions [308]. company [1546, 1376]. Compare [1764]. Comparing [16, 1050, 857, 1054, 371, 1903]. Comparison [1187, 498, 50, 621, 30, 358, 271]. Compatibility [873]. competing [1691]. Competitive [412, 1546, 1863]. Competitiveness [1592]. Compilations [280]. Compiler [1949, 287]. compilers [668]. Compiling [354]. Complete [287]. Complex [1026, 775, 1617]. Complexity [724, 1240, 65, 587]. Compliance [986, 1725, 1476]. Component [373, 1542, 1894, 870]. component-based [1894]. Component-Oriented [373]. components [1567, 1396]. componets [1397]. comprehension [882]. compressed [319]. compressing [809]. compression [272, 163, 161]. Computation [1161, 380, 921, 1339, 1420, 1915, 583, 82]. Computational [724, 708, 890, 532, 1104, 753]. Computations [1126]. Computer [1203, 1263, 1221, 730, 1009, 1099, 646, 953, 1474, 1621, 3, 1112, 1406, 811, 274, 764, 80, 295, 486, 184, 628, 1026, 724, 817, 156, 958, 430, 401, 54, 1084, 1387, 490, 31, 1195, 359, 860, 71, 116, 1185, 987, 1011, 140, 1238, 553, 232, 1017, 139, 1175, 520, 1289, 35, 198, 988, 989, 1537, 1461, 55, 1870, 1840, 749, 52, 605, 20, 288, 134, 478, 1550, 1194, 765, 391, 1728, 550, 252, 754, 193, 257, 1315, 22, 1465, 143, 348, 1505, 241, 166, 1625, 1530, 753, 1855, 136, 529]. Computer-Augmented [520]. computer-based [22]. Computer-Related [1238]. computer-supported [478, 257]. Computerized [123]. computerizing [1625]. Computers [1456, 1681, 1491, 1464, 793, 994, 524, 314, 4, 1101, 489, 197, 159, 291, 1233, 1102, 339, 1169, 482, 315, 538, 1462, 141, 477, 142].

Computing [762, 1055, 980, 897, 1324, 421, 604, 951, 835, 1176, 250, 130, 208, 350, 612, 756, 391, 169, 960, 1717, 1425, 259, 418, 673, 790, 300, 237, 1096, 774, 390, 1016, 492, 606, 239, 758, 65, 192, 159, 249, 291, 406, 554, 590, 467, 623, 675, 794, 862, 915, 987, 1041, 1208, 1326, 1417, 1714, 885, 10, 242, 346, 189, 529, 121, 101, 124, 223, 198, 173, 349, 290, 389, 519, 431, 568, 779, 965, 979, 1205, 592, 1457, 1449, 1374, 1640, 1737, 115, 856, 195, 1808, 128, 722, 595, 1664, 759, 1667, 1922, 258, 1789, 113, 141, 1890]. computing [60, 601, 1097, 1649, 521, 643, 640, 1793, 1843, 1460, 632, 1207, 1291]. Compuvision [89]. concept [655, 726, 462]. Concepts [954, 1672, 783]. Conceptual [1183, 910, 1321]. concerns [1724, 1498]. Concordia [1750]. Concurrency [792, 186, 560, 562, 53]. Concurrent [97, 555, 556, 64, 1140, 1758, 561, 559, 1531]. conduct [323]. Conference [971, 14]. confidence [1837]. Configuration [206]. confining [1125]. confirm [1724, 723]. Conformance [967]. confronting [513]. Congress [827, 1602]. Congruential [1178]. connascence [371]. connected [1930, 1832]. Connecting [1702]. Connection [1213, 1221, 1217, 1218, 1215, 582, 1216, 787, 1220, 1214, 1222, 1219, 1223, 877, 1648, 1187]. Connectionist [117]. Connections [1909, 1135]. connectivity [544]. cons [1139]. Consensus [638, 914, 1579, 399]. Consequences [960]. Considerations [1824, 698, 81, 895, 1628]. Considered [1432, 844, 429]. Consistent [1721]. Consortium [1090]. Constitution [1076]. Constitutional [150]. Constraint [69, 9]. Constraints [1180]. Construction [287, 1178, 607, 887, 7, 527, 1761]. Constructivism [1409]. consultancy [1678]. consulting [1716, 1678, 577]. Consumer [1498, 1781]. containment [1125]. contemporary [1477]. Content [1260, 1500, 1587, 1611, 1583, 1925].

Content-Based [1260]. Contents [1165]. contest [981]. Context [1308, 977, 453, 616]. context-free [453]. Contextual [1337]. continuum [1664]. Contract [1241, 387]. Contradiction [1371]. contribution [1799]. Control [522, 736, 887, 1140, 1858, 201, 1550, 257, 203, 1889, 125, 202, 200]. controls [1146, 1663]. conundrum [858, 859]. conventional [1835]. Convergence [169]. Converging [595]. conversation [677, 512, 1541]. conversational [316]. conversion [84]. COOL [557]. cooperating [254]. Cooperation [257, 1205, 254]. Cooperative [1803, 617, 454, 505, 502]. coordinated [1817]. coordinating [473]. Coordination [285, 801, 257]. cop [1839]. COPA [1698]. Copyright [1869, 280, 553, 600, 806, 360, 1826]. Copyrighted [1075]. CORBA [1653, 1652, 1648, 1649, 1650]. CORBA-based [1653]. Corbató [221]. Corn [386]. Corner [1947, 1211]. Corporate [283, 760, 1453, 1301, 1131, 595, 1500, 1783]. Correction [1851, 1831, 320]. Corrections [1273]. Correlating [1695]. Correspondence [719, 311, 792, 538, 791, 566, 776, 347, 430, 54, 534, 777, 325, 621]. Corrupted [588]. Cosmetic [48]. Cosmology [1427]. Cost [1331, 743, 848, 916, 1907, 1476, 282, 1560, 1615]. Cost-Effective [743]. Cost-Justifying [1331]. costs [587]. COTS [1437]. Could [835, 223]. Council [41, 391]. Countdown [1302, 1788]. Counting [1315]. Countries [835, 1809, 1814, 544]. Country [250, 149]. Coupling [262, 335]. Course [844, 1873, 858, 859]. Courses [1054, 1580]. cover [374]. coverage [1760]. Coyote [478]. Crackers [565]. Craig [150]. Crash [1231, 1484]. Create [845]. Created [1219]. Creating [1497, 841]. creation [979, 410].

Creative [573]. creativity [688]. Credibility [1793]. Criminological [730]. Crisis [1191, 1311]. Criteria [1388, 378]. Critical [1385, 59]. criticism [1808]. Critique [325]. Critiques [936]. critiquing [309]. Cross [50, 1898]. Cross-Cultural [50]. cross-platform [1898]. Crossroads [924, 753]. crowds [1719]. Cruelty [3, 57]. Crush [1231]. Cryptanalysis [737]. Crypto [580, 706, 1007, 1365, 805]. cryptograms [717]. Cryptography [718, 388, 1114, 1204, 139, 1414, 1904]. cryptosystems [734]. CS [1269, 1638, 1673, 1174, 189]. CSCW [505, 251]. CSCW’90 [159]. CSS [1901]. Cuba [390]. Cube [8]. cues [883]. Cul [1001]. Cul-de-Sac [1001]. Cultural [50, 1512, 467]. Culture [1237, 725]. cultures [1940]. Current [346, 22, 459, 629, 96]. Curricula [189]. Curriculum [394, 153, 1174, 490, 189, 1175, 1462]. custom [1402, 1497]. Customer [837, 838, 576, 1499]. customer-centered [576]. Customer-Developer [837]. customers [841]. customization [657]. Customized [992]. Cybercommerce [1063]. CyberNetwork [1166]. Cybersickness [1355]. Cyberspace [199, 950, 1210, 1419, 951, 1042, 1496, 733]. CYC [933, 936, 77]. Cycle [206, 98, 1547].

D [1862, 354, 1121]. DACNOS [42]. Dad [648]. Dance [492]. dark [1845]. DARPA [1738]. Data [1179, 326, 1869, 1152, 1153, 1157, 1154, 1156, 1667, 1481, 1423, 280, 600, 1288, 1083, 1308, 1162, 144, 651, 1707, 1082, 1629, 1860, 272, 333, 336, 1632, 1742, 722, 642, 1494, 1482, 857, 205, 1911, 1540, 1314, 125, 368, 1827, 1714, 1483, 374, 1627, 83, 1479, 658, 1480]. Data-intensive [1667]. Database [228, 817, 335, 1158, 1243, 934, 120, 1165, 231, 229, 1595, 225, 286, 334, 84, 1444, 227, 1553].

Databases [1155, 1152, 294, 11, 273, 368, 652]. Datacycle [415]. DATAPLEX [11]. Date [1191, 1311]. David [1415]. Dead [1696]. deaf [320]. deal [1594]. Dealing [1210]. Death [834, 348]. debacle [1852]. Debatable [25, 34]. Debate [1325, 355, 151, 233, 150, 195, 805]. Debating [341, 1375]. Debugging [1160, 1281, 1283, 1284, 1280, 797, 1393]. Debunking [337]. debut [983]. decade [982, 188]. Decency [1093, 1696]. Decertification [1831]. Decimal [260]. Decision [799, 958, 1084, 621, 442, 1552, 41, 1499, 384, 114, 1825, 271]. decision-making [1825]. Decision-Theoretic [799]. decisions [1540, 1759]. Declarative [203]. decode [717]. decoding [43]. Decrypting [340]. deep [1919, 1086, 1343, 1344]. Defeated [1093]. Defective [420]. Defense [466, 404, 1151]. Defensive [1836]. Define [1426]. defining [1728]. Definition [1491]. Degree [1175]. Dehumanizing [742]. Delay [6, 1760]. Deletion [347, 155]. Delivering [217, 1316]. Delivery [1332, 413, 1403, 409, 475]. Demand [135, 841]. Demo [1089]. Democracy [1225, 496]. democratic [602]. Demographics [1011]. Demonstration [436]. Denial [735]. Denning [151]. dense [269]. dense-map [269]. Department [140]. Departmentalization [61]. depend [1524]. dependability [586]. Dependable [443]. dependence [357]. dependencies [911]. Deployment [970]. Derived [347]. Deriving [911]. Desert [432]. Design [307, 716, 1702, 498, 747, 1400, 889, 439, 994, 1333, 615, 878, 1330, 992, 775, 90, 1054, 362, 110, 266, 1027, 900, 649, 919, 879, 1337, 578, 1799, 908, 1896, 511, 1837, 286, 1139, 1570, 679, 886, 84, 505, 613, 906, 576, 703, 1143, 691, 193, 1671, 1672, 254, 750, 94, 496, 881, 364, 1495, 546, 1398, 574, 96, 1757]. Designer [50]. Designers [284, 1377, 514].

Designing [1140, 875, 890, 563, 533, 1322, 254, 1894, 474, 884, 880, 888, 619, 882, 854]. Designs [728, 93]. Desktop [686]. desktops [1466]. Detecting [260, 1202]. detection [1870, 1840, 1841]. Determinants [281]. Develop [919]. Developed [250]. Developer [837]. developers [608]. Developing [1173, 920, 1809, 1819, 1814, 1397, 544]. Development [1090, 1598, 1385, 999, 1389, 970, 156, 837, 714, 579, 1451, 373, 987, 1704, 1630, 1414, 1399, 740, 1459, 1898, 1617, 727, 1403, 361, 801, 178, 459, 1411, 301, 1769, 1820, 572, 1709, 1727, 850, 61, 1529]. Developments [328, 1123]. device [1844]. devices [1857]. Dexter [617, 615, 613, 614, 616, 618]. Dexter-based [617, 615]. Dextrous [1045]. Dialogue [31, 110, 1797]. Dictionary [935, 272]. Did [1001, 1607]. Dienst [814]. Difference [800]. Differences [909, 802, 1903]. Difficulties [1171]. difficulty [754]. Diffusion [1362, 695, 128, 1417, 1645, 1656]. digit [46]. Digital [1367, 827, 931, 1134, 986, 1059, 1210, 1266, 1325, 1419, 1488, 1907, 1788, 1697, 820, 825, 1215, 160, 807, 1513, 1286, 169, 829, 67, 978, 1874, 196, 808, 1248, 224, 600, 806, 824, 823, 1226, 1768, 830, 1582, 1587, 1328, 818, 1510, 17, 168, 828, 1583, 1667, 1511, 1064, 1507, 1661, 1463, 342, 433, 822]. DigitalDesk [531]. Digits [260]. Dimension [725]. Dimensioning [383]. Dimensions [1224, 1179, 1228, 1225, 1078, 1231, 1230, 1229, 1226, 1227, 1162, 1613, 532, 650]. Directed [32]. Directing [1228]. direction [391]. Directions [1088, 232, 1706]. directives [88]. dirty [1647]. disabilities [315, 314]. disability [322]. Disabled [1358, 1247, 317]. Disaster [60]. discipline [55, 1450]. Discourse [1527, 754, 472]. Discovering [852, 549]. Discovery [1152, 1158, 1521, 1918, 1912, 1916, 704, 1919, 1915, 1914, 1782, 1913, 1910]. discrete [104, 103]. Discrimination [153, 768]. discuss [1205]. Disease [1350]. Disinformation [1149]. Disorders [1351]. displays [1862]. dissemination [1522]. Distributed [997, 937, 1640, 1252, 814, 1517, 331, 1540, 1163, 1878, 1423, 1034, 242, 1538, 1140, 558, 1880, 592, 1567, 1443, 11, 1883, 145, 1405, 1924, 20, 1725, 1784, 1865, 354, 1143, 1521, 557, 1512, 564, 1841, 127]. distributed-memory [354]. Distributing [1032]. distribution [1461, 1619]. disturbance [201]. Disturbances [1356]. Diversity [769, 1796]. divisive [1808]. divorce [1621, 1932]. Do [1467, 142]. DOCASE [564]. Doctorates [135, 1387, 1673]. Doctrine [553, 600]. Document [1597, 814, 1386, 240]. Documentation [191]. Documents [1705, 1516]. DoD [651]. Does [736, 1293, 583, 1680, 1827, 1624]. Doesn’t [1112]. dollars [1611]. domain [1402, 1895]. domain-specific [1402]. Domains [1426]. Dominance [1888]. done [1804]. Don’t [1810, 1764, 1533, 266, 1293, 1837]. double [1847]. double-edged [1847]. DOWL [558]. Downs [1267, 455]. Dragging [1116]. Drawings [1101, 383]. Dream [1098, 1053, 1082, 170]. drive [1636]. Driving [1049, 1939, 318]. Dropouts [1287]. DSP [1557]. DSS [1923, 1557]. Duly [1850]. Duplicate [347, 155]. during [1437]. DVI [268]. Dynabase [623]. dynabases [327]. Dynabook [291]. Dynamic [893, 20, 1300, 1303, 1302, 1538]. dynamically [1376]. dynamics [608].

e-commerce [1876, 1889, 1753, 1745]. E-education [1658]. E-Guide [1089, 983]. Early [1427, 979]. Earning [1831]. ease [1798]. East [350]. Easy [912, 1233]. EC [40, 88, 68, 39]. Ecash [1064]. Eckert [896]. Economic [111]. economics [1740, 1877, 1609, 1614]. Economy [968]. EDGAR [1591]. edge [1770]. edged [1847].

Editing [751, 815]. Editor [1198, 160, 57, 353, 520, 1415, 1373]. Editor-in-Chief [1415, 1373]. Editorial [1486, 1576, 1603, 1676, 1762, 1786, 1806, 1829, 1848, 1867, 1886, 1905, 1926, 952, 10, 172]. Editors [32]. EDR [936, 935]. Educating [417, 605]. Education [1237, 1357, 1681, 1215, 1669, 1065, 1352, 708, 1000, 467, 849, 649, 930, 209, 469, 468, 567, 611, 757, 1017, 1058, 1293, 1471, 476, 1464, 543, 1658, 1018, 488, 786, 142, 1605, 1853]. Educational [1286, 1372, 1561]. educators [1638]. EDUCOM [169]. EER [16]. Effective [153, 743, 1619, 1824]. effectiveness [1880]. Effects [15, 1184]. efficiency [1903]. Efficient [43, 1950, 1946, 1125]. Effort [827]. Efforts [261]. Eiffel [94]. eight [608]. Einschrittcompilers [1949]. Elections [1056, 123, 1936]. Electric [1213]. Electron [331]. Electronic [1091, 149, 199, 265, 377, 445, 580, 648, 834, 950, 1173, 1380, 1331, 1329, 1065, 1070, 1589, 874, 1042, 1623, 1472, 1715, 194, 255, 1060, 280, 420, 930, 935, 1679, 1296, 1610, 170, 726, 150, 831, 1467, 1064, 1655, 1590, 1525]. electronics [129, 1780]. Element [248]. Elements [304, 426]. Elephant [492]. Eli [287]. eliminate [1124]. else [1638]. Email [804, 1266, 1873, 58, 1502, 1845, 872, 956, 955]. Embed [891]. Embedded [926, 143]. Embedding [1705, 1391]. Emergence [160, 679]. emergent [1864]. Emerging [367, 1813, 1072, 623, 1610, 1495]. emotion [689]. Empirical [1764, 333, 126]. Employee [281, 1845]. employing [1556]. Employment [1387, 1673, 1081]. empowering [604]. enabled [686]. Enabling [690, 654]. encapsulation [371]. Encoding [1507, 166]. Encrypted [144]. Encryption [326, 1005, 233, 341]. encyclopedias [815]. End [864, 612, 865, 869, 1634, 595, 310]. End-User [864, 865, 869, 595]. Ending [1297]. enemy [1935]. Energetic [835]. Enfertainment [945]. Enforce [1076]. Engine [1236]. engineer [417]. Engineering [944, 500, 1128, 1657, 1150, 1430, 1240, 491, 1390, 1303, 27, 651, 1822, 1475, 641, 352, 1621, 383, 302, 175, 1902, 505, 1555, 1614, 18, 395, 870, 1251, 652, 1686, 1422, 853, 650]. Engineers [646, 1767, 1933]. Engines [835]. English [1949, 934]. Enhance [1129, 476, 1592]. Enhanced [545]. Enhancements [1768]. Enhancing [1707, 134]. Enigma [1371]. enough [607]. Entering [982]. Enterprise [1381, 732, 1725, 1620, 1922, 1483, 374, 1649]. Enterprises [991, 1741, 1686]. Enterprising [848]. Entirely [1873]. entity [911]. entity-relationship [911]. Environment [927, 421, 1425, 920, 331, 332, 168, 1545, 1125]. Environmental [523]. Environments [1357, 1384, 1199, 885, 187, 1351, 520, 1354, 1707, 333, 1687, 852, 530, 329, 881, 1668, 786, 1543, 854]. Envision [817]. epic [1852]. Epilogue [451]. EPSS [1334]. Equal [504]. Equation [281]. Equity [849, 486]. equivalencies [84]. era [450, 1890, 310]. Error [1183, 260]. Error-based [1183]. Error-Detecting [260]. Errors [1202, 382]. Eschew [56]. Escrow [1005, 1004, 1007]. Esprit [1931]. essential [1143]. establishing [40]. estimating [282]. estimation [107, 1760, 1560]. Ethical [1172, 960, 954, 1084, 1388, 956, 1010, 607, 513]. Ethics [953, 1430, 1197, 957, 197, 518, 121, 323, 442, 1795, 1902, 1637, 508]. Ethnocritical [907]. Ethnography [1337]. EU [1518]. Europe [1537, 37]. European [38, 1572, 36]. EV [439]. EV-4 [439]. Evaluated [1061]. Evaluating [1263, 974, 1084, 1051, 1651, 1615, 1870, 1840]. Evaluation [131, 156, 739, 59, 1201, 308, 866, 345, 886, 22, 379]. event [104, 103, 1299]. Events [1037]. Ever [1233, 1082, 1559]. ever-widening [1559]. Everything [1534]. evidence [1839]. Evil [1389]. Evolution [1778, 693, 368, 268]. Evolutionary [1915, 636]. Evolvable [1777, 1775, 1774]. evolved [1780]. Evolving [1908, 1330, 1779, 1712, 462, 1543]. exact [357]. Examining [195, 1479]. example [735]. Examples [1027]. exceed [1827]. Exception [278]. Exclusion [137]. Executive [1294, 1366, 770, 444]. executives [1907]. Expectations [720]. Expecting [181]. Experience [394, 684, 1186, 1280, 332, 1142, 573, 1794]. Experiences [132, 917, 502, 712, 127, 1404]. Experiment [645, 441, 818, 178]. Experimental [646, 30]. Experiments [1779]. Expert [593, 105, 378, 30, 1554, 739, 591, 309]. exploitation [860]. Exploiting [305, 1875]. explore [1874]. Explorer [18]. Exploring [906, 870, 1860]. Explosion [1243]. export [1663]. Exportability [233]. Exposure [1347]. Express [1601]. expression [1896]. Extended [539]. extendible [64]. Extending [374]. extension [1536]. Extensions [230, 700]. extra [1410]. Extracting [1153, 119]. Extraction [975]. Eye [1109]. eyes [618].

FAA [1943]. Fab [1260]. Faces [1243]. faceted [1688, 180]. facilitate [1104]. facilitation [503]. Facilities [1247, 1459]. Factor [645, 836, 1565]. Factories [649]. Factoring [356, 242]. Factors [1685]. Faculty [1011, 1459]. fail [734, 220]. Failure [845, 1909, 723]. failures [737]. Fair [553, 600, 694]. family [669]. famine [75]. famous [1533]. Farewell [310]. Fast [311, 1, 13, 244, 1178, 1948, 246, 63, 247, 245, 85, 382, 336, 1573]. father [55]. Fault [1200, 1030, 145, 1143, 1760, 1125, 1761]. Fault-Tolerant [1200, 1030, 145, 1143]. faults [1125]. Fax [58]. Fear [1222]. Feasible [443]. Feature [167, 996].

Feature-Based [996]. features [1650]. Federal [1133, 1601]. FedEx [998]. FedStats [1514]. Feel [44, 455, 360]. fellows [1489]. Fellowship [761]. Fernando [221]. few [1117, 100]. Fewer [1244]. FidoNet [542]. Field [1810, 441, 1170, 1776, 478]. Field-programmable [1776]. fields [255]. Fifth [295, 449]. figures [1788]. File [12, 336, 20]. files [749, 620]. filing [1621]. Filling [1272]. Filtering [413, 1259, 1262, 412, 408, 414, 407, 1925, 416]. filters [411, 410]. Final [448]. Finals [981]. Financial [1051, 375]. Finding [728]. findings [1765]. Fingers [639]. Finland [1731]. Fire [707]. Firewalls [1307]. firm [1789]. First [292, 566, 844, 140, 858, 859, 1462, 258, 183]. first-course [858, 859]. first-year [1462]. Fisheye [752]. fit [1802]. Five [916, 1205]. Five-Year [916]. Flame [598]. flexibility [839]. Flexible [287, 1033, 1801]. Flogging [1535]. flowless [1573]. Fly [1016]. flying [1930]. focus [984, 1800, 1734]. Fold [24]. Followed [259]. Following [1774]. Folly [1850]. Force [189, 488]. forces [1939]. Forecast [916]. Forget [1810]. Form [566, 286]. form-based [286]. Formal [1705, 1087, 1186, 727, 1447, 178]. formalisms [857]. Former [568]. formulas [384]. formulation [370]. Forsees [235]. Fortran [355, 354, 519]. Fortress [37]. Forum [1270, 1438, 1374, 1681, 1647, 1581, 1491, 1626, 1474, 1509, 1660, 1831, 1695, 1810, 1850, 1716, 1764, 1928, 1888, 1909, 1733, 1869, 1269, 1870, 1271, 1252, 1253, 1273, 466, 1268, 1249, 1274, 1275, 1272, 1250, 1251, 1248, 518, 1393, 1276, 1136, 25, 602, 404, 388, 278, 34]. Forums [1431, 803]. forward [1435]. Fostering [1283]. foundation [465, 175, 256]. Foundations [1162]. founding [55]. Four [645, 358]. Four-Factor [645]. Fourth [909, 725, 982]. Framework [960, 367, 1051, 1899, 378, 1401, 1404, 1753, 1856, 1670, 1397, 1645, 1398, 1528]. Frameworks [1491, 1396, 1400, 1405, 1395, 1403, 1620, 1894]. Framwork [1399]. framworks [1402]. Franz [979]. Fraternity [37]. Fraud [359]. Free [522, 1773, 781, 234, 453, 1698]. Free-Hand [522]. Freedom [1764]. FreeNet [1336]. Friendly [707]. Front [437, 1731, 328, 1732]. Frontier [199, 265, 377, 445, 580, 648, 834, 950, 29, 1220, 874, 1042, 1715, 149, 1623, 1472]. frontiers [309]. Fugue [1947]. Full [1282, 1802, 166, 1519]. full-motion [166]. full-text [1519]. Fun [1099]. Function [1106, 1107, 970, 441]. functional [211, 911]. Functions [1386, 273, 230]. fund [1064]. funding [256]. furtive [1315]. fusion [252]. Future [1237, 1218, 1241, 1242, 1243, 1234, 1240, 1236, 291, 1233, 849, 1238, 1232, 1204, 1239, 1115, 389, 1235, 1941, 1323, 334, 175, 486, 1698, 391, 489, 786, 1460]. Fuzzy [631, 632, 513]. FY [66].

gains [1490]. gambling [1485]. Game [1101, 18, 1042]. Gammon [800]. gap [1551]. Garage [1218, 1734]. Garbage [1181]. Gartner [1418]. gate [1776]. Gathering [836]. GBB [1302]. GemStone [228]. Gender [768]. General [1056, 196, 621, 1523, 278]. general-purpose [1523]. generalize [292]. Generalized [5]. Generating [32, 319]. Generation [295, 708, 909, 456, 449, 209, 1017, 229, 187, 225, 1882, 1714, 1759, 1067]. Generator [13]. Generators [776, 1178, 534, 777]. generic [344]. Genetic [636, 1180]. GENOCOP [1180]. Genome [238, 1918]. Geographic [270, 1288, 1782]. Geometric [1046, 774, 995]. GeoRoute [270]. Gerasch [146]. Germany [772]. gesture [318]. Gestures [522]. Gets [1171]. getting [1884]. Ghosts [112]. Gifts [708, 757]. Gigabit [1335]. gigabyte [1860].

Giving [318]. glass [767, 1850]. Global [695, 429, 1417, 968, 883, 543, 1656]. Globalization [130]. Globalizing [1237]. go [142, 18]. Goal [1850, 1701, 1019, 1082, 1546]. Goal-Based [1019]. Goal-Oriented [1701]. goals [866, 784, 1376, 602]. Goes [1040, 1328]. Going [829]. Gold [1159, 1494]. gone [1315]. Good [1266, 1732, 607, 1117, 1748]. goods [217, 1467]. Gordon [744]. Government [693, 1661, 1938, 1116]. Governs [1342]. GPS [1782]. GPS-based [1782]. Grace [1368, 766, 310]. gradient [107]. Graduate [761]. Graham [1541]. grand [640]. Granting [1333]. Graph [1182, 657, 549]. graph-based [657]. Graphic [503]. Graphical [749, 457, 456, 752, 851]. Graphics [1100, 193, 1859, 1855]. Graphs [645]. Great [265, 290]. Grid [1426, 1424]. grossed [1862]. grossed-beam [1862]. Group [1094, 1091, 1035, 1030, 194, 1029, 1034, 725, 1033, 592, 686, 1113, 1418]. group-enabled [686]. GroupLens [1262]. Groups [760, 1734]. Groupware [132, 608, 861]. Growing [1864, 1206, 1543]. Grows [1190, 138]. Growth [697, 459]. guaranteeing [1683]. Guard [648]. Guard-All [648]. Guest [1198, 160, 353, 520, 948]. Guide [494, 339, 1936, 493, 750, 1089, 983]. guidelines [1400, 282]. Gulf [1490].

h [1000]. Habanero [1571]. Hackers [1764]. hacking [946, 150]. Hairiest [1279]. Haiti [1820]. Halpern [1373]. Hand [522]. Handling [1202]. hands [1853]. Hanging [1074]. Happen [1230]. Happenings [540]. Hardware [1778, 1465, 111, 1214, 272, 1777, 1775, 1774]. Harmful [844]. Harmonic [1185]. harmony [88]. hash [273]. Hashing [244, 246, 247, 245, 64, 63]. Hassles [872]. having [1907]. Headed [239]. Health [1633, 262, 1359, 1361, 1348, 1942, 1346]. heart [877]. heartburn [1623]. Hedonic [596]. Heeks [1851]. held [1368]. Helf [1474]. Help [481, 1391]. helps [1708]. heritage [1512]. Hermeneutic [1289]. Heroes [1891]. heterogeneous [11, 1539]. Hierarchy [73]. High [1198, 1837, 1031, 169, 1135, 490, 788, 1019, 640, 436, 568, 334, 1124, 1646, 573, 1664]. High-Assurance [1198]. High-confidence [1837]. High-Performance [1135, 436, 568, 573, 1664]. High-Speed [169, 788, 1124]. higher [1464]. Highlights [80]. highly [1758]. Highway [1022]. Hill [540]. Hindsight [1581]. History [664, 1234, 554, 542]. hitch [1491]. hoax [1715]. Holist [222]. HoloSketch [1047]. Home [1171, 1169, 1861, 1167]. HomeNet [1170]. honor [1790]. honored [1292]. Hope [1222]. Hopes [1244]. Hopper [1368, 766, 310]. Horizon [467]. Horizons [932]. Horse [863, 1885]. Horses [1580]. Horus [1033]. Host [701]. house [1535, 1367]. HPC [1663]. HPCC [235, 581, 863]. HTML [986, 1901]. Hubris [1693]. Human [1213, 1221, 1217, 1218, 1215, 876, 767, 1216, 156, 836, 1220, 461, 31, 248, 1214, 1343, 1222, 1017, 1219, 1223, 1918, 238]. Human-Computer [156, 31, 1017]. Humane [1233]. humans [1858]. Hurdles [1287]. Hyper [889]. hyperdocuments [893]. Hyperform [892]. HyperIntelligence [29]. Hypermedia [1304, 886, 613, 878, 885, 879, 884, 882, 892, 875, 617, 615, 267, 616, 815, 877, 240]. hypermedia/time [240]. hypermedia/time-based [240]. hyperspace [619]. Hypertext [966, 881, 890, 614, 887, 888, 28]. HyTime [240].

I-WAY [1424, 1494]. i750 [165]. IANA [1682]. IBM [1949, 1629, 1949]. IBM/360 [1949, 1949]. ICANN [1816]. Ideas [117, 633]. Identifiers [464]. identifying [1670]. Identity [19, 1340]. Identity-Related [1340]. IEEE-CS [189]. If [1190, 1112]. II [1009]. III [70]. ill [205]. ill-structured [205]. Illinois [824]. I’m [1533]. Image [67, 1882, 1356, 1856, 1854]. images [751, 167]. Imaging [1051, 331]. Immediacy [1280]. Immersive [1423]. immunology [1406]. Impact [1242, 294, 1268, 133, 1680, 1483]. Impacts [206]. impairments [319]. Implementation [1179, 1949, 92, 1129, 22, 1937, 1757, 1949]. Implementations [13]. Implementing [1883, 990, 534, 1174, 180, 1125, 563, 297]. Implications [1208, 546]. Importance [1388]. Improve [1362, 204]. Improved [10, 584]. improvement [1318, 577]. Improvements [788, 1319, 1620]. Improving [1810, 31]. In-Memory [12]. Incident [705]. Including [1184]. incorporating [727]. Increasing [1798]. incredible [1408]. incremental [9]. Independent [745]. index [962, 961]. Indexing [1271]. India [1818, 1662, 492]. Indian [49]. indistinguishable [1854]. Individual [1181]. induction [1488, 938]. Inductive [939, 1913]. Industrial [937, 1381, 932, 1627, 1186, 1777, 625, 1545]. Industrial-strength [1627]. Industry [924, 1120, 262, 184, 925, 634, 1313, 767]. industry/academe [1313]. inevitably [1808]. Inference [1154]. Influence [1083, 182, 869]. influences [1148]. influencing [1685]. Info [778]. Informal [1325, 423]. Informatics [238]. Information [1360, 1013, 1036, 1209, 408, 1059, 1210, 1217, 975, 1564, 1888, 810, 832, 1385, 1050, 524, 413, 1164, 444, 1305, 1080, 991, 119, 985, 1268, 72, 1609, 1243, 237, 270, 954, 976, 1114, 1022, 407, 1073, 957, 816, 959, 1151, 1079, 674, 1077, 1361, 833, 458, 420, 968, 1288, 1081, 1811, 40, 412, 1082, 594, 1443, 131, 1606, 1691, 176, 1617, 36, 798, 414, 1444, 1440, 133, 1588, 1439, 813, 1591, 1494, 1482, 1518, 1264, 22, 1900, 409, 678, 1512, 1847, 1836, 1942, 1526, 1359, 416, 183, 148, 1624, 1675, 1539, 521, 411, 410, 803, 171]. Information-Filtering [413, 416]. informational [1563]. informations [1688]. informed [1540]. Informedia [820]. Infowar [1833]. Infoware [1214]. Infrastructure [1380, 1575, 933, 1073, 1427, 901, 483, 1559, 1817]. infrastructures [1813, 1469]. Inheritance [719, 1692, 1410]. Initial [259]. Initiatives [1522, 1465, 826]. Innovation [1232, 253, 1646]. Innovations [848, 1235]. Innovative [927]. Inoculating [1838]. Input [317, 319, 1857, 318]. Inquiry [1293]. Inroads [1066]. insects [1254]. Insecurity [86]. Insertion [146, 2]. insiders [1945]. insights [313]. Inspection [24, 1478, 584]. Inspections [1810, 1765, 1805]. inspired [1254]. Insurance [924, 1094, 1369]. Intangibles [1810, 403]. Integer [30]. Integers [242]. Integrated [714, 1879, 577, 1462, 1922, 687]. Integrating [1301, 1704, 593, 211]. Integration [711, 1899, 216, 1690, 1553, 574]. integrative [1322]. integrator [295]. Integrity [1411, 1549]. Intel [1193]. Intellect [876]. Intellectual [1003, 746, 968, 328, 1732]. Intelligence [517, 404, 1196, 349, 1863]. Intelligent [1100, 188, 202, 995, 676, 1020, 319, 1672, 1503, 810]. intensive [1667]. intentional [501]. Intentions [281]. interact [682]. Interacting [531]. Interaction [808, 1017, 1306, 1138, 426, 1921]. Interactions [457]. Interactive [169, 457, 331, 590, 458, 1169, 1374, 1752, 642, 167, 502]. Interconnection [23]. interest [609, 624]. interests [549, 411]. Interface [1103, 462, 156, 1233, 266, 319, 683, 215, 158, 1706, 1528]. Interfaces [456, 110, 1223, 1639, 1900, 460]. internal [2]. International [661, 897, 1490, 1662, 951, 835, 1613, 1342, 250, 130, 184, 208, 612, 695, 756, 1164, 1074, 1809, 1363, 401, 300, 1872, 390, 1016, 492, 1245, 432, 568, 913, 981, 1205, 1815, 288]. Internet [732, 1146, 1237, 1818, 1068, 1610, 1062, 704, 1409, 1490, 1662, 1066, 1473, 1724, 1815, 1718, 1764, 1909, 1128, 1362, 1611, 1497, 683, 705, 1342, 764, 695, 1065, 1544, 1168, 693, 1591, 545, 1170, 1167, 1680, 1672, 696, 1234, 1135, 1562, 874, 1816, 1813, 1485, 1708, 1364, 1069, 1307, 1874, 1820, 1814, 590, 731, 794, 1417, 1645, 1060, 1656, 1493, 1654, 894, 1293, 1435, 1817, 1498]. Internetworking [473, 541, 787]. Interoperability [1511, 179]. interpersonal [1903]. interpret [114]. Interpretation [907]. Interpreting [237, 171]. Interspace [824]. intervention [1724, 946]. Interview [221, 427]. Intranet [1041, 1304]. Introducing [562]. Introduction [555, 1198, 569, 160, 807, 251, 795, 103, 836, 541, 353, 493, 1278, 456, 90, 932, 758, 922, 449, 45, 1565, 469, 1104, 433, 520, 292, 70, 210, 1510, 662]. introductions [1137]. Intrusion [1841, 1870, 1840, 1839]. Intuitive [1825]. invaluable [1559]. Investigating [15]. Investigation [1079, 480]. Investment [933, 1338, 1477]. investments [131]. Invisible [648, 1098]. involving [1691]. IP [1067]. IQ [876]. IRIS [267]. ISDN [422]. isn’t [851]. ISO [1002]. Israel [661]. Issue [555, 1888, 541, 90, 469, 433, 520]. Issues [132, 1385, 957, 1166, 1553, 782, 529, 1138, 322, 886, 615, 461, 1348, 513]. Iterative [703]. ither [1000].

J [55]. Jack [377]. Jaguar [897]. Jam [1869]. Japan [284, 944, 635, 377, 599, 548, 39]. Jasper [476]. Java [1569, 1571, 1468, 1568, 1903, 1565, 1566, 1749]. Java-based [1749]. Jim [1947]. Jini [1843]. job [802]. Jobs [1215].

Joe [1373]. Joint [498, 189]. José [221]. journal [984, 1373]. Journals [1070, 1520]. JPEG [162, 161]. JTEC [599]. July [41]. junkyard [1494]. jurisdiction [1679]. Just [1332, 1119, 1546, 1785]. Just-in-Time [1332]. justice [1472]. Justification [157]. Justifying [1331].

K-12 [488]. K12 [543]. Kansas [1284]. Kasparov [1086, 1343, 1344]. Katz [1040]. kbit [164]. kbit/s [164]. KDD [1153]. keeps [1206]. Key [1005, 1004, 1286, 1022, 1007, 1365, 1058, 1006]. Keyboard [1227]. keystroke [19]. Kicking [1116]. Kicking-and-Screaming [1116]. kid [1734]. Kids [1024, 849]. KidSim [681]. kits [527]. Kiwis [1016]. Know [1112, 266, 1638]. Knowledge [78, 1919, 1332, 1269, 1362, 1152, 1153, 599, 526, 1914, 1705, 626, 335, 815, 1158, 933, 1910, 1023, 369, 352, 176, 18, 370, 1913, 1709, 472, 200, 1539, 574]. Knowledge-Based [599, 626, 526, 815, 176, 370, 200]. knowledge-building [472].

Labelling [1182]. LabNetwork [471]. laboratories [480]. Laboratory [331]. Landmark [1093]. landscape [1792]. Language [940, 1949, 424, 974, 909, 976, 745, 589, 45, 325, 977, 972, 558, 78, 973, 92, 562, 510, 240, 416, 303, 681, 79, 1569]. Language-Independent [745]. Languages [843, 519, 69, 285, 1542, 49, 379]. Laptop [1462, 1457]. Laptops [1602]. Large [933, 242, 1399, 1443, 273, 620]. Large-Scale [933]. LaSSIE [176]. lasts [577]. latencies [19]. Latest [1123]. Latter [1203]. Launching [450]. Laurence [819]. Law [1255, 1184, 224, 280, 873, 780, 1507]. Laws [1850, 637]. Lawsuits [44, 360]. Lax [723]. Lead [1001, 1767]. leadership [1772]. leading [1205]. Leaking [1050]. Leap [338]. Leap-year [338]. learn [1254]. Learned [918, 1909, 1898, 1405, 366, 1437]. Learner [1028, 1018]. Learner-Centered [1028, 1018]. Learning [864, 1025, 1026, 1065, 1378, 492, 929, 1027, 900, 477, 800, 485, 938, 828, 684, 1911, 1104, 658, 1463, 122]. LearningWorks [1403]. Learns [685, 209]. Lebanon [1717]. Lecture [741, 1052, 724, 426, 1053]. LEDA [774]. Legacy [1690, 1301, 651, 1755, 1555, 654, 653]. Legal [1165, 956, 729, 1475, 946, 1584]. Legally [1679, 114, 76, 224, 183, 148, 280, 553, 420, 455, 600, 746, 694, 806, 873, 968, 1165, 1075, 1297, 1624, 1871, 1696]. Legion [1194]. legitimacy [1551]. LEGOs [1888]. lemma [453]. Length [244, 246, 247, 245, 2, 63]. Less [250]. Less-Developed [250]. Lessons [918, 1909, 1598, 1405, 1591, 366, 1437, 94, 1298, 1769, 723, 1336, 1820, 670, 1344, 672]. Letter [56, 948, 87, 234, 386, 913, 965, 466, 137, 779]. Letters [57, 326]. Level [1079, 604, 179]. leverage [216]. Lexical [934]. Lexicons [973]. Liability [1850, 420, 1369]. Liberal [1175]. Librarie [1525]. Libraries [792, 807, 814, 829, 806, 828, 94, 1667, 1511, 1507, 830]. Library [1367, 827, 825, 1192, 808, 821, 824, 823, 1768, 1328, 818, 1510, 1521, 1519, 827, 820, 826, 822]. License [1474]. licensed [1391, 1624]. Licensing [1810, 1888, 1933, 736, 1767]. lies [1833]. Life [1094, 199, 206, 1025, 485, 945, 1401, 98, 348, 1547, 1855]. Life-long [485]. Lifelike [945]. Light [431]. Likelihood [107]. Limiting [1850]. Limits [1888, 1240, 1501, 459]. line [984, 1620, 119, 105, 109, 1757]. Linear [1180, 155]. Linguistic [1057]. link [1593]. linking [605]. Links [236, 837, 419, 1376, 1596]. LinkWinds [642]. Linux [1850, 1770]. LISP [210, 218, 217, 1304, 214, 215, 212]. LispView [216]. List [763, 1238]. Lists [62]. Lit [598]. Literacy [1113, 1514]. Literate [1951, 33]. Literature [817, 768].

Litigation [929]. Little [835]. live [1328]. living [1695, 1646]. Lobby [1550]. local [883, 1819, 1516, 1938]. location [521]. locking [1144, 1844]. Loebner [672]. Log [708, 849, 649, 930, 209, 142, 468, 567, 611, 757, 1017, 1058, 1293, 1605, 1853, 749]. Logic [1325, 939, 295, 294, 1231, 293, 797, 69, 1913, 296, 632]. logical [517, 1573]. LogP [1161]. Long [697, 868, 1471, 389, 1679, 485, 1578]. long-arm [1679]. Long-Term [697, 868, 389, 1578]. longer [1455]. Look [1145, 829, 44, 455, 1235, 495, 360]. Looking [516, 1233, 1545, 851]. looks [982]. loop [1819]. Loops [356]. Losses [375]. Lost [1341]. lot [1771]. Lotus [1392, 114]. Low [1560, 604]. Low-cost [1560]. low-level [604]. lower [1467]. LPWA [1721]. ly [429]. Lycos [1742].

M [687]. MAC [113, 1103]. Machine [425, 1187, 402, 1911, 938, 463, 49, 122, 582]. Machines [380, 1107, 284, 154, 354]. Made [1255, 1542, 174]. Madefast [1128]. magnetic [1314]. Maiden [931]. Mail [545, 170]. Mainframe [926, 1636]. Mainstream [1339, 1623]. mainstreaming [1420]. Maintaining [1646, 234]. Maintenance [920, 587, 307, 1461, 61]. Majors [1764]. make [1546]. makes [770]. Making [1228, 747, 1025, 818, 134, 576, 904, 1084, 861, 1859, 905, 442, 523, 1825, 1471, 662]. malcontents [1562]. Management [398, 206, 228, 569, 1379, 787, 846, 1365, 1386, 229, 306, 1783, 1597, 272, 1740, 283, 1633, 570, 571, 1322, 1572, 95, 1877, 1653, 722, 1439, 1639, 1900, 282, 847, 666, 803, 1480]. Manager [72, 205, 215]. Managerial [1628]. Managing [866, 1444, 367, 955, 1783, 1548]. Manufacture [994]. Manufacturing [991, 996, 995, 989, 990, 109, 988]. Many [1524, 650]. map [269]. Mapping [237]. Market [591, 1083, 40, 38, 88, 1264, 141].

Marketing [1494, 1493, 1495, 1874, 1498]. marketplaces [1610, 1467]. Markets [1065, 1130, 133]. Marriage [1390]. MARS [186]. Marvin [677]. mass [1497, 17]. mass-production [1497]. Massively [641, 242]. Master [136]. Match [1343]. Matching [1950, 1946, 1948]. math [476]. mathematical [906, 370]. mathematics [1110]. MathWorlds [1110]. matter [1785]. Maturity [1317, 1447]. Maurice [466]. May [1115]. Mayday [181]. maze [597]. MBONE [699]. McLuhan [862]. me [1533]. Mead [1412]. mean [1680]. Means [1119, 371, 1410]. Measurement [1108, 441, 306]. measures [772]. Measuring [51, 644, 1388, 585]. Mechanisms [1202, 92]. Media [421, 1218, 422, 1336, 709, 224, 1440, 703, 1623, 1583, 1500, 1104]. MediaView [196]. Medical [1383, 129, 786]. Medicine [1362, 1349]. medium [603]. MeDoc [1520]. Meeks [1375]. meet [1294]. Meeting [194]. meetings [726]. Meets [1353, 945, 862, 1366]. megaprogramming [400]. melding [797]. Melissa [1812]. Member [1418, 1206]. membership [322]. Memoing [296]. Memory [12, 352, 744, 354, 871, 656, 1125]. memory-based [656]. men [1637]. Mentoring [771, 770]. Message [1812, 1085]. Messages [1263]. messyware [1920]. Metacomputing [330]. Metal [213]. Metalevel [1596]. Metamanagement [1491]. Metaphor [1228, 754]. metaphorical [750]. Metasystems [1666]. Method [1183, 356, 561, 577, 46]. methodic [564]. methodical [509]. methodologies [658]. Methodology [1160, 1429, 335, 156, 878, 667, 308]. Methods [413, 1087, 740, 641, 739]. Mexico [1413]. MG [809]. Michigan [825]. Microcomputer [111, 47]. Microkernels [1127]. Microprocessor [1236, 665, 666]. microprocessors [1501]. Microscopy [331].

Microsoft [1320, 1542]. Middle [350]. middleman [1920]. Middleware [997]. Midnight [138]. might [682]. migration [20, 1652]. Military [1268]. million [1852]. Milner [427]. MIMD [354]. Mind [14, 1109]. Mine [1159, 1494]. Minimal [13, 155, 273, 2]. Mining [1155, 1152, 1157, 1154, 1156, 1917, 1911]. Minsky [677]. MIPS [352]. MIS [171, 1431, 281, 870]. Misappropriation [929]. Misleading [1831]. Missing [1276]. mission [1880, 1883, 125, 1878]. Misuse [1340]. Mixed [207]. MMX [1947, 1193]. Mobile [722, 1199, 1748, 521, 1749]. Mobility [701]. Model [997, 1161, 1317, 281, 490, 861, 1428, 879, 1175, 1692, 1447, 1542, 613, 614, 616, 370, 395, 1495, 911, 82]. Modeling [1569, 333, 1897, 1046, 1703, 1700, 515, 411, 306, 369, 365, 857, 361, 366, 368, 1690, 1321, 667, 374]. Models [16, 206, 1705, 1028, 1779, 1069, 1704, 1288, 713, 1539]. Modern [1542, 1599, 110, 1301]. modular [668]. modularity [186]. Molecular [237]. momentum [316]. Money [1064, 1471]. monitors [118]. Monumental [238]. Morocco [1473]. Mosaic [1643]. Motion [1108, 1109, 166]. Motives [1287]. motor [319]. Motorola [669, 668]. move [1636]. Movement [1351]. moves [1207]. Moving [1815]. MPEG [163]. MPP [1001, 1085]. much [1117]. Muenster [1949]. multeda [1397]. Multiagent [1738, 1742, 1735, 1739]. Multicast [7, 1030, 700, 699]. multidisciplinary [1876]. Multifarious [6]. Multilevel [177]. multilingual [1518, 1516]. Multimedia [1113, 421, 170, 971, 160, 1286, 165, 787, 1193, 196, 1020, 253, 1444, 163, 409, 1708, 477]. Multimodal [1879, 1921]. multiparty [1544]. Multiple [162, 1111, 1548]. Multiplicative [1178]. Multiprocessor [436, 81, 1125]. multiprocessors [8, 1125]. Multipurpose [1901]. multiset [428]. multithreaded [1651]. Munro [1375].

Münster [1949]. Murray [1368]. Museum [1382]. Museums [1071]. music [157]. Musical [1185]. Mutual [137, 1758]. My [872, 730, 1172, 1279, 1238]. Mysteries [112]. Myth [1831, 1255, 463]. myths [337, 1674, 1921, 1805].

N [1375]. Name [1660, 277, 269]. Names [464, 1608]. naming [510]. Napkins [1101]. Nation [235, 1445]. National [14, 233, 1818, 602, 391, 1151, 1336, 256, 1424]. Nationwide [1425, 1523]. Natural [940, 974, 976, 416, 79, 972, 78, 973]. Nature [724, 767]. Navigating [619, 28, 1663]. Navigation [1261]. Necessary [1389]. need [1436, 1871, 1834]. Needed [1491, 1376]. Needs [1357, 865, 1561, 1687, 1773]. Negative [1327]. neglected [895]. negotiate [1377]. negotiation [1547, 1747]. Neidorf [150]. Neil [1375]. nerve [219]. NEST [106]. Net [406, 915, 1508, 1835, 1929, 1735, 1930, 1228, 862]. Net.Speech [915]. Netizen [1936]. Netizens [1093]. Netscape [1898]. Network [106, 6, 544, 1082, 698, 1877, 1653, 1739, 1639, 784, 320, 543, 1942, 546, 1843]. network-centric [1843]. Networked [1737, 235, 329, 1675]. Networking [1335, 1124, 1878]. Networks [795, 1378, 169, 1259, 788, 987, 1133, 1083, 634, 58, 635, 1821, 1515, 789, 1914, 798, 288, 1550, 128, 796, 550, 7, 475, 633, 1712, 786, 603, 785, 782, 632]. Neural [635, 634, 1914, 633, 632]. Never [1297]. Never-Ending [1297]. New-Generation [187]. Newell [1412, 405]. News [1039, 1092, 1370, 1416, 1434, 1452, 1470, 1487, 1506, 1532, 1558, 1577, 1604, 1622, 1642, 1659, 1677, 1694, 1713, 1730, 1763, 1787, 1807, 1830, 1849, 1868, 1887, 1906, 1927, 169, 119, 1262, 1732]. Newspapers [169, 1874]. Newstrack [964]. Next [1291, 29, 225, 1191, 1067, 456, 1208, 1238, 1232, 1230, 1017, 229, 1212, 1043, 1714].

Next-Generation [229, 225]. NeXTstep [162]. Niche [1272]. Nicol [1415]. NII [746, 692, 624]. NIIIP [1014]. Nine [282]. Nintendo [209]. NIST [342, 343]. No [1343, 1472, 1455, 1471]. Nobody [191]. Nomadic [885]. non [1815]. non-U.S.-centric [1815]. Noncollaborative [1285]. Noncommand [460]. nonspeakers [317]. Normal [566]. North [71, 241, 951]. Not-So-Accidental [222]. Note [244, 246, 247, 245]. Notebook [1021]. Noted [1850]. Notes [311, 297, 192, 289, 1392]. Notion [719]. NSF [810, 1518, 1135, 1422]. NSF-EU [1518]. NSFNET [697]. NT [1041]. Number [13, 776, 1178, 534, 777, 1759]. Numbers [66, 1831, 928, 108]. Numerical [1180].

O [445]. Obfuscation [56]. Obituary [898, 122]. Object [555, 719, 710, 228, 372, 917, 999, 1395, 941, 335, 711, 91, 73, 920, 90, 714, 362, 1701, 921, 916, 120, 1027, 1704, 363, 919, 879, 519, 558, 97, 1148, 716, 1567, 563, 561, 1571, 556, 1640, 211, 1542, 1865, 98, 1300, 657, 559, 364, 1690, 713, 1619, 1651, 658, 96]. Object-Oriented [555, 719, 710, 372, 917, 999, 941, 335, 711, 91, 920, 90, 714, 362, 1701, 921, 120, 1027, 919, 879, 519, 1395, 558, 97, 1148, 716, 563, 561, 556, 211, 1865, 98, 657, 559, 364, 1619, 96]. object-sharing [1571]. Objectively [1509, 1432, 1389, 1784, 1865, 1620, 1484]. Objects [1256, 522, 1200, 1303, 539, 1302, 230, 1448, 396]. ObjectStore [227]. Observations [1001, 1265, 1681]. obstacles [1614]. oceanography [550]. ODA [1070]. odyssey [1893]. offer [1573]. Offers [1094, 1286]. office [530]. offices [895]. offs [1560]. often [1789]. old [1534, 100, 1605]. OMA [1649]. OMG [1649]. On-line [119, 984, 105, 1757]. On-ramp [1082]. Once [1375]. One [1949, 1392, 1832, 1471, 1435]. One-Step [1949]. Ones [1191]. ongoing [1889]. Onion [1720, 1626]. Online [812, 1197, 1070, 1071, 1934, 1723, 1443, 818, 1781, 1917]. Only [1608]. onto [1304]. Ontological [1162]. ontology [1754]. Onward [1324]. OO [918, 1491, 1810, 1405, 1924, 1784]. Open [1367, 1083, 1923, 252, 618, 168, 1729, 1769]. open-source [1729, 1769]. OpenVMS [435]. Operating [738, 1124, 1126, 389, 1123, 127]. operations [64]. Opinion [1755, 804, 1324, 1014, 844, 1309, 1364, 1339, 1289, 1692, 1728, 1674, 1711, 1827, 1504, 1804, 1903, 1884]. Opportunities [1585, 1614, 1872, 231]. Opportunity [406, 930, 504]. optical [1821]. Optimization [1758, 1180]. Orchestrating [570]. order [1377]. Organization [569, 570, 571, 470, 1392, 1376]. Organizational [1095, 597]. Organizations [1196, 1691, 1477, 133, 1864]. organized [1376]. Organizing [1379, 1802, 871]. orientation [396]. Oriented [555, 719, 710, 372, 917, 999, 941, 335, 711, 91, 920, 90, 714, 362, 1701, 373, 921, 120, 1027, 919, 879, 519, 558, 97, 1148, 716, 511, 563, 561, 556, 1395, 211, 1865, 906, 98, 657, 559, 364, 1619, 96]. origin [1771]. ORION [998]. OS/400 [918]. OSI [23]. OSPF [700]. Other [1101, 1169, 1205, 1678]. Our [1015, 846, 386, 1323, 533, 876, 1790, 847, 849]. Outlawing [1644]. outreach [1669]. outreaches [1413]. outsourced [1727]. Outsourcing [1080, 1079, 1081, 1077]. Overcoming [459, 1347]. Overload [1210, 678]. Overplanting [386]. Overview [1067, 164, 669, 283, 631, 1649]. Owns [950]. Oz [1284].

Pacific [984, 897]. paged [871]. pages [1500, 1861]. pains [1543]. palace [340]. palatable [174]. palms [1853]. Palmtop [524]. Panel [599]. Paper [1203, 531, 532]. Paradigm [91]. Paradigms [1102, 571, 1097]. paradox [594, 1612, 1613].

Parallel [1012, 1491, 1161, 334, 380, 104, 356, 1425, 921, 154, 239, 1339, 1201, 1129, 641, 1420, 51, 1144, 166, 643, 82]. Parkinson [1350]. Part [1025, 610]. participates [1414]. participation [506]. Participatory [839, 512, 603]. partnerships [1422]. party [1801]. PASCAL [1949, 1949]. PASM [243]. passing [1085]. Passwords [659, 647]. Past [1234, 291, 489]. Patching [1304]. patent [337]. Patented [76]. Patents [1277, 264, 1695, 1646]. path [2, 1774]. Patient [1363]. Pattern [986, 1950, 1946, 1143, 1639, 1948]. Pattern-Matching [1950, 1946]. Patterns [1145, 372, 695, 919, 1139, 1570, 1142, 1396, 1894, 1540, 1137]. Paul [1541]. Payoff [861]. payoffs [1726]. PC [26, 1890]. PCs [1193, 1789, 1115, 1869]. PD [497, 498, 495, 509, 499, 507, 493, 504, 494, 512]. peace [1472]. peeling [1626]. People [421, 778, 319, 314, 682]. Perceptions [955]. Percheron [863]. Perfect [715, 273]. Performance [1331, 1333, 1135, 788, 1338, 1083, 436, 568, 334, 1329, 573, 802, 1664, 51, 64, 1484, 644, 667, 358, 640]. Performance-Centered [1333]. Perils [1063]. Perlis [55]. Perplexed [339]. Persian [1490]. Persistant [1126]. Personal [72, 959, 1789, 65, 113, 192, 141, 159, 249, 291, 406, 554, 590, 467, 623, 675, 794, 862, 915, 987, 1041, 1208, 1326, 1417, 1714, 1890, 449, 499, 684]. Personalized [413, 1261, 409]. persons [315]. Perspective [1068, 941, 1158, 283, 1322, 1480]. Perspectives [661, 897, 951, 835, 1342, 250, 130, 184, 208, 612, 695, 756, 1164, 1074, 300, 390, 1016, 492, 1245, 432, 449, 101, 223, 198, 349, 290, 389, 519, 431, 568, 1490, 1662, 1809, 706, 1872, 1863, 124, 173]. Persuasive [1850, 1791, 1795, 1792]. Pervasive [1217]. Petaflops [1324]. philosophy [212]. PHOAKS [1258]. Phobias [1347]. Phoenix [12]. Physical [1353, 84, 1349]. physically [317, 1859]. Physician [1360]. Physicians [1496]. PICS [1146]. PICTIVE [504]. Picture [1312, 161]. Pile [1115]. Pioneer [122]. Pioneering [765]. pipe [528]. pipeline [1408]. piracy [1607]. Pitfalls [1063]. Place [385]. placement [269]. Plagiarism [1888, 952, 694, 1846]. Plain [580]. Plaintext [729]. Plan [1042, 1457, 831, 1151]. planning [572, 60]. Plans [1309, 1249, 656]. Plasma [921]. Platform [774, 1898, 1722]. platforms [1640, 1882]. Plowshares [661]. Plus [1206, 1205, 1418]. Pointers [1603, 1676, 1181, 1486, 1576, 1762, 1786, 1806, 1829, 1848, 1867, 1886, 1905, 1926]. Points [441]. poles [1314]. policies [597]. Policy [580, 1059, 1228, 236, 1869, 235, 1326, 660, 609, 1296, 602, 182, 1606, 718, 706, 1824, 1826]. Policy-Making [1228]. Political [288]. Politicizing [660]. Politicking [1428]. Politics [68, 1550, 1131, 908, 1134, 1248]. Polling [588]. Polychannel [17]. polygon [344]. poor [1483]. Popular [1764]. Population [1072]. Portable [1178, 291]. Porting [435]. Positions [1011]. Possible [1224, 792, 1053]. Post [22, 1890, 1223, 602]. post-Cold [602]. post-PC [1890]. Post-WIMP [1223]. POSTGRES [229]. Power [1327, 1374, 1736, 666, 1596]. Powerful [1909, 1306]. PowerPC [669, 665, 664, 663, 667, 668, 666, 662]. Pracnique [242]. Practical [1475, 1063, 272, 1161, 698, 273, 1150, 1371, 1191, 1267, 1678, 1578, 1852, 1808, 1765, 585, 102, 191, 174, 266, 376, 454, 539, 639, 357, 289, 309]. Practice [1150, 1700, 500, 1251, 1685]. Practices [494, 10, 1803, 1546, 1572, 1800, 906, 493, 22, 1796, 597]. Practitioner [494, 493]. Praises [1491]. precipitous [1920]. predict [1607]. predictability [204]. Predicted [223]. Prediction [316]. predictive [5]. Predictors [1934]. preference [644]. preferences [1722]. prefix [43]. preparation [1875]. prepare [1678]. preparedness [1943]. Preparing [1334]. Present [840, 291, 489]. presented [1124]. President [1812, 1076, 56, 948, 1891, 1699, 87, 234, 386, 779, 913, 965]. Presper [896]. Press [1764, 983, 1295]. Price [1089, 1274, 1467, 596]. Pricing [1364]. priesthood [1455]. primitives [1144]. Principles [533, 943]. print [1394]. printed [157]. priorities [1477]. Privacy [545, 1900, 1130, 1341, 959, 720, 1361, 956, 597, 955, 279, 1723, 1724, 1718, 548, 1715, 1722, 312, 546, 547, 1498]. Private [199, 789]. Probabilistic [62]. probability [797]. Probe [1106]. Problem [1888, 1026, 137, 1488, 655, 510, 74, 204]. problem-solving [204]. Problems [695, 1180, 704, 641, 1546, 298, 1899, 338, 1605, 513]. Procedure [843, 21, 53]. procedures [1376]. Proceedings [971]. Process [998, 365, 1432, 1389, 1153, 577, 848, 1799, 592, 1567, 1741, 366, 1484, 881, 301, 572, 303, 658]. Processes [775, 714, 1546, 1548, 1686]. Processing [940, 974, 976, 972, 78, 118, 1124, 336, 973, 1714, 1543]. Processor [165, 51, 124]. product [500, 1802, 1480]. production [1497]. Productivity [909, 1148, 594, 1612, 1613, 134, 830]. productivity-enhancing [134]. products [345, 1617, 686, 841, 1117, 308]. Profess [1010]. profession [1323]. Professional [723, 234, 323, 1369]. Professionals [147, 1551]. Professor [1040]. Profiling [136, 1574]. profit [1773]. Program [1094, 655, 1101, 1135, 76, 567, 1418, 1723]. programmability [186]. programmable [1776, 1503]. Programmer [1150, 1371, 1191, 1267, 102, 191, 266, 376, 454, 539, 639, 1678, 1578, 1852, 1808, 1765, 174, 173]. programmers [852, 1436, 81]. Programmiersprache [1949]. Programming [555, 719, 428, 1012, 939, 843, 1810, 1949, 1282, 294, 1951, 293, 1668, 325, 33, 1450, 97, 218, 981, 561, 69, 556, 30, 852, 1924, 295, 893, 211, 1542, 1784, 1865, 557, 657, 298, 1144, 559, 1913, 564, 1504, 851, 656, 1531, 681, 854]. Programs [1055, 761, 1176, 771, 259, 418, 673, 790, 77, 553, 1573, 1555, 93, 296, 136]. Progress [1614, 406, 1545, 650]. Project [398, 569, 1909, 825, 1070, 824, 885, 433, 1631, 570, 571, 1687, 1772, 295, 1908, 1591, 479, 572, 1918, 1751, 1931, 822, 918, 1188, 998, 238, 185, 821, 449]. Project-based [1631]. project-specific [1687]. projections [313]. Projects [1334, 1385, 1309, 1207, 495, 1647, 1578, 505, 1670, 1524, 826, 1727, 1574]. Prolog [70, 297]. Prologue [4]. Promise [916]. promotes [1514]. Proof [1182]. Property [950, 1003, 947, 746, 968, 148, 328, 1732]. proposal [1923, 343]. Proposed [1269, 342]. proprietary [1611]. pros [1139]. prospects [1082]. Protect [1075]. Protected [738]. Protecting [548, 1559, 1583, 1469, 312]. Protection [1165, 547]. protest [1414]. Protocol [701, 1923, 737]. protocols [783]. Prototypes [1118, 500]. Prototyping [106, 992, 639, 892, 506, 502]. prover [178]. provide [1369, 1436]. providers [1611, 183]. Providing [1183, 701]. Pseudo [777]. Pseudo-Random [777]. Psychometrics [1355]. Public [1591, 624, 609, 1606, 1519]. Publication [196, 1209]. publications [984, 1296]. Publishable [1203]. published [1291, 1295]. Publishing [1088, 1069, 794, 1697, 150, 831, 1901, 1525]. pumping [453]. purchase [1461]. purpose [1523]. Putting [162, 1924, 253, 158]. puzzle [340].

Qaulity [970]. Quagmire [1159]. Quality [1481, 1308, 1162, 1319, 1707, 1316, 585, 1317, 1318, 1482, 1483, 1615, 1479, 1480]. Quandary [956]. Quantum [1319]. Query [1183]. Quest [1468]. question [1438]. Questioned [581]. Questioning [1474, 1269]. Questions [907, 692, 855].

Queueing [1184]. Quick [142]. Quo [1775].

R [1121]. R&D [66, 1121, 1523]. R-D [1121]. Race [802, 1885]. Racks [376]. RADARSAT [1302]. RADARSAT-1 [1302]. railroad [395]. Rails [537]. ramp [1082]. Rampart [1032]. Random [13, 776, 1178, 108, 534, 777]. Randomized [1950, 1946]. Randomness [536, 535]. range [1802]. Rapid [994, 992, 892]. Rate [1094]. ratings [1131]. ratio [107]. Re [840, 1555]. Re-engineering [1555]. Re-Present [840]. Reaches [924, 1090]. Reactive [1384]. reactor [643]. Read [1229]. Readers [1509, 278]. readership [851]. readiness [1944]. Reading [1247, 468]. Reads [1429, 191]. Ready [1339, 1943, 1944]. Real [1256, 218, 201, 795, 1200, 1127, 125, 200, 520, 1858, 1860, 1139, 177, 523, 1568, 202, 1878, 1895, 1884, 1377]. Real-Time [1200, 218, 201, 125, 200, 1858, 1568, 202, 1878, 1895]. Real-World [795]. Realistic [380, 1856, 304]. realities [1726]. Reality [1049, 1569, 1048, 1046, 1364, 1350, 1383, 525, 170, 526, 1944, 532, 1878, 1854, 1044, 1377]. Realizing [1862]. Really [3, 1942, 1624]. Reasoning [995, 627, 1573]. reasons [1748]. Reassessing [805]. Rebuttal [152]. Receipts [1061]. receives [1412]. Rechenanlage [1949]. recodable [1844]. recognition [1540, 79]. recognizing [656]. Recommendation [1260]. Recommendations [1258]. Recommender [1263, 1257]. Reconsider [741]. Reconsidered [566, 518]. Records [1363]. Recoverable [1126]. Recovery [1199, 1006, 1842, 60]. Recruiting [769]. Redress [1681]. Reduce [1087, 678]. Reducing [204, 1759]. reduction [1575]. reengineer [654]. Reengineering [998, 1928, 1478, 1051, 1823, 1574]. refereed [1296]. reference [614]. Referral [1259]. Reflections [87, 74]. Reflects [1002, 1449].

Reform [1286, 568, 473]. Reforming [780]. Regarding [1850]. region [984]. regions [1819]. registry [1252]. Regressing [1545]. Regulation [1075, 1816]. regulations [1871, 1435]. Regulatory [959]. Rehabilitation [1349, 1352, 992]. Rehabilitative [1351]. Reinventing [1429]. Rejoinder [278]. Rekindled [355]. Related [1038, 1340, 1238]. Relational [16, 120, 652]. Relationship [1150, 911]. Relationships [1122, 1099]. Relative [1160]. Reliability [441, 126]. reliable [592]. relying [1904]. Remain [743]. Remark [791]. Remarks [326, 534]. remedies [1578]. Remedy [1433]. remember [1682, 1299]. remembers [979]. Remote [522]. Rendezvous [424]. Renewing [609]. replaces [893]. Replacing [1797]. Report [599, 812, 1014, 275, 237, 746, 232, 189, 99, 573]. Reports [811, 812]. repositories [1443]. Representation [747, 1185]. Representations [16, 904, 899, 748]. representing [908]. Republic [1474]. Request [1136, 1651]. requests [1527]. Require [736]. Requirements [1335, 836, 1684, 24, 1426, 1701, 658, 651, 525, 1437, 1685]. requires [1377]. Research [533, 1241, 1172, 1176, 391, 1243, 996, 362, 1000, 159, 649, 27, 263, 894, 581, 721, 1414, 1710, 704, 283, 1117, 870, 96]. Residential [1170]. Resource [763, 704, 1782]. Resources [764, 771, 856, 1521]. Response [1491, 705, 1253, 1275, 536, 232, 671]. Responses [936, 343]. Responsibility [1015, 207]. Responsive [530, 1354]. Restricted [670]. restrictions [1414]. results [585]. Resuscitation [208]. Retire [355]. Retrieval [1305, 976, 408, 809, 1443, 818, 798, 1526, 620, 1519, 810]. Retrieving [1441]. Retrospect [1226]. Retrospective [1188, 42, 495, 1557]. Return [986]. Returns [550]. Reusable [919, 1881, 1405, 560, 305]. Reuse [1246, 1148, 855, 1804, 180]. Reverse [650, 728, 651, 1475, 652]. Review [768, 346]. Reviewing [963]. Revised [1175, 1826]. revision [1871]. Revisited [1491, 291]. Revisiting [1313, 198]. Revolution [1236, 1226]. rhinos [1315]. Rhythms [903]. Rich [808, 1743]. Richard [1851]. riddle [1488]. Right [1229]. Righteous [1764]. Rights [1326, 968, 150, 1683, 183, 445]. Rim [897]. Ring [6]. Rise [922, 1920, 124]. Risk [190, 1944, 1575]. RISKS [299, 872, 190, 845, 1015, 1265, 793, 598, 832, 1621, 610, 622, 1290, 928, 1433, 947, 1114, 1073, 402, 588, 123, 277, 565, 429, 452, 551, 579, 647, 674, 637, 912, 963, 1149, 1163, 1057, 1038, 1177, 1087, 978, 1312, 1365, 1244, 1340, 1411, 1451, 1658, 1390, 1675, 1204, 778, 894, 1132, 1756, 129, 1670, 1693, 1505, 1485, 1602, 1641, 1847, 1729, 1828, 1925, 1945, 1805, 1712, 1904, 1885, 1866, 1785, 190, 1265, 610, 129, 402, 112, 123, 74, 222, 138, 147, 197, 261, 181, 248, 158, 375, 419, 359, 299, 385, 443, 515, 429]. RISKS [491, 537, 659, 720, 707, 755, 647, 674, 464, 207, 312, 692, 233, 729, 171, 1904]. RMM [878]. Roadblocks [1711]. Robin [427]. robot [1556]. robotic [202]. Robotics [993, 1254, 630, 1822]. Robust [1729]. Role [693, 1114, 491, 987, 1133, 638, 1610, 689, 1515, 973, 1709, 1727, 1525]. roles [828]. ROM [816, 1089]. Romania [208]. room [1264]. Rosenberg [819]. rot [1756]. Rough [756, 942, 1674, 1916]. roulette [1756]. routing [1720, 1782, 700]. Rubble [208]. Rule [1831, 269, 938]. rule-based [269]. rules [230]. runaway [1647, 1578].

s [1309, 164]. Sac [1001]. sacred [1790]. Safe [792, 12]. Safety [622, 947, 143, 59, 643, 1348]. safety-critical [59]. Sahara [951]. Saharan [432]. Salaries [1387, 1673]. same [408]. Sampling [1857]. Samuel [122]. Sandbox [1458]. Save [1042]. Saving [846, 1790, 847]. Say [950]. Scaffolded [1027]. Scalability [154, 1652]. Scalable [704, 582, 521]. Scale [933]. Scales [1111]. Scaling [27]. Scenario [985, 1019]. Scenarios [1703]. scene [1832]. Scheduling [356, 105]. Schemas [910]. Scheme [326]. Schemes [787, 1758]. scholarly [1525]. Scholars [1431, 803]. School [274, 490, 475, 849, 1019, 339, 52, 906]. schooling [470]. Schools [846, 483, 482, 489, 847, 474]. Science [1237, 1203, 236, 1474, 1241, 3, 136, 1242, 1379, 811, 764, 238, 724, 817, 1111, 430, 1243, 1234, 1240, 1387, 490, 1236, 1011, 1233, 140, 1238, 256, 1232, 232, 1239, 139, 753, 1175, 660, 1289, 634, 989, 1235, 1822, 641, 602, 55, 1621, 765, 391, 1728, 550, 1422, 1625, 1530, 1853, 1912, 988, 602, 529]. sciences [1463]. Scientific [1160, 1048, 1157, 1913, 1115, 234, 1737, 1522, 642, 480, 83]. Scientist [1009, 116]. Scientists [646, 605]. SCISOR [119]. Scope [1330, 232, 391]. Score [1185]. Screaming [1116]. scrutiny [1454]. Sea [804]. seal [1723]. seamless [703]. Search [311, 85, 2, 1440, 1549]. searches [1528]. Searching [1, 381, 1517, 1528, 382]. second [258, 80]. secondary [482]. Secret [929]. secrets [1647]. Section [569, 353, 45, 1479]. Sectors [184]. Secure [785, 521]. Securing [1062, 733, 1114, 1073]. Security [1290, 1008, 928, 720, 1307, 1204, 35, 198, 1837, 86, 1556, 1654, 1209]. Seducing [1634]. seductive [1794]. Seed [386]. Seeding [1133]. Seeing [611, 851]. Selecting [1144]. Selection [378]. Self [1474, 791, 53, 694, 1543, 302, 1391, 843, 21, 121]. Self-adaptive [1543]. Self-assessment [53, 302, 843, 21, 121]. Self-Helf [1474]. self-help [1391]. Self-Plagiarism [694]. Self-Stabilizing [791]. sell [1746]. Semantic [1526, 1539]. Semantics [910]. SEMATECH [1404]. semi [1556]. semi-autonomous [1556]. semiconductor [38, 440, 109]. Sense [77, 1630]. Sensors [1232, 1540]. separate [55]. September [1368]. sequential [562, 1760]. Serial [1129]. Series [1113, 476, 1540]. Server [1041, 346, 1546]. Service [1255, 1611, 735, 812]. Services [1091, 997, 1170, 892, 40, 990, 1436, 267]. set [1674, 1761, 1916]. sets [1860, 167, 297, 942]. settings [906]. Settlement [1061]. Seven [1748, 1309]. Seventh [274]. severe [319]. severely [317]. shaping [1606]. Share [1754]. Shared [1772, 1284, 990, 252, 1125, 549]. shared-memory [1125]. Sharing [991, 745, 1095, 589, 709, 638, 914, 781, 1002, 967, 1372, 1246, 1453, 1579, 1766, 83, 1258, 1571, 1143, 574]. Shells [378]. Shield [648]. Shift [1096]. shoot [1533]. shopping [1589]. Short [172, 1578, 717]. Short-term [1578]. shortcut [1453]. Should [76, 567, 1638, 1789, 396, 1594]. shrinking [1408]. side [969, 1845]. Sides [1325, 408]. sieve [289]. SIG [159]. SIGCHI [1090, 14]. SIGGRAPH [487, 26]. SIGKids [487]. signal [1543]. Signals [207]. Signature [342, 1758]. Signatures [729]. significance [285]. SIGPLAN [982]. Silicon [1501]. SIM Agent [1743]. SimCalc [1110]. Simple [73, 65, 1542]. Simulation [1049, 106, 108, 109, 515, 921, 728, 104, 103, 105, 640]. Simulations [1740, 1744, 1881]. simulator [669]. Singapore [1068]. single [1082]. Sins [730]. Site [1345, 1192, 1247, 847, 846, 1473, 1717, 1680, 1495, 1834, 1931]. Site-Based [846, 847]. Siteseer [1261]. Situated [524]. Situation [1264]. Situations [772]. Six [46]. Six-digit [46]. Sixteen [855]. Sixth [52]. Sketching [1047, 1151]. skills [851, 1766]. Skip [62]. Slim [1446]. Small [835, 572]. Smalltalk [924, 927, 923, 925, 922]. Smart [1096, 18]. Smartcard [738]. Smith [744]. SNA [23]. Snowbird [87]. Social [1241, 966, 869, 928, 1259, 387, 207, 601, 1138, 605, 608, 1680, 255, 1377, 1463]. Socializing [1345]. Socialware [1739]. Societal [1224, 1228, 1225, 1231, 1230, 1229, 1226, 1227]. societies [1205]. Society [637, 1939]. Socio [378]. Socio-Technical [378]. soft [632]. softbot [683]. softbot-based [683]. Software [661, 1281, 587, 206, 710, 1702, 1810, 1909, 1003, 1141, 999, 1389, 680, 1150, 1430, 1902, 970, 1317, 1024, 837, 367, 920, 143, 1240, 1779, 732, 1078, 491, 373, 1214, 239, 1390, 102, 454, 873, 649, 919, 1137, 27, 264, 1186, 1940, 1201, 187, 101, 1319, 1767, 1475, 1896, 740, 1941, 797, 1561, 1138, 1607, 607, 1320, 1898, 1621, 176, 1572, 1405, 1689, 302, 1933, 855, 727, 175, 1838, 95, 1657, 1852, 1726, 585, 268, 337, 1809, 1318, 691, 1670, 1614, 18, 361, 366, 801, 178, 1465, 586, 560, 395, 870, 188, 1411, 1729, 1690, 301, 1769, 1773]. software [59, 1251, 1397, 141, 180, 1709, 1391, 303, 1044, 1615, 1471, 853, 850, 61, 1543, 574, 173, 514]. software-based [586]. Sold [781]. soldier [1731]. Soldiers [793]. Solitons [431]. Solution [165, 344]. solutions [1899]. Solved [137]. solver [9]. Solving [1888, 1026, 1311, 204]. Some [497, 132, 74, 529, 1765, 895]. Sorting [1187, 1528]. source [1729, 1769]. South [612]. Soviet [184, 568]. Space [1111, 155, 809]. Space-Time [1111]. Spaces [421, 889, 524, 1235, 422]. Spam [1616, 1660]. span [1401]. Spatial [380, 1358, 888]. Spatially [524]. Speak [1474]. speakers [1211]. Speaking [224, 280, 553, 420, 455, 600, 746, 694, 806, 873, 968, 1165, 1075, 1297, 1679, 114, 76, 183, 148, 1624, 1871, 1696]. Special [555, 1094, 1357, 569, 541, 353, 90, 45, 433, 1479, 520]. Special-Needs [1357]. specialize [292]. specific [1402, 1687]. Specification [179, 177, 658]. Specification-level [179]. Specifications [1201, 727, 1447, 305]. Speech [950, 1247, 1698, 5, 318, 629, 79]. Speed [169, 788, 914, 809, 1124, 1834]. speeds [1690]. Spelling [320]. Spinning [1239]. spring [983]. SQL [73]. Squeeze [721]. Stabilizing [791]. Stable [728]. Staffing [1752]. Staged [998]. Stakeholders [907]. Standard [13, 164, 1085, 163, 161, 342]. standardization [1893, 1453]. Standardizing [1246]. Standards [160, 68, 745, 1095, 723, 589, 709, 638, 914, 781, 967, 1372, 1246, 1766, 341, 1695, 1646, 1585, 1002, 1453, 1579]. Stanford [821]. Starburst [230]. Starr [1699]. START [185]. State [626, 1132, 1143, 1817]. state-coordinated [1817]. statement [499]. states [1296, 150]. Statistical [1154, 1760, 1514]. Status [1207, 1872]. Staying [1930, 1832]. Step [1949, 1276, 1435]. steps [1435, 509]. Still [1343, 1339, 863, 1394, 161]. stochastic [107]. Stock [1277]. Stockholm [1294, 1366]. Stop [1660, 1932]. storage [205]. Stories [1279, 1189]. Story [66]. storyboard [502]. Strand [1174]. strangers [1445]. Strategic [732, 1457, 131]. Strategies [727, 23, 253, 1835, 1809]. Strategy [1630]. strength [1627]. Strengthening [1800]. Stress [1267, 568]. Stretching [1240]. String [1]. Strings [244, 246, 247, 245, 1948, 63]. Strong [1529, 1646]. Structural [281]. Structure [797, 775, 1083, 369, 336, 507, 619]. structure-based [619]. Structured [878, 205]. structures [8]. structuring [240, 620]. Struggle [1297]. Student [1550, 1023, 1293, 139, 1465]. Students [1520, 1028, 1019, 567, 481]. Studies [1491, 1764, 298]. study [1936, 18, 178, 126]. Studying [868]. StudySpace [885]. Style [48]. Sub [432]. Sub-Saharan [432]. Subclass [73]. Subject [1271, 961]. subjectivity [538]. Substitute [793]. Substring [311, 85]. Succeeding [842]. Success [1335, 403, 589, 1302, 1908, 802]. Successful [845, 1544, 1574]. succotash [313]. Summary [397, 189]. Supercomputer [582, 583]. supercomputers [358]. Supercomputing [1001, 236, 353]. Superconnectivity [1600].

Superhighway [832, 985, 674, 778, 1082]. Superimposing [144]. superior [1394]. Supply [135]. Supply/Demand [135]. Support [1331, 679, 1026, 284, 194, 239, 1126, 1293, 725, 1552, 890, 1124, 1322, 593, 1329, 557, 653, 1499, 1539, 773]. Supported [760, 478, 257]. Supporting [1547, 1739]. Surf [1219]. Surgeon [278]. surprising [1765]. Surveillance [551, 591]. Survey [274, 276, 629, 44, 309, 322, 52, 99]. Surveying [96]. survivability [1838]. Survivable [324]. surviving [1484]. Sustain [533]. switching [74]. sword [1847]. Swords [661]. Symbolic [1700]. Synchronous [1035, 1760]. Synergism [4]. Syntax [32]. Syntax-Directed [32]. synthesis [1857, 1856]. synthesizer [318]. System [1209, 997, 228, 738, 1047, 12, 1008, 287, 1046, 1778, 169, 868, 270, 947, 1051, 1030, 579, 1451, 196, 1126, 229, 378, 1258, 728, 1033, 317, 316, 1319, 809, 563, 1449, 1691, 226, 176, 269, 1124, 213, 268, 615, 613, 813, 1639, 739, 1521, 366, 370, 227, 557, 591, 496, 1523, 219, 127, 1377, 425, 816]. Systematic [559, 1398, 880]. systemic [1464, 473]. Systems [1360, 1263, 1198, 1217, 791, 1005, 1331, 943, 1385, 599, 160, 626, 958, 974, 214, 775, 1054, 1079, 443, 1163, 194, 1301, 1034, 1288, 1081, 47, 926, 725, 389, 1010, 1020, 1140, 1758, 651, 1755, 1148, 1399, 307, 716, 225, 283, 220, 145, 1688, 334, 30, 1870, 1840, 1322, 1738, 593, 1252, 1329, 253, 177, 20, 17, 107, 617, 1444, 1554, 133, 1739, 98, 1588, 1123, 1735, 1591, 1482, 1609, 22, 143, 586, 1437, 654, 631, 324, 871, 1499, 1481, 1448, 1882, 1359, 416, 1077, 1257, 1675, 200, 231, 309, 1543, 1864, 1538, 1529, 803, 1761]. Systems [295].

Table [621, 271]. Tables [645, 384]. tags [1539]. tailorable [1400]. tailored [318]. Take [804, 278, 396]. Taking [1015, 1277, 1767]. Tale [184, 1298]. Tales [437]. tangible [1859]. tank [1744]. tap [446, 764]. tapestry [414]. Tapping [764]. Target [874]. Task [369, 1704, 189, 488]. Task-structure [369]. Taught [1873]. Taulbee [99, 275, 276]. Tax [1190]. taxes [348]. taxonomic [1557]. Taxonomy [1005, 494, 740, 493]. TCP [788]. TCQ [5]. TD [800]. TD-Gammon [800]. Teach [567, 1822]. Teachers [1058]. Teaching [482, 1876, 3, 153, 1111, 960, 685, 1409, 828]. team [1879, 1671, 574]. Teams [102, 981, 576, 573, 1878, 572, 575]. TeamWorkStation [252]. Techies [757]. Technetronic [467]. Technical [1755, 804, 719, 311, 1324, 792, 538, 1584, 1474, 1692, 791, 566, 1171, 812, 1014, 1728, 776, 347, 844, 430, 1674, 1309, 54, 534, 1711, 1364, 1827, 1504, 1804, 1903, 777, 378, 397, 1339, 1884, 325, 621, 1289, 605, 1391, 393, 575]. Technique [24, 728, 584]. Techniques [1759, 317, 1124, 371]. Technological [1433, 432]. Technologies [1850, 1164, 736, 367, 1075, 472, 1169, 1597, 593, 1791, 444, 630, 1792, 1264, 773]. Technology [1237, 1113, 1122, 1059, 1218, 1241, 1242, 1120, 1345, 1080, 470, 1117, 1268, 367, 1243, 1119, 954, 696, 1234, 1240, 957, 1225, 1236, 452, 637, 1166, 916, 1208, 1372, 1338, 1233, 1361, 833, 849, 1238, 1232, 1118, 346, 1239, 469, 1424, 1219, 1235, 483, 1537, 1795, 315, 594, 542, 602, 131, 182, 1546, 423, 36, 1726, 391, 484, 440, 473, 193, 523, 1300, 1937, 654, 1711, 188, 488, 629, 1644, 1044, 1605, 1793, 1377, 1463, 602]. Telecom [1832]. Telecommunications [925, 1255, 780, 279, 609, 548, 320, 543, 1824]. telecommunications-use [1824]. Teledemocracy [1938]. telephone [74, 1712]. telepresence [1556]. Telepresentations [1285]. Teleputer [89]. Television [169, 590]. Tell [66]. Temporal [800]. Ten [1921, 1805]. tension [501]. Tenth [1174]. Teraflop [351]. Term [1094, 697, 868, 389, 1578]. terminology [1728]. Terrible [1231]. Test [986, 536, 967, 670, 535, 1759]. testability [716]. testable [1757]. Testbed [106, 306]. Testing [715, 710, 1870, 1840, 711, 920, 714, 1758, 1797, 1554, 739, 712, 1545, 713, 174]. Text [580, 244, 246, 247, 245, 381, 751, 726, 319, 1917, 63, 620, 1528, 1519, 382]. Their [156, 284, 285, 1853, 1377]. them [1376, 1250]. theorem [178]. Theoretic [799]. theoretical [309]. Theory [1150, 775, 1149, 1251, 1908, 1674, 1481, 1734, 1916]. Therapy [1353]. there [1804]. They’re [729]. Think [1696, 852]. Thinking [1509, 1432, 1389, 1784, 1865, 1620, 1484, 1327, 463]. Third [909, 1801]. third-party [1801]. Thoughts [1086, 448]. threat [1679]. Three [730, 1598, 1333, 23, 187, 1688]. three-faceted [1688]. Tiger [756, 1245]. Time [351, 1332, 1216, 1111, 1200, 741, 155, 218, 1858, 1860, 201, 177, 616, 1492, 948, 1540, 125, 1568, 202, 1878, 200, 1895]. time-based [240]. Times [833, 1662, 1449]. Timing [1202]. Tiny [639]. Tip [1203]. Title [1474, 1949]. TODAES [984]. Today [1192, 26, 1586]. Together [421, 690]. Token [6]. tolerance [1761]. Tolerant [1200, 1030, 145, 1143]. TOMACS [1415]. Tomorrow [1192, 1231, 675, 1235, 1586]. Tomorrows [26]. Tool [1028, 732, 304]. toolbox [1783, 507, 619]. toolkit [1743, 1032]. Tools [476, 156, 902, 159, 623, 542, 1554, 1147, 1618, 480, 1668, 1912, 308, 773]. Toolsmith [1009]. Top [872, 1783]. Total [165, 1480]. Totem [1030]. tour [393]. Touring [425]. Toys [1109]. TQM [575]. traceability [1687, 1685]. tracing [1684]. Track [1039, 1092, 1470, 1558, 1906, 1370, 1416, 1434, 1452, 1487, 1506, 1532, 1577, 1604, 1622, 1642, 1659, 1677, 1694, 1713, 1730, 1763, 1787, 1807, 1830, 1849, 1868, 1887, 1927]. Tracking [525, 697, 1417]. Tractor [649]. Trade [929, 1560]. trade-offs [1560]. Trading [352, 1754]. tradition [1422]. Traditional [110]. Traffic [1869, 787, 1827].

Training [927, 864, 865, 1880, 866, 1879, 1883, 1669, 867, 869, 1142, 577, 1882, 1878]. Transaction [118]. Transactions [1209, 1034, 1719]. Transatlantic [498]. transclusion [877]. Transfer [1122, 1120, 367, 1119, 1118, 1117, 1064]. Transformation [1270, 428]. Transforming [1379, 900]. Transis [1031]. Transition [999, 1216]. Translation [514, 49, 438]. Translations [907]. Transportation [270]. travel [1560]. Treating [1356]. Tree [621, 2, 7, 271, 336]. Trees [62, 384]. Trellis [93]. Trends [993, 1072, 1501]. Trial [1170]. trials [1584]. tribulations [1584]. Tribute [1682, 405, 896]. trip [478]. Trojan [1885]. trouble [1410]. Troubleshooting [799]. Trust [1831, 385, 1032, 1837, 1781, 1727]. TRUSTe [1723]. Trusted [345, 1842]. trustworthy [1675]. truths [1716, 1678]. tug [1889]. tug-of-war [1889]. Turbulent [833]. Turing [1052, 724, 426, 1053, 670, 741]. Turning [1895, 93]. Turnover [281]. Tutoring [1020]. TV [794]. Two [983, 1295, 13, 184, 1079, 1078, 408, 1435]. Two-Level [1079]. type [1692]. Types [1179, 1082, 230]. typical [1483]. Typographic [48]. Typology [362].

U.S [284]. U.S. [483, 1059, 1815, 182, 1176, 1168, 1250, 1422]. UART [1757]. Ubiquitous [926, 529, 1457, 1460]. UC [822]. UCLA [274, 52]. Ugly [1266, 1716, 1678]. UI [459]. UIs [461]. UK [762]. Ultimately [1203]. Ultra [775]. Ultra-Structure [775]. Ultracomputers [351]. ultrahigh [586]. UML [1892, 1897, 1893, 1894, 1895]. UMLoquent [1896]. Unbounded [1179]. Uncertainties [1166]. Uncertainty [54, 112]. Uncle [290]. Unconsummated [1390]. unconventional [1780]. Undergraduate [394, 1054]. Underlying [829]. underpinnings [547]. Understanding [145, 1794, 91, 376, 655, 1409, 383, 653, 416, 79]. Unexpected [181]. unified [877, 1655]. Unifying [91, 1539]. Union [568]. Unique [1810]. Unit [436]. unite [182, 1562]. United [150]. Units [1269]. Universal [1255, 1459]. universally [1082]. Universe [1427]. Universität [1949]. Universities [533]. University [1380, 1928, 1120, 1277, 721, 1043, 825, 1949, 1458, 816, 1823]. UNIX [126, 1771]. unknown [1731]. Unraveling [910]. unwanted [1502]. Update [1089, 256, 1641]. Updating [360]. Upgrade [1309]. upon [1377]. Ups [1267, 455]. URL [1273]. Usability [1733, 1803, 1797, 1801, 1796, 1802, 644]. usable [134]. USACM [1414]. Usage [970, 52, 274]. Use [1360, 953, 1004, 1168, 868, 284, 1084, 1233, 553, 600, 694, 977, 747, 607, 1392, 1798, 1824]. used [1147, 1942]. Useful [1153, 134]. Usenet [1262]. User [840, 864, 506, 817, 1330, 456, 865, 110, 1706, 1072, 955, 1223, 1797, 1611, 869, 1800, 1634, 1618, 595, 1683, 460, 318, 1528, 411, 1527]. User-Centered [817, 1330, 1618]. user-interface [1528]. user-tailored [318]. Users [1935, 1636, 604, 1551, 1377, 514]. Uses [162, 1866]. ushers [980]. Using [442, 1822, 453, 840, 1570, 1362, 414, 1592, 654, 1087, 464, 1448, 458, 919, 1102, 1337, 1104, 317, 522, 1322, 1862, 698, 178, 1894, 1901, 657, 502, 1913, 504, 549, 1895, 1743, 1605, 1566, 1463]. utilities [126].

Vadis [1775]. Validation [586]. Value [1418, 1697, 848, 1587, 1206, 830]. Value-added [1697]. Values [50, 959, 197]. Vanderbilt [1113]. Variable [244, 246, 247, 245, 63]. variable-length [63]. variance [204]. variety [1082]. VAX [435]. VDM [178]. vector [583]. vehicles [202]. vendor [308]. Verdict [729].

Verification [1186]. verify [1837]. Versus [914]. Very [311, 85]. Via [1873, 477]. Victory [1093]. Video [421, 840, 820, 423, 165, 1101, 1442, 164, 480, 252, 163, 168, 166, 1441]. Video-Game [1101]. videotape [1833]. Vietnam [756]. View [1003, 139, 1688, 444, 1637, 39, 1655]. Viewed [1427]. Viewing [1595, 618, 1516]. Viewpoint [1767, 1001, 730, 1607, 1932, 966, 1835, 403, 235, 1076, 1698, 743, 1277, 1646, 985, 693, 1790, 1683, 1040, 387, 1135, 1096, 846, 1298, 1375, 861, 1873, 929, 1734, 833, 1455, 1580, 1343, 1344, 1625, 139, 1327, 1811, 1663, 949]. Viewpoints [1755]. Views [752]. Village [931, 986, 1059, 1210, 1266, 1325, 1419, 1488, 1907, 1788, 1697, 1173]. VIP [701]. Virtual [1256, 1049, 1357, 1048, 1380, 1536, 1379, 1046, 991, 1381, 1378, 1363, 1382, 1376, 1196, 1045, 1428, 1044, 895, 1347, 1383, 1354, 1858, 332, 789, 1194, 1939, 1741, 523, 1671, 1882, 1136, 1878, 1686, 1884, 1569]. Virtuality [1377]. virtually [1376]. Virus [1195]. Visible [905]. vision [1194, 628]. Visit [1367]. Visual [32, 1443, 1305, 1439, 883, 1686, 332, 1440, 748, 303]. Visualization [1281, 1048, 329, 331, 458, 1862, 1689, 815, 642, 479]. Visualizing [379, 1441]. Visually [1860]. VLSI [124]. Voice [1869, 1827]. Volumes [1153]. volumetric [1862]. Voting [1937, 402]. Voyage [931]. VR [1353, 1047, 1358, 1349, 1352, 1346, 1348]. vs [1764, 1733, 150, 30, 1086, 644, 1364, 1903, 1343, 1344, 1529]. vulnerabilities [860].

W [1000]. walking [578]. War [1643, 1279, 1164, 1189, 1889, 602]. Warehouse [1156, 1707, 1629, 1632]. warehouses [1631]. Warehousing [1635, 1627]. Warfare [1888, 1151, 1249, 1836]. Warns [278]. wars [1699]. Was [1375, 140]. Washington [1833, 1889, 66, 182, 236, 1606, 1929, 1040, 1341, 1151, 1190, 1454, 1661, 1093, 540, 581, 624, 660, 721, 863, 898, 780, 279]. Wasteful [1432]. Watermarking [1419, 1587, 1585, 1582]. waters [1663]. Wave [1232]. Waves [1219]. Way [1360, 852, 1424, 1494]. Ways [1101, 611, 1605]. WCAR [763]. Weak [419, 1529]. Wearable [994, 1096]. weave [414]. web [513, 1597, 969, 702, 1897, 1733, 1304, 952, 1598, 1283, 966, 990, 1159, 1617, 811, 1721, 1752, 1168, 1588, 1591, 1259, 1901, 1516, 1711, 1706, 1499, 1495, 794, 1579, 1386, 1719, 1261, 1500, 1635, 1861, 1407, 1594]. Web-based [1617, 1499, 1594]. Webs [1239, 1539]. Webware [1290]. Week [779, 965, 980]. Where [1145, 385, 239, 759, 1315, 142]. Who [950, 1643, 1561, 1342, 1429, 967]. whole [374]. Whose [1508]. Wide [1221, 249, 1082, 969, 702, 966, 990, 1159, 811, 812]. Wide-Area [249]. Widening [913, 1559]. Wilkes [466]. Will [377, 967, 1233, 1230, 1294, 1369, 1368, 1209, 1082, 170, 220]. WIMP [1223]. Windows [930, 113, 1041]. Wings [14]. winners [1292]. Winona [1171]. Wired [1171]. wireless [722, 1819, 1930, 478]. Wiretaps [743]. wisdom [1835]. Wish [1238]. Wishes [1333]. WISs [1595, 1592, 1599, 1601, 1590]. within [1521]. Without [1050, 1146, 681, 1543]. Woes [579, 1451]. Women [762, 1207, 1474, 760, 764, 115, 771, 759, 758, 765, 1637]. won [1643]. Won’t [1076]. Word [236]. WordNet [936, 934]. Work [265, 511, 1702, 571, 510, 904, 1309, 194, 1045, 900, 901, 905, 908, 604, 1924, 253, 690, 906, 576, 257, 678, 1802, 770, 899]. Work-oriented [511]. Working [1091, 907, 1800]. Workplace [421, 768, 902, 742]. Works [1075]. workspace [252]. Workstation [1115, 596, 583]. workstations [1085]. World [1256, 715, 1221, 598, 1215, 130, 1080, 795, 1284, 520, 35, 198, 1907, 1139, 523, 1562, 141, 1377, 969, 702, 966, 990, 1159, 811].

World-Wide [1221, 969, 702, 966, 990, 1159, 811]. Worlds [1224, 1884]. worldwide [1510, 1194, 1813, 1511]. worry [1929]. Worst [985]. Worst-Case [985]. Writing [468]. wrong [1535]. WWW [880, 883, 1072].

Xday [1788]. Xerox [511]. XML [1753, 1901, 1754]. XML-based [1754]. XX [21]. XXI [53]. XXII [121]. XXIII [843].

Y2K [1788, 1626, 1850, 1943, 1725, 1872, 1641, 1828, 1805, 1875]. Yaksha [1008]. Yang [1530]. Year [1310, 259, 916, 349, 516, 1462, 338, 1488, 1476, 1309]. Year-2000 [1488, 1476, 1309]. Years [1598, 1208, 1238, 1230, 979, 1291, 1212, 1394]. Yin [1530]. Yu [326].

Zealand [1016]. zone [813].

References

Boyer:1977:FSS

[1] Robert S. Boyer and J. Strother Moore. A fast string searching algorithm. Communications of the ACM, 20(10):762– 772, October 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [1948, 85, 381].

Gerasch:1988:IAM

[2] Thomas E. Gerasch. An insertion algorithm for a minimal internal path length binary search tree. Communications of the ACM, 31(5):579–585, May 1988. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [146].

Dijkstra:1989:CRT

[3] Edsger W. Dijkstra. On the cruelty of really teaching computer science. Communications of the ACM, 32(12): 1398–1404, December 1989. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [25, 34].

Golshani:1990:PSC

[4] Forouzan Golshani and Thaddeus Regulinski. Prologue: On the synergism of computers and communication. Communications of the ACM, 33(1):8–10, January 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Marcellin:1990:GPT

[5] Michael W. Marcellin and Thomas R. Fischer. Generalized predictive TCQ of speech. Communications of the ACM, 33(1):11–19, January 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/76373.html.

Gaitonde:1990:BDM

[6] S. S. Gaitonde, Doug W. Jacobson, and A. V. Pohm. Bounding delay on a multifarious token ring network. Communications of the ACM, 33(1):20–28, January 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/76374.html.

McKinley:1990:MTC

[7] Philip K. McKinley and Jane W. S. Liu. Multicast tree construction in bus-based networks. Communications of the ACM, 33(1):29–42, January 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/76375.html.

Padmanabhan:1990:CSM

[8] Krishnan Padmanabhan. Cube structures for multiprocessors. Communications of the ACM, 33(1):43–52, January 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/76376.html.

Freeman-Benson:1990:ICS

[9] Bjorn N. Freeman-Benson, John Maloney, and Alan Borning. An incremental constraint solver. Communications of the ACM, 33(1):54–63, January 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/77531.html.

Sibley:1990:ENI

[10] Edgar H. Sibley. Editorial: New and improved computing practices. Communications of the ACM, 33(1):66–69, January 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chung:1990:DAH

[11] Chin-Wan Chung. DATAPLEX: an access to heterogeneous distributed databases. Communications of the ACM, 33(1):70–80, January 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/76377.html.

Gait:1990:PSM

[12] Jason Gait. Phoenix: a safe inmemory file system. Communications of the ACM, 33(1):81–86, January 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/76378.html.

Carta:1990:TFI

[13] David G. Carta. Two fast implementations of the “minimal standard” random number generator. Communications of the ACM, 33(1):87–88, January 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/76379.html.

Press:1990:WMS

[14] Larry Press. Wings for the mind: The 1989 SIGCHI national conference. Communications of the ACM, 33(2): 115–119, 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoadley:1990:IEC

[15] Ellen Hoadley. Investigating the effects of color. Communications of the ACM, 33(2):120–125, February 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/75578.html.

Batra:1990:CRR

[16] Dinesh Batra, Jeffrey A. Hoffler, and Robert P. Bostrom. Comparing representations with relational and EER models. Communications of the ACM, 33(2):126–139, February 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/75579.html.

Gifford:1990:PSM

[17] David K. Gifford. Polychannel systems for mass digital communications. Communications of the ACM, 33(2): 141–151, February 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/75580.html.

Kierulf:1990:SGB

[18] Anders Kierulf, Ken Chen, and Jurg Nievergelt. Smart Game Board and Go Explorer: a study in software and knowledge engineering. Communications of the ACM, 33(2):152–166, February 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/75581.html.

Joyce:1990:IAB

[19] Rick Joyce and Gopal Gupta. Identity authentication based on keystroke latencies. Communications of the ACM, 33(2):168–176, February 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/toc/Abstracts/
0001-0782/75582.html.

Gavish:1990:DFM

[20] Bezalel Gavish and Olivia R. Liu Sheng. Dynamic file migration in distributed computer systems. Communications of the ACM, 33(2):177–189, February 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/75583.html.

Rosenberg:1990:SAP

[21] J. Rosenberg, A. L. Ananda, and B. Srinivasan. Self-Assessment Procedure XX. Communications of the ACM, 33(2):190–201, February 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/75584.html.

Kumar:1990:PIE

[22] Kuldeep Kumar. Post implementation evaluation of computer-based information systems: current practices. Communications of the ACM, 33(2): 203–212, February 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/75585.html.

Tillman:1990:SOT

[23] Matthew A. Tillman and David Chi-Chung Yen. SNA and OSI: Three strategies for interconnection. Communications of the ACM, 33(2): 214–224, February 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/

0001-0782/75586.html. Computing

Reviews, Vol. 31, No. 10, October 1990.

Martin:1990:FIR

[24] Johnny Martin and W. T. Tsai. N-fold inspection: a requirements analysis technique. Communications of the ACM, 33(2):225–232, February 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/75587.html.

Bernstein:1990:AFD

[25] Danielle R. Bernstein, Joshua Stern, William Ehrich, Michael R. Dunlavey, Leonard Zettel, Bruce B. Herbert, Mark Nelson, W. F. Smyth, Bruce W. Walker, Robert M. Gordon, Stephen N. Cole, and David B. Lightstone. ACM Forum: Debatable. Communications of the ACM, 33(3):264–271, March 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [3, 34].

Press:1990:STP

[26] Larry Press. SIGGRAPH ’89 — Tomorrows’s PC today. Communications of the ACM, 33(3):274–279, 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Science:1990:SRA

[27] Computer Science and Technology Board. Scaling up: a research agenda for software engineering. Communications of the ACM, 33(3):281–293, March 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/77482.html.

Nielsen:1990:ANT

[28] Jakob Nielsen. The art of navigating through hypertext. Communications of the ACM, 33(3):296–310, March 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/77483.html.

Carlson:1990:HNF

[29] David A. Carlson and Sudha Ram. HyperIntelligence: The next frontier. Communications of the ACM, 33(3):311–321, March 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/77484.html.

Dhar:1990:IPV

[30] Vasant Dhar and Nicky Ranganathan. Integer programming vs. expert systems: an experimental comparison. Communications of the ACM, 33 (3):323–336, March 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/77485.html.

Molich:1990:IHC

[31] Rolf Molich and Jakob Nielsen. Improving a human-computer dialogue. Communications of the ACM, 33 (3):338–348, March 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/77486.html.

Arefi:1990:AGV

[32] Farahangiz Arefi, Charles E. Hughs, and David A. Workman. Automatically generating visual syntax-directed editors. Communications of the ACM, 33(3):349–360, March 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/77487.html.

VanWyk:1990:LPA

[33] Christopher J. Van Wyk. Literate programming—an assessment. Communications of the ACM, 33(3):361, 365, March 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wielgus:1990:AFM

[34] Martin Wielgus, Marvin L. Watkins, Edward Sacharuk, and Srinivasarao Damerla. ACM Forum: More debatable. Communications of the ACM, 33(4):396–397, April 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [3, 25].

Wilkes:1990:CSB

[35] Maurice V. Wilkes. Computer security in the business world. Communications of the ACM, 33(4):399–401, 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Frenkel:1990:ECI

[36] Karen A. Frenkel. The European community and information technology. Communications of the ACM, 33(4):404–410, April 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/77557.html.

Jeelof:1990:EFF

[37] Gerrit Jeelof. Europe 1992: Fraternity or fortress? Communications of the ACM, 33(4):412–416, April 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/77558.html.

Chiarodo:1990:SME

[38] Roger A. Chiarodo and Judee M. Mussehl. The semiconductor market in the European community. Communications of the ACM, 33 (4):417–423, April 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/77559.html.

Ono:1990:JVE

[39] Eiichi Ono. Japan’s view of EC ’92. Communications of the ACM, 33(4):424–425, April 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/77561.html.

Anonymous:1990:ECC

[40] Anonymous. EC commission communication on establishing an information services market. Communications of the ACM, 33(4):426– 432, April 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/77562.html.

Anonymous:1990:CDJ

[41] Anonymous. Council decision of July 26, 1988. Communications of the ACM, 33(4):434–438, April 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/77563.html.

Geihs:1990:RD

[42] Kurt Geihs and Ulf Hollberg. A retrospective on DACNOS. Communications of the ACM, 33(4):439– 448, April 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/77564.html.

Hirschberg:1990:EDP

[43] Daniel S. Hirschberg and Debra A. Lelewer. Efficient decoding of prefix codes. Communications of the ACM, 33(4):449–459, April 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/77566.html.

Samuelson:1990:SLF

[44] Pamela Samuelson and Robert J. Glushko. Survey on the look and feel lawsuits. Communications of the ACM, 33(5):483–487, 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sibley:1990:SSA

[45] Edgar H. Sibley. Special section on alphabets and language: Introduction. Communications of the ACM, 33(5):488–490, May 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/78608.html.

Qiao:1990:SDC

[46] Jinan Qiao, Yizheng Qiao, and Sanzheng Qiao. Six-digit coding method. Communications of the ACM, 33(5):491–494, May 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/78609.html.

Tayli:1990:BBM

[47] Murat Tayli and Abdulla I. Al-Salamah. Building bilingual microcomputer systems. Communications of the ACM, 33(5):495–504, May 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/78610.html.

Oman:1990:TSM

[48] Paul W. Oman and Curtis R. Cook. Typographic style is more than cosmetic. Communications of the ACM, 33(5):506–520, May 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/78611.html.

Raman:1990:ABA

[49] Subramanian Raman and Narayanan Alwar. An AI-based approach to machine translation in Indian languages. Communications of the ACM, 33(5):521–527, May 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/78612.html.

Kumar:1990:CCC

[50] Kuldeep Kumar and Niels Bjorn-Andersen. A cross-cultural comparison of IS designer values. Communications of the ACM, 33(5):528– 538, May 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/78613.html.

Karp:1990:MPP

[51] Alan H. Karp and Horace P. Flatt. Measuring parallel processor performance. Communications of the ACM, 33(5):539–543, May 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/78614.html.

Frand:1990:SAU

[52] Jason L. Frand and Julia A. Britt. Sixth annual UCLA survey of business school computer usage. Communications of the ACM, 33(5):544– 562, May 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/78615.html.

Rudolph:1990:SAP

[53] Brain A. Rudolph. Self-assessment procedure XXI: a self-assessment procedure on concurrency. Communications of the ACM, 33(5):563– 576, May 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/78616.html.

Lehman:1990:TCU

[54] M. M. Lehman. Technical correspondence: Uncertainty in computer application. Communications of the ACM, 33(5):584–586, May 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:1990:AJP

[55] Peter J. Denning. Alan J. Perlis: 1922–1990: a founding father of computer science as a separate discipline. Communications of the ACM, 33(5):604–605, May 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/214943.html.

Kocher:1990:PLE

[56] Bryan Kocher. President’s letter: Eschew obfuscation. Communications of the ACM, 33(6):625, June 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gladney:1990:LEM

[57] H. M. Gladney, M. L. Rinehart, F. Wagner, David H. Thornley, and Justin H. Coven. Letters to the editor: More on the cruelty. Communications of the ACM, 33(6):628–630, 633, June 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wilkes:1990:NEF

[58] Maurice Wilkes. Networks, email and fax. Communications of the ACM, 33(6):631–633, June 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Parnas:1990:ESC

[59] David L. Parnas, A. John van Schouwen, and Shu Po Kwan. Evaluation of safety-critical software. Communications of the ACM, 33(6):636– 648, June 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/78974.html.

Rohde:1990:DRP

[60] Renate Rohde and Jim Haskett. Disaster recovery planning for academic computing centers. Communications of the ACM, 33(6):652– 657, June 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/78975.html.

Swanson:1990:DSD

[61] E. Burton Swanson and Cynthia Mathis Beath. Departmentalization in software development and maintenance. Communications of the ACM, 33 (6):658–667, June 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/78976.html.

Pugh:1990:SLP

[62] William Pugh. Skip lists: a probabilistic alternative to balanced trees. Communications of the ACM, 33 (6):668–676, June 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/78977.html. This paper

presents skip lists, a list in which a node may have a pointer to a node some number of places ahead of it on the list. Such pointers, called “forward pointers”, therefore “skip” over intermediate nodes. A node with k forward pointers is said to be a level k node. Skip lists are probabilistic in that the level of a node is chosen randomly with the property that a node’s ith forward pointer points to the next node of level i or higher. It is shown that skips lists can efficiently implement abstract data types such as dictionaries and ordered lists in that the expected time to search for an item is O(log n).

Pearson:1990:FHV

[63] Peter K. Pearson. Fast hashing of variable-length text strings. Communications of the ACM, 33(6):677– 680, June 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-

0782/78978.html. See [1949, 244, 246, 247, 245].

Kumar:1990:COE

[64] Vijay Kumar. Concurrent operations on extendible hashing and its performance. Communications of the ACM, 33(6):681–694, June 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/78979.html.

Press:1990:PCS

[65] Larry Press. Personal computing: Simple complexity and COMDEX. Communications of the ACM, 33(7):21– 26, July 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/79205.html.

Crawford:1990:WBF

[66] Diane Crawford. From Washington – budget FY 1991: The numbers tell the R&D story. Communications of the ACM, 33(7):27–29, July 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/79206.html.

Linehan:1990:DID

[67] Thomas E. Linehan. Digital image: Digital cinema. Communications of the ACM, 33(7):30–37, July 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/79207.html.

Frenkel:1990:PSE

[68] Karen A. Frenkel. The politics of standards and the EC. Communications of the ACM, 33(7):40– 51, July 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/79208.html.

Cohen:1990:CLP

[69] Jacques Cohen. Constraint logic programming languages. Communications of the ACM, 33(7):52–68, July 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/79209.html.

Colmerauer:1990:IPI

[70] Alain Colmerauer. An introduction to Prolog III. Communications of the ACM, 33(7):69–90, July 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/79210.html.

Newborn:1990:AAN

[71] Monty Newborn and Danny Kopec. The 20th Annual ACM North American Computer Chess Championship. Communications of the ACM, 33 (7):92–104, July 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/79211.html.

Kaplan:1990:API

[72] S. Jerrold Kaplan, Mitchell D. Kapor, Edward J. Belove, Richard A. Landsman, and Todd R. Drake. AGENDA: a personal information manager. Communications of the ACM, 33(7):105– 116, July 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/79212.html.

Kung:1990:OSH

[73] Chenho Kung. Object subclass hierarchy in SQL: a simple approach. Communications of the ACM, 33 (7):117–125, July 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/79213.html.

Neumann:1990:IRS

[74] Peter G. Neumann. Inside RISKS: Some reflections on a telephone switching problem. Communications of the ACM, 33(7):154, July 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/79214.html.

Wilkes:1990:BF

[75] Maurice V. Wilkes. The bandwidth famine. Communications of the ACM, 33(8):19–21, August 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/79174.html.

Samuelson:1990:LSS

[76] Pamela Samuelson. Legally speaking: Should program algorithms be patented? Communications of the ACM, 33(8):23–27, August 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/79175.html.

Lenat:1990:CTP

[77] Douglas B. Lenat, Ramanathan V. Guha, Karen Pittman, Dexter Pratt, and Mary Shepherd. Cyc: Toward programs with common sense. Communications of the ACM, 33 (8):30–49, August 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/79176.html.

Barnett:1990:KNL

[78] Jim Barnett, Kevin Knight, Inderjeet Mani, and Elaine Rich. Knowledge and natural language processing. Communications of the ACM, 33(8):50–71, August 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/79177.html.

White:1990:NLU

[79] George M. White. Natural language understanding and speech recognition. Communications of the ACM, 33(8):72–82, August 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/79178.html.

Frenkel:1990:HSA

[80] Karen A. Frenkel. Highlights from the Second Annual Computer Bowl. Communications of the ACM, 33 (8):84–95, August 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/79179.html.

Hill:1990:CCM

[81] Mark D. Hill and James R. Larus. Cache considerations for multiprocessor programmers. Communications of the ACM, 33(8):97–102, August 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/79180.html.

Valiant:1990:BMP

[82] Leslie G. Valiant. A bridging model for parallel computation. Communications of the ACM, 33(8):103–111, August 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/79181.html.

Sterling:1990:SSD

[83] Theodor D. Sterling and James J. Weinkam. Sharing scientific data. Communications of the ACM, 33(8): 112–119, August 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/79182.html.

Gillenson:1990:PDE

[84] Mark L. Gillenson. Physical design equivalencies in database conversion. Communications of the ACM, 33 (8):120–131, August 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/79183.html.

Sunday:1990:VFS

[85] Daniel M. Sunday. A very fast substring search algorithm. Communications of the ACM, 33(8):132–142, August 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-

0782/79184.html. See also [1, 1948, 381].

Neumann:1990:IAS

[86] Peter G. Neumann. Insecurity about security? Communications of the ACM, 33(8):170, August 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/79113.html.

White:1990:PLR

[87] J. R. White. President’s letter: Reflections on Snowbird. Communications of the ACM, 33(9):19–20, September 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:1990:EDA

[88] Diane Crawford. EC directives aim for market harmony. Communications of the ACM, 33(9):21–24, September 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/84452.html.

Press:1990:CT

[89] Larry Press. Compuvision or teleputer? Communications of the ACM, 33(9):29–36, September 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/84453.html.

McGregor:1990:ISI

[90] John D. McGregor and Tim Korson. Introduction to the special issue on object-oriented design. Communications of the ACM, 33(9): 38–39, September 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/214946.html.

Korson:1990:UOO

[91] Tim Korson and John D. McGregor. Understanding object-oriented: a unifying paradigm. Communications of the ACM, 33(9):40–60, September 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/84459.html.

Jordan:1990:IBC

[92] David Jordan. Implementation benefits of C++ language mechanisms. Communications of the ACM, 33(9): 61–64, September 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/84460.html.

Kilian:1990:TTD

[93] Michael F. Kilian. Trellis: turning designs into programs. Communications of the ACM, 33(9):65–67, September 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/84462.html.

Meyer:1990:LDE

[94] Bertrand Meyer. Lessons from the design of the Eiffel libraries. Communications of the ACM, 33(9):68–88, September 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/84464.html.

Gibbs:1990:CMS

[95] Simon Gibbs, Dennis Tsichritzis, Eduardo Casais, Oscar Nierstrasz, and Xavier Pintado. Class management for software communities. Communications of the ACM, 33 (9):90–103, September 1990. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        ftp://cui.unige.ch/OO-articles/
                                        classManagement.ps.Z;   http://
                                        www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/84525.html.

Wirfs-Brock:1990:SCR

[96] Rebecca J. Wirfs-Brock and Ralph E. Johnson. Surveying current research in object-oriented design. Communications of the ACM, 33(9):104– 124, September 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/84526.html.

Agha:1990:COO

[97] Gul Agha. Concurrent objectoriented programming. Communications of the ACM, 33(9):125–141, September 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/84528.html.

Henderson-Sellers:1990:OOS

[98] Brian Henderson-Sellers and Julian M. Edwards. The object-oriented systems life cycle. Communications of the ACM, 33(9):142–159, September 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/84529.html.

Gries:1990:TSR

[99] David Gries and Dorothy Marsh. The 1988–89 Taulbee survey report. Communications of the ACM, 33(9): 160–169, September 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/84531.html.

Neumann:1990:FOC

[100] Peter G. Neumann. A few old coincidences. Communications of the ACM, 33(9):202, September 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/84533.html.

Wilkes:1990:CPA

[101] Maurice V. Wilkes. Computing perspectives: It’s all software, now. Communications of the ACM, 33 (10):19–21, October 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/84539.html.

Rettig:1990:PPS

[102] Marc Rettig. Practical programmer: Software teams. Communications of the ACM, 33(10):23–27, October 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/84543.html.

Heidelberger:1990:IDE

[103] Philip Heidelberger. Introduction: discrete event simulation. Communications of the ACM, 33(10):28–29, October 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/214948.html.

Fujimoto:1990:PDE

[104] Richard M. Fujimoto. Parallel discrete event simulation. Communications of the ACM, 33(10):30–53, October 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/84545.html.

Jain:1990:ESL

[105] Sanjay Jain, Karon Barber, and David Osterfeld. Expert simulation for on-line scheduling. Communications of the ACM, 33(10):54–60, October 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/84547.html.

Dupuy:1990:NNS

[106] Alexander Dupuy, Jed Schwartz, Yechiam Yemini, and David Bacon. NEST: a network simulation and prototyping testbed. Communications of the ACM, 33(10):63–74, October 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/84549.html.

Glynn:1990:LRG

[107] Peter W. Glynn. Likelihood ratio gradient estimation for stochastic systems. Communications of the ACM, 33(10):75–84, October 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/84552.html.

LEcuyer:1990:RNS

[108] Pierre L’Ecuyer. Random numbers for simulation. Communications of the ACM, 33(10):85–97, October 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/84555.html.

Miller:1990:SSM

[109] David J. Miller. Simulation of a semiconductor manufacturing line. Communications of the ACM, 33(10): 98–108, October 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/84556.html.

Nielsen:1990:TDD

[110] Jakob Nielsen. Traditional dialogue design applied to modern user interfaces. Communications of the ACM, 33 (10):109–118, October 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/84559.html.

Lynch:1990:EAM

[111] Brian D. Lynch, H. Rarghav Rao, and Winston T. Lin. Economic analysis of microcomputer hardware. Communications of the ACM, 33(10): 119–129, October 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/84561.html.

Neumann:1990:IRG

[112] Peter G. Neumann. Inside RISKS: Ghosts, mysteries, and uncertainty. Communications of the ACM, 33 (10):170, October 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/84563.html.

Press:1990:PCW

[113] Larry Press. Personal computing: Windows, DOS and the MAC. Communications of the ACM, 33(11): 19–26, November 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/92783.html.

Samuelson:1990:LSH

[114] Pamela Samuelson. Legally speaking: how to interpret the Lotus decision (and how not to). Communications of the ACM, 33(11): 27–33, November 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/92784.html.

Frenkel:1990:WC

[115] Karen A. Frenkel. Women and computing. Communications of the ACM, 33 (11):34–46, November 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/92756.html.

Pearl:1990:BCS

[116] Amy Pearl, Martha E. Pollack, Eve Riskin, Becky Thomas, Elizabeth Wolf, and Alice Wu. Becoming a computer scientist. Communications of the ACM, 33(11):47–57, November 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/92757.html.

Knight:1990:CIA

[117] Kevin Knight. Connectionist ideas and algorithms. Communications of the ACM, 33(11):58–74, November 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/92764.html.

Bernstein:1990:TPM

[118] Philip A. Bernstein. Transaction processing monitors. Communications of the ACM, 33(11):75–86, November 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/92767.html.

Jacobs:1990:SEI

[119] Paul S. Jacobs and Lisa F. Rau. SCISOR: Extracting information from on-line news. Communications of the ACM, 33(11):88–97, November 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/92769.html.

Premerlani:1990:OOR

[120] William J. Premerlani, Michael R. Blaha, James E. Rumbaugh, and Thomas A. Varwig. An objectoriented relational database. Communications of the ACM, 33(11): 99–109, November 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/92772.html.

Weiss:1990:XSA

[121] Eric A. Weiss. The XXII Self-Assessment: The ethics of computing. Communications of the ACM, 33(11): 110–132, November 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/92780.html.

Wiederhold:1990:OAS

[122] Gio Wiederhold, John McCarthy, and Ed Feigenbaum. Obituary: Arthur Samuel (1901–1990). pioneer in machine learning. Communications of the ACM, 33(11):137–138, November 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:1990:IRR

[123] Peter G. Neumann. Inside RISKS: Risks in computerized elections. Communications of the ACM, 33(11):170, November 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/92787.html.

Wilkes:1990:CPR

[124] Maurice V. Wilkes. Computing perspectives: the rise of the VLSI processor. Communications of the ACM, 33(12):16–17, 44, December 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/96286.html.

Muratore:1990:RTD

[125] John F. Muratore, Troy A. Heindel, Terri B. Murphy, Arthur Rasmussen, and Robert Z. McFarland. Realtime data acquisition at mission control. Communications of the ACM, 33 (12):18–31, December 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/96277.html.

Miller:1990:ESR

[126] Barton P. Miller, Lars Fredriksen, and Bryan So. An empirical study of the reliability of UNIX utilities. Communications of the ACM, 33(12): 32–44, December 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL ftp:

//grilled.cs.wisc.edu/technical_
papers/fuzz-revisited.ps;   ftp:
//grilled.cs.wisc.edu/technical_
papers/fuzz.ps;   http://www.acm.
org/pubs/toc/Abstracts/0001-0782/

96279.html. This is a fascinating paper on what happens when random input streams are fed into important UNIX utilities on several commercial UNIX systems. In some cases, the tests were able to crash the entire operating system. In 1995, a (sadly, unpublished) followup study showed that many of the failures diagnosed in 1990 still had not been repaired in the commercial systems, and that the GNU implementations were generally more robust. Both 1990 and 1995 papers, and the fuzz-generating software, are available at the authors’ FTP

                                        site at ftp://grilled.cs.wisc.edu/
                                        technical_papers/fuzz.ps and ftp:
                                        //grilled.cs.wisc.edu/technical_
                                        papers/fuzz-revisited.ps.

Tanenbaum:1990:EAD

[127] Andrew S. Tanenbaum, Robbert van Renesse, Hans van Staveren, Gregory J. Sharp, Sape J. Mullender, Jack Jansen, and Guido van Rossum. Experiences with the Amoeba distributed operating system. Communications of the ACM, 33(12):46–63, December 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/96281.html.

Gurbaxani:1990:DCN

[128] Vijay Gurbaxani. Diffusion in computing networks: the case of BIT-NET. Communications of the ACM, 33 (12):65–75, December 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/96283.html.

Jacky:1990:IRR

[129] Jonathan Jacky. Inside RISKS: risks in medical electronics. Communications of the ACM, 33(12):138, December 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/96290.html.

Goodman:1991:IPG

[130] Seymour E. Goodman. International perspectives: The globalization of computing: Perspectives on a changing world. Communications of the ACM, 34(1):19–21, January 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Clemons:1991:ESI

[131] Eric K. Clemons. Evaluation of strategic investments in information technology. Communications of the ACM, 34(1):22–36, January 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/99985.html.

Ellis:1991:GSI

[132] C. A. Ellis, S. J. Gibbs, and G. L. Rein. Groupware: Some issues and experiences. Communications of the ACM, 34(1):38–58, January 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/99987.html.

Gurbaxani:1991:IIS

[133] Vijay Gurbaxani and Seungjin Whang. The impact of information systems on organizations and markets. Communications of the ACM, 34(1):59–73, January 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/99990.html.

Gould:1991:MUU

[134] John D. Gould, Stephen J. Boies, and Clayton Lewis. Making usable, useful, productivity-enhancing computer applications. Communications of the ACM, 34(1):74–85, January 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/99993.html.

Jarvenpaa:1991:SDD

[135] Sirkka L. Jarvenpaa, Blake Ives, and Gordon B. Davis. Supply/demand of IS doctorates in the 1990s. Communications of the ACM, 34(1):86–99, January 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/99996.html.

Eerkes:1991:PCS

[136] Gary L. Eerkes. Profiling Computer Science Master’s Programs. Communications of the ACM, 34(1): 100–109, January 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/99999.html.

Lamport:1991:MEP

[137] Leslie Lamport. The mutual exclusion problem has been solved (letter). Communications of the ACM, 34 (1):110–113, January 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:1991:IRCa

[138] Peter G. Neumann. Inside RISKS: The clock grows at midnight. Communications of the ACM, 34(1):170, January 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/100000.html.

Steinberg:1991:VSV

[139] Steve Steinberg. Viewpoint: a student’s view of cryptography in computer science. Communications of the ACM, 34(2):15–17, February 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rice:1991:FCS

[140] John R. Rice. And the first computer science department was? Communications of the ACM, 34(2):19, February 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Press:1991:PCPa

[141] Larry Press. Personal computing: personal computers and the world software market. Communications of the ACM, 34(2):23–28, February 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Soloway:1991:LEQ

[142] Elliot Soloway. Log on education: Quick, where do the computers go? Communications of the ACM, 34(2): 29–33, February 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/102797.html.

Leveson:1991:SSE

[143] Nancy G. Leveson. Software safety: in embedded computer systems. Communications of the ACM, 34(2):34–46, February 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/102799.html.

Yu:1991:SED

[144] K. W. Yu and Tong Lai Yu. Superimposing encrypted data. Communications of the ACM, 34(2):48–54, February 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/102800.html. See [326].

Cristian:1991:UFT

[145] Flaviu Cristian. Understanding faulttolerant distributed systems. Communications of the ACM, 34(2):56–78, February 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/102801.html.

Richard:1991:GIA

[146] Isabelle Richard. On Gerasch’s insertion algorithm. Communications of the ACM, 34(2):79–80, February 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [2].

Neumann:1991:IRCb

[147] Peter G. Neumann. Inside RISKS: Certifying professionals. Communications of the ACM, 34(2):130, February 1991. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/102798.html.

Samuelson:1991:LSI

[148] Pamela Samuelson. Legally speaking: is information property? Communications of the ACM, 34(3):15– 18, March 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/102875.html.

Barlow:1991:EFC

[149] John P. Barlow. Electronic frontier: coming into the country. Communications of the ACM, 34(3):19– 21, March 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/102876.html.

Denning:1991:USV

[150] Dorothy E. Denning, Donn B. Parker, Steven Levy, Eugene Spafford, Paula Hawthorn, Marc Rotenberg, J. J. Buck BloomBecker, and Richard Stallman. The United States vs. Craig Neidorf: a debate on electronic publishing, Constitutional rights and hacking. Communications of the ACM, 34(3):22–43, March 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/102869.html.

Parker:1991:CDD

[151] Donn D. Parker, Stephen Levy, Eugene Spafford, Paula Hawthorn, Marc Rotenberg, J. J. Buck BloomBecker, and Richard Stallman. Colleagues debate Denning’s comments. Communications of the ACM, 34(3):33–41, March 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:1991:DR

[152] Dorothy E. Denning. Denning’s rebuttal. Communications of the ACM, 34(3):42–43, March 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gries:1991:TCD

[153] David Gries. Teaching calculation and discrimination: a more effective curriculum. Communications of the ACM, 34(3):44–55, March 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/102870.html. See letters

[278].

Nussbaum:1991:SPM

[154] Daniel Nussbaum and Anant Agarwal. Scalability of parallel machines. Communications of the ACM, 34(3):56–61, March 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/102871.html.

Teuhola:1991:MSA

[155] Jukka Teuhola and Lutz Wegner. Minimal space, average linear time duplicate deletion. Communications of the ACM, 34(3):62–73, March 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/102872.html.

Hix:1991:HCI

[156] Deborah Hix and Robert S. Schulman. Human-computer interface development tools: A methodology for their evaluation. Communications of the ACM, 34(3):74–87, March 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/102873.html.

Blostein:1991:JPM

[157] Dorothea Blostein and Lippold Haken. Justification of printed music. Communications of the ACM, 34(3):88– 99, March 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/102874.html.

Neumann:1991:IRP

[158] Peter G. Neumann. Inside RISKS: putting on your best interface. Communications of the ACM, 34(3):138, March 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/102877.html.

Press:1991:PCPb

[159] Larry Press. Personal computing: Personal computers as research tools: SIG CHI and CSCW’90. Communications of the ACM, 34(4):19– 25, April 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/103087.html.

Fox:1991:GEI

[160] Edward A. Fox. Guest editor’s introduction: Standards and the emergence of digital multimedia systems. Communications of the ACM, 34 (4):26–29, April 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/103086.html. See [273].

Wallace:1991:JSP

[161] Gregory K. Wallace. The JPEG still picture compression standard. Communications of the ACM, 34 (4):30–44, April 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/103089.html.

Cockroft:1991:NPJ

[162] Greg Cockroft and Leo Hourvitz. NeXTstep: Putting JPEG to multiple uses. Communications of the ACM, 34(4):45, 116, April 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/103097.html.

LeGall:1991:MVC

[163] Didier Le Gall. MPEG: a video compression standard for multimedia applications. Communications of the ACM, 34(4):46–58, April 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/103090.html.

Liou:1991:OKV

[164] Ming Liou. Overview of the p × 64 kbit/s video coding standard. Communications of the ACM, 34(4):59– 63, April 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/103091.html.

Harney:1991:IVP

[165] Kevin Harney, Mike Keith, Gary Lavelle, Lawrence D. Ryan, and Daniel J. Stark. The i750 video processor: a total multimedia solution. Communications of the ACM, 34(4):64–78, April 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/103092.html.

Sijstermans:1991:CFM

[166] Frans Sijstermans and Jan van der Meer. CD-I full-motion video encoding on a parallel computer. Communications of the ACM, 34(4):81– 91, April 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/103093.html.

Lippman:1991:FSI

[167] Andrew Lippman. Feature sets for interactive images. Communications of the ACM, 34(4):92–102, April 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/103094.html.

Liebhold:1991:TOE

[168] Michael Liebhold and Eric M. Hoffert. Toward an open environment for digital video. Communications of the ACM, 34(4):103–112, April 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/103095.html.

Hoffert:1991:DNS

[169] Eric M. Hoffert and Greg Gretsch. The digital news system at EDUCOM: a convergence of interactive computing, newspapers, television and high-speed networks. Communications of the ACM, 34(4):113–116, April 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/103096.html.

Borenstein:1991:MEM

[170] Nathaniel S. Borenstein. Multimedia electronic mail: will the dream become a reality? Communications of the ACM, 34(4):117–119, April 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/103098.html.

Neumann:1991:IRI

[171] Peter G. Neumann. Inside Risks: interpreting (MIS)information. Communications of the ACM, 34(4):144, April 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/103088.html.

Denning:1991:SCE

[172] Peter J. Denning. Short citedness (editorial). Communications of the ACM, 34(5):17–19, May 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wilkes:1991:CPS

[173] Maurice V. Wilkes. Computing perspectives: software and the programmer. Communications of the ACM, 34(5):23–24, May 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/103169.html.

Rettig:1991:PPT

[174] Marc Rettig. Practical programmer: testing made palatable. Communications of the ACM, 34(5):25– 29, May 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/103170.html.

Freeman:1991:BFF

[175] Peter A. Freeman and Marie-Claude Gaudel. Building a foundation for the future of software engineering. Communications of the ACM, 34(5):30–33, May 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/103168.html.

Devanbu:1991:LKB

[176] Premkumar Devanbu, Ronald J. Brachman, Peter G. Selfridge, and Bruce W. Ballard. LaSSIE: a knowledge-based software information system. Communications of the ACM, 34(5):34–49, May 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/103172.html.

Gabrielian:1991:MSR

[177] Armen Gabrielian and Matthew K. Franklin. Multilevel specification of real time systems. Communications of the ACM, 34(5):50–60, May 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/103173.html.

Lafontaine:1991:EFS

[178] Christine Lafontaine, Yves Ledru, and Pierre-Yves Schobbens. An experiment in formal software development: using the B theorem prover on a VDM case study. Communications of the ACM, 34(5):62–71, 87, May 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/103174.html.

Wileden:1991:SLI

[179] Jack C. Wileden, Alexander L. Wolf, William R. Rosenblatt, and Peri L. Tarr. Specification-level interoperability. Communications of the ACM, 34(5):72–87, May 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/103175.html.

Prieto-Diaz:1991:IFC

[180] Rubén Prieto-Díaz. Implementing faceted classification for software reuse. Communications of the ACM, 34(5):88–97, May 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/103176.html.

Neumann:1991:IRE

[181] Peter G. Neumann. Inside RISKS: Expecting the unexpected mayday. Communications of the ACM, 34(5): 128, May 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/103171.html.

Crawford:1991:WCU

[182] Diane Crawford. From Washington: CEOs unite to influence U.S. technology policy. Communications of the ACM, 34(6):15–18, June 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/103702.html.

Samuelson:1991:LSF

[183] Pamela Samuelson. Legally speaking: First Amendment rights for information providers? Communications of the ACM, 34(6):19–23, June 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/103703.html.

Goodman:1991:IPS

[184] S. E. Goodman and W. K. McHenry. International perspectives: The Soviet computer industry: a tale of two sectors. Communications of the ACM, 34(6):25–28, June 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/122192.html.

Kotov:1991:PS

[185] Vadim E. Kotov. Project START. Communications of the ACM, 34 (6):30–31, June 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/103706.html.

Kotov:1991:CMP

[186] Vadim E. Kotov. Concurrency + modularity + programmability = MARS. Communications of the ACM, 34(6):32–45, June 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/103707.html.

Tyugu:1991:TNG

[187] Enn Tyugu. Three new-generation software environments. Communications of the ACM, 34(6):46–59, June 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/103708.html. The system PRIZ

has been under development at the Institute of Cybernetics, Tallinn, Estonia since mid-seventies. This article describes three environments of different sophistication based on this system: ExpertPRIZ, a simple expert system shell; C-PRIZ, a language integrating imperative programming (C); and NUT (New UTopist; Utopist was the initial programming language of PRIZ), an object-oriented environment. Data elements in Utopist (attributes of objects) can be bound by both interand intra-object relations. These relations are used by a propositional theorem prover to generate a proof that it is possible to compute a certain datum given some other data. This (constructive) proof is used to synthesize a program which computes the datum. So PRIZ has a quite non-standard logic component: compile-time proof and program generation instead of run-time clause resolution. These technique are called Propositional Logic Programming and Structural Program Synthesis. NUT is a prototype-based (as opposed to class-based) language: any object can be used as a template (type) for a new object (and of course as a component of a new object). Limited polymorphism is supported through a generic type (any). It is notable that some of the modern programming paradigms (constraint programming, object-orientation, employing logic in computation) have been considered so early.

Narinyani:1991:IST

[188] Alexander S. Narin’yani. Intelligent software technology for the new decade. Communications of the ACM, 34(6):60–67, June 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/103709.html.

Tucker:1991:SAI

[189] Allen B. Tucker, Bruce H. Barnes, Robert M. Aiken, Keith Barker, Kim B. Bruce, J. Thomas Cain, Susan E. Conry, Gerald L. Engel, Richard G. Epstein, Doris K. Lidtke, Michael C. Mulder, Jean B. Rogers, Eugene H. Spafford, and A. Joe Turner. A summary of the ACM/IEEE-CS joint curriculum task force report: Computing curricula 1991. Communications of the ACM, 34(6):68–84, June 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/103710.html.

Charette:1991:IRR

[190] Robert N. Charette. Inside RISKS: Risks with risk analysis. Communications of the ACM, 34(6):106, June 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/103705.html.

Rettig:1991:PPN

[191] Marc Rettig. Practical programmer: Nobody reads documentation. Communications of the ACM, 34(7):19– 24, July 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/105788.html.

Press:1991:PCN

[192] Larry Press. Personal computing: Notes from COMDEX. Communications of the ACM, 34(7):25– 29, July 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/105789.html.

Kerlow:1991:ADC

[193] Isaac V. Kerlow. Art + design + computer graphics technology. Communications of the ACM, 34(7):30– 39, July 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/105786.html.

Nunamaker:1991:EMS

[194] J. F. Nunamaker, Alan R. Dennis, Joseph S. Valacich, Douglas R. Vogel, and Joey F. George. Electronic meeting systems to support group work. Communications of the ACM, 34(7):40–61, July 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/105793.html.

George:1991:ECC

[195] Joey F. George and John L. King. Examining the computing and centralization debate. Communications of the ACM, 34(7):62–72, July 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/105796.html.

Phillips:1991:MGM

[196] Richard L. Phillips. MediaView: a general multimedia digital publication system. Communications of the ACM, 34(7):74–83, July 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/105799.html.

Neumann:1991:IRCc

[197] Peter G. Neumann. Inside RISKS: Computers, ethics and values. Communications of the ACM, 34(7):106, July 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/105791.html.

Wilkes:1991:CPR

[198] Maurice V. Wilkes. Computing perspectives: Revisiting computer security in the business world. Communications of the ACM, 34(8):19– 22, August 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/108519.html.

Barlow:1991:EFP

[199] John Perry Barlow. Electronic frontier: Private life in cyberspace. Communications of the ACM, 34(8):23– 25, August 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/108520.html.

Schoppers:1991:RTK

[200] Marcel Schoppers. Real-time knowledgebased control systems. Communications of the ACM, 34(8):26–30, August 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/108517.html.

Chandrasekaran:1991:RTD

[201] B. Chandrasekaran, R. Bhatnagar, and D. D. Sharma. Real-time disturbance control. Communications of the ACM, 34(8):32–47, August 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/108524.html.

Payton:1991:IRT

[202] David W. Payton and Thomas E. Bihari. Intelligent real-time control of robotic vehicles. Communications of the ACM, 34(8):48–63, August 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/108527.html.

Kohn:1991:DCA

[203] Wolf Kohn. Declarative control architecture. Communications of the ACM, 34(8):64–79, August 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/108530.html.

Paul:1991:RPS

[204] C. J. Paul, Anurag Acharya, Bryan Black, and Jay K. Strosnider. Reducing problem-solving variance to improve predictability. Communications of the ACM, 34(8):80–93, August 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/108531.html.

Kottemann:1991:SAM

[205] Jeffrey E. Kottemann, Michael D. Gordon, and Jack W. Stott. A storage and access manager for ill-structured data. Communications of the ACM, 34(8):94–103, August 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/108534.html.

Bersoff:1991:ILC

[206] Edward H. Bersoff and Alan M. Davis. Impacts of life cycle models on software configuration management. Communications of the ACM, 34(8):104–118, August 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/108537.html.

Rosenberg:1991:IRM

[207] Ronni Rosenberg. Inside RISKS: Mixed signals about social responsibility. Communications of the ACM, 34(8):146, August 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/108521.html.

Goodman:1991:IPU

[208] Seymour E. Goodman. International perspectives: From under the rubble: Computing and the resuscitation of Romania. Communications of the ACM, 34(9):19–22, September 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/114715.html.

Soloway:1991:LEH

[209] Elliot Soloway. Log on education: How the Nintendo generation learns. Communications of the ACM, 34(9): 23–26, 95, September 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/114717.html.

Foderaro:1991:LI

[210] John Foderaro. LISP: introduction. Communications of the ACM, 34(9):27, September 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/114670.html.

Gabriel:1991:CIO

[211] Richard P. Gabriel, Jon L. White, and Daniel G. Bobrow. CLOS: integrating object-oriented and functional programming. Communications of the ACM, 34(9):28–38, September 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/114671.html.

Sinclair:1991:PL

[212] Kenneth H. Sinclair and David A. Moon. The philosophy of Lisp. Communications of the ACM, 34(9): 40–47, September 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/133119.html.

Gajek:1991:MS

[213] Oliver Gajek. The Metal system. Communications of the ACM, 34(9): 46–47, September 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/114673.html.

Layer:1991:LS

[214] D. Kevin Layer and Chris Richardson. Lisp systems in the 1990s. Communications of the ACM, 34(9): 48–57, September 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/114674.html.

McKay:1991:CCL

[215] Scott McKay. CLIM: the Common Lisp interface manager. Communications of the ACM, 34(9):58–59, September 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/114675.html.

Muller:1991:LLT

[216] Hans Muller. LispView: leverage through integration. Communications of the ACM, 34(9):59–61, September 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/126333.html.

Barber:1991:DGL

[217] Richard Barber and George Imlah. Delivering the goods with Lisp. Communications of the ACM, 34(9):

61–63, September 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/114677.html.

Allard:1991:RTP

[218] James R. Allard and Lowell B. Hawkinson. Real-time programming in Common Lisp. Communications of the ACM, 34(9):64–69, September 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/114679.html.

Stewart:1991:BNS

[219] Rocky L. Stewart. Biosphere 2 nerve system. Communications of the ACM, 34(9):69–71, September 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/114680.html.

Corbato:1991:BSW

[220] Fernando J. Corbató. On building systems that will fail. Communications of the ACM, 34(9):72–81, September 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/114686.html.

Frenkel:1991:IFJ

[221] Karen A. Frenkel. An interview with Fernando José Corbató. Communications of the ACM, 34(9):82–90, September 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/114688.html.

Neumann:1991:IRA

[222] Peter G. Neumann. Inside RISKS: The not-so-accidental holist. Communications of the ACM, 34(9):122, September 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/214895.html.

Wilkes:1991:CPC

[223] Maurice V. Wilkes. Computing perspectives: Could it have been predicted? Communications of the ACM, 34(10):19–21, October 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/125288.html.

Samuelson:1991:LSD

[224] Pamela Samuelson. Legally speaking: Digital media and the law. Communications of the ACM, 34 (10):23–28, October 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/125289.html.

Cattell:1991:NGD

[225] R. G. G. Cattell. Next-generation database systems. Communications of the ACM, 34(10):30–33, October 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/125228.html.

Deux:1991:S

[226] O. Deux et al. The O_{2} system. Communications of the ACM, 34 (10):34–48, October 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/125238.html.

Lamb:1991:ODS

[227] Charles Lamb, Gordon Landis, Jack Orenstein, and Dan Weinreb. The ObjectStore database system. Communications of the ACM, 34(10):50– 63, October 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/125244.html.

Butterworth:1991:GOD

[228] Paul Butterworth, Allen Otis, and Jacob Stein. The GemStone object database management system. Communications of the ACM, 34 (10):64–77, October 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/125254.html.

Stonebraker:1991:PNG

[229] Michael Stonebraker and Greg Kemnitz. The POSTGRES next-generation database management system. Communications of the ACM, 34(10):78– 92, October 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/125262.html.

Lohman:1991:ESO

[230] Guy M. Lohman, Bruce Lindsay, Hamid Pirahesh, and K. Bernhard Schiefer. Extensions to Starburst: objects, types, functions, and rules. Communications of the ACM, 34(10): 94–109, October 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/125266.html.

Silberschatz:1991:DSA

[231] Avi Silberschatz, Michael Stonebraker, and Jeff Ullman. Database systems: achievements and opportunities. Communications of the ACM, 34 (10):110–120, October 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/125272.html.

Simons:1991:RAR

[232] Barbara Simons, Dennis J. Frailey, A. Joe Turner, Stuart H. Zweben, and Peter J. Denning. Report: An ACM response: The scope and directions for computer science. Communications of the ACM, 34(10): 121–131, October 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/125276.html.

Weissman:1991:IRN

[233] Clark Weissman. Inside RISKS: a national debate on encryption exportability. Communications of the ACM, 34(10):162, October 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/125294.html.

White:1991:PLM

[234] John R. White. President’s letter: Maintaining free access to scientific and professional activities. Communications of the ACM, 34(11):13– 14, 16, November 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gore:1991:VHP

[235] Senator Albert Gore, Jr. Viewpoint: HPCC policy champion forsees networked nation. Communications of the ACM, 34(11):15–16, November 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:1991:WSP

[236] Diane Crawford. From Washington: Supercomputing policy links science and the ‘C’ word. Communications of the ACM, 34(11): 25–28, November 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/125512.html.

Lander:1991:RCM

[237] Eric S. Lander, Robert Langridge, and Damian M. Saccocio. A report on computing in molecular biology: Mapping and interpreting biological information. Communications of the ACM, 34(11):32–39, November 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/125491.html.

Frenkel:1991:MSA

[238] Karen A. Frenkel. A monumental science adventure: The Human Genome Project and informatics. Communications of the ACM, 34(11): 40–51, November 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/125492.html.

Pancake:1991:SSP

[239] Cherri M. Pancake. Software support for parallel computing: Where are we headed? Communications of the ACM, 34(11):52–64, November 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/125493.html.

Newcomb:1991:HHT

[240] Steven R. Newcomb, Neill A. Kipp, and Victoria T. Newcomb. The “Hy-Time”: hypermedia/time-based document structuring language. Communications of the ACM, 34(11): 67–83, November 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/125495.html.

Newborn:1991:ANA

[241] Monty Newborn and Danny Kopec. The 21st ACM North American computer chess championship. Communications of the ACM, 34(11): 84–92, November 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/125497.html.

Silverman:1991:MDC

[242] Robert D. Silverman. Massively distributed computing and factoring large integers: a pracnique. Communications of the ACM, 34(11): 94–103, November 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/125504.html.

Lumpp:1991:CCA

[243] James E. Lumpp, Jr., Samuel A. Fineberg, Wayne G. Nation, Thomas L. Casavant, Edward C. Bronson, Howard Jay Siegel, Pierre H. Pero, Thomas Schwederski, and Dan C. Marinescu. CAPS: a coding aid for PASM. Communications of the ACM, 34(11): 104–117, November 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/125507.html.

Dittmer:1991:NFH

[244] I. Dittmer. Note on fast hashing of variable length text strings. Communications of the ACM, 34(11):118, November 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Points out that Pearson’s hashing algorithm [63] was discovered fourteen years earlier by this author [1949]. See also comments in [245, 246, 247].

Savoy:1991:NFH

[245] Jacques Savoy. Note on fast hashing of variable length text strings. Communications of the ACM, 34(11): 118–119, November 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Suggests an improvement to Pearson’s hashing algorithm [63] that avoids secondary clustering. Exhibits a key set for which Pearson’s algorithm produces alarming clustering. See also comments in [244, 246, 247], and early work in [1949].

Litsios:1991:NFH

[246] James Litsios. Note on fast hashing of variable length text strings. Communications of the ACM, 34(11): 119–120, November 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Suggests a simple extension of Pearson’s hashing algorithm [63] that supports non-character data. See also comments in [244, 245, 247], and early work in [1949].

Pearson:1991:NFH

[247] Peter Pearson. Note on fast hashing of variable length text strings. Communications of the ACM, 34(11):120, November 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Responds to several comments [244, 246, 245] on the author’s earlier paper [63], and early work [1949].

Neumann:1991:IRH

[248] Peter G. Neumann. Inside RISKS: The human element. Communications of the ACM, 34(11):150, November 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/125514.html.

Press:1991:PCW

[249] Larry Press. Personal computing: Wide-area collaboration. Communications of the ACM, 34(12):21–24, December 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/125403.html.

Goodman:1991:IPC

[250] Seymour E. Goodman. International perspectives: Computing in a less-developed country. Communications of the ACM, 34(12):25–29, December 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/125408.html.

Grudin:1991:CI

[251] Jonathan Grudin. CSCW introduction. Communications of the ACM, 34 (12):30–34, December 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/125320.html.

Ishii:1991:TOS

[252] Hiroshi Ishii and Naomi Miyake. Toward an open shared workspace: computer and video fusion approach of TeamWorkStation. Communications of the ACM, 34(12):37–50, December 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/125321.html.

Francik:1991:PIW

[253] Ellen Francik, Susan Ehrlich Rudman, Donna Cooper, and Stephen Levine. Putting innovation to work: adoption strategies for multimedia communication systems. Communications of the ACM, 34(12):52–63, December 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/125322.html.

Kyng:1991:DCC

[254] Morten Kyng. Designing for cooperation: cooperating in design. Communications of the ACM, 34(12): 65–73, December 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/125323.html.

Perin:1991:ESF

[255] Constance Perin. Electronic social fields in bureaucracies. Communications of the ACM, 34(12): 74–82, December 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/125324.html.

Rosenberg:1991:UNS

[256] Laurence C. Rosenberg. Update on National Science Foundation funding of the “Collaboratory”. Communications of the ACM, 34(12):83, December 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/125392.html.

Kling:1991:CCC

[257] Rob Kling. Cooperation, coordination and control in computer-supported work. Communications of the ACM, 34 (12):83–88, December 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/125396.html.

Norman:1991:CCC

[258] Donald A. Norman. Collaborative computing: collaboration first, computing second. Communications of the ACM, 34(12):88–90, December 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/125398.html.

LaMalva:1991:APC

[259] Patrick M. LaMalva and David R. Reyes-Guerra. Accredited programs in computing, followed by year of initial accreditation. Communications of the ACM, 34(12):91–96, December 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Peterson:1991:MED

[260] W. Wesley Peterson. More errordetecting decimal digits. Communications of the ACM, 34(12):110–113, December 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:1991:IRCd

[261] Peter G. Neumann. Inside RISKS: Collaborative efforts. Communications of the ACM, 34(12):162, December 1991. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/125411.html.

Frailey:1991:AHA

[262] Frailey, Turner, and Zweben. Adequacy and health of academic-industry coupling. Communications of the ACM, 34(??):??, ???? 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Simons:1991:BRA

[263] Simons. Building a research agenda. Communications of the ACM, 34(??): ??, ???? 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stallman:1992:ASP

[264] R. Stallman and S. Garfinkel. Against software patents. Communications of the ACM, 35(1):17, January 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Barlow:1992:EFG

[265] John Perry Barlow. Electronic frontier: The great work. Communications of the ACM, 35(1):25–28, January 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/129625.html.

Rettig:1992:PPI

[266] Marc Rettig. Practical programmer: Interface design when you don’t know how. Communications of the ACM, 35(1):29–34, January 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/129626.html.

Haan:1992:IHS

[267] Bernard J. Haan, Paul Kahn, Victor A. Riley, James H. Coombs, and Norman K. Meyrowitz. IRIS hypermedia services. Communications of the ACM, 35(1):36–51, January 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/129618.html.

Green:1992:EDS

[268] James L. Green. The evolution of DVI system software. Communications of the ACM, 35(1):52–67, January 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/129619.html.

Doerschler:1992:RBS

[269] Jeffrey S. Doerschler and Herbert Freeman. A rule-based system for dense-map name placement. Communications of the ACM, 35(1):68–79, January 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/129620.html.

Lapalme:1992:GGI

[270] Guy Lapalme, Jean-Marc Rousseau, Suzanne Chapleau, Michel Cormier, Pierre Cossette, and Serge Roy. GeoRoute: a geographic information system for transportation applications. Communications of the ACM, 35(1):80–88, January 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/129624.html.

Subramanian:1992:CDT

[271] Girish H. Subramanian, John Nosek, Sankaran P. Raghunathan, and Santosh S. Kanitkar. A comparison of the decision table and tree. Communications of the ACM, 35(1):89–94, January 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/129621.html.

Bunton:1992:PDM

[272] Suzanne Bunton and Gaetano Borriello. Practical dictionary management for hardware data compression. Communications of the ACM, 35(1):95–104, January 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/129622.html.

Fox:1992:PMP

[273] Edward A. Fox, Lenwood S. Heath, Qi Fan Chen, and Amjad M. Daoud. Practical minimal perfect hash functions for large databases. Communications of the ACM, 35(1):105–121, January 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/129623.html. This is the first

published algorithm for computing minimal perfect hash functions for lists of millions of words; previous algorithms were computationally infeasible for more than a few hundred words.

Frand:1992:SAU

[274] Jason L. Frand and Julia A. Britt. Seventh Annual UCLA Survey of Business School Computer Usage. Communications of the ACM, 35(1): 123–131, January 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/129628.html.

Gries:1992:TR

[275] D. Gries and D. Marsh. The 1989-1990 Taulbee report. Communications of the ACM, 35(1):132, January 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gries:1992:TS

[276] David Gries and Dorothy Marsh. The 1989–90 Taulbee survey. Communications of the ACM, 35(1):133–143, January 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/129629.html.

Neumann:1992:IRWa

[277] Peter G. Neumann. Inside risks: What’s in a name? Communications of the ACM, 35(1):186, January 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/129627.html.

Marquart:1992:AFR

[278] R. Gary Marquart, Anthony Fedanzo, Robert V. Binder, C. J. Paul, and David Ibbetson. ACM Forum: Readers take exception to rejoinder: And the Surgeon General warns . . .. Communications of the ACM, 35(2):22, 26, February 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [153].

Wolinsky:1992:WPT

[279] Carol Wolinsky and James Sylvester. From Washington: Privacy in the telecommunications age. Communications of the ACM, 35(2):23–25, February 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/129638.html.

Samuelson:1992:LSC

[280] Pamela Samuelson. Legally speaking: Copyright law and electronic compilations of data. Communications of the ACM, 35(2):27–32, February 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/129639.html.

Igbaria:1992:DME

[281] Magid Igbaria and Jeffrey H. Greenhaus. Determinants of MIS employees’ turnover intentions: A structural equation model. Communications of the ACM, 35(2):34–49, February 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/129631.html.

Lederer:1992:NMG

[282] Albert L. Lederer and Jayesh Prasad. Nine management guidelines for better cost estimating. Communications of the ACM, 35(2):51–59, February 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/129632.html.

Clark:1992:CSM

[283] Thomas D. Clark, Jr. Corporate systems management: an overview and research perspective. Communications of the ACM, 35(2):61–75, February 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/129633.html.

Liker:1992:DTM

[284] Jeffrey K. Liker, Mitchell Fleischer, Mitsuo Nagamachi, and Michael S. Zonnevylle. Designers and their machines: CAD use and support in the U.S. and japan. Communications of the ACM, 35(2):77–95, February 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/129634.html.

Gelernter:1992:CLT

[285] David Gelernter and Nicholas Carriero. Coordination languages and their significance. Communications of the ACM, 35(2):97–107, February 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/129635.html.

Choobineh:1992:FBA

[286] Joobin Choobineh, Michael V. Mannino, and Veronica P. Tseng. A form-based approach for database analysis and design. Communications of the ACM, 35(2):108–120, February 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/129636.html.

Gray:1992:ECF

[287] Robert W. Gray, Vincent P. Heuring, Steven P. Levi, Anthony M. Sloane, and William M. Waite. Eli: a complete, flexible compiler construction system. Communications of the ACM, 35(2): 121–130, February 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/129637.html.

Goodman:1992:PAI

[288] Seymour E. Goodman. Political activity and international computer networks. Communications of the ACM, 35(2):174, February 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/129640.html.

Quesada:1992:ANP

[289] Antonio R. Quesada, Paul Pritchard, and Richard E. James, III. Additional notes on a practical sieve algorithm. Communications of the ACM, 35(3):11–14, March 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wilkes:1992:CPC

[290] Maurice V. Wilkes. Computing perspectives: Charles Babbage: The great uncle of computing? Communications of the ACM, 35(3):15– 16, 21, March 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/214839.html.

Press:1992:PCD

[291] Larry Press. Personal computing: Dynabook revisited: Portable computers past, present and future. Communications of the ACM, 35(3):25– 32, March 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/214840.html.

Cohen:1992:FST

[292] Jacques Cohen. First specialize, then generalize: introduction. Communications of the ACM, 35(3):34–39, March 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/214842.html.

Robinson:1992:LLP

[293] J. A. Robinson. Logic and logic programming. Communications of the ACM, 35(3):40–65, March 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/

                                        pubs/toc/Abstracts/0001-0782/131296.
                                        html.

Grant:1992:ILP

[294] John Grant and Jack Minker. The impact of logic programming on databases. Communications of the ACM, 35(3):66–81, March 1992. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/131297.html.

Furukawa:1992:LPI

[295] Koichi Furukawa. Logic programming as the integrator of the Fifth Generation Computer Systems project. Communications of the ACM, 35 (3):82–92, March 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/131298.html.

Warren:1992:MLP

[296] David S. Warren. Memoing for logic programs. Communications of the ACM, 35(3):93–111, March 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/toc/Abstracts/0001-0782/131299.
                                        html.

Munakata:1992:NIS

[297] Toshinori Munakata. Notes on implementing sets in Prolog. Communications of the ACM, 35(3):112– 120, March 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/131300.html.

Linn:1992:CCS

[298] Marcia C. Linn and Michael J. Clancy. The case for case studies of programming problems. Communications of the ACM, 35(3):121–132, March 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/131301.html.

Neumann:1992:IRI

[299] Peter G. Neumann. Inside RISKS: inside “RISKS of RISKS”. Communications of the ACM, 35(3):160, March 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/214843.html.

LaRovere:1992:IPC

[300] Renata L. La Rovere and S. E. Goodman. International perspectives: Computing in the Brazilian Amazon. Communications of the ACM, 35(4):21–24, April 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/129862.html.

Norman:1992:ASD

[301] Ronald J. Norman and Gene Forte. Automating the software development process: CASE in the 90’s. Communications of the ACM, 35(4):27, April 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/129853.html.

Forte:1992:SAS

[302] Gene Forte and Ronald J. Norman. A self-assessment by the software engineering community. Communications of the ACM, 35(4):28– 32, April 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/129854.html.

Shepard:1992:VSP

[303] Terry Shepard, Steve Sibbald, and Colin Wortley. A visual software process language. Communications of the ACM, 35(4):37–44, April 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/129855.html.

Huff:1992:ERC

[304] Clifford C. Huff. Elements of a realistic CASE tool adoption budget. Communications of the ACM, 35(4):45–54, April 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/129856.html.

Maiden:1992:ERS

[305] Neil A. Maiden and Alistair G. Sutcliffe. Exploiting reusable specifications through analogy. Communications of the ACM, 35(4):55–64, April 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/129857.html.

Tate:1992:CTM

[306] Graham Tate, June Verner, and Ross Jeffery. CASE: a testbed for modeling, measurement and management. Communications of the ACM, 35(4):65–72, April 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/129858.html.

Baxter:1992:DMS

[307] Ira D. Baxter. Design maintenance systems. Communications of the ACM, 35(4):73–89, April 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/129859.html.

Vessey:1992:EVP

[308] Iris Vessey, Sirkka L. Jarvenpaa, and Noam Tractinsky. Evaluation of vendor products: CASE tools as methodology companions. Communications of the ACM, 35(4):90– 105, April 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/129860.html.

Silverman:1992:SEC

[309] Barry G. Silverman. Survey of expert critiquing systems: practical and theoretical frontiers. Communications of the ACM, 35(4):106– 127, April 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/129861.html.

Sammet:1992:FGH

[310] Jean Sammet. Farewell to Grace Hopper: end of an era! Communications of the ACM, 35(4):128– 131, April 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/214846.html.

Baeza-Yates:1992:TCN

[311] Ricardo Baeza-Yates, Fred T. Krogh, Bernard Ziegler, Peter R. Sibbald, and Daniel M. Sunday. Technical correspondence: Notes on a very fast substring search algorithm. Communications of the ACM, 35(4):132–137, April 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rotenberg:1992:IRP

[312] Marc Rotenberg. Inside RISKS: Protecting privacy. Communications of the ACM, 35(4):164, April 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/129863.html.

Rettig:1992:SPI

[313] Marc Rettig. A succotash of projections and insights. Communications of the ACM, 35(5):25–30, May 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/129886.html.

Glinert:1992:CPD

[314] Ephraim P. Glinert and Bryant W. York. Computers and people with disabilities. Communications of the ACM, 35(5):32–35, May 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/129876.html.

Brown:1992:ATC

[315] Carl Brown. Assistive technology computers and persons with disabilities. Communications of the ACM, 35(5):36–45, May 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/129877.html.

Alm:1992:PCM

[316] Norman Alm, John L. Arnott, and Alan F. Newell. Prediction and conversational momentum in an augmentative communication system. Communications of the ACM, 35(5):46– 57, May 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/129878.html.

Alm:1992:IAT

[317] Norman Alm, John L. Arnott, and Alan F. Newell. Input acceleration techniques for severely physically disabled nonspeakers using a communication system. Communications of the ACM, 35(5):54–55, May 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/129879.html.

Pausch:1992:GCC

[318] Randy Pausch and Ronald D. Williams. Giving CANDY to children: usertailored gesture input driving an articulator-based speech synthesizer. Communications of the ACM, 35(5): 58–66, May 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/129880.html.

Demasco:1992:GTC

[319] Patrick W. Demasco and Kathleen F. McCoy. Generating text from compressed input: an intelligent interface for people with severe motor impairments. Communications of the ACM, 35(5):68–78, May 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/129881.html.

Kukich:1992:SCT

[320] Karen Kukich. Spelling correction for the telecommunications network for the deaf. Communications of the ACM, 35(5):80–90, May 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/129882.html.

Dipner:1992:AA

[321] Randy Dipner. ADA and the ACM. Communications of the ACM, 35(5): 90, May 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/129883.html.

Davies:1992:AMS

[322] Daniel K. Davies and Randy W. Dipner. ACM membership survey of disability and disability issues. Communications of the ACM, 35(5):91–93, May 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/129884.html.

Anderson:1992:ACE

[323] Ronald E. Anderson. ACM code of ethics and professional conduct. Communications of the ACM, 35(5):94– 99, May 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/129885.html.

Neumann:1992:SS

[324] Peter G. Neumann. Survivable systems. Communications of the ACM, 35(5):130, May 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/129887.html.

Tichy:1992:TCC

[325] Walter F. Tichy, Michael Philippsen, Phil Hatcher, and James L. Frankel. Technical correspondence: a critique of the programming language C*. Communications of the ACM, 35(6):21–24, June 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Boyd:1992:RDE

[326] Colin Boyd and Tong Lai Yu. Remarks on a data encryption scheme of Yu and Yu (letters). Communications of the ACM, 35(6):24–25, June 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [144].

Press:1992:CD

[327] Larry Press. Collective dynabases. Communications of the ACM, 35 (6):26–32, June 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/129898.html.

Samuelson:1992:DIP

[328] Pamela Samuelson, Michel Denber, and Robert J. Glushko. Developments on the intellectual property front. Communications of the ACM, 35(6):33–39, June 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/129899.html.

Hart:1992:VNE

[329] John C. Hart. Visualization in networked environments. Communications of the ACM, 35(6):42–43, June 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/129889.html.

Smarr:1992:M

[330] Larry Smarr and Charles E. Catlett. Metacomputing. Communications of the ACM, 35(6):44–52, June 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/toc/Abstracts/0001-0782/129890.
                                        html.

Mercurio:1992:DLI

[331] Philip J. Mercurio, T. T. Elvins, Stephen J. Young, Philip S. Cohen, Kevin R. Fall, and Mark H. Ellisman. The distributed laboratory: An interactive visualization environment for electron microscopy and 3D imaging. Communications of the ACM, 35(6):54–63, June 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/129891.html.

Cruz-Neira:1992:CAV

[332] Carolina Cruz-Neira, Daniel J. Sandin, Thomas A. DeFanti, Robert V. Kenyon, and John C. Hart. The CAVE: audio visual experience automatic virtual environment. Communications of the ACM, 35(6):64– 72, June 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/129892.html.

Carlbom:1992:MAE

[333] Ingrid Carlbom, William M. Hsu, Gudrun Klinker, Richard Szeliski, Keith Waters, Michael Doyle, Jim Gettys, Kristen M. Harris, Thomas M. Levergood, Ricky Palmer, Larry Palmer, Marc Picart, Demetri Terzopoulos, David Tonnesen, Michael Vannier, and Greg Wallace. Modeling and analysis of empirical data in collaborative environments. Communications of the ACM, 35(6):74–84, June 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/129893.html.

DeWitt:1992:PDS

[334] David DeWitt and Jim Gray. Parallel database systems: the future of high performance database systems. Communications of the ACM, 35(6):85–98, June 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/129894.html.

Higa:1992:OOM

[335] Kunihiko Higa, Mike Morrison, Joline Morrison, and Olivia R. Liu Sheng. An object-oriented methodology for knowledge base/database coupling. Communications of the ACM, 35(6):99–113, June 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/129895.html.

Ferguson:1992:BTD

[336] David E. Ferguson. Bit-Tree: a data structure for fast file processing. Communications of the ACM, 35(6):114–120, June 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/129896.html.

Heckel:1992:DSP

[337] Paul Heckel. Debunking the software patent myths. Communications of the ACM, 35(6):121–140, June 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/129897.html.

Neumann:1992:LYP

[338] Peter G. Neumann. Leap-year problems. Communications of the ACM, 35(6):162, June 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/129900.html.

Soloway:1992:BCY

[339] Elliot Soloway. Buying computers for your school: a guide for the perplexed. Communications of the ACM, 35(7):21–24, July 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/129909.html.

Barlow:1992:DPP

[340] John Perry Barlow. Decrypting the puzzle palace. Communications of the ACM, 35(7):25–31, July 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/129910.html.

Anonymous:1992:DES

[341] Anonymous. Debating encryption standards. Communications of the ACM, 35(7):32–34, July 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/129903.html.

NIST:1992:DSS

[342] National Institute of Standards and Technology (NIST). The Digital Signature Standard, proposed by NIST. Communications of the ACM, 35(7):36–40, July 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/129904.html.

Rivest:1992:RNP

[343] Ronald L. Rivest, Martin E. Hellman, John C. Anderson, and John W. Lyons. Responses to NIST’s proposal. Communications of the ACM, 35(7):41–54, July 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/129905.html.

Vatti:1992:GSP

[344] Bala R. Vatti. A generic solution to polygon clipping. Communications of the ACM, 35(7):56– 63, July 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/129906.html.

Chokhani:1992:TPE

[345] Santosh Chokhani. Trusted products evaluation. Communications of the ACM, 35(7):64–76, July 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/129907.html.

Sinha:1992:CSC

[346] Alok Sinha. Client-server computing: Current technology review. Communications of the ACM, 35(7):77– 98, July 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/129908.html.

Hesselink:1992:TCD

[347] Wim Hesselink, Jan Jongejan, Jukka Teuhola, and Lutz Wegner. Technical correspondence: Duplicate deletion derived. Communications of the ACM, 35(7):99–107, July 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:1992:ACL

[348] Peter G. Neumann. Aggravation by computer: life, death, and taxes. Communications of the ACM, 35(7): 122, July 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/129911.html.

Wilkes:1992:CPA

[349] Maurice V. Wilkes. Computing perspectives: Artificial intelligence as the year 2000 approaches. Communications of the ACM, 35(8):17–23, August 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/135237.html.

Goodman:1992:CME

[350] S. E. Goodman and J. D. Green. Computing in the Middle East. Communications of the ACM, 35(8):21– 24, August 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/135236.html.

Bell:1992:UTB

[351] C. Gordon Bell. Ultracomputers: a teraflop before its time. Communications of the ACM, 35(8):26–47, August 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/135227.html.

Creecy:1992:TMM

[352] Robert H. Creecy, Brij M. Masand, Stephen J. Smith, and David L. Waltz. Trading MIPS and memory for knowledge engineering. Communications of the ACM, 35(8):48–64, August 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/135228.html.

Kennedy:1992:GEI

[353] Ken Kennedy. Guest editor’s introduction to the special section on supercomputing. Communications of the ACM, 35(8):65, August 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/135229.html.

Hiranandani:1992:CFD

[354] Seema Hiranandani, Ken Kennedy, and Chau-Wen Tseng. Compiling Fortran D for MIMD distributed-memory machines. Communications of the ACM, 35(8):66–80, August 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/135230.html.

Cann:1992:RFD

[355] David Cann. Retire Fortran? A debate rekindled. Communications of the ACM, 35(8):81–89, August 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/135231.html.

Hummel:1992:FMS

[356] Susan Flynn Hummel, Edith Schonberg, and Lawrence E. Flynn. Factoring: a method for scheduling parallel loops. Communications of the ACM, 35(8):90–101, August 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/135232.html.

Pugh:1992:PAE

[357] William Pugh. A practical algorithm for exact array dependence analysis. Communications of the ACM, 35 (8):102–114, August 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/135233.html.

Simmons:1992:PCF

[358] Margaret L. Simmons, Harvey J. Wasserman, Olaf M. Lubeck, Christopher Eoyang, Raul Mendez, Hiroo Harada, and Misako Ishiguro. A performance comparison of four supercomputers. Communications of the ACM, 35(8):116–124, August 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/135234.html.

Neumann:1992:IRF

[359] Peter G. Neumann. Inside RISKS: Fraud by computer. Communications of the ACM, 35(8):154, August 1992. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/135238.html.

Samuelson:1992:UCL

[360] Pamela Samuelson. Updating the copyright look and feel lawsuits. Communications of the ACM, 35(9): 25–31, September 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/131009.html.

Korson:1992:AMS

[361] Timothy D. Korson and Vijay K. Vaishnavi. Analysis and modeling in software development. Communications of the ACM, 35(9):32–34, September 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/135225.html.

Monarchi:1992:RTO

[362] David E. Monarchi and Gretchen I. Puhr. A research typology for object-oriented analysis and design. Communications of the ACM, 35(9): 35–47, September 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/130995.html.

Rubin:1992:OBA

[363] Kenneth S. Rubin and Adele Goldberg. Object behavior analysis. Communications of the ACM, 35(9): 48–62, September 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/130996.html.

Nerson:1992:AOO

[364] Jean-Marc Nerson. Applying objectoriented analysis and design. Communications of the ACM, 35(9):63–74, September 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/130997.html.

Curtis:1992:PM

[365] Bill Curtis, Marc I. Kellner, and Jim Over. Process modeling. Communications of the ACM, 35(9):75–90, September 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/130998.html.

Krasner:1992:LLS

[366] Herb Krasner, Jim Terrel, Adam Linehan, Paul Arnold, and William H. Ett. Lessons learned from a software process modeling system. Communications of the ACM, 35(9):91– 100, September 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/130999.html.

Korson:1992:MES

[367] Timothy D. Korson and Vijay K. Vaishnavi. Managing emerging software technologies: A technology transfer framework. Communications of the ACM, 35(9):101–111, September 1992. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/131000.html.

Navathe:1992:EDM

[368] Shamkant B. Navathe. Evolution of data modeling for databases. Communications of the ACM, 35(9):112– 123, September 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/131001.html.

Chandrasekaran:1992:TSA

[369] B. Chandrasekaran, Todd R. Johnson, and Jack W. Smith. Taskstructure analysis for knowledge modeling. Communications of the ACM, 35 (9):124–137, September 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/131002.html.

Krishnan:1992:KBM

[370] Ramayya Krishnan, Xiaoping Li, and David Steier. A knowledge-based mathematical model formulation system. Communications of the ACM, 35 (9):138–146, September 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/131003.html.

Page-Jones:1992:CTM

[371] Meilir Page-Jones. Comparing techniques by means of encapsulation and connascence. Communications of the ACM, 35(9):147–151, September 1992. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/131004.html.

Coad:1992:OOP

[372] Peter Coad. Object-oriented patterns. Communications of the ACM, 35(9): 152–159, September 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/131006.html.

Nierstrasz:1992:COS

[373] Oscar Nierstrasz, Simon Gibbs, and Dennis Tsichritzis. Componentoriented software development. Communications of the ACM, 35(9):160– 165, September 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/131005.html.

Scheer:1992:EDM

[374] August-Wilhelm Scheer and Alexander Hars. Extending data modeling to cover the whole enterprise. Communications of the ACM, 35(9): 166–172, September 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/131007.html.

Neumann:1992:IRAa

[375] Peter G. Neumann. Inside RISKS: Accidental financial losses. Communications of the ACM, 35(9):194, September 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/131008.html.

Rettig:1992:PPR

[376] Marc Rettig. Practical programmer: Hat racks for understanding. Communications of the ACM, 35 (10):21–24, October 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/135247.html.

Barlow:1992:EFW

[377] John Perry Barlow. Electronic frontier: Will Japan jack in? Communications of the ACM, 35(10):27–29, October 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/135248.html.

Stylianou:1992:SCE

[378] Anthony C. Stylianou, Gregory R. Madey, and Robert D. Smith. Selection criteria for expert system shells: A socio-technical framework. Communications of the ACM, 35 (10):30–48, October 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/135240.html.

Touretzky:1992:VEA

[379] David S. Touretzky and Peter Lee. Visualizing evaluation in applicative languages. Communications of the ACM, 35(10):49–59, October 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/135241.html.

Feldman:1992:SMM

[380] Yosee Feldman and Ehud Shapiro. Spatial machines: a more realistic approach to parallel computation. Communications of the ACM, 35(10):60–73, October 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/135242.html.

Baeza-Yates:1992:NAT

[381] Ricardo Baeza-Yates and Gaston H. Gonnet. A new approach to text searching. Communications of the ACM, 35(10):74–82, October 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/135243.html. This paper de-

scribes a new linear-time string search algorithm that can handle limited regular-expression pattern matching without backtracking. See also [1948], [1], [1946], [85], and [382].

Wu:1992:FTS

[382] Sun Wu and Udi Manber. Fast text searching allowing errors. Communications of the ACM, 35(10):83–91, October 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/135244.html. This algorithm

in this paper is implemented in the agrep program, publicly available via ANONYMOUS FTP to cs.arizona. edu in the agrep subdirectory. See also [381].

Dori:1992:DAT

[383] Dov Dori. Dimensioning analysis: toward automatic understanding of engineering drawings. Communications of the ACM, 35(10): 92–103, October 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/135245.html.

Palvia:1992:TTF

[384] Shailendra C. Palvia and Steven R. Gordon. Tables, trees and formulas in decision analysis. Communications of the ACM, 35(10): 104–113, October 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/135246.html.

Neumann:1992:IRWb

[385] Peter G. Neumann. Inside RISKS: Where to place trust. Communications of the ACM, 35(10):138, October 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/135249.html.

Zweben:1992:PLW

[386] Stuart Zweben. President’s letter: We overplanting our seed corn? Communications of the ACM, 35(11):15–16, November 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Likins:1992:VBS

[387] Peter Likins. Viewpoint: a breach of the social contract. Communications of the ACM, 35(11):17–18, 111, November 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Knowles:1992:AFC

[388] Brad Knowles, Roger Schlafly, Grant D. Schultz, Lynn Zelvin, Paul Heckel, and E. Robert Yoches. ACM Forum: Comments on cryptography. Communications of the ACM, 35(11):19– 24, 112, November 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wilkes:1992:CPL

[389] Maurice V. Wilkes. Computing perspectives. the long-term future of operating systems. Communications of the ACM, 35(11):23– 24, 112, November 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/138855.html.

Mesher:1992:IPC

[390] G. M. Mesher, R. O. Briggs, S. E. Goodman, L. I. Press, and J. M. Snyder. International perspectives: Cuba, communism, and computing. Communications of the ACM, 35(11):27– 29, 112, November 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/138856.html.

Hartmanis:1992:CFC

[391] Juris Hartmanis. Computing the future: committee to assess the scope and direction of computer science and technology for the National Research Council. Communications of the ACM, 35 (11):30–40, November 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/138858.html.

Brosgol:1992:A

[392] Benjamin M. Brosgol. Ada. Communications of the ACM, 35(11): 41–42, November 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/138845.html.

Schonberg:1992:TTA

[393] Edmond Schonberg, Mark Gerhardt, and Charlene Hayden. A technical tour of Ada. Communications of the ACM, 35(11):43–52, November 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/138846.html.

Feldman:1992:AEU

[394] Michael B. Feldman. Ada experience in the undergraduate curriculum. Communications of the ACM, 35 (11):53–67, November 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/138847.html.

McCormick:1992:MRA

[395] John W. McCormick. A model railroad for Ada and software engineering. Communications of the ACM, 35 (11):68–70, November 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/138848.html.

Rosen:1992:WOS

[396] J. P. Rosen. What orientation should Ada objects take? Communications of the ACM, 35(11): 71–76, November 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/138849.html.

Taft:1992:ATS

[397] S. Tucker Taft. Ada 9X: a technical summary. Communications of the ACM, 35(11):77–82, November 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/138850.html.

Anderson:1992:APM

[398] Chris Anderson. Ada 9X project management. Communications of the ACM, 35(11):83–84, November 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/138851.html.

Ploedereder:1992:BCA

[399] Erhard Ploedereder. Building consensus for Ada 9X. Communications of the ACM, 35(11):85–88, November 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/138852.html.

Wiederhold:1992:TM

[400] Gio Wiederhold, Peter Wegner, and Stefano Ceri. Toward megaprogramming. Communications of the ACM, 35 (11):89–99, November 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/138853.html.

Kopec:1992:AAI

[401] Danny Kopec, Monty Newborn, and Mike Valvo. The 22nd Annual ACM International Computer Chess Championship. Communications of the ACM, 35(11):100–110, November 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/138854.html.

Mercuri:1992:IRV

[402] Rebecca Mercuri. Inside RISKS: Voting machine risks. Communications of the ACM, 35(11):138, November 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/138857.html.

Gersting:1992:VIS

[403] Judith L. Gersting. Viewpoint: Intangibles for success. Communications of the ACM, 35(12):11–12, December 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hayes:1992:AFD

[404] Patrick J. Hayes, Gordon S. Novak, Jr., and Wendy G. Lehnert. ACM Forum: In defense of artificial intelligence. Communications of the ACM, 35(12):13–16, December 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Traub:1992:TAN

[405] Joseph F. Traub. A tribute to Allen Newell: 1927–1992. Communications of the ACM, 35(12):20, December 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/214847.html.

Press:1992:PCN

[406] Larry Press. Personal computing: The net: Progress and opportunity. Communications of the ACM, 35(12): 21–25, December 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/138871.html.

Loeb:1992:IF

[407] Shoshana Loeb and Douglas Terry. Information filtering. Communications of the ACM, 35(12):26–28, December 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/138860.html.

Belkin:1992:IFI

[408] Nicholas J. Belkin and W. Bruce Croft. Information filtering and information retrieval: two sides of the same coin? Communications of the ACM, 35 (12):29–38, December 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/138861.html.

Loeb:1992:APD

[409] Shoshana Loeb. Architecting personalized delivery of multimedia information. Communications of the ACM, 35 (12):39–47, December 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/138862.html.

Stevens:1992:ACI

[410] Curt Stevens. Automating the creation of information filters. Communications of the ACM, 35(12):48, December 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/138863.html.

Stadnyk:1992:MUI

[411] Irene Stadnyk and Robert Kass. Modeling users’ interests in information filters. Communications of the ACM, 35 (12):49–50, December 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/138864.html.

Baclace:1992:CAI

[412] Paul E. Baclace. Competitive agents for information filtering. Communications of the ACM, 35(12):50, December 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/138865.html.

Foltz:1992:PID

[413] Peter W. Foltz and Susan T. Dumais. Personalized information delivery: An analysis of information-filtering methods. Communications of the ACM, 35 (12):51–60, December 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/138866.html.

Goldberg:1992:UCF

[414] David Goldberg, David Nichols, Brian M. Oki, and Douglas Terry. Using collaborative filtering to weave an information tapestry. Communications of the ACM, 35(12):61–70, December 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/138867.html.

Bowen:1992:DA

[415] T. F. Bowen, G. Gopal, G. Herman, T. Hickey, K. C. Lee, W. H. Mansfield, J. Raitz, and A. Weinrib. The Datacycle architecture. Communications of the ACM, 35(12): 71–81, December 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/138868.html.

Ram:1992:NLU

[416] Ashwin Ram. Natural language understanding for information-filtering systems. Communications of the ACM, 35 (12):80–81, December 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/138869.html.

Denning:1992:ENE

[417] Peter J. Denning. Educating a new engineer. Communications of the ACM, 35(12):82–97, December 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/138870.html.

LaMalva:1992:APC

[418] Patrick M. LaMalva and David R. Reyes-Guerra. Accredited programs in computing. Communications of the ACM, 35(12):98–101, December 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:1992:IRAb

[419] Peter G. Neumann. Inside RISKS: Avoiding weak links. Communications of the ACM, 35(12):146, December 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/138872.html.

Samuelson:1993:LSL

[420] Pamela Samuelson. Legally speaking: Liability for defective electronic information. Communications of the ACM, 36(1):21–26, January 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/151234.html.

Bly:1993:MWM

[421] Sara A. Bly, Steve R. Harrison, and Susan Irwin. Multimedia in the workplace: Media spaces: Bringing people together in a video, audio, and computing environment. Communications of the ACM, 36(1):28–47, January 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/151235.html.

Lyles:1993:MSB

[422] Bryan Lyles. Media spaces and broadband ISDN. Communications of the ACM, 36(1):46–47, January 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/151236.html.

Fish:1993:VTI

[423] Robert S. Fish, Robert E. Kraut, Robert W. Root, and Ronald E. Rice. Video as a technology for informal communication. Communications of the ACM, 36(1):48–61, January 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/151237.html.

Hill:1993:RLA

[424] Ralph D. Hill, Tom Brinck, John F. Patterson, Steven L. Rohall, and Wayne T. Wilner. The Rendezvous language and architecture. Communications of the ACM, 36(1):62–67, January 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/151238.html.

Arango:1993:TMS

[425] M. Arango, L. Bahler, P. Bates, M. Cochinwala, D. Cohrs, R. Fish, G. Gopal, N. Griffeth, G. E. Herman, T. Hickey, K. C. Lee, W. E. Leland, C. Lowery, V. Mak, J. Patterson, L. Ruston, M. Segal, R. C. Sekar, M. P. Vecchi, A. Weinrib, and S.-Y. Wuu. Touring Machine System. Communications of the ACM, 36(1):68–77, January 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/151239.html.

Milner:1993:EIT

[426] Robin Milner. Elements of interaction: Turing Award Lecture. Communications of the ACM, 36(1):78–89, January 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/151240.html.

Frenkel:1993:IRM

[427] Karen A. Frenkel. An interview with Robin Milner. Communications of the ACM, 36(1):90–97, January 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/151241.html.

Banatre:1993:PMT

[428] Jean-Pierre Banâtre and Daniel Le Métayer. Programming by multiset transformation. Communications of the ACM, 36(1):98–111, January 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/151242.html.

Neumann:1993:IRRa

[429] Peter G. Neumann. Inside RISKS: Risks considered global(ly). Communications of the ACM, 36(1):154, January 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/151243.html.

Kling:1993:TCB

[430] Rob Kling, Peter Wegner, John R. Rice, and Eric A. Weiss. Technical correspondence: Broadening computer science. Communications of the ACM, 36(2):15–20, February 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wilkes:1993:CPL

[431] Maurice V. Wilkes. Computing perspectives: Light amplifiers and solitons. Communications of the ACM, 36(2):21–23, February 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/151221.html.

Odedra:1993:IPS

[432] M. Odedra, M. Bennett, S. Goodman, and M. Lawrie. International perspectives: Sub-Saharan Africa: A technological desert. Communications of the ACM, 36(2):25–29, February 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/151222.html.

Supnik:1993:ISI

[433] Robert M. Supnik. Introduction to special issue on Digital’s Alpha chip project. Communications of the ACM, 36(2):30–32, February 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/151223.html.

Sites:1993:AAA

[434] Richard L. Sites. Alpha AXP architecture. Communications of the ACM, 36(2):33–44, February 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/151226.html.

Kronenberg:1993:POV

[435] Nancy Kronenberg, Thomas R. Benson, Wayne M. Cardoza, Ravindran Jagannathan, and Benjamin J. Thomas. Porting OpenVMS from VAX to Alpha AXP. Communications of the ACM, 36(2):45–53, February 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/151224.html.

Thacker:1993:ADU

[436] Charles P. Thacker, David G. Conroy, and Lawrence C. Stewart. The Alpha demonstration unit: a highperformance multiprocessor. Communications of the ACM, 36(2):55–67, February 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/151225.html.

Katz:1993:TF

[437] Ralph Katz. Tales from the front. Communications of the ACM, 36(2): 67, February 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sites:1993:BT

[438] Richard L. Sites, Anton Chernoff, Matthew B. Kirk, Maurice P. Marks, and Scott G. Robinson. Binary translation. Communications of the ACM, 36(2):69–81, February 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/151227.html.

Dobberpuhl:1993:DE

[439] Daniel Dobberpuhl and Kevin Fielding. Design of EV-4. Communications of the ACM, 36(2):82, February 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/151228.html.

Hollingsworth:1993:AST

[440] Richard Hollingsworth. Advanced semiconductor technology. Communications of the ACM, 36(2):83, February 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/151229.html.

Kemerer:1993:RFP

[441] Chris F. Kemerer. Reliability of function points measurement: a field experiment. Communications of the ACM, 36(2):85–97, February 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/151230.html.

Anderson:1993:UNA

[442] Ronald E. Anderson, Deborah G. Johnson, Donald Gotterbarn, and Judith Perrolle. Using the new ACM code of ethics in decision making. Communications of the ACM, 36(2): 98–107, February 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/151231.html.

Neumann:1993:IRD

[443] Peter G. Neumann. Inside RISKS: Are dependable systems feasible? Communications of the ACM, 36(2):146, February 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/151232.html.

Grover:1993:ITE

[444] Varun Grover and Martin Goslar. Information technologies for the 1990s: the executives’ view. Communications of the ACM, 36(3):17–19, March 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/153548.html.

Barlow:1993:EFB

[445] John Perry Barlow. Electronic frontier: Bill O’ Rights. Communications of the ACM, 36(3):21–23, March 1993. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/153551.html.

Denning:1993:TT

[446] Dorothy E. Denning. To tap or not to tap. Communications of the ACM, 36(3):24–33, March 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/153523.html.

Godwin:1993:C

[447] Mike Godwin, William A. Bayse, Marc Rotenberg, Lewis M. Branscomb, Anne M. Branscomb, Ronald L. Rivest, Andrew Grosso, and Gary T. Marx. Commentators. Communications of the ACM, 36(3):34–41, March 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/153526.html.

Denning:1993:FT

[448] Dorothy E. Denning. Final thoughts. Communications of the ACM, 36 (3):42–44, March 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/153532.html.

Shapiro:1993:FGP

[449] Ehud Shapiro and David H. D. Warren. The Fifth Generation Project: Personal perspectives: Introduction. Communications of the ACM, 36 (3):46–49, March 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/153534.html.

Fuchi:1993:LNE

[450] Kazuhiro Fuchi, Robert Kowalski, Koichi Furukawa, Kazunori Ueda, Ken Kahn, Takashi Chikayama, and Evan Tick. Launching the new era. Communications of the ACM, 36(3):49–100, March 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/153541.html.

Shapiro:1993:E

[451] Ehud Shapiro and David H. D. Warren. Epilogue. Communications of the ACM, 36(3):100–101, March 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/153547.html.

Neumann:1993:IRRb

[452] Peter G. Neumann. Inside risks: Risks of technology. Communications of the ACM, 36(3):130, March 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Billington:1993:UCF

[453] David Billington. Using the contextfree pumping lemma. Communications of the ACM, 36(4):21, 81, April 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/153590.html.

Rettig:1993:PPC

[454] Marc Rettig. Practical programmer: Cooperative software. Communications of the ACM, 36(4):23– 28, April 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/153593.html.

Samuelson:1993:LSU

[455] Pamela Samuelson. Legally speaking: The ups and downs of look and feel. Communications of the ACM, 36(4):29–35, April 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/153595.html.

Mandelkern:1993:GUI

[456] David Mandelkern. Graphical user interfaces: The next generation: Introduction. Communications of the ACM, 36(4):36–39, April 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/153574.html.

Johnson:1993:ABI

[457] Jeff A. Johnson, Bonnie A. Nardi, Craig L. Zarmer, and James R. Miller. ACE: Building interactive graphical interactions. Communications of the ACM, 36(4):40–55, April 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/153576.html.

Robertson:1993:IVU

[458] George G. Robertson, Stuart K. Card, and Jock D. Mackinlay. Information visualization using 3D interactive animation. Communications of the ACM, 36(4):56–71, April 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/153577.html.

Morse:1993:OCG

[459] Alan Morse and George Reynolds. Overcoming current growth limits in UI development. Communications of the ACM, 36(4):72–81, April 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/153579.html.

Nielsen:1993:NUI

[460] Jakob Nielsen. Noncommand user interfaces. Communications of the ACM, 36(4):82–99, April 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/153582.html.

Marcus:1993:HCI

[461] Aaron Marcus. Human communications issues in advanced UIs. Communications of the ACM, 36(4):101– 109, April 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/153670.html.

Grudin:1993:IEC

[462] Jonathan Grudin. Interface: an evolving concept. Communications of the ACM, 36(4):110–119, April 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/153585.html.

Martin:1993:MAT

[463] C. Dianne Martin. The myth of the awesome thinking machine. Communications of the ACM, 36(4):120– 133, April 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/153587.html.

Norman:1993:IRU

[464] Donald A. Norman. Inside RISKS: Using names as identifiers. Communications of the ACM, 36(4):154, April 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/214834.html.

Beach:1993:BF

[465] Gary Beach. Building a foundation. Communications of the ACM, 36(5):13–14, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/155050.html.

Grosch:1993:FDM

[466] Herb Grosch. Forum: In defense of Maurice Wilkes (letter). Communications of the ACM, 36(5):16, May 1993.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Press:1993:PCT

[467] Larry Press. Personal computing: Technetronic education: Answers on the cultural horizon. Communications of the ACM, 36(5):17– 22, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/155051.html.

Soloway:1993:LER

[468] Elliot Soloway. Log on education: Reading and writing in the 21st century. Communications of the ACM, 36(5):23–27, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/155052.html.

Soloway:1993:IIT

[469] Elliot Soloway. Introduction to the issue on technology in education. Communications of the ACM, 36(5):28–30, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/155053.html.

Hawkins:1993:TOS

[470] Jan Hawkins. Technology and the organization of schooling. Communications of the ACM, 36(5):30– 35, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/155054.html.

Ruopp:1993:L

[471] Dick Ruopp and Shahaf Gal. The LabNetwork. Communications of the ACM, 36(5):35–36, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/155055.html.

Scardamalia:1993:TKB

[472] Marlene Scardamalia and Carl Bereiter. Technologies for knowledgebuilding discourse. Communications of the ACM, 36(5):37–41, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/155056.html.

Hunter:1993:ICT

[473] Beverly Hunter. Internetworking: coordinating technology for systemic reform. Communications of the ACM, 36(5):42–46, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/155057.html.

Pearlman:1993:DNA

[474] Robert Pearlman. Designing the new American schools. Communications of the ACM, 36(5):46–49, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/155058.html.

Newman:1993:SND

[475] Denis Newman. School networks: delivery or access. Communications of the ACM, 36(5):49–51, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/155059.html.

Barron:1993:TEM

[476] Brigid Barron and Ronald J. Kantor. Tools to enhance math education: the Jasper series. Communications of the ACM, 36(5):52–54, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/155060.html.

Schank:1993:LMC

[477] Roger C. Schank. Learning via multimedia computers. Communications of the ACM, 36(5):54–56, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/155061.html.

Grant:1993:WCC

[478] Wayne C. Grant. Wireless Coyote: a computer-supported field trip. Communications of the ACM, 36(5):57– 59, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/155062.html.

Pea:1993:CVP

[479] Roy D. Pea. The collaborative visualization project. Communications of the ACM, 36(5):60–63, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/155063.html.

Rubin:1993:VLT

[480] Andee Rubin. Video laboratories: tools for scientific investigation. Communications of the ACM, 36(5):64– 65, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/155064.html.

Braun:1993:HAS

[481] Ludwig Braun. Help for all the students. Communications of the ACM, 36(5):66–69, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/155065.html.

Becker:1993:TAC

[482] Henry Jay Becker. Teaching with and about computers in secondary schools. Communications of the ACM, 36(5):69–72, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/155066.html.

Anderson:1993:TIU

[483] Ronald E. Anderson. The technology infrastructure of U.S. schools. Communications of the ACM, 36(5):72– 73, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/155067.html.

Hawkins:1993:AT

[484] Jan Hawkins, John Frederiksen, Allan Collins, Dorothy Bennett, and Elizabeth Collins. Assessment and technology. Communications of the ACM, 36(5):74–76, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/155068.html.

Edwards:1993:LLL

[485] Carol Edwards. Life-long learning. Communications of the ACM, 36(5):76–78, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/155069.html.

George:1993:CEF

[486] Yolanda S. George, Shirley M. Malcolm, and Laura Jeffers. Computer equity for the future. Communications of the ACM, 36(5):78– 81, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/155070.html.

Conn:1993:SS

[487] Coco Conn. SIGKids at SIG-GRAPH. Communications of the ACM, 36(5):81–82, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/155071.html.

Perl:1993:ATF

[488] Teri Perl. ACM task force on K-12 education and technology. Communications of the ACM, 36(5):82– 83, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/155072.html.

Lidtke:1993:CSP

[489] Doris K. Lidtke and David Moursund. Computers in schools: past, present, and how we can change the future. Communications of the ACM, 36(5):84–87, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/155073.html.

Merritt:1993:AMH

[490] Susan M. Merritt, Charles J. Buen, Philip East, Darlene Grantham, Charles Rice, Viera K. Proulx, Gerry Segal, and Carol Wolf. ACM model high school computer science curriculum. Communications of the ACM, 36(5):87–90, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/155074.html.

Neumann:1993:IRRc

[491] Peter G. Neumann. Inside RISKS: The role of software engineering. Communications of the ACM, 36(5): 114, May 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/214836.html.

Nidumolu:1993:IPC

[492] S. R. Nidumolu and S. E. Goodman. International perspectives: Computing in India: An Asian elephant learning to dance. Communications of the ACM, 36(6):15–22, June 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/163261.html.

Kuhn:1993:ITP

[493] Sarah Kuhn and Michael J. Muller. Introduction with taxonomy of PD practices: a brief practitioner’s guide. Communications of the ACM, 36 (6):24–28, June 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/163263.html.

Muller:1993:TPP

[494] Michael J. Muller, Daniel M. Wildman, and Ellen A. White. Taxonomy of PD practices: a brief practitioner’s guide. Communications of the ACM, 36(6):26–27, June 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Clement:1993:RLP

[495] Andrew Clement and Peter Van den Besselaar. A retrospective look at PD projects. Communications of the ACM, 36(6):29–37, June 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/163264.html.

Miller:1993:SDD

[496] Steven E. Miller. From system design to democracy. Communications of the ACM, 36(6):38–39, June 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/214813.html.

Bjerknes:1993:SPA

[497] Gro Bjerknes. Some PD advice. Communications of the ACM, 36(6): 39, June 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/214815.html.

Carmel:1993:PJA

[498] Erran Carmel, Randall D. Whitaker, and Joey F. George. PD and joint application design: A Transatlantic comparison. Communications of the ACM, 36(6):40–48, June 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/163265.html.

Greenbaum:1993:PPS

[499] Joan Greenbaum. PD a personal statement. Communications of the ACM, 36(6):47, June 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/214816.html.

Anderson:1993:EPC

[500] William L. Anderson and William T. Crocca. Engineering practice and codevelopment of product prototypes. Communications of the ACM, 36 (6):49–56, June 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/163266.html.

Euchner:1993:BIT

[501] Jim Euchner and Patricia Sachs. The benefits of intentional tension.

Communications of the ACM, 36(6): 53, June 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/214817.html.

Madsen:1993:EUC

[502] Kim Halskov Madsen and Peter H. Aiken. Experiences using cooperative interactive storyboard prototyping. Communications of the ACM, 36(6):57–64, June 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/163268.html.

Crane:1993:GF

[503] Darlene Crane. Graphic facilitation. Communications of the ACM, 36(6):64–65, June 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/214820.html.

Muller:1993:EOP

[504] Michael J. Muller, Daniel M. Wildman, and Ellen A. White. “equal opportunity” PD using PICTIVE. Communications of the ACM, 36(6):64– 66, June 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/214818.html.

Gronbaek:1993:CCC

[505] Kaj Grønbæk, Morten Kyng, and Preben Mogensen. CSCW challenges: cooperative design in engineering projects. Communications of the ACM, 36(6):67–77, June 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/163272.html.

Harker:1993:UPP

[506] Susan Harker. User participation in prototyping. Communications of the ACM, 36(6):77–81, June 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/214822.html.

Kensing:1993:PST

[507] Finn Kensing and Andreas Munk-Madsen. PD: structure in the toolbox. Communications of the ACM, 36(6):78–85, June 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/163278.html.

Mumford:1993:EA

[508] Enid Mumford. The ETHICS approach. Communications of the ACM, 36(6):82, June 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/214824.html.

Floyd:1993:SMA

[509] Christiane Floyd. STEPS: a methodical approach to PD. Communications of the ACM, 36(6):83, June 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/214826.html.

Katzenberg:1993:WLA

[510] Barbara Katzenberg and Peter Piela. Work language analysis and the naming problem. Communications of the ACM, 36(6):86–92, June 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/163286.html.

Blomberg:1993:WOD

[511] Jeanette Blomberg, Denise McLaughlin, and Lucy Suchman. Workoriented design at Xerox. Communications of the ACM, 36(6):91, June 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/214828.html.

Novick:1993:PCP

[512] David Novick and Eleanor Wynn. Participatory conversation in PD. Communications of the ACM, 36(6): 93, June 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/214830.html.

Wagner:1993:WFP

[513] Ina Wagner. A web of fuzzy problems: confronting the ethical issues. Communications of the ACM, 36(6):94–101, June 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/163290.html. Analyses

inter-human relations and problems related to the task of scheduling rooms and persons having to do with surgery in a large hospital.

Williams:1993:TBS

[514] Marian G. Williams and Vivienne Begg. Translation between software designers and users. Communications of the ACM, 36(6):102– 103, June 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/214831.html.

Neumann:1993:IRM

[515] Peter G. Neumann. Inside RISKS: Modeling and simulation. Communications of the ACM, 36(6):124, June 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/163293.html.

Brown:1993:LTY

[516] Judith R. Brown. Looking toward the year 2000. Communications of the ACM, 36(7):13–14, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/159548.html.

Angell:1993:ILB

[517] Ian O. Angell. Intelligence: logical or biological. Communications of the ACM, 36(7):15–16, 119, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/159550.html.

Steele:1993:FCE

[518] Guy L. Steele, Jr., Danny Hillis, Richard Stallman, Gerald J. Sussman, Marvin Minsky, John McCarthy, John Backus, Fernando Corbató, and Ronald E. Anderson. Forum: Code of ethics reconsidered. Communications of the ACM, 36(7):17–20, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wilkes:1993:CPF

[519] Maurice V. Wilkes. Computing perspectives: From Fortran and Algol to object-oriented languages. Communications of the ACM, 36(7):21– 23, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/159553.html.

Wellner:1993:GEI

[520] Pierre Wellner, Wendy Mackay, and Rich Gold. Guest editors’ introduction to the special issue on computeraugmented environments: Back to the real world. Communications of the ACM, 36(7):24–27, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/159555.html.

Spreitzer:1993:SSM

[521] Michael Spreitzer and Marvin Theimer. Scalable, secure, mobile computing with location information. Communications of the ACM, 36(7):27, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/159558.html.

Baudel:1993:CRC

[522] Thomas Baudel and Michel Beaudouin-Lafon. Charade: Remote control of objects using free-hand gestures. Communications of the ACM, 36(7):28–35, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/159562.html.

Krueger:1993:ETM

[523] Myron W. Krueger. Environmental technology: making the real world virtual. Communications of the ACM, 36(7):36–37, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/159563.html.

Fitzmaurice:1993:SIS

[524] George W. Fitzmaurice. Situated information spaces and spatially aware palmtop computers. Communications of the ACM, 36(7):38– 49, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/159566.html.

Azuma:1993:TRA

[525] Ronald Azuma. Tracking requirements for augmented reality. Communications of the ACM, 36(7):50– 51, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/159581.html.

Feiner:1993:KBA

[526] Steven Feiner, Blair Macintyre, and Dorée Seligmann. Knowledge-based augmented reality. Communications of the ACM, 36(7):53–62, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/159587.html.

Resnick:1993:BCK

[527] Mitchel Resnick. Behavior construction kits. Communications of the ACM, 36(7):64–71, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/159593.html.

Gold:1993:P

[528] Rich Gold. This is not a pipe. Communications of the ACM, 36(7): 72, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/159598.html.

Weiser:1993:SCS

[529] Mark Weiser. Some Computer Science issues in ubiquitous computing. Communications of the ACM, 36(7):74–84, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/159617.html.

Elrod:1993:ROE

[530] Scott Elrod, Gene Hall, Rick Costanza, Michael Dixon, and Jim Des Rivières.

Responsive office environments. Communications of the ACM, 36(7):84– 85, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/159626.html.

Wellner:1993:IPD

[531] Pierre Wellner. Interacting with paper on the DigitalDesk. Communications of the ACM, 36(7):86– 96, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/159630.html.

Mackay:1993:ARA

[532] Wendy Mackay, Gilles Velay, Kathy Carter, Chaoying Ma, and Daniele Pagani. Augmenting reality: adding computational dimensions to paper. Communications of the ACM, 36(7):96– 97, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/159638.html.

Denning:1993:DNP

[533] Peter J. Denning. Designing new principles to sustain research in our Universities. Communications of the ACM, 36(7):99–104, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/159640.html.

Marsaglia:1993:TCR

[534] George Marsaglia. Technical correspondence: Remarks on choosing and implementing random number generators. Communications of the ACM, 36(7):105–108, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sullivan:1993:ATR

[535] Stephen J. Sullivan. Another test for randomness. Communications of the ACM, 36(7):108, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Park:1993:ATR

[536] Stephen K. Park, Keith W. Miller, and Paul K. Stockmeyer. Another test for randomness: Response. Communications of the ACM, 36(7):108–110, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:1993:IRRd

[537] Peter G. Neumann. Inside RISKS: RISKS on the rails. Communications of the ACM, 36(7):130, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/214812.html.

Colonna:1993:TCS

[538] Jean-François Colonna. Technical correspondence: The subjectivity of computers. Communications of the ACM, 36(8):15–18, August 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rettig:1993:PPE

[539] Marc Rettig, Gary Simons, and John Thomson. Practical programmer: Extended objects. Communications of the ACM, 36(8):19–24, August 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/163413.html.

Weingarten:1993:WHH

[540] Fred W. Weingarten. From Washington: Happenings on the hill. Communications of the ACM, 36 (8):25–27, August 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/163417.html.

Ishida:1993:ISI

[541] Haruhisa Ishida and Lawrence H. Landweber. Introduction to the special issue on internetworking. Communications of the ACM, 36(8):28– 30, August 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/163382.html.

Bush:1993:FTT

[542] Randy Bush. FidoNet: technology, tools, and history. Communications of the ACM, 36(8):31–35, August 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/163383.html.

Murray:1993:KNG

[543] Janet Murray. K12 network: global education through telecommunications. Communications of the ACM, 36 (8):36–41, August 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/163386.html.

Sadowsky:1993:NCD

[544] George Sadowsky. Network connectivity for developing countries. Communications of the ACM, 36 (8):42–47, August 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/163388.html.

Kent:1993:IPE

[545] Stephen T. Kent. Internet privacy enhanced mail. Communications of the ACM, 36(8):48–60, August 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/163390.html.

Rotenberg:1993:CPI

[546] Marc Rotenberg. Communications privacy: implications for network design. Communications of the ACM, 36(8):61–68, August 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/163393.html.

Tuerkheimer:1993:UPP

[547] Frank M. Tuerkheimer. The underpinnings of privacy protection. Communications of the ACM, 36 (8):69–73, August 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/163394.html.

Hiramatsu:1993:PTP

[548] Tsuyoshi Hiramatsu. Protecting telecommunications privacy in Japan. Communications of the ACM, 36 (8):74–77, August 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/163400.html.

Schwartz:1993:DSI

[549] Michael F. Schwartz and David C. M. Wood. Discovering shared interests using graph analysis. Communications of the ACM, 36(8):78–89, August 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/163402.html.

Hesse:1993:RSC

[550] Bradford W. Hesse, Lee S. Sproull, Sara B. Kiesler, and John P. Walsh. Returns to science: computer networks in oceanography. Communications of the ACM, 36(8):90–101, August 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/163409.html.

Neumann:1993:IRRe

[551] Peter G. Neumann. Inside risks: Risks of surveillance. Communications of the ACM, 36(8):122, August 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/163420.html.

Hoffman:1993:CC

[552] Lance J. Hoffman. Clipping Clipper. Communications of the ACM, 36 (9):15–17, September 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/162691.html.

Samuelson:1993:LSC

[553] Pamela Samuelson. Legally speaking: Computer programs and copyright’s fair use doctrine. Communications of the ACM, 36(9):19–25, September 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/162693.html.

Press:1993:PCB

[554] Larry Press. Personal computing: Before the Altair: The history of personal computing. Communications of the ACM, 36(9):27–33, September 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/162697.html.

Anonymous:1993:ISI

[555] Anonymous. Introduction to the special issue on concurrent object-oriented programming. Communications of the ACM, 36(9):34, September 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cohen:1993:COO

[556] Jacques Cohen. Concurrent objectoriented programming. Communications of the ACM, 36(9):35–36, September 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/214809.html.

Lea:1993:CSS

[557] Rodger Lea, Christian Jacquemot, and Eric Pillevesse. COOL: system support for distributed programming. Communications of the ACM, 36(9): 37–46, September 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/162699.html.

Achauer:1993:DDO

[558] Bruno Achauer. The DOWL distributed object-oriented language. Communications of the ACM, 36(9): 48–55, September 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/162703.html.

Meyer:1993:SCO

[559] Bertrand Meyer. Systematic concurrent object-oriented programming. Communications of the ACM, 36(9): 56–80, September 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/162705.html.

Lohr:1993:CAR

[560] Klaus-Peter Löhr. Concurrency annotations for reusable software. Communications of the ACM, 36(9):81–89, September 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/162708.html.

Caromel:1993:TMO

[561] Denis Caromel. Toward a method of object-oriented concurrent programming. Communications of the ACM, 36 (9):90–102, September 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/162711.html.

Karaorman:1993:ICS

[562] Murat Karaorman and John Bruno. Introducing concurrency to a sequential language. Communications of the ACM, 36(9):103–115, September 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/162715.html.

Campbell:1993:DIC

[563] Roy H. Campbell, Nayeem Islam, David Raila, and Peter Madany. Designing and implementing Choices: an object-oriented system in C++. Communications of the ACM, 36(9): 117–126, September 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/162717.html.

Muhlhauser:1993:DMA

[564] Max Mühlhäuser, Wolfgang Gerteis, and Lutz Heuser. DOCASE: a methodic approach to distributed programming. Communications of the ACM, 36(9):127–138, September 1993. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/162721.html.

Neumann:1993:IRA

[565] Peter G. Neumann. Inside risks: Animal crackers. Communications of the ACM, 36(9):162, September 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/163427.html.

Dekleva:1993:TCF

[566] Sasa Dekleva and Stanley Martens. Technical correspondence: First normal form reconsidered. Communications of the ACM, 36(10):19–20, October 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Soloway:1993:LES

[567] Elliot Soloway. Log on education: Should we teach students to program? Communications of the ACM, 36(10):21–24, October 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/164061.html.

Wolcott:1993:IPU

[568] P. Wolcott and S. E. Goodman. International perspectives: Under the stress of reform: High-performance computing in the former Soviet Union. Communications of the ACM, 36 (10):25–29, October 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/164063.html.

Constantine:1993:ISS

[569] Larry L. Constantine and Lucy A. D. Lockwood. Introduction to the special section on project organization and management. Communications of the ACM, 36(10):30–32, October 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Constantine:1993:OPO

[570] Larry L. Constantine and Lucy A. D. Lockwood. Orchestrating project organization and management. Communications of the ACM, 36(10):31– 32, October 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/163434.html.

Constantine:1993:WOP

[571] Larry L. Constantine. Work organization: paradigms for project management and organization. Communications of the ACM, 36(10):35–43, October 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/163435.html.

Rettig:1993:PPD

[572] Marc Rettig and Gary Simons. A project planning and development process for small teams. Communications of the ACM, 36(10):45–55, October 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/163440.html.

Hyman:1993:CCH

[573] Risa B. Hyman. Creative chaos in highperformance teams: an experience report. Communications of the ACM, 36(10):57–60, October 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/163444.html.

Walz:1993:ISD

[574] Diane B. Walz, Joyce J. Elam, and Bill Curtis. Inside a software design team: knowledge acquisition, sharing, and integration. Communications of the ACM, 36(10):63–77, October 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/163447.html.

Zultner:1993:TTT

[575] Richard E. Zultner. TQM for technical teams. Communications of the ACM, 36(10):79–91, October 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/163577.html.

Holtzblatt:1993:MCC

[576] Karen Holtzblatt and Hugh Beyer. Making customer-centered design work for teams. Communications of the ACM, 36(10):92–103, October 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/164050.html.

Hutchings:1993:PIL

[577] Tony Hutchings, Michael G. Hyde, David Marca, and Lou Cohen. Process improvement that lasts: an integrated training and consulting method. Communications of the ACM, 36(10): 104–113, October 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/164057.html.

Zahniser:1993:DWA

[578] Richard A. Zahniser. Design by walking around. Communications of the ACM, 36(10):114–123, October 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/164060.html.

Neumann:1993:IRS

[579] Peter G. Neumann. Inside risks: System development woes. Communications of the ACM, 36(10):146, October 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/164065.html.

Barlow:1993:EFP

[580] John Perry Barlow. The electronic frontier: a plain text on crypto policy. Communications of the ACM, 36 (11):21–26, November 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/163378.html.

Weingarten:1993:WHR

[581] Fred W. Weingarten. From Washington: HPCC research questioned. Communications of the ACM, 36(11): 27–29, November 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/163379.html.

Hillis:1993:CCM

[582] W. Daniel Hillis and Lewis W. Tucker. The CM-5 Connection Machine: a scalable supercomputer. Communications of the ACM, 36(11):31–40, November 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/163361.html.

Thomborson:1993:DYW

[583] Clark D. Thomborson. Does your workstation computation belong on a vector supercomputer? Communications of the ACM, 36(11): 41–49, November 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/163363.html.

Knight:1993:IIT

[584] John C. Knight and E. Ann Myers. An improved inspection technique. Communications of the ACM, 36(11):51–61, November 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/163366.html.

Grady:1993:PRM

[585] Robert B. Grady. Practical results from measuring software quality. Communications of the ACM, 36 (11):62–68, November 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/163369.html.

Littlewood:1993:VUD

[586] Bev Littlewood and Lorenzo Strigini. Validation of ultrahigh dependability for software-based systems. Communications of the ACM, 36(11): 69–80, November 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/163373.html.

Banker:1993:SCM

[587] Rajiv D. Banker, Srikant M. Datar, Chris F. Kemerer, and Dani Zweig. Software complexity and maintenance costs. Communications of the ACM, 36 (11):81–94, November 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/163375.html.

Mercuri:1993:IRC

[588] Rebecca Mercuri. Inside risks: Corrupted polling. Communications of the ACM, 36(11):122, 94, November 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/163380.html.

Rada:1993:SSS

[589] Roy Rada. Sharing standards: Standards: The language for success. Communications of the ACM, 36(12): 17–23, December 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/163351.html.

Press:1993:PCI

[590] Larry Press. Personal computing: The Internet and interactive television. Communications of the ACM, 36 (12):19–23, December 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/163352.html.

Lucas:1993:MES

[591] Henry C. Lucas, Jr. Market expert surveillance system. Communications of the ACM, 36(12):27–34, December 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/163301.html.

Birman:1993:PGA

[592] Kenneth P. Birman. The process group approach to reliable distributed computing. Communications of the ACM, 36(12):37–53, December 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/163303.html.

El-Najdawi:1993:ESS

[593] M. K. El-Najdawi and Anthony C. Stylianou. Expert support systems: integrating AI technologies. Communications of the ACM, 36(12): 55–65, December 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/163306.html.

Brynjolfsson:1993:PPI

[594] Erik Brynjolfsson. The productivity paradox of information technology. Communications of the ACM, 36 (12):66–77, December 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/163309.html.

Kappelman:1993:CEU

[595] Leon A. Kappelman, John P. Thompson, and Ephraim R. McLean. Converging end-user and corporate computing. Communications of the ACM, 36(12):79–92, December 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/163314.html.

Rao:1993:HPA

[596] H. Raghav Rao and Brian D. Lynch. Hedonic price analysis of workstation attributes. Communications of the ACM, 36(12):95–102, December 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/163348.html.

Smith:1993:PPP

[597] H. Jeff Smith. Privacy policies and practices: inside the organizational maze. Communications of the ACM, 36 (12):104–122, December 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/163349.html.

Denning:1993:IRW

[598] Peter J. Denning. Inside risks: a world lit by flame. Communications of the ACM, 36(12):170, December 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/163353.html.

Feigenbaum:1994:KBS

[599] Edward A. Feigenbaum, Peter E. Friedland, Bruce B. Johnson, H. Penny Nii, Herbert Schnorr, Howard E. Shrobe, and Robert S. Engelmore. Knowledge-based systems in Japan (report of the JTEC panel). Communications of the ACM, 37(1):17–20, January 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/175232.html.

Samuelson:1994:LSC

[600] Pamela Samuelson. Legally speaking: Copyright’s fair use doctrine and digital data. Communications of the ACM, 37(1):21–27, January 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/175233.html.

Schuler:1994:SC

[601] Doug Schuler. Social computing. Communications of the ACM, 37 (1):28–29, January 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/175223.html.

Chapman:1994:NFS

[602] Gary Chapman. The National Forum on Science and Technology Goals: building a democratic, post-Cold War science and technology policy. Communications of the ACM, 37(1):30–37, January 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/175224.html.

Schuler:1994:CNB

[603] Doug Schuler. Community networks: building a new participatory medium. Communications of the ACM, 37 (1):38–51, January 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/175225.html.

Clement:1994:CWE

[604] Andrew Clement. Computing at work: empowering action by “low-level users”. Communications of the ACM, 37(1):52–63, January 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/175226.html.

Friedman:1994:ECS

[605] Batya Friedman and Peter H. Kahn, Jr. Educating computer scientists: linking the social and the technical. Communications of the ACM, 37(1):64–70, January 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/175227.html.

Nissenbaum:1994:CA

[606] Helen Nissenbaum. Computing and accountability. Communications of the ACM, 37(1):72–80, January 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/175228.html.

Collins:1994:HGG

[607] W. Robert Collins, Keith W. Miller, Bethany J. Spielman, and Phillip Wherry. How good is good enough?: an ethical analysis of software construction and use. Communications of the ACM, 37(1):81–91, January 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/175229.html.

Grudin:1994:GSD

[608] Jonathan Grudin. Groupware and social dynamics: eight challenges for developers. Communications of the ACM, 37(1):92–105, January 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/175230.html.

Anonymous:1994:RCP

[609] Anonymous. Renewing the commitment to a public interest telecommunications policy. Communications of the ACM, 37(1):106–108, January 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/175231.html.

Dorsett:1994:IRR

[610] Robert D. Dorsett. Inside RISKS: Risks in aviation (part 1 of 2). Communications of the ACM, 37 (1):154, January 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/175234.html.

Soloway:1994:LEW

[611] Elliot Soloway. Log on education: Ways of seeing. Communications of the ACM, 37(2):15–20, February 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/175244.html.

Goodman:1994:IPC

[612] S. E. Goodman. International perspectives: Computing in south africa: An end to “apartness”? Communications of the ACM, 37(2):21–25, February 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/175245.html.

Gronbaek:1994:HSD

[613] Kaj Grønbæk and Randall H. Trigg. Hypermedia system design applying the Dexter model. Communications of the ACM, 37(2):26–29, February 1994. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/175236.html.

Halasz:1994:DHR

[614] Frank Halasz and Mayer Schwartz. The Dexter hypertext reference model. Communications of the ACM, 37(2): 30–39, February 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/175237.html.

Gronbaek:1994:DID

[615] Kaj Grønbæk and Randall H. Trigg. Design issues for a Dexter-based hypermedia system. Communications of the ACM, 37(2):40–49, February 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/175238.html.

Hardman:1994:AHM

[616] Lynda Hardman, Dick C. A. Bulterman, and Guido van Rossum. The Amsterdam hypermedia model: adding time and context to the Dexter model. Communications of the ACM, 37(2): 50–62, February 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/175239.html.

Gronbaek:1994:CHS

[617] Kaj Grønbæk, Jens A. Hem, Ole L. Madsen, and Lennert Sloth. Cooperative hypermedia systems: a Dexter-based architecture. Communications of the ACM, 37(2):64–74, February 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/175240.html.

Leggett:1994:VDO

[618] John J. Leggett and John L. Schnase. Viewing Dexter with open eyes. Communications of the ACM, 37(2):76–86, February 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/175241.html.

Rivlin:1994:NHD

[619] Ehud Rivlin, Rodrigo Botafogo, and Ben Shneiderman. Navigating in hyperspace: designing a structure-based toolbox. Communications of the ACM, 37(2):87–96, February 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/175242.html.

Salton:1994:ASR

[620] Gerard Salton, James Allan, and Chris Buckley. Automatic structuring and retrieval of large text files. Communications of the ACM, 37(2): 97–108, February 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/175243.html.

Vanthienen:1994:TCM

[621] Jan Vanthienen, Girish Subramanian, and John Nosek. Technical correspondence: a more general comparison of the decision table and tree.

Communications of the ACM, 37(2): 109–113, February 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dorsett:1994:IRS

[622] Robert D. Dorsett. Inside risks: Safety in the air. Communications of the ACM, 37(2):146, February 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/175246.html.

Press:1994:PCE

[623] Larry Press. Personal computing: Emerging Dynabase tools. Communications of the ACM, 37(3):11–16, March 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/175260.html.

Weingarten:1994:WPI

[624] Fred W. Weingarten. From Washington: Public interest and the NII. Communications of the ACM, 37 (3):17–19, March 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/175261.html.

Munakata:1994:CIA

[625] Toshinori Munakata. Commercial and industrial AI. Communications of the ACM, 37(3):23–26, March 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/175248.html.

Hayes-Roth:1994:SKB

[626] Frederick Hayes-Roth and Neil Jacobstein. The state of knowledge-based systems. Communications of the ACM, 37(3):26–39, March 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/175249.html.

Allen:1994:CBR

[627] Bradley P. Allen. Case-based reasoning: business applications. Communications of the ACM, 37(3):40– 42, March 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/175250.html.

Grimson:1994:CVA

[628] W. E. L. Grimson and J. L. Mundy. Computer vision applications. Communications of the ACM, 37(3):44– 51, March 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/175251.html.

Rudnicky:1994:SCS

[629] Alexander I. Rudnicky, Alexander G. Hauptmann, and Kai-Fu Lee. Survey of current speech technology. Communications of the ACM, 37 (3):52–57, March 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/175252.html.

Kanade:1994:NTA

[630] Takeo Kanade, Michael L. Reed, and Lee E. Weiss. New technologies and applications in robotics. Communications of the ACM, 37 (3):58–67, March 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/175253.html.

Munakata:1994:FSO

[631] Toshinori Munakata and Yashvant Jani. Fuzzy systems: an overview. Communications of the ACM, 37 (3):68–76, March 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/175254.html.

Zadeh:1994:FLN

[632] Lotfi A. Zadeh. Fuzzy logic, neural networks, and soft computing. Communications of the ACM, 37 (3):77–84, March 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/175255.html.

Rumelhart:1994:BIN

[633] David E. Rumelhart, Bernard Widrow, and Michael A. Lehr. The basic ideas in neural networks. Communications of the ACM, 37(3):87–92, March 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/175256.html.

Widrow:1994:NNA

[634] Bernard Widrow, David E. Rumelhart, and Michael A. Lehr. Neural networks: Applications in industry, business, and science. Communications of the ACM, 37(3):93–105, March 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/175257.html.

Asakawa:1994:NNJ

[635] Kazuo Asakawa and Hideyuki Takagi. Neural networks in Japan. Communications of the ACM, 37 (3):106–112, March 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/175258.html.

Goldberg:1994:GEA

[636] David E. Goldberg. Genetic and evolutionary algorithms come of age. Communications of the ACM, 37 (3):113–119, March 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/175259.html.

Neumann:1994:IRT

[637] Peter G. Neumann. Inside risks: Technology, laws, and society. Communications of the ACM, 37(3):138, March 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/175262.html.

Rada:1994:SSR

[638] Roy Rada, George S. Carson, and Christopher Haynes. Sharing standards: The role of consensus. Communications of the ACM, 37(4):15– 16, April 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/175286.html.

Rettig:1994:PPP

[639] Marc Rettig. Practical programmer: Prototyping for tiny fingers. Communications of the ACM, 37 (4):21–27, April 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/175288.html.

Tentner:1994:HPC

[640] Adrian M. Tentner. High performance computing grand challenges in simulation. Communications of the ACM, 37(4):28–29, April 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/175278.html.

Camp:1994:MPM

[641] W. J. Camp, S. J. Plimpton, B. A. Hendrickson, and R. W. Leland. Massively parallel methods for engineering and science problems. Communications of the ACM, 37(4):30– 41, April 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/175279.html.

Jacobson:1994:LIS

[642] Allan S. Jacobson, Andrew L. Berkin, and Martin N. Orton. LinkWinds: interactive scientific data analysis and visualization. Communications of the ACM, 37(4):42–52, April 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/175280.html.

Tentner:1994:APC

[643] A. M. Tentner, R. N. Blomquist, P. L. Garner, E. M. Gelbard, K. C. Gross, M. Minkoff, R. A. Valentin, and T. R. Canfield. Advances in parallel computing for reactor analysis and safety. Communications of the ACM, 37(4):54–64, April 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/175281.html.

Nielsen:1994:MUP

[644] Jakob Nielsen and Jonathan Levy. Measuring usability: preference vs. performance. Communications of the ACM, 37(4):66–75, April 1994. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/175282.html.

Coll:1994:GTF

[645] Richard A. Coll, Joan H. Coll, and Ganesh Thakur. Graphs and tables: a four-factor experiment. Communications of the ACM, 37(4):76– 86, April 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/175283.html.

CSTB:1994:ACE

[646] Computer Science and Telecommunications Board. Academic careers for experimental computer scientists and engineers. Communications of the ACM, 37(4):87–90, April 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/175285.html.

Neumann:1994:IRRa

[647] Peter G. Neumann. Inside RISKS: Risks of passwords. Communications of the ACM, 37(4):126, April 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/175289.html.

Barlow:1994:EFD

[648] John Perry Barlow. Electronic frontier: Dad’s invisible guard-all shield. Communications of the ACM, 37(5): 15–17, May 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/175307.html.

Schank:1994:LET

[649] Roger C. Schank. Log on education: Tractor factories and research in software design. Communications of the ACM, 37(5):19–21, May 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/175309.html.

Waters:1994:REP

[650] Richard G. Waters and Elliot Chikofsky. Reverse engineering: progress along many dimensions. Communications of the ACM, 37(5):22–25, May 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/175291.html.

Aiken:1994:DLS

[651] Peter Aiken, Alice Muntz, and Russ Richards. DoD legacy systems: reverse engineering data requirements. Communications of the ACM, 37(5): 26–41, May 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/175292.html.

Premerlani:1994:ARE

[652] William J. Premerlani and Michael R. Blaha. An approach for reverse engineering of relational databases. Communications of the ACM, 37(5):42– 49, May 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/175293.html.

Ning:1994:ASL

[653] Jim Q. Ning, Andre Engberts, and W. Voytek Kozaczynski. Automated support for legacy code understanding. Communications of the ACM, 37(5):50–57, May 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/175295.html.

Markosian:1994:UET

[654] Lawrence Markosian, Philip Newcomb, Russell Brand, Scott Burson, and Ted Kitzmiller. Using an enabling technology to reengineer legacy systems. Communications of the ACM, 37(5):58–70, May 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/175297.html.

Biggerstaff:1994:PUC

[655] Ted J. Biggerstaff, Bharat G. Mitbander, and Dallas E. Webster. Program understanding and the concept assignment problem. Communications of the ACM, 37(5):72–82, May 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/175300.html.

Quilici:1994:MBA

[656] Alex Quilici. A memory-based approach to recognizing programming plans. Communications of the ACM, 37(5):84–93, May 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/175301.html.

Lieberherr:1994:AOO

[657] Karl J. Lieberherr, Ignacio Silva-Lepe, and Cun Xiao. Adaptive object-oriented programming using graph-based customization. Communications of the ACM, 37(5):94– 101, May 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/175303.html.

Vessey:1994:RSL

[658] Iris Vessey and Sue A. Conger. Requirements specification: learning object, process, and data methodologies. Communications of the ACM, 37(5):102–113, May 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/175305.html.

Neumann:1994:IRA

[659] Peter G. Neumann. Inside RISKS: Alternative passwords. Communications of the ACM, 37(5):146, May 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/175311.html.

Weingarten:1994:WPS

[660] Fred W. Weingarten. From Washington: The politicizing of science policy. Communications of the ACM, 37(6):13–15, June 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/175219.html.

Ariav:1994:IPI

[661] G. Ariav and S. E. Goodman. International perspectives: Israel: Of swords and software plowshares. Communications of the ACM, 37(6):17– 21, June 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/175220.html.

Ullah:1994:MPI

[662] Nasr Ullah and Philip K. Brownfield. The making of the PowerPC: introduction. Communications of the ACM, 37(6):22, June 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/175209.html.

Moore:1994:PA

[663] Charles R. Moore and Russell C. Stanphill. The PowerPC alliance. Communications of the ACM, 37(6):25– 27, June 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/175210.html.

Diefendorff:1994:HPA

[664] Keith Diefendorff. History of the PowerPC architecture. Communications of the ACM, 37(6):28–33, June 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/175211.html.

Burgess:1994:PM

[665] Brad Burgess, Nasr Ullah, Peter Van Overen, and Deene Ogden. The PowerPC 603 microprocessor. Communications of the ACM, 37(6):34– 42, June 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/175212.html.

Suessmith:1994:PMP

[666] Brad W. Suessmith and George Paap, III. PowerPC 603 microprocessor power management. Communications of the ACM, 37(6):43– 46, June 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/175213.html.

Poursepanj:1994:PPM

[667] Ali Poursepanj. The PowerPC performance modeling methodology. Communications of the ACM, 37(6):47– 55, June 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/175214.html.

Shipnes:1994:MAM

[668] Julie Shipnes and Mike Phillip. A modular approach to Motorola PowerPC compilers. Communications of the ACM, 37(6):56–63, June 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/175215.html.

Anderson:1994:OMP

[669] William Anderson. An overview of Motorola’s PowerPC simulator family. Communications of the ACM, 37(6):64–69, June 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/175216.html.

Shieber:1994:LRT

[670] Stuart M. Shieber. Lessons from a restricted Turing test. Communications of the ACM, 37(6):70– 78, June 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/175217.html.

Loebner:1994:R

[671] Hugh Gene Loebner. In response. Communications of the ACM, 37 (6):79–82, June 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/175218.html.

Shieber:1994:LL

[672] Stuart M. Shieber. On Loebner’s lessons. Communications of the ACM, 37(6):83–84, June 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/175604.html.

LaMalva:1994:APC

[673] Patrick M. LaMalva and George D. Peterson. Accredited programs in computing. Communications of the ACM, 37(6):85–88, June 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:1994:IRRb

[674] Peter G. Neumann. Inside RISKS: Risks on the information superhighway. Communications of the ACM, 37(6):114, June 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/175221.html.

Press:1994:PCT

[675] Larry Press. Personal computing: Tomorrow’s campus. Communications of the ACM, 37(7):13– 17, July 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/176806.html.

Riecken:1994:IA

[676] Doug Riecken. Intelligent agents. Communications of the ACM, 37(7):18– 21, July 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/176790.html.

Minsky:1994:CMM

[677] Marvin Minsky and Doug Riecken. A conversation with Marvin Minsky about agents. Communications of the ACM, 37(7):22–29, July 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/176791.html.

Maes:1994:ARW

[678] Pattie Maes. Agents that reduce work and information overload. Communications of the ACM, 37(7):30– 40, July 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/176792.html.

Edmonds:1994:SCD

[679] Ernest A. Edmonds, Linda Candy, Rachel Jones, and Bassel Soufi. Support for collaborative design: agents and emergence. Communications of the ACM, 37(7):41–47, July 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/176793.html.

Genesereth:1994:SA

[680] Michael R. Genesereth and Steven P. Ketchpel. Software agents. Communications of the ACM, 37(7):48– 53, July 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/176794.html.

Smith:1994:KPA

[681] David Canfield Smith, Allen Cypher, and Jim Spohrer. KidSim: programming agents without a programming language. Communications of the ACM, 37(7):54–67, July 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/176795.html.

Norman:1994:HMP

[682] Donald A. Norman. How might people interact with agents. Communications of the ACM, 37(7):68– 71, July 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/176796.html.

Etzioni:1994:SBI

[683] Oren Etzioni and Daniel Weld. A softbot-based interface to the Internet. Communications of the ACM, 37(7):72–76, July 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/176797.html.

Mitchell:1994:ELP

[684] Tom M. Mitchell, Rich Caruana, Dayne Freitag, John McDermott, and David Zabowski. Experience with a learning personal assistant. Communications of the ACM, 37(7):80– 91, July 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/176798.html.

Selker:1994:CTA

[685] Ted Selker. COACH: a teaching agent that learns. Communications of the ACM, 37(7):92–99, July 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/176799.html.

Greif:1994:DAG

[686] Irene Greif. Desktop agents in group-enabled products. Communications of the ACM, 37(7):100– 105, July 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/176800.html.

Riecken:1994:MAI

[687] Doug Riecken. M: an architecture of integrated agents. Communications of the ACM, 37(7):106– 116, July 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/176801.html.

Boden:1994:AC

[688] Margaret A. Boden. Agents and creativity. Communications of the ACM, 37(7):117–121, July 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/176802.html.

Bates:1994:REB

[689] Joseph Bates. The role of emotion in believable agents. Communications of the ACM, 37(7):122– 125, July 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/176803.html.

Guha:1994:EAW

[690] R. V. Guha and Douglas B. Lenat. Enabling agents to work together. Communications of the ACM, 37 (7):126–142, July 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/176804.html.

Kautz:1994:BDS

[691] Henry A. Kautz, Bart Selman, and Michael Coen. Bottom-up design of software agents. Communications of the ACM, 37(7):143–146, July 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/176805.html.

Simons:1994:IRQ

[692] Barbara Simons. Inside RISKS: Questions about the NII. Communications of the ACM, 37(7):170, 146, July 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/176807.html.

Kahn:1994:VRG

[693] Robert E. Kahn. Viewpoint: The role of government in the evolution of the Internet. Communications of the ACM, 37(8):15–19, August 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/179729.html.

Samuelson:1994:LSS

[694] Pamela Samuelson. Legally speaking: Self-plagiarism or fair use? Communications of the ACM, 37 (8):21–25, August 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/179731.html.

Goodman:1994:IPG

[695] S. E. Goodman, L. I. Press, S. R. Ruth, and A. M. Rutkowski. International perspectives: The global diffusion of the Internet: Patterns and problems. Communications of the ACM, 37(8):27–31, August 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/179733.html.

Leiner:1994:IT

[696] Barry M. Leiner. Internet technology. Communications of the ACM, 37(8):32–33, August 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/179611.html.

Claffy:1994:TLT

[697] Kimberly C. Claffy, Hans-Werner Braun, and George C. Polyzos. Tracking long-term growth of the NSFNET. Communications of the ACM, 37 (8):34–45, August 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/179616.html.

Eidnes:1994:PCN

[698] Håvard Eidnes. Practical considerations for network addressing using CIDR. Communications of the ACM, 37(8):46–53, August 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/179622.html.

Eriksson:1994:MMB

[699] Hans Eriksson. MBONE: the multicast backbone. Communications of the ACM, 37(8):54–60, August 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/179627.html.

Moy:1994:MRE

[700] John Moy. Multicast routing extensions for OSPF. Communications of the ACM, 37(8):61–66, August 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/179654.html.

Teraoka:1994:VPP

[701] Fumio Teraoka, Keisuke Uehara, Hideki Sunahara, and Jun Murai. VIP: a protocol providing host mobility. Communications of the ACM, 37(8):67–75, August 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/179657.html.

Berners-Lee:1994:WWW

[702] Tim Berners-Lee, Robert Cailliau, Ari Luotonen, Henrik Frystyk Nielsen, and Arthur Secret. The World-Wide Web. Communications of the ACM, 37(8):76–82, August 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/179671.html.

Ishii:1994:IDS

[703] Hiroshi Ishii, Minoru Kobayashi, and Kazuho Arita. Iterative design of seamless collaboration media. Communications of the ACM, 37(8):83– 97, August 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/179687.html.

Bowman:1994:SIR

[704] C. Mic Bowman, Peter B. Danzig, Udi Manber, and Michael F. Schwartz. Scalable Internet resource discovery: research problems and approaches. Communications of the ACM, 37 (8):98–107, August 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/179704.html.

Fithen:1994:CIR

[705] Katherine Fithen and Barbara Fraser. CERT incident response and the Internet. Communications of the ACM, 37(8):108–114, August 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/179721.html.

Landau:1994:CPP

[706] Susan Landau, Stephen Kent, Clinton C. Brooks, Scott Charney, Dorothy E. Denning, Whitfield Diffie, Anthony Lauck, Douglas Miller, Peter G. Neumann, and David L. Sobel. Crypto policy perspectives. Communications of the ACM, 37(8): 115–121, August 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/179726.html.

Neumann:1994:IRF

[707] Peter G. Neumann. Inside RISKS: Friendly fire. Communications of the ACM, 37(8):138, August 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/toc/Abstracts/0001-0782/179742.
html.

Kafai:1994:LEC

[708] Yasmin Kafai and Elliot Soloway. Log on education: Computational gifts for the Barney generation. Communications of the ACM, 37(9):19–22, September 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rada:1994:SSN

[709] Roy Rada and George S. Carson. Sharing standards: The new media. Communications of the ACM, 37(9): 23–25, September 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/214782.html. Discusses

the problem of the multiplicity of document markup formats, and the role that SGML and OpenDoc might play in providing a common intermediate format for document exchange.

Binder:1994:OOS

[710] Robert V. Binder. Object-oriented software testing. Communications of the ACM, 37(9):28–29, September 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/182988.html.

Jorgensen:1994:OOI

[711] Paul C. Jorgensen and Carl Erickson. Object-oriented integration testing. Communications of the ACM, 37 (9):30–38, September 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/182989.html.

Murphy:1994:ECC

[712] Gail C. Murphy, Paul Townsend, and Pok Sze Wong. Experiences with cluster and class testing. Communications of the ACM, 37(9):39–47, September 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/182990.html.

Poston:1994:ATO

[713] Robert M. Poston. Automated testing from object models. Communications of the ACM, 37(9):48–58, September 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/184074.html.

McGregor:1994:IOO

[714] John D. McGregor and Timothy D. Korson. Integrated object-oriented testing and development processes. Communications of the ACM, 37(9): 59–77, September 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/184075.html.

Arnold:1994:TPW

[715] Thomas R. Arnold and William A. Fuson. Testing “in A perfect world”. Communications of the ACM, 37(9): 78–86, September 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/184076.html.

Binder:1994:DTO

[716] Robert V. Binder. Design for testability in object-oriented systems. Communications of the ACM, 37(9):87– 101, September 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/184077.html.

Hart:1994:DSC

[717] George W. Hart. To decode short cryptograms. Communications of the ACM, 37(9):102–108, September 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/184078.html.

Hoffman:1994:CP

[718] Lance J. Hoffman, Faraz A. Ali, Steven L. Heckler, and Ann Huybrechts. Cryptography policy. Communications of the ACM, 37(9):109– 117, September 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/184079.html.

Baclawski:1994:TCN

[719] Kenneth Baclawski and Bipin Indurkhya. Technical correspondence: The notion of inheritance in objectoriented programming. Communications of the ACM, 37(9):118–119, September 1994. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/184080.html.

Neumann:1994:IRE

[720] Peter G. Neumann. Inside RISKS: Expectations of security and privacy. Communications of the ACM, 37(9):138, September 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/184081.html.

Weingarten:1994:WUR

[721] Fred W. Weingarten. From Washington: University research in a squeeze. Communications of the ACM, 37 (10):15–17, October 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/194315.html.

Imielinski:1994:MWC

[722] Tomasz Imieliński and B. R. Badrinath. Mobile wireless computing: challenges in data management. Communications of the ACM, 37(10):18– 28, October 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/194317.html.

Oz:1994:WPS

[723] Effy Oz. When professional standards are lax: The CONFIRM failure and its lessons. Communications of the ACM, 37(10):29–36, October 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/194319.html.

Hartmanis:1994:TAL

[724] Juris Hartmanis. Turing Award Lecture: On computational complexity and the nature of computer science. Communications of the ACM, 37(10):37–43, October 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/214781.html.

Watson:1994:CFD

[725] Richard T. Watson, Teck Hua Ho, and K. S. Raman. Culture: a fourth dimension of group support systems. Communications of the ACM, 37(10):44–55, October 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/194320.html.

Chen:1994:ACC

[726] H. Chen, P. Hsu, R. Orwig, L. Hoopes, and J. F. Nunamaker. Automatic concept classification of text from electronic meetings. Communications of the ACM, 37(10):56–73, October 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/194322.html.

Fraser:1994:SIF

[727] Martin D. Fraser, Kuldeep Kumar, and Vijay K. Vaishnavi. Strategies for incorporating formal specifications in software development. Communications of the ACM, 37 (10):74–86, October 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/194399.html.

Wild:1994:FSS

[728] Rosemary H. Wild and Joseph J. Pignatiello, Jr. Finding stable system designs: a reverse simulation technique. Communications of the ACM, 37(10):87–98, October 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/194405.html.

Wright:1994:IRV

[729] Benjamin Wright. Inside RISKS: The verdict on plaintext signatures: They’re legal. Communications of the ACM, 37(10):122, October 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/175203.html.

BloomBecker:1994:VMT

[730] Jay J. Buck BloomBecker. Viewpoint: My three computer criminological sins. Communications of the ACM, 37(11):15–16, November 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/188283.html.

Press:1994:CI

[731] Larry Press. Commercialization of the Internet. Communications of the ACM, 37(11):17–21, November 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/188286.html.

Nejmeh:1994:IST

[732] Brian A. Nejmeh. Internet: a strategic tool for the software enterprise. Communications of the ACM, 37(11): 23–27, November 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/188288.html.

Ganesan:1994:SC

[733] Ravi Ganesan and Ravi Sandhu. Securing cyberspace. Communications of the ACM, 37(11):28–31, November 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/188289.html.

Anderson:1994:WCF

[734] Ross J. Anderson. Why cryptosystems fail. Communications of the ACM, 37 (11):32–40, November 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/188291.html.

Needham:1994:DSE

[735] Roger M. Needham. Denial of service: an example. Communications of the ACM, 37(11):42–46, November 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/188294.html.

Hauser:1994:DLR

[736] Ralf C. Hauser. Does licensing require new access control technologies? Communications of the ACM, 37 (11):48–55, November 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/188295.html.

Simmons:1994:CPF

[737] Gustavus J. Simmons. Cryptanalysis and protocol failures. Communications of the ACM, 37(11): 56–65, November 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/188298.html.

Clark:1994:BSP

[738] Paul C. Clark and Lance J. Hoffman. BITS: A Smartcard protected operating system. Communications of the ACM, 37(11):66–70, November 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/188371.html.

Kirani:1994:EES

[739] Shekhar H. Kirani, Imran A. Zualkernan, and Wei-Tek Tsai. Evaluation of expert system testing methods. Communications of the ACM, 37 (11):71–81, November 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/188373.html.

Blum:1994:TSD

[740] Bruce I. Blum. A taxonomy of software development methods. Communications of the ACM, 37(11): 82–94, November 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/188377.html.

Stearns:1994:TAL

[741] Richard Edwin Stearns. Turing award lecture: It’s time to reconsider time. Communications of the ACM, 37(11): 95–99, November 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/188379.html.

Grosch:1994:DW

[742] Herb Grosch. Dehumanizing the workplace. Communications of the ACM, 37(11):122, November 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/188381.html.

Hanson:1994:VCW

[743] Robin Hanson. Viewpoint: Can wiretaps remain cost-effective? Communications of the ACM, 37(12): 13–15, December 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/198367.html.

Frenkel:1994:MGS

[744] Karen Frenkel. In memory: Gordon Smith 1919–1994. Communications of the ACM, 37(12):16, December 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/198368.html.

Moore:1994:SSL

[745] James W. Moore, David Emery, and Roy Rada. Sharing standards: Language-independent standards. Communications of the ACM, 37(12):17–20, December 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/198369.html.

Samuelson:1994:LSN

[746] Pamela Samuelson. Legally speaking: The NII intellectual property report. Communications of the ACM, 37 (12):21–27, December 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/198372.html.

Carroll:1994:MUD

[747] John M. Carroll. Making use a design representation. Communications of the ACM, 37(12):28–35, December 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/198374.html.

Lohse:1994:CVR

[748] Gerald L. Lohse, Kevin Biolsi, Neff Walker, and Henry H. Rueter. A classification of visual representations. Communications of the ACM, 37(12): 36–49, December 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/198376.html.

Eick:1994:GAC

[749] Stephen G. Eick, Michael C. Nelson, and Jeffery D. Schmidt. Graphical analysis of computer log files. Communications of the ACM, 37(12): 50–56, December 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/198378.html.

Madsen:1994:GMD

[750] Kim Halskov Madsen. A guide to metaphorical design. Communications of the ACM, 37(12):57–62, December 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/198381.html.

Bagley:1994:EIT

[751] Steven C. Bagley and Gary E. Kopec. Editing images of text. Communications of the ACM, 37(12):63–72, December 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/198382.html.

Sarkar:1994:GFV

[752] Manojit Sarkar and Marc H. Brown. Graphical fisheye views. Communications of the ACM, 37(12):73–84, December 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/198384.html.

Stevenson:1994:SCS

[753] D. E. Stevenson. Science, computational science, and computer science: at a crossroads. Communications of the ACM, 37(12):85–96, December 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/198386.html.

Johnson:1994:MDC

[754] Gerald J. Johnson. Of metaphor and the difficulty of computer discourse. Communications of the ACM, 37(12): 97–102, December 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/198388.html.

Neumann:1994:IRI

[755] Peter G. Neumann. Inside RISKS: Inside “inside RISKS.”. Communications of the ACM, 37(12):154, December 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goodman:1995:IPC

[756] S. E. Goodman and L. I. Press. International perspectives: Computing in Vietnam: An Asian tiger in the rough. Communications of the ACM, 38(1):11–16, January 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/204867.html.

Soloway:1995:LEB

[757] Elliot Soloway. Log on education: Beware, techies bearing gifts. Communications of the ACM, 38 (1):17–24, January 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/204871.html.

Pearl:1995:WCI

[758] Amy Pearl. Women in computing: Introduction. Communications of the ACM, 38(1):26–28, January 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/204873.html.

Klawe:1995:WCW

[759] Maria Klawe and Nancy Leveson. Women in computing: where are we now? Communications of the ACM, 38(1):29–35, January 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/204874.html.

Diwanji:1995:CSW

[760] Pavani Diwanji. Corporate supported women’s groups. Communications of the ACM, 38(1):36, January 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Baylor:1995:GFP

[761] Sandra Johnson Baylor. Graduate fellowship programs. Communications of the ACM, 38(1):37–42, January 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Adam:1995:WCU

[762] Alison Adam. Women and computing in the UK. Communications of the ACM, 38(1):43, January 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Downey:1995:ANR

[763] Laura L. Downey. Announcing a new resource: The WCAR list. Communications of the ACM, 38(1):43–44, January 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Freeman:1995:TTI

[764] Elisabeth Freeman and Susanne Hupfer. TAP: Tapping Internet resources for women in computer science. Communications of the ACM, 38(1):44, January 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gurer:1995:PWC

[765] Denise W. Gürer. Pioneering women in computer science. Communications of the ACM, 38(1):45–54, January 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/204875.html.

Borg:1995:GHC

[766] Anita Borg and Telle Whitney. The Grace Hopper celebration. Communications of the ACM, 38(1):50–51, January 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hemenway:1995:HNG

[767] Kathleen Hemenway. Human nature and the glass ceiling in industry. Communications of the ACM, 38(1):55–62, January 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/204878.html.

Isaacs:1995:GDW

[768] Ellen Isaacs. Gender discrimination in the workplace: a literature review. Communications of the ACM, 38(1):58–59, January 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Camp:1995:DR

[769] Tracy Camp. Diversity recruiting. Communications of the ACM, 38(1): 61, January 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pfleeger:1995:EMW

[770] Shari Lawrence Pfleeger and Norma Mertz. Executive mentoring: what makes it work? Communications of the ACM, 38(1):63–73, January 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/204880.html.

Goyal:1995:MRP

[771] Amita Goyal. Mentoring resources and programs for women. Communications of the ACM, 38(1):66–67, January 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Oechtering:1995:SAM

[772] Veronika Oechtering and Roswitha Behnke. Situations and advancement measures in Germany. Communications of the ACM, 38(1):75–82, January 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/204881.html.

Vessey:1995:CTC

[773] Iris Vessey and Ajay Paul Sravanapudi. CASE tools as collaborative support technologies. Communications of the ACM, 38(1):83–95, January 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/204882.html.

Mehlhorn:1995:LPC

[774] Kurt Mehlhorn and Stefan Näher. LEDA: a platform for combinatorial and geometric computing. Communications of the ACM, 38(1):96–102, January 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/204889.html.

Long:1995:USD

[775] Jeffrey G. Long and Dorothy E. Denning. Ultra-structure: a design theory for complex systems and processes. Communications of the ACM, 38(1):103–120, January 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/204892.html.

Herring:1995:TCR

[776] Charles Herring and Julian I. Palmore. Technical correspondence: Random number generators are chaotic. Communications of the ACM, 38(1): 121–122, January 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/204895.html.

Ross:1995:TCP

[777] D. E. Ross. Technical correspondence: Pseudo-random number generators for a calculator. Communications of the ACM, 38(1):122–124, January 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/204896.html.

Shneiderman:1995:IRI

[778] Ben Shneiderman. Inside risks: The info superhighway: For the people. Communications of the ACM, 38(1):162, January 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/204898.html.

Zweben:1995:PLA

[779] Stuart Zweben. President’s letter: ACM Computing Week ’95. Communications of the ACM, 38(2): 13–14, February 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/204827.html.

Weingarten:1995:WRT

[780] Fred W. Weingarten. From Washington: Reforming telecommunications law. Communications of the ACM, 38(2):19–22, February 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/204828.html.

Rada:1995:SSS

[781] Roy Rada and John Berg. Sharing standards: Standards: Free or sold? Communications of the ACM, 38(2):23–27, February 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/204829.html.

Vetter:1995:ICA

[782] Ronald J. Vetter and David H. C. Du. Issues and challenges in ATM networks. Communications of the ACM, 38(2):28–29, February 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/204830.html.

Vetter:1995:ACA

[783] Ronald J. Vetter. ATM concepts, architectures, and protocols. Communications of the ACM, 38(2):30–38, February 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/204831.html.

Kim:1995:ANG

[784] B. G. Kim and P. Wang. ATM network: goals and challenges. Communications of the ACM, 38(2):39–44, February 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/204832.html.

Stevenson:1995:SCA

[785] Daniel Stevenson, Nathan Hillery, and Greg Byrd. Secure communications in ATM networks. Communications of the ACM, 38(2):45–52, February 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/204844.html.

Schnepf:1995:BFM

[786] James A. Schnepf, David H. C. Du, E. Russell Ritenour, and Aaron J. Fahrmann. Building future medical education environments over ATM networks. Communications of the ACM, 38(2):54–69, February 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/204845.html.

Iwata:1995:ACT

[787] A. Iwata, N. Mori, C. Ikeda, H. Suzuki, and M. Ott. ATM connection and traffic management schemes for multimedia Internetworking. Communications of the ACM, 38(2):72–89, February 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/204846.html.

Perloff:1995:ITP

[788] Michael Perloff and Kurt Reiss. Improvements to TCP performance in high-speed ATM networks. Communications of the ACM, 38(2):90–100, February 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/204849.html.

Fotedar:1995:AVP

[789] Shivi Fotedar, Mario Gerla, Paola Crocetti, and Luigi Fratta. ATM virtual private networks. Communications of the ACM, 38(2):101–109, February 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/204852.html.

LaMalva:1995:APC

[790] Patrick M. LaMalva and George D. Peterson. Accredited programs in computing. Communications of the ACM, 38(2):110–114, February 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Debest:1995:TCR

[791] Xavier A. Debest. Technical correspondence: Remark about self-stabilizing systems. Communications of the ACM, 38(2):115–117, February 1995. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Buhr:1995:TCS

[792] Peter A. Buhr. Technical correspondence: Are safe concurrency libraries possible. Communications of the ACM, 38(2):117–120, February 1995. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Demchak:1995:IRC

[793] C. C. Demchak and S. E. Goodman. Inside risks: Computers as substitute soldiers. Communications of the ACM, 38(2):154, February 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/204854.html.

Press:1995:PCI

[794] Larry Press. Personal computing: The Internet is not TV: Web publishing. Communications of the ACM, 38(3):17–23, March 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://info.cern.ch/
hypertext/Web/Daemon/Overview.html;
 http://www.acm.org/pubs/toc/Abstracts/
0001-0782/203332.html.

Heckerman:1995:RWA

[795] David Heckerman, Abe Mamdani, and Michael P. Wellman. Real-world applications of Bayesian networks: Introduction. Communications of the ACM, 38(3):24–26, March 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/203334.html.

Heckerman:1995:BN

[796] David Heckerman and Michael P. Wellman. Bayesian networks. Communications of the ACM, 38(3):27– 30, March 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/203336.html.

Burnell:1995:SCM

[797] Lisa Burnell and Eric Horvitz. Structure and chance: melding logic and probability for software debugging. Communications of the ACM, 38(3): 31–41, 57, March 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/203338.html.

Fung:1995:ABN

[798] Robert Fung and Brendan Del Favero. Applying Bayesian networks to information retrieval. Communications of the ACM, 38(3):42–48, March 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/203340.html.

Heckerman:1995:DTT

[799] David Heckerman, John S. Breese, and Koos Rommelse. Decisiontheoretic troubleshooting. Communications of the ACM, 38(3):49–57, March 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/203341.html.

Tesauro:1995:TDL

[800] Gerald Tesauro. Temporal difference learning and TD-gammon. Communications of the ACM, 38(3):58– 68, March 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/203343.html.

Kraut:1995:CSD

[801] Robert E. Kraut and Lynn A. Streeter. Coordination in software development. Communications of the ACM, 38(3):69–81, March 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/203345.html.

Igbaria:1995:RDJ

[802] Magid Igbaria and Wayne M. Wormley. Race differences in job performance and career success. Communications of the ACM, 38(3):82– 92, March 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/203347.html.

Walstrom:1995:FMI

[803] Kent A. Walstrom, Bill C. Hardgrave, and Rick L. Wilson. Forums for management information systems scholars. Communications of the ACM, 38(3):93–107, March 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/203348.html.

Arensburger:1995:TOT

[804] Andrew Arensburger and Azriel Rosenfeld. Technical opinion: To take arms against a sea of email. Communications of the ACM, 38(3):108– 109, March 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/203350.html.

Neumann:1995:RCD

[805] Peter G. Neumann. Reassessing the crypto debate. Communications of the ACM, 38(3):138, March 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/203352.html.

Samuelson:1995:LSC

[806] Pamela Samuelson. Legally speaking: Copyright and digital libraries. Communications of the ACM, 38(4): 15–21, 110, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/205324.html.

Fox:1995:DLI

[807] Edward A. Fox, Robert M. Akscyn, Richard K. Furuta, and John J. Leggett. Digital libraries: Introduction. Communications of the ACM, 38(4):22–28, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/205325.html.

Rao:1995:RID

[808] Ramana Rao, Jan O. Pedersen, Marti A. Hearst, Jock D. Mackinlay, Stuart K. Card, Larry Masinter, Per-Kristian Halvorsen, and George G. Robertson. Rich interaction in the digital library. Communications of the ACM, 38(4):29– 39, April 1995. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/205326.html.

Bell:1995:MRS

[809] Timothy C. Bell, Alistair Moffat, Ian H. Witten, and Justin Zobel. The MG retrieval system: compressing for space and speed. Communications of the ACM, 38(4):41– 42, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/205327.html.

Croft:1995:NCI

[810] W. Bruce Croft. NSF Center for Intelligent Information Retrieval. Communications of the ACM, 38(4):42–43, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://ciir.cs.umass.
edu/;   http://www.acm.org/pubs/
toc/Abstracts/0001-0782/205328.
html.

Fox:1995:WWW

[811] Edward A. Fox. World-Wide Web and computer science reports. Communications of the ACM, 38 (4):43–44, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/cs-tr.cs.cornell.edu/;   http:/
/fas.sfu.ca/1/projects/EPiCS/CS-
TechReports;   http://harvest.cs.
colorado.edu/brokers/cstech/query.
html;   http://www.acm.org/pubs/
toc/Abstracts/0001-0782/205329.
html;   http://www.cs.cmu.edu:8001/
afs/cs.cmu.edu/user/jblythe/Mosaic/
cs-reports.html; http://www.cs.

indiana.edu/cstr/search;   http:
//www.cs.odu.edu/WATERS/WATERS-
GS.html;   http://www.rdt.monash.
edu.au/tr/siteslist.html.

French:1995:WAT

[812] James C. French, Edward A. Fox, Kurt Maly, and Alan L. Selman. Wide Area Technical Report Service: Technical reports online. Communications of the ACM, 38(4):45, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/205330.html;   http://
www.cs.odu.edu/WATERS/WATERS-GS.
html.

Kacmar:1995:IZS

[813] Charles Kacmar and Dean Jue. The information zone system. Communications of the ACM, 38(4):46–47, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/205332.html.

Lagoze:1995:DAD

[814] Carl Lagoze and James R. Davis. Dienst: An architecture for distributed document libraries. Communications of the ACM, 38(4):47, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://cs-tr.cs.
cornell.edu/;   http://www.acm.
org/pubs/toc/Abstracts/0001-0782/
205331.html.

Huser:1995:KBE

[815] Christoph Hüser, Klaus Reichenberger, Lothar Rostek, and Norbert A. Streitz. Knowledge-based editing and visualization for hypermedia encyclopedias. Communications of the ACM, 38(4):49–51, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/205333.html.

Merrill:1995:UCC

[816] Deane Merrill, Nathan Parker, Fredric Gey, and Chris Stuber. The University of California CD-ROM Information System. Communications of the ACM, 38(4):51–52, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//cedr.lbl.gov/cdrom/doc/cdrom.
html;   http://www.acm.org/pubs/
toc/Abstracts/0001-0782/205334.
html.

Heath:1995:EUC

[817] Lenwood S. Heath, Deborah Hix, Lucy T. Nowell, William C. Wake, Guillermo A. Averboch, Eric Labow, Scott A. Guyer, Dennis J. Brueni, Robert K. France, Kaushal Dalal, and Edward A. Fox. Envision: a usercentered database for computer science literature. Communications of the ACM, 38(4):52–53, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://ei.cs.vt.edu/EIproj.html.

Entlich:1995:MDL

[818] Richard Entlich, Lorrin Garson, Michael Lesk, Lorraine Normore, Jan Olsen, and Stuart Weibel. Making a digital library: the chemistry online retrieval experiment. Communications of the ACM, 38(4):54, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/205335.html.

Olson:1995:ALR

[819] Gary M. Olson. An appreciation of Laurence Rosenberg. Communications of the ACM, 38(4):55, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/205336.html.

Christel:1995:IDV

[820] M. Christel, T. Kanade, M. Mauldin, R. Reddy, M. Sirbu, S. Stevens, and H. Wactlar. Informedia Digital Video Library. Communications of the ACM, 38(4):57–58, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://fuzine.mt.
cs.cmu.edu/im/informedia.html;
http://www.acm.org/pubs/toc/Abstracts/
0001-0782/205337.html.

SDLG:1995:SDL

[821] The Stanford Digital Libraries Group. The Stanford Digital Library Project. Communications of the ACM, 38 (4):59–60, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www-diglib.stanford.edu/diglib;
http://www.acm.org/pubs/toc/Abstracts/
0001-0782/205338.html.

Wilensky:1995:UBD

[822] Robert Wilensky. UC Berkeley’s Digital Library project. Communications of the ACM, 38(4):60, April 1995. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://elib.

cs.berkeley.edu/;   http://www.
acm.org/pubs/toc/Abstracts/0001-
0782/205339.html.

Smith:1995:ADL

[823] Terence R. Smith and James Frew. Alexandria digital library. Communications of the ACM, 38(4):61–62, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://alexandria.sdc.ucsb.
edu/;   http://www.acm.org/pubs/
toc/Abstracts/0001-0782/205340.
html.

Schatz:1995:BII

[824] Bruce Schatz. Building the interspace: The Illinois digital library project. Communications of the ACM, 38(4):62–63, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/205341.html;   http://
www.grainger.uiuc.edu/dli.

Crum:1995:UMD

[825] Laurie Crum. University of Michigan digital library project. Communications of the ACM, 38(4):63– 64, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/205342.html;   http://www.
sils.umich.edu/UMDL/HomePage.html.

Purday:1995:BLI

[826] Jonathan Purday. The British Library’s initiatives for access projects. Communications of the ACM, 38 (4):65–66, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL gopher:

//portico.bl.uk/;   http://www.
acm.org/pubs/toc/Abstracts/0001-
0782/205343.html.

Becker:1995:LCD

[827] Herbert S. Becker. Library of Congress digital library effort. Communications of the ACM, 38(4):66, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/205344.html;   http://www.
loc.gov/.

Marchionini:1995:RDL

[828] Gary Marchionini and Hermann Maurer. The roles of digital libraries in teaching and learning. Communications of the ACM, 38(4):67– 75, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/205345.html.

Levy:1995:GDL

[829] David M. Levy and Catherine C. Marshall. Going digital: a look at assumptions underlying digital libraries. Communications of the ACM, 38(4):77–84, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/205346.html.

Wiederhold:1995:DLV

[830] Gio Wiederhold. Digital libraries, value, and productivity. Communications of the ACM, 38(4):85–96, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/205347.html.

Denning:1995:AEP

[831] Peter J. Denning and Bernard Rous. The ACM electronic publishing plan. Communications of the ACM, 38 (4):97–109, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/205348.html.

Denning:1995:IRI

[832] Peter J. Denning. Inside risks: Information superhighway 2015. Communications of the ACM, 38(4):138, April 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/205349.html.

Robbins:1995:VTT

[833] Stuart Robbins. Viewpoint: Turbulent times for information technology. Communications of the ACM, 38(5):15–16, May 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/203357.html.

Barlow:1995:EFD

[834] John Perry Barlow. Electronic frontier: Death from above. Communications of the ACM, 38(5):17– 20, May 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/203358.html.

Dedrick:1995:IPL

[835] J. L. Dedrick, S. E. Goodman, and K. L. Kraemer. International perspectives: Little engines that could: Computing in small energetic countries. Communications of the ACM, 38(5):21–26, May 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/203359.html.

Holtzblatt:1995:RGH

[836] Karen Holtzblatt and Hugh R. Beyer. Requirements gathering: The human factor: Introduction. Communications of the ACM, 38(5):30– 32, May 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/203361.html.

Keil:1995:CDL

[837] Mark Keil and Erran Carmel. Customerdeveloper links in software development. Communications of the ACM, 38(5):33–44, May 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/203363.html.

Beyer:1995:AC

[838] Hugh R. Beyer and Karen Holtzblatt. Apprenticing with the customer. Communications of the ACM, 38(5):45– 52, May 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/203365.html.

Kjaer:1995:PAF

[839] Arne Kjær and Kim Halskov Madsen. Participatory analysis of flexibility. Communications of the ACM, 38(5):53–60, May 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/203367.html.

Brun-Cottan:1995:UVR

[840] Françoise Brun-Cottan and Patricia Wall. Using video to re-present the user. Communications of the ACM, 38(5):61–71, May 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/203368.html.

Hutchings:1995:CPC

[841] Anthony F. Hutchings and Steve T. Knox. Creating products customers demand. Communications of the ACM, 38(5):72–80, May 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/203370.html.

Allen:1995:SCC

[842] C. Dennis Allen. Succeeding as a clandestine change agent. Communications of the ACM, 38(5):81– 86, May 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/203371.html.

Clifton:1995:SAP

[843] Mitchell H. Clifton. Self-Assessment Procedure XXIII: Programming languages. Communications of the ACM, 38(5):89–97, May 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/203378.html.

Johnson:1995:TOC

[844] L. F. Johnson. Technical opinion: C in the first course considered harmful. Communications of the ACM, 38(5):99–101, May 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/203373.html.

Charette:1995:IRH

[845] Robert N. Charette. Inside risks: How to create a successful failure. Communications of the ACM, 38(5): 122, May 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/203375.html.

Moursund:1995:VSB

[846] David G. Moursund. Viewpoint: Sitebased management saving our schools. Communications of the ACM, 38(6): 13–15, June 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moursund:1995:SBM

[847] David G. Moursund. Site-based management: saving our schools. Communications of the ACM, 38(6):15– 17, June 1995. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/203243.html.

Nejmeh:1995:EIP

[848] Brian A. Nejmeh. Enterprising innovations: Process cost and value analysis. Communications of the ACM, 38(6):19–24, June 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/203246.html.

Rockman:1995:LES

[849] Saul Rockman. Log on education: In school or out: Technology, equity, and the future of our kids. Communications of the ACM, 38(6):25–29, June 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/203247.html.

Stacy:1995:CSD

[850] Webb Stacy. Cognition and software development. Communications of the ACM, 38(6):31, June 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/203250.html.

Petre:1995:WLI

[851] Marian Petre. Why looking isn’t always seeing: readership skills and graphical programming. Communications of the ACM, 38(6):33–44, June 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/203251.html.

Dumas:1995:DWP

[852] Joseph Dumas and Paige Parsons. Discovering the way programmers think about new programming environments. Communications of the ACM, 38(6):45–56, June 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/203253.html.

Stacy:1995:CBS

[853] Webb Stacy and Jean MacMillian. Cognitive bias in software engineering. Communications of the ACM, 38(6):57–63, June 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/203256.html.

Winograd:1995:PEE

[854] Terry Winograd. From programming environments to environments for designing. Communications of the ACM, 38(6):65–74, June 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/203259.html.

Frakes:1995:SQA

[855] William B. Frakes and Christopher J. Fox. Sixteen questions about software reuse. Communications of the ACM, 38(6):75–87, 112, June 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/203260.html.

Gagliano:1995:AAC

[856] Ross A. Gagliano, Martin D. Fraser, and Mark E. Schaefer. Auction allocation of computing resources. Communications of the ACM, 38(6):88– 102, June 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/203262.html.

Kim:1995:CDM

[857] Young-Gul Kim and Salvatore T. March. Comparing data modeling formalisms. Communications of the ACM, 38(6):103–115, June 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/203265.html.

Abelson:1995:FCC

[858] Hal Abelson, Kim Bruce, Andy van Dam, Brian Harvey, Allen Tucker, and Peter Wegner. The first-course conundrum. Communications of the ACM, 38(6):116–117, June 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/203266.html.

Horwitz:1995:FCC

[859] Susan Horwitz, Kenneth Appel, Theresa Cuprak, David Kay, Christopher Nevison, Leon Schram, Mark Stehlik, and Owen Astrachan. The first-course conundrum, why change? Communications of the ACM, 38 (6):117–118, June 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/203270.html.

Neumann:1995:CVE

[860] Peter G. Neumann. Computer vulnerabilities: exploitation or avoidance. Communications of the ACM, 38(6):138, June 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/203273.html.

Nosek:1995:VMG

[861] John Nosek and Garth Shephard. Viewpoint: Making groupware payoff: The British model. Communications of the ACM, 38(7):11– 13, July 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/214784.html.

Press:1995:PCM

[862] Larry Press. Personal computing: McLuhan meets the Net. Communications of the ACM, 38(7):15– 20, July 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/214787.html.

Weingarten:1995:WHS

[863] Fred Weingarten. From Washington: The HPCC: Still a Percheron horse. Communications of the ACM, 38(7):21–23, July 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/214789.html.

Compeau:1995:EUT

[864] Deborah Compeau, Lorne Olfman, Maung Sei, and Jane Webster. Enduser training and learning. Communications of the ACM, 38(7):24– 26, July 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/214791.html.

Nelson:1995:AEU

[865] R. Ryan Nelson, Ellen M. Whitener, and Henry H. Philcox. The assessment of end-user training needs. Communications of the ACM, 38(7):27– 39, July 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/214793.html.

Carroll:1995:MEG

[866] John M. Carroll and Mary Beth Rosson. Managing evaluation goals for training. Communications of the ACM, 38(7):40–48, July 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/214796.html.

Fitzgerald:1995:CTB

[867] Edmond P. Fitzgerald and Aileen Cater-Steel. Champagne training on a beer budget. Communications of the ACM, 38(7):49–60, July 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/214798.html.

Kay:1995:SLT

[868] Judy Kay and Richard C. Thomas. Studying long-term system use. Communications of the ACM, 38(7):61– 69, July 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/214799.html.

Galletta:1995:SIE

[869] Dennis F. Galletta, Manju Ahuja, Amir Hartman, Thompson Teo, and A. Graham Peace. Social influence and enduser training. Communications of the ACM, 38(7):70–79, July 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/214800.html.

Morrison:1995:ESE

[870] Joline Morrison and Joey F. George. Exploring the software engineering component in MIS research. Communications of the ACM, 38(7):80– 91, July 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/214802.html.

Newman:1995:OAP

[871] Gary Newman. Organizing arrays for paged memory systems. Communications of the ACM, 38(7):93– 103, 108–110, July 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/214804.html.

Agre:1995:IRM

[872] Philip E. Agre. Inside risks: My top 10 Email hassles. Communications of the ACM, 38(7):122, July 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/214806.html.

Samuelson:1995:LSS

[873] Pamela Samuelson. Legally speaking: Software compatibility and the law. Communications of the ACM, 38(8):15–22, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/208373.html.

Meeks:1995:EFT

[874] Brock N. Meeks. Electronic frontier: Target: Internet. Communications of the ACM, 38(8):23–25, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/208375.html.

Bieber:1995:DHA

[875] Michael Bieber and Tomás Isakowitz. Designing hypermedia applications. Communications of the ACM, 38 (8):26–29, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/208345.html.

Engelbart:1995:TAH

[876] Douglas C. Engelbart. Toward augmenting the human intellect and boosting our collective IQ. Communications of the ACM, 38(8):30, 32–33, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/208352.html.

Nelson:1995:HCH

[877] Theodor Holm Nelson. The heart of connection: hypermedia unified by transclusion. Communications of the ACM, 38(8):31–33, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/208353.html.

Isakowitz:1995:RMS

[878] Tomás Isakowitz, Edward A. Stohr, and P. Balasubramanian. RMM: a methodology for structured hypermedia design. Communications of the ACM, 38(8):34–44, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/208346.html.

Schwabe:1995:OOH

[879] Daniel Schwabe and Gustavo Rossi. The object-oriented hypermedia design model. Communications of the ACM, 38(8):45–46, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/208354.html.

Balasubramanian:1995:SAD

[880] V. Balasubramanian, Bang Min Ma, and Joonhee Yoo. A systematic approach to designing a WWW application. Communications of the ACM, 38(8):47–48, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/208355.html.

Nanard:1995:HDE

[881] Jocelyne Nanard and Marc Nanard. Hypertext design environments and the hypertext design process. Communications of the ACM, 38(8):49– 56, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/208347.html.

Thuring:1995:HCD

[882] Manfred Thüring, Jörg Hannemann, and Jörg M. Haake. Hypermedia and cognition: designing for comprehension. Communications of the ACM, 38(8):57–66, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/208348.html.

Kahn:1995:VCL

[883] Paul Kahn. Visual cues for local and global coherence in the WWW. Communications of the ACM, 38 (8):67–69, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/208356.html.

Streitz:1995:DHC

[884] Norbert A. Streitz. Designing hypermedia: a collaborative activity.

Communications of the ACM, 38 (8):70–71, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/208357.html.

Schnase:1995:SPC

[885] John L. Schnase, Edward L. Cunnius, and S. Bruce Dowton. The StudySpace project: Collaborative hypermedia in nomadic computing environments. Communications of the ACM, 38(8):72–73, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/208363.html.

Garzotto:1995:HDA

[886] Franca Garzotto, Luca Mainetti, and Paolo Paolini. Hypermedia design, analysis, and evaluation issues. Communications of the ACM, 38(8):74– 86, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/208349.html.

Johnson:1995:CHC

[887] Scott Johnson. Control for hypertext construction. Communications of the ACM, 38(8):87, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/208365.html.

Marshall:1995:SHD

[888] Catherine C. Marshall and Frank M. Shipman, III. Spatial hypertext: designing for change. Communications of the ACM, 38(8):88–97, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/208350.html.

Dieberger:1995:DHS

[889] Andreas Dieberger and Jay D. Bolter. On the design of hyper “spaces”. Communications of the ACM, 38(8):98, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/208366.html.

Bieber:1995:DHS

[890] Michael Bieber and Charles Kacmar. Designing hypertext support for computational applications. Communications of the ACM, 38(8):99–107, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/208351.html.

Davis:1995:EEL

[891] Hugh Davis. To embed or not to embed . . .. Communications of the ACM, 38(8):108–109, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/208367.html.

Wiil:1995:HRP

[892] Uffe K. Wiil. Hyperform: Rapid prototyping of hypermedia services. Communications of the ACM, 38(8): 109–111, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/208369.html.

Furuta:1995:DHA

[893] Richard Furuta and David Stotts. Dynamic hyperdocuments: authoring replaces programming. Communications of the ACM, 38(8):111–112, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/208370.html.

Snyder:1995:IRR

[894] Joel M. Snyder. Inside risks: Research on the Internet. Communications of the ACM, 38(8):130, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/208377.html.

Snizek:1995:VOS

[895] William E. Snizek. Virtual offices: some neglected considerations. Communications of the ACM, 38(9): 15–17, September 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/223252.html.

Wilkes:1995:TPE

[896] Maurice V. Wilkes. A tribute to Presper Eckert. Communications of the ACM, 38(9):20–22, September 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/223253.html.

Baeza-Yates:1995:IPC

[897] R. A. Baeza-Yates, D. A. Fuller, J. A. Pino, and S. E. Goodman. International perspectives: Computing in Chile: The jaguar of the Pacific Rim? Communications of the ACM, 38(9): 23–28, September 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/223254.html.

Weingarten:1995:WOA

[898] Fred W. Weingarten. From Washington: Obituary for an agency. Communications of the ACM, 38(9): 29–32, September 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/223256.html.

Suchman:1995:RW

[899] Lucy Suchman. Representations of work. Communications of the ACM, 38(9):33–35, September 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/223257.html.

Sachs:1995:TWC

[900] Patricia Sachs. Transforming work: Collaboration, learning, and design. Communications of the ACM, 38(9): 36–44, September 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/223258.html.

Star:1995:WI

[901] Susan Leigh Star and Geoffrey C. Bowker. Work and infrastructure. Communications of the ACM, 38(9):41, September 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jordan:1995:TW

[902] Brigitte Jordan, Ron Goldman, and Patricia Sachs. Tools for the workplace. Communications of the ACM, 38(9):42, September 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wild:1995:RC

[903] Helga Wild, Chris Darrouzet, Ted Kahn, and Susan U. Stucky. Rhythms of collaboration. Communications of the ACM, 38(9):45, September 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/223260.html.

Kyng:1995:MRW

[904] Morten Kyng. Making representations work. Communications of the ACM, 38(9):46–55, September 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/223261.html.

Suchman:1995:MWV

[905] Lucy Suchman. Making work visible. Communications of the ACM, 38 (9):56–64, September 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/223263.html.

Hall:1995:EDO

[906] Rogers Hall. Exploring design oriented mathematical practices in school and work settings. Communications of the ACM, 38(9):62–64, September 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/223264.html.

Muller:1995:EQW

[907] Michael J. Muller. Ethnocritical questions for working with translations, interpretation and the stakeholders. Communications of the ACM, 38(9): 64–65, September 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/223265.html.

Bannon:1995:PDR

[908] Liam J. Bannon. The politics of design: representing work. Communications of the ACM, 38(9):66–68, September 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/223267.html.

Klepper:1995:TFG

[909] Robert Klepper and Douglas Bock. Third and fourth generation language productivity differences. Communications of the ACM, 38(9):69–79, September 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/223268.html.

Papazoglou:1995:USC

[910] M. P. Papazoglou. Unraveling the semantics of conceptual schemas. Communications of the ACM, 38(9): 80–94, September 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/223275.html.

Ram:1995:DFD

[911] Sudha Ram. Deriving functional dependencies from the entity-relationship model. Communications of the ACM, 38(9):95–107, September 1995. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/toc/Abstracts/
0001-0782/223277.html.

Neumann:1995:IRRa

[912] Peter G. Neumann. Inside risks: Risks of easy answers. Communications of the ACM, 38(9):130, September 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zweben:1995:PLW

[913] Stuart Zweben. President’s letter: Widening international boundaries. Communications of the ACM, 38(10):15–16, October 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rada:1995:SSC

[914] Roy Rada. Sharing standards: Consensus versus speed. Communications of the ACM, 38(10):21–23, October 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Press:1995:PCN

[915] Larry Press. Personal computing: Net.Speech: Audio comes to the net. Communications of the ACM, 38 (10):25–31, October 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pancake:1995:PCO

[916] Cherri M. Pancake. The promise and the cost of object technology: A five-year forecast. Communications of the ACM, 38(10):33–49, October 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fayad:1995:OOE

[917] Mohamed E. Fayad and Wei-Tek Tsai. Object-oriented experiences. Communications of the ACM, 38(10):50–53, October 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berg:1995:LLO

[918] William Berg, Marshall Cline, and Mike Girou. Lessons learned from the OS/400 OO Project. Communications of the ACM, 38(10):54–64, October 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schmidt:1995:UDP

[919] Douglas C. Schmidt. Using design patterns to develop reusable object-oriented communication software. Communications of the ACM, 38(10):65–74, October 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kung:1995:DOO

[920] David Kung, Jerry Gao, Pei Hsia, Yasufumi Toyoshima, Chris Chen, Young-Si Kim, and Young-Kee Song. Developing an object-oriented software testing and maintenance environment. Communications of the ACM, 38(10):75– 87, October 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Norton:1995:OOP

[921] Charles D. Norton, Boleslaw K. Szymanski, and Viktor K. Decyk. Objectoriented parallel computation for plasma simulation. Communications of the ACM, 38(10):88–100, October 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shan:1995:ISR

[922] Yen-Ping Shan. Introduction: Smalltalk on the rise. Communications of the ACM, 38(10):102–104, October 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goldberg:1995:WS

[923] Adele Goldberg. Why Smalltalk? Communications of the ACM, 38(10): 105–107, October 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Adamczyk:1995:SRC

[924] Jim Adamczyk. Smalltalk reaches crossroads in the insurance industry. Communications of the ACM, 38(10): 107–109, October 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Grochow:1995:STI

[925] Jerrold M. Grochow. Smalltalk in the telecommunications industry. Communications of the ACM, 38(10):110– 111, October 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Thomas:1995:UAE

[926] Dave Thomas. Ubiquitous applications: Embedded systems to mainframe. Communications of the ACM, 38(10):112–114, October 1995. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Auer:1995:STI

[927] Ken Auer. Smalltalk training: As innovative as the environment. Communications of the ACM, 38(10):115– 117, October 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Garfinkel:1995:IRR

[928] Simson L. Garfinkel. Inside risks: Risks of social security numbers. Communications of the ACM, 38(10):146, October 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rachlin:1995:VLL

[929] Robert D. Rachlin. Viewpoint: Learning from litigation: Trade secret misappropriation. Communications of the ACM, 38(11):15–17, November 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/219720.html.

Shneiderman:1995:LEW

[930] Ben Shneiderman, Maryann Alavi, Kent Norman, and Ellen Yu Borkowski. Log on education: Windows of opportunity in electronic classrooms. Communications of the ACM, 38(11): 19–24, November 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/219725.html.

Berghel:1995:DVM

[931] Hal Berghel. Digital village: Maiden voyage. Communications of the ACM, 38(11):25–27, November 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/223056.html.

Munakata:1995:INH

[932] Toshinori Munakata. Introduction: New horizons in commercial and industrial AI. Communications of the ACM, 38(11):28–31, November 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/219734.html.

Lenat:1995:CLS

[933] Douglas B. Lenat. CYC: a large-scale investment in knowledge infrastructure. Communications of the ACM, 38 (11):33–38, November 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/219745.html.

Miller:1995:WLD

[934] George A. Miller. WordNet: a lexical database for English. Communications of the ACM, 38(11): 39–41, November 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/219748.html.

Yokoi:1995:EED

[935] Toshio Yokoi. The EDR electronic dictionary. Communications of the ACM, 38(11):42–44, November 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/219752.html.

Lenat:1995:CWE

[936] Douglas B. Lenat, George A. Miller, and Toshio Yokoi. CYC, WordNet, and EDR: Critiques and responses. Communications of the ACM, 38(11): 45–48, November 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/219757.html.

Chaib-draa:1995:IAD

[937] B. Chaib-draa. Industrial applications of distributed AI. Communications of the ACM, 38(11): 49–53, November 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/219761.html.

Langley:1995:AML

[938] Pat Langley and Herbert A. Simon. Applications of machine learning and rule induction. Communications of the ACM, 38(11):54–64, November 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/219768.html.

Bratko:1995:AIL

[939] Ivan Bratko and Stephen Muggleton. Applications of inductive logic programming. Communications of the ACM, 38(11):65–70, November 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/219771.html.

Church:1995:CAN

[940] Kenneth W. Church and Lisa F. Rau. Commercial applications of natural language processing. Communications of the ACM, 38(11): 71–79, November 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/219778.html.

Harmon:1995:OOA

[941] Paul Harmon. Object-oriented AI: a commercial perspective. Communications of the ACM, 38(11): 80–86, November 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/219783.html.

Pawlak:1995:RS

[942] Zdzislaw Pawlak, Jerzy Grzymala-Busse, Roman Slowinski, and Wojciech Ziarko. Rough sets. Communications of the ACM, 38(11): 88–95, November 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/219791.html.

Ditto:1995:PAC

[943] William Ditto and Toshinori Munakata. Principles and applications of chaotic systems. Communications of the ACM, 38(11): 96–102, November 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/219797.html.

Aihara:1995:CEJ

[944] Kazuyuki Aihara and Ryu Katayama. Chaos engineering in Japan. Communications of the ACM, 38(11): 103–107, November 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/219801.html.

Maes:1995:ALM

[945] Pattie Maes. Artificial life meets enfertainment: Lifelike autonomous agents. Communications of the ACM, 38(11): 108–114, November 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/219808.html.

Coldwell:1995:AAT

[946] R. A. Coldwell. Australian attitudes toward legal intervention into hacking. Communications of the ACM, 38(11):

115–117, November 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/219812.html.

Leveson:1995:IRS

[947] Nancy G. Leveson. Inside risks: Safety as a system property. Communications of the ACM, 38(11):146, November 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/219816.html.

Lynch:1995:GPL

[948] Dan Lynch and Bert Herzog. Guest President’s letter: a time for celebration. Communications of the ACM, 38 (12):11–12, December 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/219665.html.

Zurk:1995:VWA

[949] Lisa M. Zurk, Barbara Simons, Rebecca Parsons, and Dawn Cohen. Viewpoint: When an advantage is not an advantage. Communications of the ACM, 38(12):17–18, December 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/219667.html.

Barlow:1995:EFP

[950] John Perry Barlow. Electronic frontier: Property and speech: Who owns what you say in cyberspace? Communications of the ACM, 38(12):

19–22, December 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/219673.html.

Danowitz:1995:IPC

[951] A. K. Danowitz, Y. Nassef, and S. E. Goodman. International perspectives: Cyberspace across the Sahara: Computing in North Africa. Communications of the ACM, 38(12): 23–28, December 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/219674.html.

Denning:1995:EPW

[952] Peter J. Denning. Editorial: Plagiarism in the Web. Communications of the ACM, 38(12):29, December 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/219675.html.

Conger:1995:ECU

[953] Sue Conger and Karen D. Loch. Ethics and computer use. Communications of the ACM, 38(12): 30–32, December 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/219676.html.

Laudon:1995:ECI

[954] Kenneth C. Laudon. Ethical concepts and information technology. Communications of the ACM, 38(12): 33–39, December 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/219677.html.

Weisband:1995:MUP

[955] Suzanne P. Weisband and Bruce A. Reinig. Managing user perceptions of Email privacy. Communications of the ACM, 38(12):40–47, December 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/219678.html.

Sipior:1995:ELQ

[956] Janice C. Sipior and Burke T. Ward. The ethical and legal quandary of Email privacy. Communications of the ACM, 38(12):48–54, December 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/219679.html.

Mason:1995:AEI

[957] Richard O. Mason. Applying ethics to information technology issues. Communications of the ACM, 38(12): 55–57, December 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/219681.html.

Johnson:1995:ACD

[958] Deborah G. Johnson and John M. Mulvey. Accountability and computer decision systems. Communications of the ACM, 38(12):58–64, December 1995. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/219682.html.

Milberg:1995:VPI

[959] Sandra J. Milberg, Sandra J. Burke, H. Jeff Smith, and Ernest A. Kallman. Values, personal information privacy, and regulatory approaches. Communications of the ACM, 38(12): 65–74, December 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/219683.html.

Huff:1995:CCF

[960] Chuck Huff and C. Dianne Martin. Computing consequences: a framework for teaching ethical computing. Communications of the ACM, 38 (12):75–84, December 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/219687.html.

Anonymous:1995:CAS

[961] Anonymous. Communications of the ACM 1995 subject index. Communications of the ACM, 38(12): 85–90, December 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/219690.html.

Anonymous:1995:CAA

[962] Anonymous. Communications of the ACM 1995 author index. Communications of the ACM, 38(12): 91–99, December 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/219693.html.

Neumann:1995:IRRb

[963] Peter G. Neumann. Inside risks: Reviewing the risks archives. Communications of the ACM, 38(12):138, December 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/219694.html.

Fox:1996:N

[964] Robert Fox. Newstrack. Communications of the ACM, 39(1):9–10, January 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zweben:1996:PLC

[965] Stuart Zweben. President’s letter: Computing week ’96. Communications of the ACM, 39(1):11–12, January 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Erickson:1996:VWW

[966] Thomas Erickson. Viewpoint: The World-Wide Web as social hypertext. Communications of the ACM, 39(1):15–17, January 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/227582.html.

Rada:1996:SSW

[967] Roy Rada. Sharing standards: Who will test conformance? Communications of the ACM, 39(1):19–22, January 1996. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/227583.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
234175.html.

Samuelson:1996:LSI

[968] Pamela Samuelson. Legally speaking: Intellectual property rights and the global information economy. Communications of the ACM, 39(1):23–28, January 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/227584.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
234176.html.

Berghel:1996:CSW

[969] Hal Berghel. The client’s side of the World-Wide Web. Communications of the ACM, 39(1):30–40, January 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/227585.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
234177.html.

Haag:1996:QFD

[970] Stephen Haag, M. K. Raja, and L. L. Schkade. Qaulity function deployment: Usage in software development. Communications of the ACM, 39(1):41–49, January 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/227586.html;   http:/
/www.acm.org/pubs/toc/Abstracts/
cacm/234178.html.

Cheyney:1996:TMC

[971] Matthew Cheyney, Peter Gloor, Donald B. Johnson, Fillia Makedon, James Matthews, and Panagiotis Takis Metaxas. Toward multimedia conference proceedings. Communications of the ACM, 39(1):50–59, January 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/227587.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
234179.html.

Wilks:1996:NLP

[972] Yorick Wilks. Natural language processing. Communications of the ACM, 39(1):60–62, January 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/227588.html;   http:/
/www.acm.org/pubs/toc/Abstracts/
cacm/234180.html.

Guthrie:1996:RLN

[973] Louise Guthrie, James Pustejovsky, Yorick Wilks, and Brian M. Slator. The role of Lexicons in natural language processing. Communications of the ACM, 39(1):63–72, January 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/227589.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
234204.html.

King:1996:ENL

[974] Margaret King. Evaluating natural language processing systems. Communications of the ACM, 39(1):73–79, January 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/227590.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
234208.html.

Cowie:1996:IE

[975] Jim Cowie and Wendy Lehnert. Information extraction. Communications of the ACM, 39(1):80–91, January 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/227591.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
234209.html.

Lewis:1996:NLP

[976] David D. Lewis and Karen Sparck Jones. Natural language processing for information retrieval. Communications of the ACM, 39(1):92–101, January 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/227592.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
234210.html.

Wiebe:1996:LUC

[977] Janyce Wiebe, Graeme Hirst, and Diane Horton. Language use in context. Communications of the ACM, 39 (1):102–111, January 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/227593.html;   http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/234212.html.

Neumann:1996:RDC

[978] Peter G. Neumann. Risks in digital commerce. Communications of the ACM, 39(1):154, January 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/toc/Abstracts/cacm/234214.
html.

Alt:1996:FAR

[979] Franz Alt. Franz Alt remembers the early years of computing and the creation of ACM. Communications of the ACM, 39(2 (supplement)): 1, 5, 6, February 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). ACM Member-Net.

Anonymous:1996:CWU

[980] Anonymous. Computing week 96 ushers in ACM’s 50th anniversary. Communications of the ACM, 39(2 (supplement)):1–2, February 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). ACM Member-Net.

Anonymous:1996:FTI

[981] Anonymous. Finals teams: 1996 international collegiate programming contest. Communications of the ACM, 39(2 (supplement)):3, February 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). ACM MemberNet.

Anonymous:1996:EFD

[982] Anonymous. Entering its fourth decade, SIGPLAN looks ahead. Communications of the ACM, 39(2 (supplement)):4, February 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). ACM Member-Net.

Anonymous:1996:AGD

[983] Anonymous. ACM E-guide to debut this spring. two new books from ACM Press books. Communications of the ACM, 39(2 (supplement)):7, February 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). ACM MemberNet.

Anonymous:1996:NLJ

[984] Anonymous. New on-line journal. TO-DAES begins publications. new focus added: Asia Pacific region. Communications of the ACM, 39(2 (supplement)):8, February 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). ACM Member-Net.

Johnson:1996:VIS

[985] Jeff Johnson. Viewpoint: The information superhighway: A worst-case scenario. Communications of the ACM, 39(2):15–17, February 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:1996:DVH

[986] Hal Berghel. Digital village: HTML compliance and the return of the test pattern. Communications of the ACM, 39(2):19–22, February 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/230799.html.

Press:1996:PCR

[987] Larry Press. Personal computing: The role of computer networks in development. Communications of the ACM, 39(2):23–30, February 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/230800.html.

Woxny:1996:CSM

[988] Michael J. Woxny and William C. Regli. Computer science in manufacturing. Communications of the ACM, 39(2):32–33, February 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/230801.html.

Wozny:1996:CSM

[989] Michael J. Wozny and William C. Regli. Computer science in manufacturing. Communications of the ACM, 39(2):33, February 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Erkes:1996:ISM

[990] J. W. Erkes, K. B. Kenny, J. W. Lewis, B. D. Sarachan, M. W. Sobolewski, and R. N. Sum, Jr. Implementing shared manufacturing services on the World-Wide Web. Communications of the ACM, 39(2):34–45, February 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/230802.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
230802.html.

Hardwick:1996:SMI

[991] Martin Hardwick, David L. Spooner, Tom Rando, and K. C. Morris. Sharing manufacturing information in virtual enterprises. Communications of the ACM, 39(2):46–54, February 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/230803.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
230803.html.

Kumar:1996:RDP

[992] Vijay Kumar, Ruzena Bajcsy, William Harwin, and Patrick Harker. Rapid design and prototyping of customized rehabilitation aids. Communications of the ACM, 39(2):55–61, February 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/230804.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
230804.html.

Bekey:1996:CTR

[993] George A. Bekey. Commentary: Trends in robotics. Communications of the ACM, 39(2):62, February 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/230805.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
230805.html.

Finger:1996:RDM

[994] S. Finger, M. Terk, E. Subrahmanian, C. Kasabach, F. Prinz, D. P.

Siewiorek, A. Smailagic, J. Stivoric, and L. Weiss. Rapid design and manufacture of wearable computers. Communications of the ACM, 39(2): 63–70, February 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/230806.html;   http:/
/www.acm.org/pubs/toc/Abstracts/
cacm/230806.html.

Requicha:1996:GRI

[995] Aristides A. G. Requicha. Geometric reasoning for intelligent manufacturing. Communications of the ACM, 39(2):71–76, February 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/230807.html;   http:/
/www.acm.org/pubs/toc/Abstracts/
cacm/230807.html.

Mantyla:1996:CFB

[996] Martti Mäntylä, Dana Nau, and Jami Shah. Challenges in feature-based manufacturing research. Communications of the ACM, 39(2):77–85, February 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/230808.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
230808.html.

Bernstein:1996:MMD

[997] Philip A. Bernstein. Middleware: a model for distributed system services. Communications of the ACM, 39(2):86–98, February 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/230809.html;   http:/
                                        /www.acm.org/pubs/toc/Abstracts/
                                        cacm/230809.html.

Candler:1996:OPS

[998] James W. Candler, Prashant C. Palvia, Jane D. Thompson, and Steven M. Zeltmann. The ORION Project: Staged business process reengineering at FedEx. Communications of the ACM, 39(2):99–107, February 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/230810.html;   http://www.
                                        acm.org/pubs/toc/Abstracts/cacm/
                                        230810.html.

Fayad:1996:TOO

[999] Mohamed E. Fayad, Wei-Tek Tsai, and Milton L. Fulghum. Transition to object-oriented software development. Communications of the ACM, 39 (2):108–121, February 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/230811.html;   http:/
                                        /www.acm.org/pubs/toc/Abstracts/
                                        cacm/230811.html.

Neumann:1996:WHI

[1000] Peter G. Neumann and Peter J. Denning. W(h)ither research and education? Communications of the ACM, 39(2):154, February 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/230813.html;   http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/230813.html.

Bell:1996:VOS

[1001] Gordon Bell. Viewpoint: 1995 observations on supercomputing alternatives: Did the MPP bandwagon lead to a culde-sac? Communications of the ACM, 39(3):11–15, March 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/227235.html.

Rada:1996:SSI

[1002] Roy Rada. Sharing standards: ISO 9000 reflects the best in standards. Communications of the ACM, 39 (3):17–20, March 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/227236.html.

Davis:1996:NVI

[1003] Randall Davis, Pamela Samuelson, Mitchell Kapor, and Jerome Reichman. A new view of intellectual property and software. Communications of the ACM, 39(3):21–30, March 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/227237.html;   http:/
/www.acm.org/pubs/toc/Abstracts/
cacm/227237.html.

Ganesan:1996:HUK

[1004] Ravi Ganesan. How to use key escrow. Communications of the ACM, 39(3):33, March 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/227238.html;   http:/
/www.acm.org/pubs/toc/Abstracts/
cacm/227238.html.

Denning:1996:TKE

[1005] Dorothy E. Denning and Dennis K. Branstad. A taxonomy for key escrow encryption systems. Communications of the ACM, 39(3):34–40, March 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/227239.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
227239.html.

Walker:1996:CKR

[1006] Stephen T. Walker, Steven B. Lipner, Carl M. Ellison, and David M. Balenson. Commercial key recovery. Communications of the ACM, 39(3):41–47, March 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/227240.html;   http:/
/www.acm.org/pubs/toc/Abstracts/
cacm/227240.html.

Maher:1996:CBK

[1007] David Paul Maher. Crypto backup and key escrow. Communications of the ACM, 39(3):48–53, March 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/227241.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
227241.html.

Ganesan:1996:YSS

[1008] Ravi Ganesan. The Yaksha security system. Communications of the ACM, 39(3):55–60, March 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/227242.html;   http:/
/www.acm.org/pubs/toc/Abstracts/
cacm/227242.html.

Brooks:1996:CST

[1009] Frederick P. Brooks, Jr. The computer scientist as toolsmith II. Communications of the ACM, 39(3):61– 68, March 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/227243.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
227243.html.

Wood-Harper:1996:HWP

[1010] A. T. Wood-Harper, Steve Corder, J. R. G. Wood, and Heather Watson. How we profess: The ethical systems analyst. Communications of the ACM, 39(3):69–77, March 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/227244.html;   http:/
/www.acm.org/pubs/toc/Abstracts/
cacm/227244.html.

Ralston:1996:DCF

[1011] Anthony Ralston. The demographics of candidates for faculty positions in computer science. Communications of the ACM, 39(3):78– 84, March 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/227245.html;   http://www.
                                        acm.org/pubs/toc/Abstracts/cacm/
                                        227245.html.

Blelloch:1996:PPA

[1012] Guy E. Blelloch. Programming parallel algorithms. Communications of the ACM, 39(3):85–97, March 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/227246.html;   http://www.
                                        acm.org/pubs/toc/Abstracts/cacm/
                                        227246.html.

Anonymous:1996:IAa

[1013] Anonymous. Information for authors. Communications of the ACM, 39(3):98–99, March 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goldschmidt:1996:TOR

[1014] Art Goldschmidt. Technical opinion: Report on NIIIP. Communications of the ACM, 39(3):100–103, March 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/227247.html.

Charette:1996:TRO

[1015] Robert N. Charette. Taking responsibility for our risks. Communications of the ACM, 39(3):122, March 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/227599.html;   http://www.

acm.org/pubs/toc/Abstracts/cacm/
227599.html.

Myers:1996:IPC

[1016] Michael D. Myers. International perspectives: Can Kiwis fly? computing in New Zealand. Communications of the ACM, 39(4):11–15, April 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/227212.html.

Soloway:1996:LEN

[1017] Elliot Soloway and Amanda Pryor.

Log on education: The next generation in human-computer interaction. Communications of the ACM, 39(4):16–18, April 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/227213.html.

Norman:1996:LCE

[1018] Donald A. Norman and James C.

Spohrer. Learner-centered education. Communications of the ACM, 39(4):24–27, April 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/227215.html;   http:/
/www.acm.org/pubs/toc/Abstracts/
cacm/227215.html.

Schank:1996:GBS

[1019] Roger C. Schank and Alex Kass. A goal-based scenario for high school students. Communications of the ACM, 39(4):28–29, April 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/227216.html;   http:/
                                        /www.acm.org/pubs/toc/Abstracts/
                                        cacm/227216.html.

Woolf:1996:IMT

[1020] Beverly Park Woolf. Intelligent multimedia tutoring systems. Communications of the ACM, 39(4):30– 31, April 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/227217.html.

Edelson:1996:CN

[1021] Daniel C. Edelson, Roy D. Pea, and Louis M. Gomez. The Collaboratory Notebook. Communications of the ACM, 39(4):32–33, April 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/227218.html;   http://www.
                                        acm.org/pubs/toc/Abstracts/cacm/
                                        227218.html.

Linn:1996:KIH

[1022] Marcia C. Linn. Key to the information highway. Communications of the ACM, 39(4):34–35, April 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/227219.html;   http:/
                                        /www.acm.org/pubs/toc/Abstracts/
                                        cacm/227219.html.

Scardamalia:1996:SCA

[1023] Marlene Scardamalia and Carl Bereiter. Student communities for the advancement of knowledge. Communications of the ACM, 39(4):36– 37, April 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/227220.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
227220.html.

Kafai:1996:SKK

[1024] Yasmin B. Kafai. Software by kids for kids. Communications of the ACM, 39(4):38–39, April 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/227221.html;   http:/
/www.acm.org/pubs/toc/Abstracts/
cacm/227221.html.

Eden:1996:MLP

[1025] Hal Eden, Mike Eisenberg, Gerhard Fischer, and Alexander Repenning. Making learning a part of life. Communications of the ACM, 39(4):40–42, April 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/227222.html;   http:/
/www.acm.org/pubs/toc/Abstracts/
cacm/227222.html.

Guzdial:1996:CSL

[1026] M. Guzdial, J. Kolodner, C. Hmelo, H. Narayanan, D. Carlson, N. Rappin, R. Hübscher, J. Turns, and W. Newstetter. Computer support for learning through complex problem solving. Communications of the ACM, 39(4):43–45, April 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/227600.html;   http:/
                                        /www.acm.org/pubs/toc/Abstracts/
                                        cacm/227600.html.

Rosson:1996:SEL

[1027] Mary Beth Rosson and John M. Carroll. Scaffolded examples for learning object-oriented design. Communications of the ACM, 39(4):46– 47, April 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/227223.html;   http://www.
                                        acm.org/pubs/toc/Abstracts/cacm/
                                        227223.html.

Jackson:1996:LCT

[1028] Shari L. Jackson, Steven J. Stratford, Joseph Krajcik, and Elliot Soloway. A learner-centered tool for students building models. Communications of the ACM, 39(4):48– 49, April 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/227224.html;   http://www.
                                        acm.org/pubs/toc/Abstracts/cacm/
                                        227224.html.

Powell:1996:GC

[1029] David Powell. Group communication. Communications of the ACM, 39(4):50–53, April 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/227225.html;   http:/
                                        /www.acm.org/pubs/toc/Abstracts/
                                        cacm/227225.html.

Moser:1996:TFT

[1030] Louise E. Moser, P. M. Melliar-Smith, Deborah A. Agarwal, Ravi K. Budhia, and Colleen A. Lingley-Papadopoulos. Totem: a fault-tolerant multicast group communication system. Communications of the ACM, 39(4):54–63, April 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/227226.html;   http:/
/www.acm.org/pubs/toc/Abstracts/
cacm/227226.html.

Dolev:1996:TAH

[1031] Danny Dolev and Dalia Malki. The Transis approach to high availability cluster communication. Communications of the ACM, 39(4):64– 70, April 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/227227.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
227227.html.

Reiter:1996:DTR

[1032] Michael K. Reiter. Distributing trust with the Rampart Toolkit. Communications of the ACM, 39(4):71– 74, April 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/227228.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
227228.html.

vanRenesse:1996:HFG

[1033] Robbert van Renesse, Kenneth P. Birman, and Silvano Maffeis. Horus: a flexible group communication system. Communications of the ACM, 39(4):76–83, April 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        cacm/227229.html.

Schiper:1996:GCT

[1034] André Schiper and Michael Raynat. From group communication to transactions in distributed systems. Communications of the ACM, 39(4):84–87, April 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/227230.html;   http:/
                                        /www.acm.org/pubs/toc/Abstracts/
                                        cacm/227230.html.

Cristian:1996:SAG

[1035] Flaviu Cristian. Synchronous and asynchronous group communication. Communications of the ACM, 39 (4):88–97, April 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        0001-0782/227231.html.

Anonymous:1996:IAb

[1036] Anonymous. Information for authors. Communications of the ACM, 39(4):98–99, April 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:1996:CE

[1037] Anonymous. Calendar of events. Communications of the ACM, 39(4):100– 101, 107–109, April 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:1996:IRRa

[1038] Peter G. Neumann. Inside risks: a risks related bookshelf. Communications of the ACM, 39(4):122, April 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:1996:NTa

[1039] Robert Fox. News track. Communications of the ACM, 39(5):11–12, May 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Katz:1996:VPK

[1040] Randy H. Katz. Viewpoint: Professor Katz goes to Washington. Communications of the ACM, 39(5):13–17, May 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Press:1996:PCW

[1041] Larry Press. Personal computing:

Windows NT as a personal or intranet server. Communications of the ACM, 39(5):19–23, May 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/229460.html.

Meeks:1996:EFG

[1042] Brock N. Meeks. Electronic frontier: The game plan to save cyberspace. Communications of the ACM, 39(5):24–26, May 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/229461.html.

Denning:1996:UNC

[1043] Peter J. Denning. The university’s next challenges. Communications of the ACM, 39(5):27–31, May 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/229462.html.

Singh:1996:VRS

[1044] Gurminder Singh, Steven K. Feiner, and Daniel Thalmann. Virtual reality: software and technology. Communications of the ACM, 39(5):35–36, May 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/toc/Abstracts/cacm/229463.
html.

Poston:1996:DVW

[1045] Timothy Poston and Luis Serra. Dextrous virtual work. Communications of the ACM, 39(5):37–45, May 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/toc/Abstracts/cacm/229464.
html.

Green:1996:GMA

[1046] Mark Green and Sean Halliday. A geometric modeling and animation system for virtual reality. Communications of the ACM, 39(5):46–53, May 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/toc/Abstracts/cacm/229465.
html.

Deering:1996:HVS

[1047] Michael F. Deering. The HoloSketch VR sketching system. Communications of the ACM, 39(5):54–61, May 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/toc/Abstracts/cacm/229466.
html.

Bryson:1996:VRS

[1048] Steve Bryson. Virtual reality in scientific visualization. Communications of the ACM, 39(5):62–71, May 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/toc/Abstracts/cacm/229467.
html.

Bayarri:1996:VRD

[1049] Salvador Bayarri, Marcos Fernandez, and Mariano Perez. Virtual reality for driving simulation. Communications of the ACM, 39(5):72–76, May 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/toc/Abstracts/cacm/229468.
html.

Fagin:1996:CIL

[1050] Ronald Fagin, Moni Naor, and Peter Winkler. Comparing information without leaking it. Communications of the ACM, 39(5):77–85, May 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/229469.html.

Lucas:1996:RFE

[1051] Henry C. Lucas, Jr., Donald J. Berndt, and Greg Truman. A reengineering framework for evaluating a financial imaging system. Communications of the ACM, 39(5):86–96, May 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/toc/Abstracts/cacm/229470.
                                        html.

Feigenbaum:1996:TAL

[1052] Edward A. Feigenbaum. Turing Award Lecture: How the “what” becomes the “how”. Communications of the ACM, 39(5):97–104, May 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        cacm/229471.html.

Reddy:1996:TAL

[1053] Raj Reddy. Turing Award Lecture: To dream the possible dream. Communications of the ACM, 39(5):105–112, May 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/toc/Abstracts/cacm/233436.
                                        html.

McLeod:1996:CUC

[1054] Raymond McLeod, Jr. Comparing undergraduate courses in systems analysis and design. Communications of the ACM, 39(5):113–121, May 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/toc/Abstracts/cacm/229472.
                                        html.

Anonymous:1996:APC

[1055] Anonymous. Accredited programs in computing. Communications of the ACM, 39(5):122–127, May 1996. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:1996:AGE

[1056] Anonymous. ACM general elections. Communications of the ACM, 39(5):128–135, May 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:1996:IRL

[1057] Peter G. Neumann. Inside risks: Linguistic risks. Communications of the ACM, 39(5):154, May 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Soloway:1996:LET

[1058] Elliot Soloway. Log on education: Teachers are the key. Communications of the ACM, 39(6):11– 14, June 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/228504.html.

Berghel:1996:DVU

[1059] Hal Berghel. Digital village: U.S.

technology policy in the information age. Communications of the ACM, 39(6):15–18, June 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/228506.html.

Pyle:1996:ECI

[1060] Raymond Pyle. Electronic commerce and the Internet. Communications of the ACM, 39(6):22–23, June 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-

0782/228507.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
228507.html.

Sivori:1996:ERS

[1061] John R. Sivori. Evaluated receipts and settlement at Bell Atlantic. Communications of the ACM, 39(6):24–28, June 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/228508.html.

Bhimani:1996:SCI

[1062] Anish Bhimani. Securing the commercial Internet. Communications of the ACM, 39(6):29–35, June 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/228509.html;   http:/
/www.acm.org/pubs/toc/Abstracts/
cacm/228509.html.

Borenstein:1996:PPP

[1063] Nathaniel S. Borenstein. Perils and pitfalls of practical cybercommerce. Communications of the ACM, 39(6):36–44, June 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/228510.html.

Panurach:1996:MEC

[1064] Patiwat Panurach. Money in electronic commerce: digital cash, electronic fund transfer, and Ecash. Communications of the ACM, 39(6):45– 50, June 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-

0782/228512.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
228512.html.

Hamalainen:1996:EML

[1065] Matti Hämäläinen, Andrew B. Whinston, and Svetlana Vishik. Electronic markets for learning: Education brokerages on the Internet. Communications of the ACM, 39(6):51–58, June 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/228513.html;   http:/
/www.acm.org/pubs/toc/Abstracts/
cacm/228513.html.

Chon:1996:II

[1066] Kilnam Chon. Internet inroads. Communications of the ACM, 39(6):59– 60, June 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/228515.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
228515.html.

Hinden:1996:ING

[1067] Robert M. Hinden. IP Next Generation overview. Communications of the ACM, 39(6):61–71, June 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/228517.html;   http:/
/www.acm.org/pubs/toc/Abstracts/
cacm/228517.html.

Ang:1996:CIS

[1068] Peng Hwa Ang and Berlinda Nadarajan. Censorship and the Internet: A Singapore perspective. Communications of the ACM, 39(6):72– 78, June 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/228520.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
228520.html.

OReilly:1996:PMI

[1069] Tim O’Reilly. Publishing models for Internet commerce. Communications of the ACM, 39(6):79– 86, June 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/228522.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
228522.html.

Kirstein:1996:COP

[1070] Peter Kirstein and Goli Montasser-Kohsari. The C-ODA project: Online access to electronic journals. Communications of the ACM, 39(6):87– 99, June 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/228523.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
228523.html.

Mannoni:1996:BMO

[1071] Bruno Mannoni. Bringing museums online. Communications of the ACM, 39(6):100–105, June 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/228524.html;   http:/
/www.acm.org/pubs/toc/Abstracts/
cacm/228524.html.

Pitkow:1996:ETW

[1072] James E. Pitkow and Colleen M.

Kehoe. Emerging trends in the WWW user population. Communications of the ACM, 39(6):106– 108, June 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/228525.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
228525.html.

Lunt:1996:IRS

[1073] Teresa F. Lunt. Inside risks: Securing the information infrastructure. Communications of the ACM, 39(6):130, June 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Greenberg:1996:IPB

[1074] L. T. Greenberg and S. E. Goodman. International perspectives: Is Big Brother hanging by his bootstraps? Communications of the ACM, 39(7):11–16, July 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/233978.html.

Samuelson:1996:LSR

[1075] Pamela Samuelson. Legally speaking: Regulation of technologies to protect copyrighted works. Communications of the ACM, 39(7):17–24, July 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/toc/Abstracts/cacm/233980.
html.

Grosso:1996:VWW

[1076] Andrew Grosso. Viewpoint: Why won’t the President enforce the Constitution? Communications of the ACM, 39(7):25–26, July 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        cacm/233982.html.

Rao:1996:ISO

[1077] H. Raghav Rao, Kichan Nam, and A. Chaudhury. Information systems outsourcing. Communications of the ACM, 39(7):27–28, July 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        cacm/233984.html.

Nelson:1996:TDS

[1078] Paul Nelson, William Richmond, and Abraham Seidmann. Two dimensions of software acquisition. Communications of the ACM, 39(7):29–35, July 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/toc/Abstracts/cacm/233986.
                                        html.

Nam:1996:TLI

[1079] Kichan Nam, S. Rajagopalan, H. Raghav Rao, and A. Chaudhury. A two-level investigation of information systems outsourcing. Communications of the ACM, 39(7):36–44, July 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        cacm/233989.html.

Gurbaxani:1996:NWI

[1080] Vijay Gurbaxani. The new world of information technology outsourcing. Communications of the ACM, 39(7):45–46, July 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/233992.html.

Slaughter:1996:EOI

[1081] Sandra A. Slaughter and Soon Ang.

Employment outsourcing in information systems. Communications of the ACM, 39(7):47–54, July 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/233994.html.

Bell:1996:RPI

[1082] Gordon Bell and Jim Gemmell. Onramp prospects for the information superhighway dream: will the goal of a single, universally available network carrying a wide variety of data types ever be achieved? Communications of the ACM, 39(7):55–61, July 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/toc/Abstracts/cacm/233995.
html.

Streeter:1996:HOD

[1083] Lynn A. Streeter, Robert E. Kraut, Henry C. Lucas, Jr., and Laurence Caby. How open data networks influence business performance and market structure. Communications of the ACM, 39(7):62–73, July 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/233998.html.

Loch:1996:EED

[1084] Karen D. Loch and Sue Conger. Evaluating ethical decision making and computer use. Communications of the ACM, 39(7):74–83, July 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/233999.html.

Dongarra:1996:MPS

[1085] Jack J. Dongarra, Steve W. Otto, Marc Snir, and David Walker. A message passing standard for MPP and workstations. Communications of the ACM, 39(7):84–90, July 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/234000.html.

Munakata:1996:TDB

[1086] Toshinori Munakata. Thoughts on Deep Blue vs. Kasparov. Communications of the ACM, 39(7):91–92, July 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/toc/Abstracts/cacm/234001.
html.

Neumann:1996:IRU

[1087] Peter G. Neumann. Inside risks: Using formal methods to reduce risks. Communications of the ACM, 39(7):114, July 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/toc/Abstracts/cacm/234595.
html.

Anonymous:1996:NPD

[1088] Anonymous. New publishing directions for ACM. Communications of the ACM, 39(8 (supplement)):3, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). ACM MemberNet.

Anonymous:1996:NGC

[1089] Anonymous. New E-Guide CD-ROM demo. price update. Communications of the ACM, 39(8 (supplement)):4, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). ACM MemberNet.

Anonymous:1996:SDC

[1090] Anonymous. SIGCHI Development Consortium reaches out. Communications of the ACM, 39(8 (supplement)):7, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). ACM Member-Net.

Anonymous:1996:ESW

[1091] Anonymous. Electronic services working group. Communications of the ACM, 39(8 (supplement)):8, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). ACM MemberNet.

Fox:1996:NTb

[1092] Robert Fox. News track. Communications of the ACM, 39(8):11–12, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sieger:1996:WCD

[1093] Jonah Sieger. From Washington:

Communications decency act is defeated: Landmark victory for netizens. Communications of the ACM, 39(8):13–15, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/232018.html.

Anonymous:1996:AGI

[1094] Anonymous. ACM group insurance program offers term life insurance at special rate. Communications of the ACM, 39(8 (supplement)):14–15, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). ACM MemberNet.

Moore:1996:SSO

[1095] James W. Moore and Roy Rada. Sharing standards: Organizational badge collecting. Communications of the ACM, 39(8):17–21, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/232020.html.

Mann:1996:VSC

[1096] Steve Mann. Viewpoint: Smart clothing: The shift to wearable computing. Communications of the ACM, 39(8):23–24, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/232021.html.

Selker:1996:NPC

[1097] Ted Selker. New paradigms for computing. Communications of the ACM, 39(8):28–30, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/232022.html.

Brown:1996:DID

[1098] John Seely Brown. To dream the invisible dream. Communications of the ACM, 39(8):30, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/232024.html;   http:/
/www.acm.org/pubs/toc/Abstracts/
cacm/232024.html.

Bushnell:1996:RBF

[1099] Nolan Bushnell. Relationships between fun and the computer business. Communications of the ACM, 39(8):31–37, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/232025.html;   http:/
/www.acm.org/pubs/toc/Abstracts/
cacm/232025.html.

Lieberman:1996:IG

[1100] Henry Lieberman. Intelligent graphics. Communications of the ACM, 39(8):38–48, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/232026.html;   http:/
/www.acm.org/pubs/toc/Abstracts/
cacm/232026.html.

Kahn:1996:DNV

[1101] Ken Kahn. Drawings on napkins, video-game animation, and other ways to program computers. Communications of the ACM, 39(8):49–59, August 1996. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/232028.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
232028.html.

Selker:1996:NPU

[1102] Ted Selker. New paradigms for using computers. Communications of the ACM, 39(8):60–69, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/232030.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
232030.html.

Gentner:1996:AMI

[1103] Don Gentner and Jakob Nielson. The Anti-Mac interface. Communications of the ACM, 39(8):70–82, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/232032.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
232032.html.

Soloway:1996:IUC

[1104] Elliot Soloway and Amanda Pryor. Introduction: Using computational media to facilitate learning. Communications of the ACM, 39(8):83, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/232033.html.

Rubin:1996:CTC

[1105] Andee Rubin, Scott Bresnahan, and Ted Ducas. Cartwheeling through CamMotion. Communications of the ACM, 39(8):84–85, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/232035.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
232035.html.

Confrey:1996:FP

[1106] Jere Confrey and Alan Maloney.

Function probe. Communications of the ACM, 39(8):86–87, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/toc/Abstracts/0001-0782/232036.
html;   http://www.acm.org/pubs/
toc/Abstracts/cacm/232036.html.

Feurzeig:1996:FM

[1107] Wallace Feurzeig and John Richards.

Function machines. Communications of the ACM, 39(8):88–90, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/232038.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
232038.html.

Cappo:1996:MM

[1108] Marge Cappo and Kathy Darling.

Measurement in motion. Communications of the ACM, 39(8):91–93, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/232039.html;   http://www.
                                        acm.org/pubs/toc/Abstracts/cacm/
                                        232039.html.

Schwartz:1996:MTE

[1109] Judah L. Schwartz. Motion toys for eye and mind. Communications of the ACM, 39(8):94–96, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/232040.html;   http://www.
                                        acm.org/pubs/toc/Abstracts/cacm/
                                        232040.html.

Roschelle:1996:SMM

[1110] Jeremy Roschelle and James J. Kaput. SimCalc MathWorlds for the mathematics of change. Communications of the ACM, 39(8):97–99, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/toc/Abstracts/cacm/232041.
                                        html.

Horwitz:1996:TSM

[1111] Paul Horwitz, Eric Neumann, and Joyce Schwartz. Teaching science at multiple space-time scales. Communications of the ACM, 39(8):100–102, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/0001-
                                        0782/232042.html;   http://www.
                                        acm.org/pubs/toc/Abstracts/cacm/
                                        232042.html.

Dillenbourg:1996:WIC

[1112] Pierre Dillenbourg and J. A. Self. What if the computer doesn’t know the answer? Communications of the ACM, 39(8):103–105, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
0001-0782/232044.html;   http:/
/www.acm.org/pubs/toc/Abstracts/
cacm/232044.html.

Anonymous:1996:MLS

[1113] Anonymous. A multimedia literacy series: The Cognition and Technology Group at Vanderbilt. Communications of the ACM, 39(8):106–109, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-
0782/232046.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
232046.html.

Lin:1996:IRC

[1114] Herbert Lin. Inside risks: Cryptography’s role in securing information. Communications of the ACM, 39(8):131, August 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sterling:1996:SWF

[1115] Thomas Sterling. The scientific workstation of the future may be a pile of PCs. Communications of the ACM, 39 (9):11–12, September 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/234461.html.

Meeks:1996:DKS

[1116] Brock N. Meeks. Dragging a kickingand-screaming Government into the 21st Century. Communications of the ACM, 39(9):13–15, September 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/toc/Abstracts/cacm/234462.
html.

Isaacs:1996:TTM

[1117] Ellen A. Isaacs and John C. Tang. Technology transfer: so much research, so few good products. Communications of the ACM, 39(9):22–25, September 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/toc/Abstracts/cacm/234463.
html.

Scholtz:1996:TTT

[1118] Jean Scholtz. Technology transfer through prototypes. Communications of the ACM, 39(9):26–27, September 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/toc/Abstracts/cacm/234464.
html.

Kuchinsky:1996:TMM

[1119] Allan Kuchinsky. Transfer means more than just technology. Communications of the ACM, 39(9):28–29, September 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/toc/Abstracts/cacm/234465.
html.

Foley:1996:TTU

[1120] Jim Foley. Technology transfer from university to industry. Communications of the ACM, 39(9):30–31, September 1996. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/cacm/
234466.html.

Johnson:1996:RDR

[1121] Jeff Johnson. R-D, not R&D.

Communications of the ACM, 39(9): 32–34, September 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/234467.html.

Bennett:1996:BRT

[1122] John L. Bennett. Building relationships for technology transfer. Communications of the ACM, 39(9): 35–36, September 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/234468.html.

Islam:1996:LDO

[1123] Nayeem Islam and Roy H. Campbell.

Latest developments in operating systems. Communications of the ACM, 39 (9):38–40, September 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/234469.html.

Druschel:1996:OSS

[1124] Peter Druschel. Operating system support for high-speed communication: techniques to eliminate processing bottlenecks in high-speed networking are presented. Communications of the ACM, 39(9):41–51, September 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/

                                        pubs/toc/Abstracts/cacm/234470.
                                        html.

Rosenblum:1996:IEF

[1125] Mendel Rosenblum, John Chapin, Dan Teodosiu, Scott Devine, Tirthankar Lahiri, and Anoop Gupta. Implementing efficient fault containment for multiprocessors: confining faults in a shared-memory multiprocessor environment. Communications of the ACM, 39(9):52–61, September 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/toc/Abstracts/cacm/234471.
                                        html.

Rosenberg:1996:OSS

[1126] John Rosenberg, Alan Dearle, David Hulse, Anders Lindström, and Stephen Norris. Operating system support for persistant and recoverable computations. Communications of the ACM, 39(9):62–69, September 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        cacm/234472.html.

Liedtke:1996:TRM

[1127] Jochen Liedtke. Toward real microkernels. Communications of the ACM, 39 (9):70–77, September 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        cacm/234473.html.

Cutkosky:1996:MCE

[1128] Mark R. Cutkosky, Jay M. Tenenbaum, and Jay Glicksman. Madefast: Collaborative engineering over the Internet. Communications of the ACM, 39 (9):78–87, September 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/234474.html.

Vishkin:1996:CPA

[1129] Uzi Vishkin. Can parallel algorithms enhance serial implementation? Communications of the ACM, 39(9): 88–91, September 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/234475.html.

Laudon:1996:MP

[1130] Kenneth C. Laudon. Markets and privacy. Communications of the ACM, 39 (9):92–104, September 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/234476.html.

Baker:1996:CPC

[1131] Richard Baker. The corporate politics of CMM ratings. Communications of the ACM, 39(9):105–106, September 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/toc/Abstracts/cacm/234477.
html.

Weinstein:1996:IRB

[1132] Lauren Weinstein. Inside risks: Behind the state of the art. Communications of the ACM, 39(9):130, September 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/toc/Abstracts/cacm/234478.
html.

Press:1996:SNF

[1133] Larry Press. Seeding networks: The federal role. Communications of the ACM, 39(10):11–18, October 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/toc/Abstracts/cacm/234979.
                                        html.

Berghel:1996:DP

[1134] Hal Berghel. Digital politics. Communications of the ACM, 39(10):19– 25, October 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/cacm/
                                        236158.html.

Luker:1996:VNN

[1135] Mark Luker. Viewpoint: NSF’s new program for high-performance Internet connections. Communications of the ACM, 39(10):27–28, October 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/toc/Abstracts/cacm/236160.
                                        html.

Rada:1996:AFR

[1136] Roy Rada, Anne Mette Hass, Don Herbison-Evans, John Waterworth, and Magnus Kempe. ACM forum: Request for a virtual IT college. Communications of the ACM, 39(10):29–35, October 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        cacm/236162.html.

Schmidt:1996:SPI

[1137] Douglas C. Schmidt, Mohamed Fayad, and Ralph E. Johnson. Software patterns: introductions. Communications of the ACM, 39(10):36–39, October 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/toc/Abstracts/cacm/234982.
html.

Cockburn:1996:ISI

[1138] Alistair Cockburn. The interaction of social issues and software architecture. Communications of the ACM, 39(10):40–46, October 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/234983.html.

Cline:1996:PCA

[1139] Marshall P. Cline. The pros and cons of adopting and applying design patterns in the real world. Communications of the ACM, 39(10):47–49, October 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/toc/Abstracts/cacm/234984.
html.

Aarsten:1996:DCD

[1140] Amund Aarsten, Davide Brugali, and Giuseppe Menga. Designing concurrent and distributed control systems. Communications of the ACM, 39(10):50–58, October 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/236168.html.

Fayad:1996:ASA

[1141] Mohamed Fayad and Marshall P. Cline. Aspects of software adaptability. Communications of the ACM, 39 (10):58–59, October 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        cacm/236170.html.

Goldfedder:1996:TEP

[1142] Brandon Goldfedder and Linda Rising. A training experience with patterns. Communications of the ACM, 39(10):60–64, October 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        cacm/236171.html.

Islam:1996:EDP

[1143] Nayeem Islam and Murthy Devarakonda. An essential design pattern for fault-tolerant distributed state sharing. Communications of the ACM, 39(10):65–74, October 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        cacm/236172.html.

McKenney:1996:SLP

[1144] Paul E. McKenney. Selecting locking primitives and parallel programming. Communications of the ACM, 39(10):75–82, October 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

                                        /www.acm.org/pubs/toc/Abstracts/
                                        cacm/236174.html.

Anonymous:1996:WLP

[1145] Anonymous. Where to look for patterns. Communications of the ACM, 39(10):82, October 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Resnick:1996:PIA

[1146] Paul Resnick and James Miller. PICS:

Internet access controls without censorship. Communications of the ACM, 39(10):87–93, October 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/234989.html;   http://www.
acm.org/pubs/toc/Abstracts/cacm/
236175.html.

Iivari:1996:WCT

[1147] Juhani Iivari. Why are case tools not used? Communications of the ACM, 39(10):94–103, October 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/236183.html.

Basili:1996:HRI

[1148] Victor R. Basili, Lionel C. Briand, and Walcélio L. Melo. How reuse influences productivity in object-oriented systems. Communications of the ACM, 39 (10):104–116, October 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/
cacm/236184.html.

Neumann:1996:IRDa

[1149] Peter G. Neumann. Inside risks:

Disinformation theory. Communications of the ACM, 39(10):138, October 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/toc/Abstracts/cacm/234992.
                                        html.

Glass:1996:PPR

[1150] Robert L. Glass. Practical programmer: The relationship between theory and practice in software engineering. Communications of the ACM, 39 (11):11–13, November 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Munro:1996:WSN

[1151] Neil Munro. From Washington: Sketching a National Information Warfare Defense Plan. Communications of the ACM, 39(11):15–17, November 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fayyad:1996:DMK

[1152] Usama Fayyad and Ramasamy Uthurusamy. Data mining and knowledge discovery in databases. Communications of the ACM, 39(11):24–27, November 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fayyad:1996:KPE

[1153] Usama Fayyad, Gregory Piatetsky-Shapiro, and Padhraic Smyth. The KDD process for extracting useful knowledge from volumes of data. Communications of the ACM, 39(11):27–34, November 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glymour:1996:SID

[1154] Clark Glymour, David Madigan, Daryl Pregibon, and Padhraic Smyth. Statistical inference and data mining. Communications of the ACM, 39(11): 35–41, November 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brachman:1996:MBD

[1155] Ronald J. Brachman, Tom Khabaza, Willi Kloesgen, Gregory Piatetsky-Shapiro, and Evangelos Simoudis. Mining business databases. Communications of the ACM, 39(11):42–48, November 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Inmon:1996:DWD

[1156] W. H. Inmon. The data warehouse and data mining. Communications of the ACM, 39(11):49–50, November 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fayyad:1996:MSD

[1157] Usama Fayyad, David Haussler, and Paul Stolorz. Mining scientific data. Communications of the ACM, 39(11): 51–57, November 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Imielinski:1996:DPK

[1158] Tomasz Imielinski and Heikki Mannila. A database perspective on knowledge discovery. Communications of the ACM, 39(11):58–64, November 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Etzioni:1996:WWW

[1159] Oren Etzioni. The World-Wide Web: Quagmire or gold mine? Communications of the ACM, 39(11):65–77, November 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Abramson:1996:RDN

[1160] David Abramson, Ian Foster, John Michalakes, and Rok Socič. Relative debugging: a new methodology for debugging scientific applications. Communications of the ACM, 39 (11):69–77, November 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.cit.gu.edu.au/~davida/guard.

html. Uses IBM Data Explorer to illustrate the computational differences between two versions of the same program, using the Guard (Griffith University Relative Debugger) system to automatically carry out run-time monitoring and comparison of results of two implementations of the same program, possibly written in different programming languages, and possibly running on different computer architectures at different geographic locations.

Culler:1996:LPM

[1161] David E. Culler, Rachard M. Karp, David Patterson, Abhijit Sahay, Eunice E. Santos, Klaus Erik Schauser, Ramesh Subramonian, and Thorsten von Eicken. LogP: a practical model of parallel computation. Communications of the ACM, 39(11):78–85, November 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wand:1996:ADQ

[1162] Yair Wand and Rachard Y. Wang. Anchoring data quality dimensions in ontological foundations. Communications of the ACM, 39(11):86–95, November 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:1996:IRDb

[1163] Peter G. Neumann. Inside risks:

Distributed systems have distributed risks. Communications of the ACM, 39(11):130, November 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goodman:1996:IPW

[1164] S. E. Goodman. International perspectives: War, information technologies, and international asymmetries. Communications of the ACM, 39(12): 11–15, December 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samuelson:1996:LSL

[1165] Pamela Samuelson. Legally speaking: Legal protection for database contents. Communications of the ACM, 39 (12):17–23, December 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Noll:1996:CTI

[1166] A. Michael Noll. CyberNetwork technology: Issues and uncertainties. Communications of the ACM, 39(12):27–31, 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kraut:1996:I

[1167] Robert Kraut. The Internet@Home. Communications of the ACM, 39(12): 32–35, December 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoffman:1996:IWU

[1168] Donna L. Hoffman, William D. Kalsbeek, and Thomas P. Novak. Internet and Web use in the U.S. Communications of the ACM, 39(12):36–46, December 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Venkatesh:1996:COI

[1169] Alladi Venkatesh. Computers and other interactive technologies for the home. Communications of the ACM, 39(12):47–54, December 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kraut:1996:HFT

[1170] Robert Kraut, William Scherlis, Tridas Mukhopadhyay, Jane Manning, and Sara Kiesler. The HomeNet field trial of residential Internet services. Communications of the ACM, 39 (12):55–63, December 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Franzke:1996:WGW

[1171] Marita Franzke and Anne McClard. Winona gets wired: Technical difficulties in the home. Communications of the ACM, 39(12):64–66, December 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Duncan:1996:MRE

[1172] George T. Duncan. Is my research ethical? Communications of the ACM, 39 (12):67–68, December 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Carroll:1996:DBE

[1173] John M. Carroll and Mary Beth Rosson. Developing the Blacksburg Electronic Village. Communications of the ACM, 39(12):69–74, December 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Martin:1996:ITS

[1174] C. Dianne Martin, Chuck Huff, Donald Gotterbarn, and Keith Miller. Implementing a tenth strand in the CS curriculum. Communications of the ACM, 39(12):75–84, December 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Walker:1996:RMC

[1175] Henry M. Walker and G. Michael Schneider. A revised model curriculum for a liberal arts degree in computer science. Communications of the ACM, 39(12):85–95, December 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Geist:1996:CRP

[1176] Robert Geist, Madhu Chetuparambil, Stephen Hedetniemi, and A. Joe Turner. Computing research programs in the U.S. Communications of the ACM, 39(12):96–99, December 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:1996:IRRb

[1177] Peter G. Neumann. Inside risks: Risks of anonymity. Communications of the ACM, 39(12):162, December 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kemp:1996:CFP

[1178] C. David Kemp. The construction of fast portable multiplicative congruential random number generators. Communications of the ACM, 39(12es): 163–166, 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
1996-39-12es/a163-kemp/a163-kemp.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/1996-39-
12es/a163-kemp/.

Barrero:1996:IAD

[1179] Alejandro Barrero. Implementation of abstract data types with arrays of unbounded dimensions. Communications of the ACM, 39(12es): 167–174, 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1996-39-12es/a167-barrero/.

Michalewicz:1996:GGA

[1180] Zbigniew Michalewicz and Cezary Z. Janikow. GENOCOP: a genetic algorithm for numerical optimization problems with linear constraints. Communications of the ACM, 39(12es): 175–201, 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
1996-39-12es/a175-michalewicz/a175-

michalewicz.pdf;   http://www.acm.
org/pubs/citations/journals/cacm/
1996-39-12es/a175-michalewicz/.

Pearlmutter:1996:GCP

[1181] Barak A. Pearlmutter. Garbage collection with pointers to individual cells. Communications of the ACM, 39(12es): 202–206, 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1996-39-12es/a202-pearlmutter/

.

Rojas:1996:GLP

[1182] Raúl Rojas. A graph labelling proof of the backpropagation algorithm. Communications of the ACM, 39(12es): 207–215, 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1996-39-12es/a207-rojas/.

Chu:1996:EBC

[1183] Wesley W. Chu, Kuorong Chiang, Chih-Cheng Hsu, and Henrick Yau. An error-based conceptual clustering method for providing approximate query answers. Communications of the ACM, 39(12es):216–230, 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://dblp.uni-trier.de/
db/journals/cacm/ChuCHY96.html;
http://www.acm.org/pubs/citations/
journals/cacm/1996-39-12es/a216-
chu/.

Nelson:1996:IQE

[1184] Randolph D. Nelson. Including queueing effects in Amdahl’s law. Communications of the ACM, 39(12es):

231–238, 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1996-39-12es/a231-nelson/.

Prather:1996:HAC

[1185] Ronald E. Prather. Harmonic analysis from the computer representation of a musical score. Communications of the ACM, 39(12es): 239–255, 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1996-39-12es/a239-prather/.

Staskausas:1996:EFV

[1186] Mark G. Staskausas. An experience in the formal verification of industrial software. Communications of the ACM, 39(12es):256– 272, 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1996-39-12es/a256-staskauskas/

.

Blelloch:1996:CSA

[1187] Guy E. Blelloch, Charles E. Leiserson, Bruce M. Maggs, C. Greg Plaxton, Stephen J. Smith, and Marco Zagha. A comparison of sorting algorithms for the Connection Machine CM-2. Communications of the ACM, 39(12es):273– 297, 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Borenstein:1996:CAP

[1188] Nathaniel Borenstein. CMU’s Andrew Project: a retrospective. Communications of the ACM, 39(12es):

298–311, 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
1996-39-12es/a298-borenstein/a298-
borenstein.pdf;   http://www.acm.
org/pubs/citations/journals/cacm/
1996-39-12es/a298-borenstein/.

Morris:1996:WSA

[1189] James H. Morris. War stories from Andrew. Communications of the ACM, 39(12es):312–321, 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/articles/journals/
cacm/1996-39-12es/a312-morris/a312-
morris.pdf;   http://www.acm.org/
pubs/citations/journals/cacm/1996-
39-12es/a312-morris/.

Munro:1997:WII

[1190] Neil Munro. From Washington:

If it grows, tax it. Communications of the ACM, 40(1):11–13, January 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-1/p11-munro/.

Glass:1997:PPN

[1191] Robert L. Glass. Practical programmer: The next date crisis and the ones after that. Communications of the ACM, 40(1):15–17, January 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-1/p15-glass/.

Fox:1997:STL

[1192] Robert Fox. On site: Tomorrow’s library today. Communications of the ACM, 40(1):20–21, January 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-1/p20-fox/.

Peleg:1997:IMM

[1193] Alex Peleg, Sam Wilkie, and Uri Weiser. Intel MMX for multimedia PCs. Communications of the ACM, 40(1):24–38, January 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-1/p25-peleg/.  See

also Blinn’s comments [1947] about MMX instruction set deficiencies.

Grimshaw:1997:LVW

[1194] Andrew S. Grimshaw and William A. Wulf. The Legion vision of a worldwide virtual computer. Communications of the ACM, 40(1):39–45, January 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-1/p39-grimshaw/.

Nachenberg:1997:CVA

[1195] Carey Nachenberg. Computer virus: Antivirus coevolution. Communications of the ACM, 40(1):46–51, January 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-1/p46-nachenberg/.

OLeary:1997:AIV

[1196] Daniel E. O’Leary, Daniel Kuokka, and Robert Plant. Artificial intelligence and virtual organizations. Communications of the ACM, 40(1):52–59, January 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-1/p52-o_leary/.

Johnson:1997:EO

[1197] Deborah G. Johnson. Ethics online. Communications of the ACM, 40(1):60–65, January 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/1997-40-1/p60-johnson/.

Bhattacharya:1997:GEI

[1198] Sourav Bhattacharya, Akira Onoma, and Farokh Bastani. Guest editor’s introduction: High-assurance systems. Communications of the ACM, 40(1): 67, January 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-1/p67-bhattacharya/

.

Neves:1997:ARM

[1199] Nuno Neves and W. Kent Fuchs.

Adaptive recovery for mobile environments. Communications of the ACM, 40(1):68–74, January 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/1997-
40-1/p68-neves/.

Kim:1997:FTR

[1200] K. H. Kim and Chittur Subbaraman. Fault-tolerant real-time objects. Communications of the ACM, 40(1):75–82, January 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-1/p75-kim/.

Tsai:1997:PES

[1201] Jeffrey J. P. Tsai, Bing Li, and Eric Y. T. Juan. Parallel evaluation of software architecture specifications. Communications of the ACM, 40(1):83–86, January 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-1/p83-tsai/.

Stewart:1997:MDH

[1202] David B. Stewart and Pradeep K. Khosla. Mechanisms for detecting and handling timing errors. Communications of the ACM, 40(1):87–93, January 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-1/p87-stewart/.

Alpert:1997:UPC

[1203] Sherman R. Alpert and Richard B. Lam. The ultimately publishable computer science paper for the latter ’90s: a tip for authors. Communications of the ACM, 40(1):94, January 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). This parody paper manages to use Java and JavaScript more than a dozen times.

Schneier:1997:IRC

[1204] Bruce Schneier. Inside risks: Cryptography, security, and the future. Communications of the ACM, 40 (1):138, January 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.counterpane.com/csf.html.

Anonymous:1997:APF

[1205] Anonymous. ACM plus five other leading computing societies met to discuss international cooperation. Communications of the ACM, 40(2S):3–??, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:1997:MVP

[1206] Anonymous. Member value plus keeps on growing. Communications of the ACM, 40(2S):5–??, February 1997. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:1997:ACS

[1207] Anonymous. ACM Committee on the Status of Women in Computing moves ahead with new projects. Communications of the ACM, 40(2S):8–??, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Press:1997:PCTa

[1208] Larry Press. Personal computing:

Technology in bloom: Implications for the next 50 years. Communications of the ACM, 40(2):11–17, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p11-press/.

Anonymous:1997:NAT

[1209] Anonymous. New ACM Transactions on Information and System Security will begin publication. Communications of the ACM, 40(2S):14–??, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:1997:DVC

[1210] Hal Berghel. Digital village: Cyberspace 2000: Dealing with information overload. Communications of the ACM, 40(2):19–24, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p19-berghel/.

Myhrvold:1997:ASC

[1211] Nathan Myhrvold. ACM 97: speakers corner. Communications of the ACM, 40(2):27–28, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/1997-
40-2/p27-myhrvold/.

Crawford:1997:NY

[1212] Diane Crawford. The next 50 years. Communications of the ACM, 40(2): 29, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p29-crawford/.

Bell:1997:HCB

[1213] Gordon Bell. The human connection: The body electric. Communications of the ACM, 40(2):30–32, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p31-bell/.

OReilly:1997:HCH

[1214] Tim O’Reilly. The human connection: Hardware, software, and infoware. Communications of the ACM, 40(2):33–34, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/1997-40-2/p33-o_reilly/.

Dyson:1997:HCE

[1215] Esther Dyson. The human connection: Education and jobs in the digital world. Communications of the ACM, 40(2):35–36, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p35-dyson/.

Hillis:1997:HCT

[1216] Danny Hillis. The human connection: a time of transition. Communications of the ACM, 40(2):37–39, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p37-hillis/.

Birnbaum:1997:HCP

[1217] Joel Birnbaum. The human connection: Pervasive information systems. Communications of the ACM, 40(2):40–41, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/1997-40-2/p40-birnbaum/.

Davis:1997:HCG

[1218] Marc Davis. The human connection: Garage cinema and the future of media technology. Communications of the ACM, 40(2):42–48, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p42-davis/.

Tsichritzis:1997:HCH

[1219] Dennis Tsichritzis. The human connection: How to surf the technology waves we created. Communications of the ACM, 40(2):49–54, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p49-tsichritzis/.

Lanier:1997:HCF

[1220] Jaron Lanier. The human connection: The frontier between us. Communications of the ACM, 40(2):55–56, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p55-lanier/.

Berners-Lee:1997:HCW

[1221] Tim Berners-Lee. The human connection: World-wide computer. Communications of the ACM, 40(2):57–58, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p57-berners-lee/.

Shneiderman:1997:HCB

[1222] Ben Shneiderman. The human connection: Between hope and fear. Communications of the ACM, 40(2):59–62, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p59-shneiderman/.

vanDam:1997:HCP

[1223] Andries van Dam. The human connection: Post-WIMP user interfaces. Communications of the ACM, 40(2):63–67, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p63-van_dam/.

Barlow:1997:SDB

[1224] John Perry Barlow. Societal dimensions: The best of all possible worlds. Communications of the ACM, 40(2):68–74, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/1997-
40-2/p69-barlow/.

Meeks:1997:SDB

[1225] Brock N. Meeks. Societal dimensions: Better democracy through technology. Communications of the ACM, 40(2):75–78, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/1997-
40-2/p75-meeks/.

Sterling:1997:SDD

[1226] Bruce Sterling. Societal dimensions: The digital revolution in retrospect. Communications of the ACM, 40(2): 79, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p79-sterling/.

Talbott:1997:SDA

[1227] Steve Talbott. Societal dimensions: Asleep at the keyboard. Communications of the ACM, 40(2):80–82, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p80-talbott/.

Berman:1997:SDD

[1228] Jerry Berman and Daniel J. Weitzner. Societal dimensions: Directing policymaking beyond the Net’s metaphor. Communications of the ACM, 40(2): 83–84, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/1997-40-2/p83-berman/.

Stallman:1997:SDR

[1229] Richard Stallman. Societal dimensions: The right to read. Communications of the ACM, 40(2):85–87, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p85-stallman/.

Selker:1997:SDW

[1230] Ted Selker. Societal dimensions: What will happen in the next 50 years? Communications of the ACM, 40(2):88–89, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p88-selker/.

Nelson:1997:SDC

[1231] Theodor Holm Nelson. Societal dimensions: Crush and crash: Logic of a terrible tomorrow. Communications of the ACM, 40(2):90–91, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p90-nelson/.

Saffo:1997:SFT

[1232] Paul Saffo. The science of future technology: Sensors: The next wave of innovation. Communications of the ACM, 40(2):92–97, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p92-saffo/.

Raskin:1997:SFT

[1233] Jef Raskin. The science of future technology: Looking for a humane interface: Will computers ever become easy to use? Communications of the ACM, 40(2):98–101, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p98-raskin/.

Leiner:1997:SFT

[1234] Barry M. Leiner, Vinton G. Cerf, David D. Clark, Robert E. Kahn, Leonard Kleinrock, Daniel C. Lynch, Jon Postel, Lawrence G. Roberts, and Stephen S. Wolff. The science of future technology: The past and future history of the Internet. Communications of the ACM, 40(2):102–108, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p102-leiner/.

Wulf:1997:SFT

[1235] Wm. A. Wulf. The science of future technology: Look in the spaces for tomorrow’s innovations. Communications of the ACM, 40(2):109–111, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p109-wulf/.

Moore:1997:SFT

[1236] Gordon E. Moore. The science of future technology: The microprocessor: Engine of the technology revolution. Communications of the ACM, 40(2):112–114, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p112-moore/.

Adam:1997:SFT

[1237] Nabil Adam, Baruch Awerbuch, Jacob Slonim, Peter Wegner, and Yelena Yesha. The science of future technology: Globalizing business, education, culture through the Internet. Communications of the ACM, 40(2):115–121, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p115-adam/.

Rosen:1997:SFT

[1238] Charlie Rosen. The science of future technology: My computer-related wish list for the next 50 years. Communications of the ACM, 40(2):122–124, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p122-rosen/.

Slingers:1997:SFT

[1239] Sachem Web Slingers. The science of future technology: Spinning webs into the 21st century. Communications of the ACM, 40(2):125–128, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p125-sachem_web_
slingers/.

Leveson:1997:SFT

[1240] Nancy G. Leveson. The science of future technology: Software engineering: Stretching the limits of complexity. Communications of the ACM, 40(2):129–131, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-2/p129-leveson/.

Denning:1997:SFT

[1241] Peter J. Denning. The science of future technology: a new social contract for research. Communications of the ACM, 40(2):132–134, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-2/p132-denning/.

Farber:1997:SFT

[1242] David J. Farber. The science of future technology: Communications technology and its impact by 2010. Communications of the ACM, 40(2):135–138, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-2/p135-farber/.

Korth:1997:SFT

[1243] Henry F. Korth and Abraham Silberschatz. The science of future technology: Database research faces the information explosion. Communications of the ACM, 40(2):139–142, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p139-korth/.

Neumann:1997:IRH

[1244] Peter G. Neumann. Inside risks:

Hopes for fewer risks. Communications of the ACM, 40(2):170, February 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-2/p170-neumann/.

ORiain:1997:IPB

[1245] Seán ÓRiain. International perspectives: The birth of a Celtic tiger. Communications of the ACM, 40(3):11– 16, March 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-3/p11-oriain/.

Rada:1997:SSS

[1246] Roy Rada and James Moore. Sharing standards: Standardizing reuse. Communications of the ACM, 40(3):19– 23, March 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-3/p19-rada/.

Friedlander:1997:SSF

[1247] Carl Friedlander. On site: Speech facilities for the reading disabled. Communications of the ACM, 40(3):24– 25, March 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-3/p24-friedlander/.

Sallach:1997:FDP

[1248] David Sallach and Andrew Oram. Forum: Digital politics. Communications of the ACM, 40(3):26–27, March 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-3/p26-sallach/.

Jones:1997:FIB

[1249] Michael G. Jones. Forum: Inside the Beltway warfare plans. Communications of the ACM, 40(3):27– 28, March 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-3/p27-jones/.

Pemberton:1997:FUT

[1250] Steven Pemberton. Forum: U.S. and them. Communications of the ACM, 40(3):28, March 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

                                        www.acm.org/pubs/citations/journals/
                                        cacm/1997-40-3/p28-pemberton/.

Phoha:1997:FTP

[1251] Vir V. Phoha. Forum: Theory and practice in software engineering. Communications of the ACM, 40(3):28, March 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-3/p28-phoha/.

Fedanzo:1997:FDS

[1252] Anthony Fedanzo. Forum: Distributed systems registry. Communications of the ACM, 40(3):29, March 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-3/p29-fedanzo/.

Glass:1997:FR

[1253] Robert L. Glass. Forum: Response.

Communications of the ACM, 40(3): 29, March 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-3/p29-glass/.

Beer:1997:BIA

[1254] Randall D. Beer, Roger D. Quinn, Hillel J. Chiel, and Roy E. Ritzmann. Biologically inspired approaches to robotics: what can we learn from insects? Communications of the ACM, 40(3):30–38, March 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-3/p31-beer/.

Mueller:1997:UST

[1255] Milton Mueller. Universal service and the telecommunications act: Myth made law. Communications of the ACM, 40(3):39–47, March 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-3/p39-mueller/.

Aliaga:1997:VOR

[1256] Daniel G. Aliaga. Virtual objects in the real world. Communications of the ACM, 40(3):49–54, March 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-3/p49-aliaga/.

Resnick:1997:RS

[1257] Paul Resnick and Hal R. Varian. Recommender systems. Communications of the ACM, 40(3):56–58, March 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-3/p56-resnick/.

Terveen:1997:PSS

[1258] Loren Terveen, Will Hill, Brian Amento, David McDonald, and Josh Creter. PHOAKS: a system for sharing recommendations. Communications of the ACM, 40(3):59–62, March 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-3/p59-terveen/.

Kautz:1997:RWC

[1259] Henry Kautz, Bart Selman, and Mehul Shah. Referral Web: Combining social networks and collaborative filtering. Communications of the ACM, 40(3):63–65, March 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-3/p63-kautz/.

Balabanovic:1997:FCB

[1260] Marko Balabanovíc and Yoav Shoham. Fab: Content-based, collaborative recommendation. Communications of the ACM, 40(3):66–72, March 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/1997-
40-3/p66-balabanovic/.

Rucker:1997:SPN

[1261] James Rucker and Marcos J. Polanco.

Siteseer: Personalized navigation for the Web. Communications of the ACM, 40(3):73–76, March 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/1997-
40-3/p73-rucker/.

Konstan:1997:GAC

[1262] Joseph A. Konstan, Bradley N. Miller, David Maltz, Jonathan L. Herlocker, Lee R. Gordon, and John Riedl. GroupLens: Applying collaborative filtering to Usenet news. Communications of the ACM, 40(3):77–87, March 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-3/p77-konstan/.

Avery:1997:RSE

[1263] Christopher Avery and Richard Zeckhauser. Recommender systems for evaluating computer messages. Communications of the ACM, 40(3):88– 89, March 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-3/p88-avery/.

Koumpis:1997:SRA

[1264] Adamantios Koumpis. Situation room analysis in the information technologies market. Communications of the ACM, 40(3):90–92, March 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-3/p90-koumpis/.

Cook:1997:IRO

[1265] Richard I. Cook. Inside risks: Observations on RISKS and risks. Communications of the ACM, 40(3):122, March 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:1997:DVE

[1266] Hal Berghel. Digital village: Email: The good, the bad, and the ugly. Communications of the ACM, 40(4):11– 15, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-4/p11-berghel/.

Glass:1997:PPU

[1267] Robert L. Glass. Practical programmer: The ups and downs of programmer stress. Communications of the ACM, 40(4):17–19, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-4/p17-glass/.

Johnson:1997:FMI

[1268] Jeff Johnson. Forum: The military impact of information technology.

Communications of the ACM, 40(4): 20, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-4/p20-johnson/.

Davis:1997:FQP

[1269] Ronald L. Davis and Roger W. Webster. Forum: Questioning proposed CS knowledge units. Communications of the ACM, 40(4):20–21, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cowan:1997:FAT

[1270] Crispin Cowan. Forum: Amazing transformation. Communications of the ACM, 40(4):21, April 1997. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Exner:1997:FSI

[1271] Frank Exner. Forum: The subject of indexing. Communications of the ACM, 40(4):21, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Criticizes the Communications’ ‘1996 Subject Index’ as just an extended table of contents.

Patel:1997:FFN

[1272] Mayur Patel. Forum: Filling a niche. Communications of the ACM, 40(4):21–22, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goldstein:1997:FUC

[1273] Barry Goldstein. Forum: URL corrections. Communications of the ACM, 40(4):22, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Criticizes the incorrect insertion of hyphens in URLs that are split across lines.

Lingle:1997:FWP

[1274] Michael Lingle. Forum: What price books. Communications of the ACM, 40(4):22, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nachenberg:1997:FR

[1275] Carey Nachenberg. Forum: Response. Communications of the ACM, 40(4):22, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Yun:1997:FMS

[1276] James K. Yun. Forum: Missing step. Communications of the ACM, 40(4):22, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hsu:1997:VTS

[1277] Richard C. Hsu. Viewpoint: Taking stock for university patents. Communications of the ACM, 40(4):23– 25, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-4/p23-hsu/.

Lieberman:1997:I

[1278] Henry Lieberman. Introduction. Communications of the ACM, 40(4):26– 29, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-4/p26-lieberman/.

Eisenstadt:1997:MHB

[1279] Marc Eisenstadt. My hairiest bug war stories. Communications of the ACM, 40(4):30–37, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-4/p30-eisenstadt/.

Ungar:1997:DEI

[1280] David Ungar, Henry Lieberman, and Christopher Fry. Debugging and the experience of immediacy. Communications of the ACM, 40(4):38– 43, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-4/p38-ungar/.

Baecker:1997:SVD

[1281] Ron Baecker, Chris DiGiano, and Aaron Marcus. Software visualization for debugging. Communications of the ACM, 40(4):44–54, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-4/p44-baecker/.

Fry:1997:PAF

[1282] Christopher Fry. Programming on an already full brain. Communications of the ACM, 40(4):55–64, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-4/p55-fry/.

Domingue:1997:FDC

[1283] John Domingue and Paul Mulholland. Fostering debugging communities on the Web. Communications of the ACM, 40(4):65–71, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-4/p65-domingue/.

Smith:1997:KOC

[1284] Randall B. Smith, Mario Wolczko, and David Ungar. From Kansas to Oz: Collaborative debugging when a shared world breaks. Communications of the ACM, 40(4):72–78, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-4/p72-smith/.

Gemmell:1997:NTC

[1285] D. James Gemmell and C. Gordon Bell. Noncollaborative telepresentations come of age. Communications of the ACM, 40(4):79– 89, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-4/p79-gemmell/.

Hardaway:1997:DMO

[1286] Don Hardaway and Richard P. Will. Digital multimedia offers key to educational reform. Communications of the ACM, 40(4):90–96, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-4/p90-will/.

Katz:1997:MHD

[1287] James Katz and Philip Aspden. Motives, hurdles, and dropouts. Communications of the ACM, 40(4):97– 102, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-4/p97-katz/.

Shekhar:1997:DMG

[1288] Shashi Shekhar, Mark Coyle, Brajesh Goyal, Duen-Ren Liu, and Shyamsundar Sarkar. Data models in geographic information systems. Communications of the ACM, 40(4):103– 111, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-4/p103-shekhar/.

West:1997:TOH

[1289] Dave West. Technical opinion:

Hermeneutic computer science. Communications of the ACM, 40(4):115– 116, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-4/p115-west/.

Felten:1997:IRW

[1290] Edward W. Felten. Inside risks:

Webware security. Communications of the ACM, 40(4):130, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/

                                        pubs/citations/journals/cacm/1997-
                                        40-4/p130-felten/.

Anonymous:1997:BCN

[1291] Anonymous. Beyond Calculation: The Next 50 years of Computing is published. Communications of the ACM, 40(5 (supplement)):2–??, May 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:1997:AWH

[1292] Anonymous. Award winners are honored at 1997 ACM awards ceremony. Communications of the ACM, 40(5 (supplement)):4–??, May 1997. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Soloway:1997:LED

[1293] Elliot Soloway and Raven Wallace. Log on education: Does the Internet support student inquiry? don’t ask. Communications of the ACM, 40(5):11– 16, May 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-5/p11-soloway/.

Anonymous:1997:AECa

[1294] Anonymous. ACM Executive Committee will meet in Stockholm. Communications of the ACM, 40(5 (supplement)):13–??, May 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:1997:TNB

[1295] Anonymous. Two new books are published by ACM Press. Communications of the ACM, 40(5 (supplement)):15–??, May 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:1997:ASP

[1296] Anonymous. ACM states policy on refereed electronic publications. Communications of the ACM, 40(5 (supplement)):16–??, May 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samuelson:1997:LSN

[1297] Pamela Samuelson. Legally speaking:

The never-ending struggle for balance. Communications of the ACM, 40(5): 17–21, May 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-5/p17-samuelson/.

Mowshowitz:1997:VLC

[1298] Abbe Mowshowitz. Viewpoint: Lessons from a cautionary tale. Communications of the ACM, 40(5):23–25, May 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-5/p23-mowshowitz/.

House:1997:AER

[1299] Charles H. House. ACM 97: an event to remember. Communications of the ACM, 40(5):31–34, May 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-5/p31-house/.

Laddaga:1997:DOT

[1300] Robert Laddaga and James Veitch. Dynamic object technology. Communications of the ACM, 40(5):36– 38, May 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-5/p36-laddaga/.

Robertson:1997:ILS

[1301] Paul Robertson. Integrating legacy systems with modern corporate applications. Communications of the ACM, 40(5):39–46, May 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/1997-40-5/p39-robertson/.

Corkill:1997:CSD

[1302] Daniel D. Corkill. Countdown to success: dynamic objects, GBB, and RADARSAT-1. Communications of the ACM, 40(5):48–58, May 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-5/p48-corkill/.

Phillips:1997:DOE

[1303] Robert E. Phillips. Dynamic objects for engineering automation. Communications of the ACM, 40(5):59– 65, May 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-5/p59-phillips/.

Davies:1997:PWC

[1304] Byron Davies and Victoria Bryan Davies. Patching onto the Web: Common Lisp hypermedia for the Intranet. Communications of the ACM, 40(5): 66–69, May 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-5/p66-davies/.

Gupta:1997:VIR

[1305] Amarnath Gupta and Ramesh Jain.

Visual information retrieval. Communications of the ACM, 40(5):70– 79, May 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-5/p70-gupta/.

Wegner:1997:WIM

[1306] Peter Wegner. Why interaction is more powerful than algorithms. Communications of the ACM, 40(5):80– 91, May 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-5/p80-wegner/.

Oppliger:1997:ISF

[1307] Rolf Oppliger. Internet security:

Firewalls and beyond. Communications of the ACM, 40(5):92–102, May 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-5/p92-oppliger/.

Strong:1997:DQC

[1308] Diane M. Strong, Yang W. Lee, and Richard Y. Wang. Data quality in context. Communications of the ACM, 40(5):103–110, May 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-5/p103-strong/.

Lefkon:1997:TOS

[1309] Dick Lefkon. Technical opinion(s): Seven work plans for Year-2000 upgrade projects. Communications of the ACM, 40(5):111– 113, May 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-5/p111-lefkon/.

Brigham:1997:Y

[1310] Georges Brigham. The year 2000. Communications of the ACM, 40(5):113– 115, May 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-5/p113-brigham/.

Wagner:1997:SDC

[1311] Robert A. Wagner. Solving the date crisis. Communications of the ACM, 40(5):115–117, May 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

                                        www.acm.org/pubs/citations/journals/
                                        cacm/1997-40-5/p115-wagner/.

Neumann:1997:IRB

[1312] Peter G. Neumann. Inside risks: The big picture. Communications of the ACM, 40(5):136, May 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/1997-
40-5/p136-neumann/.

Glass:1997:RIA

[1313] Robert L. Glass. Revisiting the industry/academe communication chasm. Communications of the ACM, 40(6): 11–13, June 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-6/p11-glass/.

Munro:1997:MPD

[1314] Neil Munro. The magnetic poles of data collection. Communications of the ACM, 40(6):17–19, June 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-6/p17-munro/.

Kraus:1997:CFB

[1315] Jeff Kraus. Counting furtive beasts by computer: where have all the rhinos gone? Communications of the ACM, 40(6):20–23, June 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/1997-
40-6/p20-kraus/.

Fox:1997:QAI

[1316] Christopher Fox and William Frakes.

The quality approach: is it delivering? Communications of the ACM, 40(6): 24–29, June 1997. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-6/p24-fox/.

Herbsleb:1997:SQC

[1317] James Herbsleb, David Zubrow, Dennis Goldenson, Will Hayes, and Mark Paulk. Software quality and the capability maturity model. Communications of the ACM, 40(6):30– 40, June 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-6/p30-herbsleb/.

Hollenbach:1997:CQS

[1318] Craig Hollenbach, Ralph Young, Al Pflugrad, and Doug Smith. Combining quality and software improvement. Communications of the ACM, 40(6): 41–45, June 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-6/p41-hollenbach/.

Arthur:1997:QIS

[1319] Lowell Jay Arthur. Quantum improvements in software system quality. Communications of the ACM, 40(6):46– 52, June 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-6/p46-arthur/.

Cusumano:1997:HMB

[1320] Michael A. Cusumano and Richard W. Selby. How Microsoft builds software. Communications of the ACM, 40(6): 53–61, June 1997. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-6/p53-cusumano/.

Parsons:1997:CCC

[1321] Jeffrey Parsons and Yair Wand. Choosing classes in conceptual modeling. Communications of the ACM, 40(6): 63–69, June 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-6/p63-parsons/.

Dutta:1997:DMS

[1322] Soumitra Dutta, Berend Wierenga, and Arco Dalebout. Designing management support systems using an integrative perspective. Communications of the ACM, 40(6):70–79, June 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-6/p70-dutta/.

Dahlbom:1997:FOP

[1323] Bo Dahlbom and Lars Mathiassen.

The future of our profession. Communications of the ACM, 40(6):80– 89, June 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-6/p80-dahlbom/.

Bailey:1997:TOO

[1324] David H. Bailey. Technical opinion:

Onward to petaflops computing. Communications of the ACM, 40(6):90– 92, June 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-6/p90-bailey/.

Berghel:1997:DVI

[1325] Hal Berghel. Digital village: Informal logic on both sides of the CDA debate. Communications of the ACM, 40(7): 11–15, July 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-7/p11-berghel/.

Press:1997:PCTb

[1326] Larry Press. Personal computing: Toward an access rights policy. Communications of the ACM, 40(7):17– 21, July 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-7/p17-press/.

Talin:1997:VPN

[1327] Talin. Viewpoint: The power of negative thinking. Communications of the ACM, 40(7):25–27, July 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-7/p25-talin/.

Denning:1997:ADL

[1328] Peter J. Denning. The ACM digital library goes live. Communications of the ACM, 40(7):28–29, July 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-7/p28-denning/.

Fischer:1997:EPS

[1329] Olivier Fischer and Richard Horn.

Electronic performance support systems. Communications of the ACM, 40(7):31–32, July 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/1997-40-7/p31-fischer/.

Karat:1997:ESU

[1330] John Karat. Evolving the scope of user-centered design. Communications of the ACM, 40(7):33–38, July 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-7/p33-karat/.

Desmarais:1997:CJE

[1331] Michel C. Desmarais, Richard Leclair, Jean-Yves Fiset, and Hichem Talbi. Cost-justifying electronic performance support systems. Communications of the ACM, 40(7):39–48, July 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-7/p39-desmarais/.

Cole:1997:JTK

[1332] Kevin Cole, Olivier Fischer, and Phyllis Saltzman. Just-in-time knowledge delivery. Communications of the ACM, 40(7):49–53, July 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-7/p49-cole/.

Gery:1997:GTW

[1333] Gloria Gery. Granting three wishes through performance-centered design. Communications of the ACM, 40(7): 54–59, July 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-7/p54-gery/.

Benko:1997:PEP

[1334] Shelia Benko and Shelley Webster. Preparing for EPSS projects. Communications of the ACM, 40(7):60– 63, July 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-7/p60-benko/.

Chatterjee:1997:RSG

[1335] Samir Chatterjee. Requirements for success in gigabit networking. Communications of the ACM, 40(7):64– 73, July 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-7/p64-chatterjee/.

Patrick:1997:MLN

[1336] Andrew S. Patrick. Media lessons from the National Capital FreeNet. Communications of the ACM, 40(7):74– 80, July 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-7/p74-patrick/.

Simonsen:1997:UEC

[1337] Jesper Simonsen and Finn Kensing. Using ethnography in contextual design. Communications of the ACM, 40(7):82–88, July 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/1997-
40-7/p82-simonsen/.

Rai:1997:TIB

[1338] Arun Rai, Ravi Patnayakuni, and Nainika Patnayakuni. Technology investment and business performance. Communications of the ACM, 40(7): 89–97, July 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-7/p89-rai/.

Talia:1997:TOP

[1339] Domenico Talia. Technical opinion: Parallel computation still not ready for the mainstream. Communications of the ACM, 40(7):98– 99, July 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-7/p98-talia/.

Neumann:1997:IRIa

[1340] Peter G. Neumann. Inside risks:

Identity-related misuse. Communications of the ACM, 40(7):112, July 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-7/p112-neumann/.

Meeks:1997:WPL

[1341] Brock N. Meeks. From Washington:

Privacy lost, anytime, anywhere. Communications of the ACM, 40(8):11– 13, August 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-8/p11-meeks/.

Foster:1997:IPW

[1342] Will A. Foster, Anthony M. Rutkowski, and Seymour E. Goodman. International perspectives: Who governs the Internet? Communications of the ACM, 40(8):15–20, August 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-8/p15-foster/.

Seirawan:1997:VKV

[1343] Yasser Seirawan. Viewpoint: Kasparov vs. Deep Blue: The aftermath still no match for the human brain. Communications of the ACM, 40(8):21– 22, August 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-8/p21-seirawan/.

Simon:1997:VKV

[1344] Herbert A. Simon and Toshinori Munakata. Viewpoint: Kasparov vs. Deep Blue: AI lessons. Communications of the ACM, 40(8):21, 23–25, August 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-8/p21-seirawan/.

Fox:1997:SSA

[1345] Robert Fox. On site: Socializing around arcade technology. Communications of the ACM, 40(8):26–28, August 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-8/p26-fox/.

Strickland:1997:VHC

[1346] Dorothy Strickland. VR and health care. Communications of the ACM, 40(8):32, August 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/1997-40-8/p32-strickland/.

Strickland:1997:OPV

[1347] Dorothy Strickland, Larry Hodges, Max North, and Suzanne Weghorst. Overcoming phobias by virtual exposure. Communications of the ACM, 40(8):34–39, August 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/1997-40-8/p34-strickland/.

Viire:1997:HSI

[1348] Erik Viire. Health and safety issues for VR. Communications of the ACM, 40(8):40–41, August 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/1997-
40-8/p40-viire/.

Greenleaf:1997:AVP

[1349] Walter J. Greenleaf. Applying VR to physical medicine and rehabilitation. Communications of the ACM, 40(8):42–46, August 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/1997-40-8/p42-greenleaf/.

Weghorst:1997:ARP

[1350] Suzanne Weghorst. Augmented reality and Parkinson’s disease. Communications of the ACM, 40(8):47– 48, August 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-8/p47-weghorst/.

Wann:1997:REA

[1351] John P. Wann, Simon K. Rushton, Martin Smyth, and David Jones. Rehabilitative environments for attention and movement disorders. Communications of the ACM, 40(8):49– 52, August 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-8/p49-wann/.

Inman:1997:VER

[1352] Dean P. Inman, Ken Loge, and John Leavens. VR education and rehabilitation. Communications of the ACM, 40(8):53–58, August 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-8/p53-inman/.

Bowman:1997:VMP

[1353] Tim Bowman. VR meets physical therapy. Communications of the ACM, 40(8):59–60, August 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/1997-40-8/p59-bowman/.

Zeltzer:1997:RVE

[1354] David Zeltzer and Rita K. Addison. Responsive virtual environments. Communications of the ACM, 40(8):61– 64, August 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-8/p61-zeltzer/.

Stanney:1997:PC

[1355] Kay M. Stanney and Robert S.

Kennedy. The psychometrics of cybersickness. Communications of the ACM, 40(8):66–68, August 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/1997-
40-8/p66-stanney/.

Riva:1997:TBI

[1356] Giuseppe Riva, Luca Melis, and Mirco Bolzoni. Treating bodyimage disturbances. Communications of the ACM, 40(8):69–71, August 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-8/p69-riva/.

Brown:1997:VES

[1357] David J. Brown, Steven Kerr, and John R. Wilson. Virtual environments in special-needs education. Communications of the ACM, 40(8):72– 75, August 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-8/p72-brown/.

Foreman:1997:VSA

[1358] Nigel Foreman, Paul Wilson, and Danae Stanton. VR and spatial awareness in disabled children. Communications of the ACM, 40(8):76–77, August 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-8/p76-foreman/.

Raghupathi:1997:HCI

[1359] W. Raghupathi. Health care information systems. Communications of the ACM, 40(8):80–82, August 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-8/p80-raghupathi/.

Anderson:1997:CWP

[1360] James G. Anderson. Clearing the way for physicians’ use of clinical information systems. Communications of the ACM, 40(8):83–90, August 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-8/p83-anderson/.

Rindfleisch:1997:PIT

[1361] Thomas C. Rindfleisch. Privacy, information technology, and health care. Communications of the ACM, 40(8):93–100, August 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/1997-40-8/p93-rindfleisch/.

Detmer:1997:UII

[1362] William M. Detmer and Edward H.

Shortliffe. Using the Internet to improve knowledge diffusion in medicine. Communications of the ACM, 40(8): 101–108, August 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/1997-40-8/p101-detmer/.

Kilman:1997:ICB

[1363] David G. Kilman and David W.

Forslund. An international collaboratory based on virtual patient records. Communications of the ACM, 40(8):110–117, August 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-8/p110-kilman/.

Noll:1997:TOI

[1364] A. Michael Noll. Technical opinion:

Internet pricing vs. reality. Communications of the ACM, 40(8):118–121, August 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-8/p118-noll/.

Neumann:1997:IRC

[1365] Peter G. Neumann. Inside risks:

Crypto key management. Communications of the ACM, 40(8):136, August 1997. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-8/p136-neumann/.

Anonymous:1997:AECb

[1366] Anonymous. ACM Executive Committee meets in Stockholm. Communications of the ACM, 40(9):3–??, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:1997:VAD

[1367] Anonymous. Visit ACM’s Digital Library Open House. Communications of the ACM, 40(9):5–??, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:1997:GMH

[1368] Anonymous. The Grace Murray Hopper celebration will be held in September. Communications of the ACM, 40 (9):6–??, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:1997:AWP

[1369] Anonymous. ACM will provide professional liability insurance. Communications of the ACM, 40(9):7–??, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:1997:NTa

[1370] Robert Fox. News track. Communications of the ACM, 40(9):9–10, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-9/p9-fox/.

Glass:1997:PPC

[1371] Robert L. Glass. Practical programmer: Cobol: a contradiction and an enigma. Communications of the ACM, 40(9):11–13, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-9/p11-glass/.

Rada:1997:SSE

[1372] Roy Rada and James R. Schoening. Sharing standards: Educational technology standards. Communications of the ACM, 40(9):15–18, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-9/p15-rada/.

Anonymous:1997:JHN

[1373] Anonymous. Joe Halpern is the new Editor-in-Chief of Journal of the ACM. Communications of the ACM, 40(9): 15–??, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:1997:FPI

[1374] Diane Crawford. Forum: The power of interactive computing. Communications of the ACM, 40(9):20–23, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-9/p20-crawford/.

Munro:1997:VDW

[1375] Neil Munro and Brock N. Meeks. Viewpoint: Debating (what once was) the CDA. for: Neil Munro; against: Brock N. Meeks. Communications of the ACM, 40(9):25–28, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-9/p25-munro/.

Mowshowitz:1997:VOV

[1376] Abbe Mowshowitz. Virtual organization: a virtually organized company dynamically links its business goals with the procedures needed to achieve them. Communications of the ACM, 40(9):30–37, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-9/p30-mowshowitz/.

Turoff:1997:VAU

[1377] Murray Turoff. Virtuality: Allowing users to negotiate an agreedupon reality, virtuality requires system designers to anticipate how their technology alters the social order of the real world. Communications of the ACM, 40(9):38–43, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-9/p38-turoff/.

Hiltz:1997:ALN

[1378] Starr Roxanne Hiltz and Barry Wellman. Asynchronous learning networks as a virtual classroom. Communications of the ACM, 40(9):44–49, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-9/p44-hiltz/.

Faucheux:1997:HVO

[1379] Claude Faucheux. How virtual organizing is transforming management science. Communications of the ACM, 40(9):50–55, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-9/p50-faucheux/.

Chellappa:1997:EIV

[1380] Ramnath Chellappa, Anitesh Barua, and Andrew B. Whinston. An electronic infrastructure for a virtual university. Communications of the ACM, 40(9):56–58, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-9/p56-chellappa/.

Hardwick:1997:IVE

[1381] Martin Hardwick and Richard Bolton.

The industrial virtual enterprise. Communications of the ACM, 40(9):59–60, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-9/p59-hardwick/.

Mannoni:1997:VM

[1382] Bruno Mannoni. A virtual museum. Communications of the ACM, 40(9): 61–62, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

                                        www.acm.org/pubs/citations/journals/
                                        cacm/1997-40-9/p61-mannoni/.

Zajtchuk:1997:MAV

[1383] Russ Zajtchuk and Richard M. Satava. Medical applications of virtual reality. Communications of the ACM, 40(9):63–64, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-9/p63-zajtchuk/.

Cooperstock:1997:RE

[1384] Jeremy R. Cooperstock, Sidney S. Fels, William Buxton, and Kenneth C. Smith. Reactive environments. Communications of the ACM, 40(9):65–73, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-9/p65-cooperstock/.

Ewusi-Mensah:1997:CIA

[1385] Kweku Ewusi-Mensah. Critical issues in abandoned information systems development projects. Communications of the ACM, 40(9):74–80, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-9/p74-ewusi-mensah/

.

Rein:1997:CDM

[1386] Gail L. Rein, Daniel L. McCue, and Judith A. Slein. A case for document management functions on the Web. Communications of the ACM, 40(9):81–89, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-9/p81-rein/.

Maisel:1997:ESR

[1387] Herbert Maisel and Catherine Gaddy.

Employment and salaries of recent doctorates in computer science. Communications of the ACM, 40(9):90–93, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-9/p90-maisel/.

Pearson:1997:MIE

[1388] J. Michael Pearson, Leon Crosby, and J. P. Shim. Measuring the importance of ethical behavior criteria. Communications of the ACM, 40(9):94–100, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-9/p94-pearson/.

Fayad:1997:TOS

[1389] Mohamed E. Fayad. Thinking objectively: Software development process: A necessary evil. Communications of the ACM, 40(9):101–103, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-9/p101-fayad/.

Parnas:1997:IRS

[1390] David Lorge Parnas. Inside risks: Software engineering: An unconsummated marriage. Communications of the ACM, 40(9):128, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-9/p128-parnas/.

Samuelson:1997:ETS

[1391] Pamela Samuelson. Embedding technical self-help in licensed software. Communications of the ACM, 40(10):13– 17, October 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-10/p13-samuelson/.

Lai:1997:OOU

[1392] Tung Lai Lai and Efraim Turban. One organization’s use of Lotus Notes. Communications of the ACM, 40(10):19–21, October 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-10/p19-lai/.

Webber:1997:FMD

[1393] David R. R. Webber. Forum: More on debugging. Communications of the ACM, 40(10):23–24, October 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-10/p23-????/.

Hsu:1997:AYP

[1394] Richard C. Hsu and William E.

Mitchell. After 400 years, print is still superior. Communications of the ACM, 40(10):27–28, October 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-10/p27-hsu/.

Fayad:1997:OOA

[1395] Mohamed Fayad and Douglas C.

Schmidt. Object-oriented application frameworks. Communications of the ACM, 40(10):32–38, October 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-10/p32-fayad/.

Johnson:1997:FCP

[1396] Ralph E. Johnson. Frameworks = (components + patterns). Communications of the ACM, 40(10):39–42, October 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-10/p39-johnson/.

Posnak:1997:AFD

[1397] Edward J. Posnak, R. Greg Lavender, and Harrick M. Vin. An adaptive framework for developing multeda software componets. Communications of the ACM, 40(10):43–47, October 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-10/p43-posnak/.

Schmid:1997:SFD

[1398] Han Albrecht Schmid. Systematic framework design. Communications of the ACM, 40(10):48–51, October 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-10/p48-schmid/.

Baumer:1997:FDL

[1399] Dirk Bäumer, Guido Gryczan, Rolf Knoll, Carola Linienthal, Dirk Riehle, and Heinz Züllighoven. Framwork development for large systems. Communications of the ACM, 40(10):53– 59, October 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-10/p53-baumer/.

Demeyer:1997:DGT

[1400] Serge Demeyer, Theo Dirk Meijlar, and Oscar Nierstrasz. Design guidelines for “tailorable” frameworks. Communications of the ACM, 40(10):60–64, October 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-10/p60-demeyer/.

Brugali:1997:FLS

[1401] Davide Brugali, Giusee Menga, and Amund Aarsten. The framework life span. Communications of the ACM, 40(10):65–68, October 1997. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-10/p65-brugali/.

Codenie:1997:CAD

[1402] Wim Codenie, Koen De Hondt, Patrick Steyaert, and Arlette Vercammen. From custom applications to domain-specific framworks. Communications of the ACM, 40(10):71–77, October 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-10/p71-codenie/.

Goldberg:1997:LDD

[1403] Adele Goldberg, Steven T. Abell, and David Leibs. The LearningWorks development and delivery frameworks. Communications of the ACM, 40(10):78–81, October 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-10/p78-goldberg/.

Doscher:1997:SEC

[1404] David Doscher and Robert Hodges.

SEMATECH’s experiences with the CIM framework. Communications of the ACM, 40(10):82–84, October 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-10/p82-doscher/.

Fayad:1997:LLB

[1405] Mohamed E. Fayad and Douglas C.

Schmidt. Lessons learned building reusable OO frameworks for distributed software. Communications of the ACM, 40(10):85–87, October 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-10/p85-fayad/.

Forrest:1997:CI

[1406] Stephane Forrest, Steven A. Hofmeyr, and Anil Somayaji. Computer immunology. Communications of the ACM, 40(10):88–96, October 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/journals/cacm/1997-
                                        40-10/p88-forrest/.

Weintraub:1997:AWW

[1407] Annette Weintraub. Art on the Web, the Web as art. Communications of the ACM, 40(10):97–102, October 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-10/p97-weintraub/.

Camp:1997:ISP

[1408] Tracy Camp. The incredible shrinking pipeline. Communications of the ACM, 40(10):103–110, October 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1997-40-10/p103-camp/.

Brandt:1997:CTU

[1409] D. Scott Brandt. Constructivism: teaching for understanding of the Internet. Communications of the ACM, 40 (10):112–117, October 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/1997-40-10/p112-brandt/.

Ryant:1997:WIM

[1410] Ivan Ryant. Why inheritance means extra trouble. Communications of the ACM, 40(10):118–119, October 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-10/p118-ryant/.

Neumann:1997:IRIb

[1411] Peter G. Neumann. Inside risks: Integrity in software development. Communications of the ACM, 40(10):144, October 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1997-40-10/p144-neumann/.

Anonymous:1997:CMR

[1412] Anonymous. Carver Mead receives the ACM Allen Newell Award. Communications of the ACM, 40(11(S)):2–??, November 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:1997:AOM

[1413] Anonymous. ACM outreaches to Mexico. Communications of the ACM, 40 (11(S)):4–??, November 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:1997:UPP

[1414] Anonymous. USACM participates in protest against restrictions on cryptography research and development. Communications of the ACM, 40(11(S)): 5–??, November 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:1997:DNN

[1415] Anonymous. David Nicol is the new Editor-in-Chief of TOMACS. Communications of the ACM, 40(11(S)):7–??, November 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:1997:NTb

[1416] Robert Fox. News track. Communications of the ACM, 40(11):9–10, November 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1997-40-11/p9-fox/.

Press:1997:PCTc

[1417] Larry Press. Personal computing: Tracking the global diffusion of the Internet. Communications of the ACM, 40(11):11–17, November 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1997-40-
                                        11/p11-press/.

Anonymous:1997:AMV

[1418] Anonymous. ACM Member Value Plus program adds the Gartner Group. Communications of the ACM, 40 (11(S)):16–??, November 1997. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:1997:DVW

[1419] Hal Berghel. Digital village — watermarking cyberspace. Communications of the ACM, 40(11):19–24, November 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1997-40-
11/p19-berghel/.

Crawford:1997:MPC

[1420] Diane Crawford. The mainstreaming of parallel computation. Communications of the ACM, 40(11):25–26, November 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1997-40-
11/p25-crawford/.

Smarr:1997:TC

[1421] Larry Smarr. Toward the 21st century. Communications of the ACM, 40(11):28–32, November 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1997-40-
11/p28-smarr/.

Smith:1997:NPT

[1422] Philip M. Smith. The NSF partnerships and the tradition of U.S. science and engineering. Communications of the ACM, 40(11):34–37, November 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/

citations/journals/cacm/1997-40-
11/p34-smith/.

Reed:1997:DDI

[1423] Daniel A. Reed, Roscoe C. Giles, and Charles E. Catlett. Distributed data and immersive collaboration. Communications of the ACM, 40(11):38–48, November 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1997-40-11/p38-reed/.

Stevens:1997:WNT

[1424] Rick Stevens, Paul Woodward, Tom DeFanti, and Charlie Catlett. From the I-WAY to the National Technology Grid. Communications of the ACM, 40(11):50–60, November 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1997-40-
11/p50-stevens/.

Kennedy:1997:NPC

[1425] Ken Kennedy, Charles F. Bender, John W. D. Connolly, John L. Hennessy, Mary K. Vernon, and Larry Smarr. A nationwide parallel computing environment. Communications of the ACM, 40(11):62–72, November 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1997-40-
11/p62-kennedy/.

McRae:1997:HAD

[1426] Gregory J. McRae. How application domains define requirements for the grid. Communications of the ACM, 40(11):74–83, November 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1997-40-
11/p74-mcrae/.

Ostriker:1997:CEU

[1427] Jeremiah P. Ostriker and Michael L.

Norman. Cosmology of the early universe viewed through the new infrastructure. Communications of the ACM, 40(11):84–94, November 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1997-40-
11/p84-ostriker/.

Romm:1997:TVP

[1428] Celia T. Romm and Nava Pliskin.

Toward a virtual politicking model. Communications of the ACM, 40(11): 95–100, November 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//www.acm.org:80/pubs/citations/
journals/cacm/1997-40-11/p95-romm/

.

Hidding:1997:RMW

[1429] Gezinus J. Hidding. Reinventing methodology: Who reads it and why? Communications of the ACM, 40(11):102–109, November 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1997-40-
11/p102-hidding/.

Gotterbarn:1997:SEC

[1430] Don Gotterbarn, Keith Miller, and Simon Rogerson. Software engineering code of ethics. Communications of the ACM, 40(11):110–118, November 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1997-40-
                                        11/p110-gotterbarn/.

Hardgrave:1997:FMS

[1431] Bill C. Hardgrave and Kent A. Walstrom. Forums for MIS scholars. Communications of the ACM, 40(11): 119–124, November 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        journals/cacm/1997-40-11/p119-hardgrave/

.

Fayad:1997:TOP

[1432] Mohamed E. Fayad and Mauri Laitinen. Thinking objectively

— process assessment considered wasteful. Communications of the ACM, 40(11):125–128, November 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1997-40-
                                        11/p125-fayad/.

Ladkin:1997:IRR

[1433] Peter B. Ladkin. Inside risks — risks of technological remedy. Communications of the ACM, 40(11):160, November 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1997-40-
                                        11/p160-ladkin/.

Fox:1997:NTc

[1434] Robert Fox. News track. Communications of the ACM, 40(12):9–10, December 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1997-40-12/p9-fox/.

Tan:1997:CNI

[1435] Zixiang Alex Tan, Milton Mueller, and Will Foster. China’s new Internet regulations: two steps forward, one step back. Communications of the ACM, 40(12):11–16, December 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1997-40-
12/p11-tan/.

Glass:1997:HBP

[1436] Robert L. Glass. How best to provide the services IS programmers need. Communications of the ACM, 40(12):17–19, December 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1997-40-
12/p17-glass/.

Maiden:1997:LLD

[1437] Neil A. M. Maiden, Cornelious Ncube, and Andrew Moore. Lessons learned during requirements acquisition for COTS systems. Communications of the ACM, 40(12):21–25, December 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1997-40-
12/p21-maiden/.

Crawford:1997:FCQ

[1438] Diane Crawford. Forum: Cobol in question. Communications of the ACM, 40(12):27–29, December 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1997-40-
                                        12/p27-crawford/.

Jain:1997:VIM

[1439] Ramesh Jain. Visual information management. Communications of the ACM, 40(12):30–32, December 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1997-40-
                                        12/p30-jain/.

Gupta:1997:SIV

[1440] Amarnath Gupta, Simone Santini, and Ramesh Jain. In search of information in visual media. Communications of the ACM, 40(12):34–42, December 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1997-40-
                                        12/p34-gupta/.

Yeo:1997:RVV

[1441] Boon-Lock Yeo and Minerva M. Yeung. Retrieving and visualizing video. Communications of the ACM, 40(12):43–52, December 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org:80/pubs/citations/journals/
                                        cacm/1997-40-12/p43-yeo/.

Lienhart:1997:VA

[1442] Rainer Lienhart, Silvia Pfeiffer, and Wolfgang Effelsberg. Video abstracting. Communications of the ACM, 40(12):54–62, December 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1997-40-
12/p54-lienhart/.

Chang:1997:VIR

[1443] Shih-Fu Chang, John R. Smith, Mandis Beigi, and Ana Benitez. Visual information retrieval from large distributed online repositories. Communications of the ACM, 40(12):63–71, December 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1997-40-
12/p63-chang/.

Grosky:1997:MMI

[1444] William I. Grosky. Managing multimedia information in database systems. Communications of the ACM, 40(12):72–80, December 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1997-40-
12/p72-grosky/.

Katz:1997:NS

[1445] James E. Katz and Philip Aspden. A nation of strangers? Communications of the ACM, 40(12):81–86, December 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1997-40-
12/p81-katz/.

Franz:1997:SB

[1446] Michael Franz and Thomas Kistler. Slim binaries. Communications of the ACM, 40(12):87–94, December 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1997-40-
                                        12/p87-franz/.

Fraser:1997:FSM

[1447] Martin D. Fraser and Vijay K. Vaishnavi. A formal specifications maturity model. Communications of the ACM, 40(12):95–103, December 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1997-40-
                                        12/p95-fraser/.

Parsons:1997:UOS

[1448] Jeffrey Parsons and Yair Wand. Using objects for systems analysis. Communications of the ACM, 40(12):104–110, December 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org:80/pubs/citations/journals/
                                        cacm/1997-40-12/p104-parsons/.

Coulter:1997:ACC

[1449] Neal Coulter. ACM’s computing classification system reflects changing times. Communications of the ACM, 40(12):111–112, December 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1997-40-
                                        12/p111-coulter/.

Ward:1997:PD

[1450] Nigel Ward. Programming as a discipline? Communications of the ACM, 40(12):113–114, December 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1997-40-
12/p113-ward/.

Neumann:1997:IRS

[1451] Peter G. Neumann. Inside risks: System development woes. Communications of the ACM, 40(12):160, December 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1997-40-12/p160-neumann/.

Fox:1998:NTa

[1452] Robert Fox. News track. Communications of the ACM, 41(1):9–10, January 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
1/p9-fox/.

Rada:1998:SSCa

[1453] Roy Rada. Sharing standards: Corporate shortcut to standardization. Communications of the ACM, 41(1):11–15, January 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1998-41-1/p11-rada/.

Munro:1998:WAC

[1454] Neil Munro. From Washington: Access charges under scrutiny. Communications of the ACM, 41(1):17–19, January 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
1/p17-munro/.

Sabelli:1998:VWN

[1455] Nora H. Sabelli. Viewpoint: We are no longer a priesthood. Communications of the ACM, 41(1):20–21, January 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
1/p20-sabelli/.

Burg:1998:CAC

[1456] Jennifer J. Burg and Stan J. Thomas. Computers across campus. Communications of the ACM, 41(1):22–25, January 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1998-41-1/p22-burg/.

Brown:1998:SPU

[1457] David G. Brown, Jennifer J. Burg, and Jay L. Dominick. A strategic plan for ubiquitous laptop computing. Communications of the ACM, 41(1):26–35, January 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1998-41-1/p26-brown/.

MacDougall:1998:AUS

[1458] Glenn MacDougall. Acadia University’s “Sandbox”. Communications of the ACM, 41(1):32–33, January 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
1/p32-macdougall/.

Candiotti:1998:CUA

[1459] Alan Candiotti and Neil Clarke. Combining universal access with faculty development and academic facilities. Communications of the ACM, 41(1):36–41, January 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
1/p36-candiotti/.

Weiser:1998:FUC

[1460] Mark Weiser. The future of ubiquitous computing on campus. Communications of the ACM, 41(1):41–42, January 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
1/p41-weiser/.

Biros:1998:CCP

[1461] Janice Biros. Changes in computer purchase, distribution, and maintenance. Communications of the ACM, 41(1):43–44, January 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
1/p43-biros/.

Kiaer:1998:LCI

[1462] Lynn Kiaer, David Mutchler, and Jeffrey Froyd. Laptop computers in an integrated first-year curriculum. Communications of the ACM, 41(1):45–49, January 1998. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org:80/pubs/citations/journals/
                                        cacm/1998-41-1/p45-kiaer/.

Watters:1998:DAU

[1463] Carolyn Watters, Marshall Conley, and Cynthia Alexander. The digital agora: using technology for learning in the social sciences. Communications of the ACM, 41(1):50–57, January 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        1/p50-watters/.

Deden:1998:CSC

[1464] Ann Deden. Computers and systemic change in higher education. Communications of the ACM, 41(1):58–63, January 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1998-41-1/p58-deden/.

LeBlanc:1998:HSC

[1465] Richard J. LeBlanc, Jr. and Steven L. Teal. Hardware and software choices for student computer initiatives. Communications of the ACM, 41(1):64–69, January 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org:80/pubs/citations/journals/
                                        cacm/1998-41-1/p64-leblanc/.

Grier:1998:CD

[1466] Samuel L. Grier and Larry W. Bryant. The case for desktops. Communications of the ACM, 41(1):70–71, January 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
1/p70-grier/.

Lee:1998:DEM

[1467] Ho Guen Lee. Do electronic marketplaces lower the price of goods? Communications of the ACM, 41(1):73–80, January 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1998-41-1/p73-lee/.

Cohen:1998:QJ

[1468] Sonya Freeman Cohen. Quest for Java. Communications of the ACM, 41(1):81–83, January 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
1/p81-cohen/.

Neumann:1998:PI

[1469] Peter G. Neumann. Protecting the infrastructures. Communications of the ACM, 41(1):128, January 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
1/p128-neumann/.

Fox:1998:NTb

[1470] Robert Fox. News track. Communications of the ACM, 41(2):9–11, February 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/

citations/journals/cacm/1998-41-
2/p9-fox/.

Soloway:1998:LEN

[1471] Elliot Soloway. Long on education: no one is making money in education software. Communications of the ACM, 41(2):11–16, February 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
2/p11-soloway/.

Meeks:1998:EFN

[1472] Brock N. Meeks. Electronic frontier: no justice, no peace. Communications of the ACM, 41(2):17–20, February 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
2/p17-meeks/.

Clark:1998:SIC

[1473] Jan Guynes Clark and Vincent S. Lai. On site: Internet comes to Morocco. Communications of the ACM, 41(2):21–26, February 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
2/p21-guynes_clark/.

Crawford:1998:FQT

[1474] Diane Crawford, Larry Brunelle, Morton Grosser, Danielle Bernstein, George Mallard, Louis A. Russ, and Magnus Kempe. Forum: Questioning technical self-helf; women and computer science; title by license; citizens of the republic speak. Communications of the ACM, 41(2):24–26, February 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
2/p24-crawford/.

Behrens:1998:PLA

[1475] Brian C. Behrens and Reuven R.

Levary. Practical legal aspects of software reverse engineering. Communications of the ACM, 41(2):27–29, February 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1998-41-2/p27-behrens/.

Kappelman:1998:CCY

[1476] Leon A. Kappelman, Darla Fent, Kellie B. Keeling, and Victor Prybutok. Calculating the cost of Year-2000 compliance. Communications of the ACM, 41(2):30–39, February 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
2/p30-kappelman/.

Grover:1998:IPC

[1477] Varun Grover, James T. C. Teng, and Kirk D. Fiedler. IS investment priorities in contemporary organizations. Communications of the ACM, 41(2):40–48, February 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
2/p40-grover/.

Johnson:1998:RI

[1478] Philip M. Johnson. Reengineering inspection. Communications of the ACM, 41(2):49–53, February 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        2/p49-johnson/.

Tayi:1998:SSE

[1479] Giri Kumar Tayi and Donald P. Ballou. Special section: examining data quality. Communications of the ACM, 41(2):54–57, February 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        2/p54-kumar_tayi/.

Wang:1998:PPT

[1480] Richard Y. Wang. A product perspective on total data quality management. Communications of the ACM, 41(2):58–65, February 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        2/p58-wang/.

Orr:1998:DQS

[1481] Ken Orr. Data quality and systems theory. Communications of the ACM, 41(2):66–71, February 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        journals/cacm/1998-41-2/p66-orr/

.

Kaplan:1998:ADQ

[1482] David Kaplan and Ramayya Krishnan.

Assessing data quality in accounting information systems. Communications of the ACM, 41(2):72–78, February 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
2/p72-kaplan/.

Redman:1998:IPD

[1483] Thomas C. Redman. The impact of poor data quality on the typical enterprise. Communications of the ACM, 41(2):79–82, February 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
2/p79-redman/.

Laitinen:1998:TOS

[1484] Mauri Laitinen and Mohamed E.

Fayad. Thinking objectively: surviving a process performance crash. Communications of the ACM, 41(2):83–86, February 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1998-41-2/p83-laitinen/.

Neumann:1998:IRI

[1485] Peter G. Neumann. Inside risks: Internet gambling. Communications of the ACM, 41(2):112, February 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
2/p112-neumann/.

Crawford:1998:EPa

[1486] Diane Crawford. Editorial pointers. Communications of the ACM, 41(3):5, March 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1998-41-3/p5-crawford/.

Fox:1998:NTc

[1487] Robert Fox. News track. Communications of the ACM, 41(3):9–11, March 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        3/p9-fox/.

Berghel:1998:DVY

[1488] Hal Berghel. Digital village: the Year-2000 problem and the new riddle of induction. Communications of the ACM, 41(3):13–17, March 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        3/p13-berghel/.

Crawford:1998:AF

[1489] Diane Crawford. ACM fellows. Communications of the ACM, 41(3):18, March 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1998-41-3/p18-crawford/.

Burkhart:1998:IPIa

[1490] Grey E. Burkhart and Seymour E. Goodman. International perspectives: the Internet gains acceptance in the Persian Gulf. Communications of the ACM, 41(3):19–25, March 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
3/p19-burkhart/.

Crawford:1998:FPC

[1491] Diane Crawford, Jimmy A. Castro, Raman Kannan, Mark C. Kilby, Mohamed Fayad, Earl F. Glynn, Tracy Camp, Mike Yaphe, Abbe Mowshowitz, and Domenico Talla. Forum: Praises come with a hitch; OO frameworks; response; more studies needed; response; metamanagement definition; response; parallel computers revisited. Communications of the ACM, 41(3):26–29, March 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1998-41-3/p26-crawford/.

House:1998:TC

[1492] Charles H. House. It’s time for a change. Communications of the ACM, 41(3):31, March 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
3/p31-house/.

Rao:1998:MI

[1493] H. Raghav Rao, A. F. Salam, and Brian DosSantos. Marketing and the Internet. Communications of the ACM, 41(3):32–34, March 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//www.acm.org:80/pubs/citations/

                                        journals/cacm/1998-41-3/p32-rao/

.

Kannan:1998:MIW

[1494] P. K. Kannan, Ai-Mei Chang, and Andrew B. Whinston. Marketing information on the I-Way data junkyard or information gold mine? Communications of the ACM, 41(3):35–43, March 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        3/p35-kannan/.

Palmer:1998:EMW

[1495] Jonathan W. Palmer and David A. Griffith. An emerging model of Web site design for marketing. Communications of the ACM, 41(3):44–51, March 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        3/p44-palmer/.

Eder:1998:PC

[1496] Lauren B. Eder and Marvin E. Darter. Physicians in cyberspace. Communications of the ACM, 41(3):52–54, March 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        3/p52-eder/.

Elofson:1998:CCM

[1497] Greg Elofson and William N. Robinson. Creating a custom mass-production channel on the Internet. Communications of the ACM, 41(3):56–62, March 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
3/p56-elofson/.

Wang:1998:CPC

[1498] Huaiqing Wang, Matthew K. O. Lee, and Chen Wang. Consumer privacy concerns about Internet marketing. Communications of the ACM, 41(3):63–70, March 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
3/p63-wang/.

OKeefe:1998:WBC

[1499] Robert M. O’Keefe and Tim McEachern. Web-based customer decision support systems. Communications of the ACM, 41(3):71–76, March 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
3/p71-o_keefe/.

Salam:1998:CCW

[1500] A. F. Salam, H. R. Rao, and C. C.

Pegels. Content of corporate Web pages as advertising media. Communications of the ACM, 41(3):76–77, March 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
3/p76-salam/.

Bohr:1998:STL

[1501] Mark Bohr. Silicon trends and limits for advanced microprocessors. Communications of the ACM, 41(3):80–87, March 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1998-41-3/p80-bohr/.

Hall:1998:HAU

[1502] Robert J. Hall. How to avoid unwanted email. Communications of the ACM, 41(3):88–95, March 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        3/p88-hall/.

Nardi:1998:CPI

[1503] Bonnie A. Nardi, James R. Miller, and David J. Wright. Collaborative, programmable intelligent agents. Communications of the ACM, 41(3):96–104, March 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1998-41-3/p96-nardi/.

Nosek:1998:TOC

[1504] John T. Nosek. Technical opinion: the case for collaborative programming. Communications of the ACM, 41(3):105–108, March 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        3/p105-nosek/.

Neumann:1998:IRC

[1505] Peter G. Neumann. Inside risks: Are computer addictive? Communications of the ACM, 41(3):128, March 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
3/p128-neumann/.

Fox:1998:NTd

[1506] Robert Fox. News track. Communications of the ACM, 41(4):11–12, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
4/p11-fox/.

Samuelson:1998:ELD

[1507] Pamela Samuelson. Encoding the law into digital libraries. Communications of the ACM, 41(4):13–18, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
4/p13-samuelson/.

Borenstein:1998:WNI

[1508] Nathaniel S. Borenstein. Whose net is it anyway? Communications of the ACM, 41(4):19–20, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
4/p19-borenstein/.

Crawford:1998:FRT

[1509] Diane Crawford. Forum: Readers “thinking objectively”. Communications of the ACM, 41(4):22–27, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
4/p22-crawford/.

Fox:1998:TWD

[1510] Edward A. Fox and Gary Marchionini. Toward a worldwide digital library: introduction. Communications of the ACM, 41(4):28–32, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
4/p28-fox/.

Paepcke:1998:IDL

[1511] Andreas Paepcke, Chen-Chuan K. Chang, Terry Winograd, and Héctor García-Molina. Interoperability for digital libraries worldwide. Communications of the ACM, 41(4):33–42, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
4/p33-paepcke/.

Moen:1998:ADC

[1512] William E. Moen. Accessing distributed cultural heritage information. Communications of the ACM, 41(4): 44–48, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1998-41-4/p44-moen/.

Gladney:1998:DAA

[1513] Henry M. Gladney, Fred Mintzer, Fabio Schiattarella, Julián Bescós, and Martin Treu. Digital access to antiquities. Communications of the ACM, 41(4): 49–57, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1998-41-4/p49-gladney/.

Dippo:1998:FPS

[1514] Cathryn S. Dippo. FedStats promotes statistical literacy. Communications of the ACM, 41(4):58–60, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1998-41-4/p58-dippo/; http:
//www.fedstats.gov.

Cowan:1998:NRC

[1515] D. D. Cowan, C. I. Mayfield, F. W.

Tompa, and W. Gasparini. New role for community networks. Communications of the ACM, 41(4):61–63, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
4/p61-cowan/.

Maeda:1998:VMD

[1516] Akira Maeda, Myriam Dartois, Takehisa Fujita, Tetsuo Sakaguchi, Shigeo Sugimoto, and Koichi Tabata. Viewing multilingual documents on your local Web browser. Communications of the ACM, 41(4):64–65, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
4/p64-maeda/.

Leong:1998:DCB

[1517] M. K. Leong, L. Cao, and Y. Lu.

Distributed Chinese bibliographic searching. Communications of the ACM, 41(4):66–68, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        4/p66-leong/.

Klavans:1998:NEM

[1518] Judith L. Klavans and Peter Scha`‘uble. NSF-EU multilingual information access. Communications of the ACM, 41(4):69–70, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        journals/cacm/1998-41-4/p69-klavans/

.

Witten:1998:PLB

[1519] Ian H. Witten, Craig Nevill-Manning, Rodger McNab, and Sally Jo Cunningham. A public library based on full-text retrieval. Communications of the ACM, 41(4):71–75, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        4/p71-witten/.

Endres:1998:SAB

[1520] Albert Endres and Norbert Fuhr. Students access books and journals through MeDoc. Communications of the ACM, 41(4):76–77, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        4/p76-endres/.

Kovacs:1998:DRW

[1521] László Kovács. Discovery of resources within a distributed library system. Communications of the ACM, 41(4): 78–79, April 1998. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1998-41-4/p78-kovacs/.

Goncalves:1998:ICS

[1522] Marcos André Gonçalves and Claudia Bauzer Medieros. Initiatives that center on scientific dissemination. Communications of the ACM, 41(4): 80–81, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1998-41-4/p80-gonccedilalves/

.

Myaeng:1998:RNG

[1523] Sung Hyon Myaeng. R&D for a nationwide general-purpose system. Communications of the ACM, 41(4):82–85, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1998-41-4/p82-myaeng/.

McKnight:1998:MPD

[1524] Cliff McKnight. Many projects that depend on collaboration. Communications of the ACM, 41(4):86–87, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
4/p86-mcknight/.

Treloar:1998:LNR

[1525] Andrew E. Treloar. Libraries’ new role in electronic scholarly publishing. Communications of the ACM, 41(4): 88–89, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org:80/pubs/citations/journals/
                                        cacm/1998-41-4/p88-treloar/.

Pejtersen:1998:SIR

[1526] Annelise Mark Pejtersen. Semantic information retrieval. Communications of the ACM, 41(4):90–92, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        4/p90-pejtersen/.

Talja:1998:DAU

[1527] Sanna Talja, Rami Heinisuo, Eija-Liisa Kasesniemi, Harri Kemainen, Sinikka Luukkainen, Kirsi Pispa, and Kalervo Järvelin. Discourse analysis of user requests. Communications of the ACM, 41(4):93–94, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        4/p93-talja/.

Shneiderman:1998:SSU

[1528] Ben Shneiderman, Donald Byrd, and W. Burce Croft. Sorting out searching: a user-interface framework for text searches. Communications of the ACM, 41(4):95–98, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        4/p95-shneiderman/.

Vessey:1998:SVW

[1529] Iris Vessey and Robert Glass. Strong vs. weak approaches to systems development. Communications of the ACM, 41(4):99–102, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
4/p99-vessey/.

Sodan:1998:YYC

[1530] A. C. Sodan. Yin and Yang in computer science. Communications of the ACM, 41(4):103–114, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
4/p103-sodan/.

Schneider:1998:CP

[1531] Fred B. Schneider. On concurrent programming. Communications of the ACM, 41(4):128, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
4/p128-schneider/.

Fox:1998:NTe

[1532] Robert Fox. News track. Communications of the ACM, 41(5):9–11, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
5/p9-fox/.

Meeks:1998:DSM

[1533] Brock N. Meeks. Don’t shoot me, I’m famous. Communications of the ACM, 41(5):13–15, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        5/p13-meeks/.

Glass:1998:EON

[1534] Robert Glass. Everything old is new again. Communications of the ACM, 41(5):17–20, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        5/p17-glass/.

Crawford:1998:FWH

[1535] Diane Crawford. Flogging the wrong house? Communications of the ACM, 41(5):22–27, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        5/p22-crawford/.

Crawford:1998:VE

[1536] Diane Crawford. Virtual extension. Communications of the ACM, 41(5):28, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1998-41-5/p28-crawford/.

Barnes:1998:ACT

[1537] Bruce H. Barnes. Assessing computer technology in Europe. Communications of the ACM, 41(5):29–32, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        5/p29-barnes/.

Veitch:1998:DDS

[1538] James Veitch and Robert Laddaga.

Distributed dynamic systems. Communications of the ACM, 41(5):34–36, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1998-41-5/p34-veitch/.

Singh:1998:UHI

[1539] Narinder Singh. Unifying heterogeneous information models: semantic tags support knowledge webs. Communications of the ACM, 41(5):37–44, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1998-41-5/p37-singh/.

Morrill:1998:DRP

[1540] Jeffrey Morrill. Distributed recognition of patterns in time series data: from sensors to informed decisions. Communications of the ACM, 41(5):45–51, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1998-41-5/p45-morrill/.

Veitch:1998:CPG

[1541] Jim Veitch. A conversation with Paul Graham. Communications of the ACM, 41(5):52–54, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
5/p52-veitch/.

Gray:1998:MLM

[1542] David N. Gray, John Hotchkiss, Seth LaForge, Andrew Shalit, and Toby Weinberg. Modern languages and Microsoft’s component object model: programming COM made simple. Communications of the ACM, 41(5):55–65, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1998-41-5/p55-gray/.

Sztipanovits:1998:SAS

[1543] Janos Sztipanovits, Gabor Karsai, and Ted Bapty. Self-adaptive software for signal processing: evolving systems in changing environments without growing pains. Communications of the ACM, 41(5):66–73, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
5/p66-sztipanovits/.

Hardman:1998:SMA

[1544] Vicky Hardman, Martina Angela Sasse, and Isidor Kouvelas. Successful multiparty audio communication over the Internet. Communications of the ACM, 41(5):74–80, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
5/p74-hardman/.

Onoma:1998:RTI

[1545] Akira K. Onoma, Wei-Tek Tsai, Mustafa Poonawala, and Hiroshi Suganuma. Regressing testing in an industrial environment: progress is attained by looking backward. Communications of the ACM, 41(5):81–86, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
5/p81-onoma/.

Duchessi:1998:CSB

[1546] Deter Duchessi and InduShobha Chengalur-Smith. Client/server benefits, problems, best practices: the goal: not just better technology but better business processes that make a company more competitive. Communications of the ACM, 41(5):87–94, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
5/p87-duchessi/.

Robinson:1998:SNL

[1547] William N. Robinson and Vecheslav Volkov. Supporting the negotiation life cycle. Communications of the ACM, 41(5):95–102, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
5/p95-robinson/.

Fayad:1998:AMM

[1548] Mohamed E. Fayad. The art of managing multiple processes. Communications of the ACM, 41(5):103–105, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
5/p103-fayad/.

Mercuri:1998:SAI

[1549] Rebecca Mercuri. In search of academic integrity. Communications of the ACM, 41(5):136, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        5/p136-mercuri/.

Grier:1998:PCC

[1550] David Alan Grier. Politics, control, and computer networks: the Chinese Student Lobby of 1989. Communications of the ACM, 41(5es):137–145, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        5es/p137-grier/.

Keable:1998:LGB

[1551] Michel Keable, Maurice Landry, and Claude Banville. The legitimacy gap between IS professionals and users. Communications of the ACM, 41 (5es):146–156, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        journals/cacm/1998-41-5es/p146-
                                        keable/.

Angehrn:1998:CBD

[1552] Albert A. Angehrn and Soumitra Dutta. Case-based decision support. Communications of the ACM, 41(5es):157–165, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        journals/cacm/1998-41-5es/p157-
                                        angehrn/.

Parent:1998:IAD

[1553] Christine Parent and Stefano Spaccapietra. Issues and approaches of database integration. Communications of the ACM, 41(5es):166–178, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
5es/p166-parent/.

Gupta:1998:ATT

[1554] Uma G. Gupta. Automatic tools for testing expert systems. Communications of the ACM, 41(5es):179–184, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1998-41-5es/p179-gupta/.

Joiner:1998:REL

[1555] J. K. Joiner and W. T. Tsai. Reengineering legacy Cobol programs. Communications of the ACM, 41 (5es):185–197, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//www.acm.org:80/pubs/citations/
journals/cacm/1998-41-5es/p185-
joiner/.

Nishiyama:1998:SAS

[1556] Takashi Nishiyama, Junji Nomura, Ryoji Nakajima, and Tetsuji Miyanoh. A semi-autonomous security robot employing telepresence. Communications of the ACM, 41(5es):198–205, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
5es/p198-nishiyama/.

Sen:1998:DDT

[1557] Arun Sen. From DSS to DSP: a taxonomic retrospective. Communications of the ACM, 41(5es):206–216, May 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        5es/p206-sen/.

Fox:1998:NTf

[1558] Robert Fox. News track. Communications of the ACM, 41(6):9–10, June 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        10/p9-fox/.

Lukasik:1998:PIE

[1559] Stephen J. Lukasik, Lawrence T. Greenberg, and Seymour E. Goodman. Protecting an invaluable and everwidening infrastructure. Communications of the ACM, 41(6):11–16, June 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        6/p11-lukasik/.

Press:1998:LCE

[1560] Larry Press. Low-cost estimation of travel trade-offs. Communications of the ACM, 41(6):17–20, June 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        6/p17-press/.

Carlisle:1998:WNE

[1561] Paul Carlisle, Mike Anderson, Ray Brown, James Mitchell, Anthony Ralston, Herb Grosch, Stewart A. Denenberg, and Rainer Schoenrank. Who needs educational software? Communications of the ACM, 41(6):22–26, June 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
articles/journals/cacm/1998-41-
6/p22-crawford/p22-crawford.pdf;
http://www.acm.org:80/pubs/citations/
journals/cacm/1998-41-6/p22-carlisle/
;   http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
6/p22-crawford/.

Madsen:1998:IMW

[1562] Wayne Madsen. Internet malcontents of the world — unite! Communications of the ACM, 41(6):27–28, June 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
6/p27-madsen/.

Morgan:1998:BIB

[1563] Christopher Morgan. Building an informational bridge to China. Communications of the ACM, 41(6):29–30, June 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
6/p29-morgen/.

Crawford:1998:IA

[1564] Diane Crawford, Thomas Lambert, Robert Fox, and Andrew Rosenbloom.

Information for authors. Communications of the ACM, 41(6):31–33, June 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
6/p31-crawford/.

Singhal:1998:JFI

[1565] Sandeep Singhal and Binh Nguyen. The Java factor: Introduction. Communications of the ACM, 41(6):34–37, June 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1998-41-6/p34-singal/.

Tyma:1998:WWU

[1566] Paul Tyma. Why are we using Java again? Communications of the ACM, 41(6):38–42, June 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
6/p38-tyma/.

Bohrer:1998:BPC

[1567] Kathy Bohrer, Verlyn Johnson, Anders Nilsson, and Bradley Rubin. Business process components for distributed object applications. Communications of the ACM, 41(6):43–48, June 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1998-41-6/p43-bohrer/.

Nilsen:1998:ART

[1568] Kelvin Nilsen. Adding real-time capabilities to Java. Communications of the ACM, 41(6):49–56, June 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
6/p49-nilsen/.

Brutzman:1998:VRM

[1569] Don Brutzman. The Virtual Reality Modeling Language and Java. Communications of the ACM, 41(6):57–64, June 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1998-41-6/p57-brutzman/.

Cooper:1998:UDP

[1570] James W. Cooper. Using design patterns. Communications of the ACM, 41(6):65–68, June 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
6/p65-cooper/.

Chabert:1998:JOS

[1571] Annie Chabert, Ed Grossman, Larry S.

Jackson, Stephen R. Pietrowiz, and Chris Seguin. Java object-sharing in Habanero. Communications of the ACM, 41(6):69–76, June 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
6/p69-chabert/.

Dutta:1998:BES

[1572] Soumitra Dutta, Luk N. Van Wassenhove, and Selvan Kulandaiswamy. Benchmarking European software management practices. Communications of the ACM, 41(6):77–86, June 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        6/p77-dutta/.

Wos:1998:POF

[1573] Larry Wos. Programs that offer fast, flowless, logical reasoning. Communications of the ACM, 41(6):87–95, June 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        6/p87-wos/.

Teng:1998:PSR

[1574] James T. C. Teng, Seung Ryul Jeong, and Varun Grover. Profiling successful reengineering projects. Communications of the ACM, 41(6):96–102, June 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        6/p96-teng/.

Lawson:1998:IRR

[1575] Harold W. Lawson. Infrastructure risk reduction. Communications of the ACM, 41(6):120, June 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        6/p120-lawson/.

Crawford:1998:EPb

[1576] Diane Crawford. Editorial pointers. Communications of the ACM, 41(7):5, July 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1998-41-7/p5-crawford/.

Fox:1998:NTg

[1577] Robert Fox. News track. Communications of the ACM, 41(7):9–11, July 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
7/p9-fox/.

Glass:1998:PPS

[1578] Robert L. Glass. Practical programmer: Short-term and long-term remedies for runaway projects. Communications of the ACM, 41(7):13–15, July 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
7/p13-glass/.

Rada:1998:SSCb

[1579] Roy Rada, Carl Cargill, and John Klensin. Sharing standards: Consensus and the Web. Communications of the ACM, 41(7):17–22, July 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
7/p17-rada/.

Schank:1998:VHC

[1580] Roger C. Schank. Viewpoint: Horses for courses. Communications of the ACM, 41(7):23–25, July 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        7/p23-schank/.

Crawford:1998:FHA

[1581] Diane Crawford. Forum: Hindsight is not always 20/20. Communications of the ACM, 41(7):26–28, July 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        7/p26-crawford/.

Yeung:1998:DW

[1582] Minerva M. Yeung. Digital watermarking. Communications of the ACM, 41(7):30–33, July 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        7/p30-yeung/.

Memon:1998:PDM

[1583] Nasir Memon and Ping Wah Wong. Protecting digital media content. Communications of the ACM, 41(7):35–43, July 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1998-41-7/p35-memon/.

Craver:1998:TTL

[1584] Scott Craver, Boon-Lock Yeo, and Minerva Yeung. Technical trials and legal tribulations. Communications of the ACM, 41(7):45–54, July 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/

citations/journals/cacm/1998-41-
7/p45-craver/.

Mintzer:1998:OWS

[1585] Fred Mintzer, Gordon W. Braudaway, and Alan E. Bell. Opportunities for watermarking standards. Communications of the ACM, 41(7):57–64, July 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
7/p57-mintzer/.

Zhao:1998:BTT

[1586] Jian Zhao, Eckhard Koch, and Chenghui Luo. In business today and tomorrow. Communications of the ACM, 41(7):66–72, 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/1998-
41-7/p67-zhao/.

Acken:1998:HWA

[1587] John M. Acken. How watermarking adds value to digital content. Communications of the ACM, 41(7):75–77, July 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1998-41-7/p75-acken/.

Isakowitz:1998:WIS

[1588] Tomás Isakowitz, Michael Bieber, and Fabio Vitali. Web information systems. Communications of the ACM, 41(7): 78–80, July 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org:80/pubs/citations/journals/
                                        cacm/1998-41-7/p78-isakowitz/.

Lohse:1998:ES

[1589] Gerald L. Lohse and Peter Spiller. Electronic shopping. Communications of the ACM, 41(7):81–87, July 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        7/p81-lohse/.

Tenenbaum:1998:WEC

[1590] Jay M. Tenenbaum. WISs and electronic commerce. Communications of the ACM, 41(7):89–90, July 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        7/p89-tenenbaum/.

Kambil:1998:PAW

[1591] Ajit Kambil and Mark Ginsberg. Public access Web information systems: Lessons from the Internet EDGAR project. Communications of the ACM, 41(7):91–97, July 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        7/p91-kambil/.

Lederer:1998:UWE

[1592] Albert L. Lederer, Dinesh A. Mirchandani, and Kenneth Sims. Using WISs to enhance competitiveness. Communications of the ACM, 41(7):94–95, 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Oinas-Kukkonen:1998:WIL

[1593] Harri Oinas-Kukkonen. What is inside a link? Communications of the ACM, 41(7):98, July 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
7/p98-oinas-kukkonen/.

vanHeck:1998:HSC

[1594] Eric van Heck and Peter Vervest. How should CIOs deal with Web-based auctions? Communications of the ACM, 41(7):99–100, July 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//www.acm.org:80/pubs/citations/
journals/cacm/1998-41-7/p99-van_
heck/.

Arocena:1998:VWD

[1595] Gustavo Arocena and Alberto Mendelzon. Viewing WISs as database applications. Communications of the ACM, 41(7):101–102, July 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
7/p101-arocena/.

Takahashi:1998:MLM

[1596] Kenji Takahashi. Metalevel links: More power to your links. Communications of the ACM, 41(7):103–105, July 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
7/p103-takahashi/.

Balasubramanian:1998:DMW

[1597] V. Balasubramanian and Alf Bashian. Document management and Web technologies. Communications of the ACM, 41(7):107–115, July 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        7/p107-balasubramanian/.

Dennis:1998:LTY

[1598] Alan R. Dennis. Lessons from three years of Web development. Communications of the ACM, 41(7):112– 113, 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1998-41-7/p112-dennis/.

Maurer:1998:MW

[1599] Hermann A. Maurer. Modern WISs. Communications of the ACM, 41(7): 114–115, 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/1998-41-7/p114-maurer/.

Turoff:1998:S

[1600] Murray Turoff and Starr Roxanne Hiltz. Superconnectivity. Communications of the ACM, 41(7):116, July 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        7/p116-turoff/.

Senn:1998:WFE

[1601] James A. Senn. WISs at Federal Express. Communications of the ACM, 41(7):117–118, July 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
7/p117-senn/.

Neumann:1998:IRL

[1602] Peter G. Neumann. Inside risks:

Laptops in Congress? Communications of the ACM, 41(7):128, July 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
7/p128-neumann/.

Crawford:1998:EPc

[1603] Diane Crawford. Editorial pointers.

Communications of the ACM, 41(8):5, August 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/articles/journals/cacm/
1998-41-9/p128-neumann/p128-neumann.
pdf;   http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
8/p5-crawford/.

Fox:1998:NTh

[1604] Robert Fox. News track. Communications of the ACM, 41(8):9–10, August 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
8/p9-fox/.

Soloway:1998:LEU

[1605] Elliot Soloway and Cathie Norris. Log on education: using technology to address old problems in new ways. Communications of the ACM, 41(8):11–18, August 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1998-41-8/p11-soloway/.

Crawford:1998:WSP

[1606] Diane Crawford. From Washington: shaping public policy in the information age. Communications of the ACM, 41(8):21–24, August 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        8/p21-crawford/.

Coldwell:1998:VDC

[1607] Roger A. Coldwell. Viewpoint: did Chuck Babbage predict software piracy? Communications of the ACM, 41(8):25–27, August 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        8/p25-coldwelll/.

Crawford:1998:ONC

[1608] Diane Crawford. Only the names have changed. Communications of the ACM, 41(8):28–31, August 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        8/p28-crawford/.

Kauffman:1998:ISE

[1609] Robert J. Kauffman and Frederick J. Riggins. Information systems and economics. Communications of the ACM, 41(8):32–34, August 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
8/p32-kauffman/.

Bakos:1998:ERE

[1610] Yannis Bakos. The emerging role of electronic marketplaces on the Internet. Communications of the ACM, 41(8):35–42, August 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
8/p35-bakos/.

Dewan:1998:ISP

[1611] Rajiv M. Dewan, Marshall L. Freimer, and Abraham Seidmann. Internet service providers, proprietary content, and the battle for users’ dollars. Communications of the ACM, 41(8):43–48, August 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1998-41-8/p43-dewan/.

Brynjolfsson:1998:BPP

[1612] Erik Brynjolfsson and Lorin Hitt. Beyond the productivity paradox. Communications of the ACM, 41(8):49–55, August 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1998-41-8/p49-brynjolfsson/.

Dewan:1998:IDP

[1613] Sanjeev Dewan and Kenneth L. Kraemer. International dimensions of the productivity paradox. Communications of the ACM, 41(8):56–62, August 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
8/p56-dewan/.

Kemerer:1998:POO

[1614] Chris F. Kemerer. Progress, obstacles, and opportunities in software engineering economics. Communications of the ACM, 41(8):63–66, August 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
8/p63-kemerer/.

Slaughter:1998:ECS

[1615] Sandra A. Slaughter, Donald E. Harter, and Mayuram S. Krishnan. Evaluating the cost of software quality. Communications of the ACM, 41(8):67–73, August 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1998-41-8/p67-slaughter/.

Cranor:1998:S

[1616] Lorre Faith Cranor and Brian A. LaMacchia. Spam! Communications of the ACM, 41(8):74–83, August 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
8/p74-cranor/.

Fielding:1998:WBD

[1617] Roy T. Fielding, E. James Whitehead, Jr., Kenneth M. Anderson, Gregory A.

Bolcer, Peyman Oreizy, and Richard N. Taylor. Web-based development of complex information products. Communications of the ACM, 41(8):84–92, August 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1998-41-8/p84-fielding/.

Jarzabek:1998:CUC

[1618] Stan Jarzabek and Riri Huang.

The case for user-centered CASE tools. Communications of the ACM, 41(8):93–99, August 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
8/p93-jarzabek/.

Purao:1998:EDO

[1619] Sandeep Purao, Hemant Jain, and Derek Narareth. Effective distribution of object-oriented applications. Communications of the ACM, 41(8):100– 108, August 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1998-41-8/p100-purao/.

Hamu:1998:TOA

[1620] David S. Hamu and Mohamed E.

Fayad. Thinking objectively: achieving bottom-line improvements with enterprise frameworks. Communications of the ACM, 41(8):110–113, August 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
8/p110-hamu/.

Denning:1998:IRC

[1621] Peter J. Denning. Inside risks: Computer science and software engineering: filing for divorce? Communications of the ACM, 41(8):128, August 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        8/p128-denning/.

Fox:1998:NTi

[1622] Robert Fox. News track. Communications of the ACM, 41(9):9–10, September 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1998-41-9/p9-fox/.

Meeks:1998:EFM

[1623] Brock N. Meeks. Electronic frontier: mainstream media heartburn. Communications of the ACM, 41(9):11–13, September 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org:80/pubs/citations/journals/
                                        cacm/1998-41-9/p11-meeks/.

Samuelson:1998:LSD

[1624] Pamela Samuelson. Legally speaking: does information really have to be licensed? Communications of the ACM, 41(9):15–20, September 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        9/p15-samuelson/.

Smith:1998:VCC

[1625] Don Smith. Viewpoint: computerizing computer science. Communications of the ACM, 41(9):21–23, September 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
9/p21-smith/.

Crawford:1998:FPY

[1626] Diane Crawford. Forum: peeling the Y2K onion. Communications of the ACM, 41(9):24–27, September 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
9/p24-crawford/.

Sen:1998:ISD

[1627] Arun Sen and Varghese S. Jacob.

Industrial-strength data warehousing. Communications of the ACM, 41(9): 28–31, September 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//www.acm.org:80/pubs/citations/
journals/cacm/1998-41-9/p28-sen/

.

Watson:1998:MC

[1628] Hugh J. Watson and Barbara J. Haley. Managerial considerations. Communications of the ACM, 41(9):32–37, September 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1998-41-9/p32-watson/.

Bontempo:1998:IDW

[1629] Charles Bontempo and George Zagelow. The IBM data warehouse architecture. Communications of the ACM, 41(9):38–48, September 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        9/p38-bontempo/.

Sigal:1998:CSD

[1630] Michael Sigal. A common sense development strategy. Communications of the ACM, 41(9):42–43, 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

                                        //dblp.uni-trier.de/db/journals/
                                        cacm/Sigal98.html.

Sutter:1998:PBW

[1631] James R. Sutter. Project-based warehouses. Communications of the ACM, 41(9):49–51, September 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        9/p49-sutter/.

Gardner:1998:BDW

[1632] Stephen R. Gardner. Building the data warehouse. Communications of the ACM, 41(9):52–60, September 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        9/p52-gardner/.

Claudio:1998:HCM

[1633] Cecilia Claudio. Health care management. Communications of the ACM, 41(9):58–59, September 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
9/p58-claudio/.

Glassey:1998:SEU

[1634] Katherine Glassey. Seducing the end user. Communications of the ACM, 41(9):62–69, September 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
9/p62-glassey/.

Scott:1998:WW

[1635] Jim Scott. Warehousing over the Web. Communications of the ACM, 41(9):64–65, September 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
9/p64-scott/.

Huber:1998:UDM

[1636] Linda Huber. Users drive move from mainframe. Communications of the ACM, 41(9):66–67, September 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
9/p66-huber/.

Kreie:1998:HMW

[1637] Jennifer Kreie and Timothy Paul Cronan. How men and women view ethics. Communications of the ACM, 41(9):70–76, September 1998.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
9/p70-kreie/.

Gal-Ezer:1998:WES

[1638] Judith Gal-Ezer and David Harel. What (else) should CS educators know? Communications of the ACM, 41(9):77–84, September 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//www.acm.org:80/pubs/citations/
journals/cacm/1998-41-9/p77-gal-
ezer/.

Keller:1998:PSN

[1639] Rudolf K. Keller, Jean Tessier, and Gregor von Bochmann. A pattern system for network management interfaces. Communications of the ACM, 41(9):86–93, September 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
9/p86-keller/.

Dogac:1998:DOC

[1640] Asuman Dogac, Cevdet Dengi, and M. Tamer Öszu. Distributed object computing platforms. Communications of the ACM, 41(9):95–103, September 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
9/p95-dogac/.

Neumann:1998:IRY

[1641] Peter G. Neumann. Inside risks: Y2K update. Communications of the ACM, 41(9):128, September 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
9/p128-neumann/.

Fox:1998:NTj

[1642] Robert Fox. News track. Communications of the ACM, 41(10):9–11, October 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:1998:WWM

[1643] Hal Berghel. Who won the Mosaic war? Communications of the ACM, 41(10):13–16, October 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
10/p13-berghel/.

Simons:1998:OT

[1644] Barbara Simons. Outlawing technology. Communications of the ACM, 41(10):17–18, October 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
10/p17-simons/.

Press:1998:IDF

[1645] Larry Press, Grey Burkhart, Will Foster, Sy Goodman, Peter Wolcott, and Jon Woodard. An Internet diffusion framework. Communications of the ACM, 41(10):21–26, October 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        10/p21-press/.

Hsu:1998:VMH

[1646] Richard C. Hsu. Viewpoint: Maintaining high living standards through innovation, strong patents. Communications of the ACM, 41(10):27–28, October 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1998-41-10/p27-hsu/.

Crawford:1998:FDS

[1647] Diane Crawford. Forum: The dirty secrets of runaway projects. Communications of the ACM, 41(10):29–33, October 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1998-41-10/p29-crawford/.

Seetharaman:1998:CC

[1648] Krishnan Seetharaman. The CORBA connection. Communications of the ACM, 41(10):34–36, October 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        10/p34-seetharaman/.

Siegel:1998:OOC

[1649] Jon Siegel. OMG overview: CORBA and the OMA in enterprise computing. Communications of the ACM, 41(10):37–43, October 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
10/p37-siegel/.

Vinoski:1998:NFC

[1650] Steve Vinoski. New features for CORBA 3.0. Communications of the ACM, 41(10):44–52, October 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
10/p44-vinoski/.

Schmidt:1998:EAM

[1651] Douglas C. Schmidt. Evaluating architectures for multithreaded object request brokers. Communications of the ACM, 41(10):54–60, October 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
10/p54-schmidt/.

Henning:1998:BMS

[1652] Michi Henning. Binding, migration, and scalability in CORBA. Communications of the ACM, 41(10):62–71, October 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1998-41-10/p62-henning/.

Haggerty:1998:BCB

[1653] Paul Haggerty and Krishnan Seetharaman. The benefits of CORBA-based network management. Communications of the ACM, 41(10):73–79, October 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:

                                        80/pubs/citations/journals/cacm/
                                        1998-41-10/p73-haggerty/.

Segev:1998:ISC

[1654] Arie Segev, Jaana Porra, and Malu Roldan. Internet security and the case of Bank of America. Communications of the ACM, 41(10):81–87, October 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        10/p81-segev/.

Riggins:1998:TUV

[1655] Frederick J. Riggins and Hyeun-Suk (Sue) Rhee. Toward a unified view of electronic commerce. Communications of the ACM, 41(10):88–95, October 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1998-41-10/p88-riggins/.

Rai:1998:HAI

[1656] Arun Rai, T. Ravichandran, and Subhashish Samaddar. How to anticipate the Internet’s global diffusion. Communications of the ACM, 41(10):97– 106, October 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org:80/pubs/citations/journals/
                                        cacm/1998-41-10/p97-rai/.

Gisselquist:1998:ES

[1657] Richard Gisselquist. Engineering in software. Communications of the ACM, 41(10):107–108, October 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
10/p107-gisselquist/.

Neumann:1998:IRR

[1658] Peter G. Neumann. Inside risks:

Risks of E-education. Communications of the ACM, 41(10):136, October 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
10/p136-neumann/.

Fox:1998:NTk

[1659] Robert Fox. News track. Communications of the ACM, 41(11):9–10, November 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1998-41-11/p9-fox/.

Crawford:1998:FSN

[1660] Diane Crawford. Forum: Stop, in the name of spam. Communications of the ACM, 41(11):11–14, November 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
11/p11-crawford/.

Schorr:1998:WDG

[1661] Herbert Schorr and Salvatore J. Stolfo.

From Washington: digital government for the 21st century. Communications of the ACM, 41(11):15–19, November 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
11/p15-schorr/.

Burkhart:1998:IPIb

[1662] Grey E. Burkhart, Seymour E. Goodman, Arun Mehta, and Larry Press. International perspectives: the Internet in India: better times ahead? Communications of the ACM, 41(11):21–26, November 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1998-41-11/p21-burkhart/.

Wolcott:1998:VHE

[1663] Peter Wolcott, Patrick Homer, and Seymour Goodman. Viewpoint: HPC export controls: navigating choppy waters. Communications of the ACM, 41(11):27–30, November 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
11/p27-wolcott/.

Karin:1998:HPC

[1664] Sidney Karin and Susan Graham. The high-performance computing continuum. Communications of the ACM, 41(11):32–35, November 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
11/p32-karin/.

Messina:1998:A

[1665] Paul Messina, David Culler, Wayne Pfeiffer, Wiliam Martin, J. Tinsley Oden, and Gary Smith. Architecture. Communications of the ACM, 41(11):36–44, November 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
11/p36-messina/.

Grimshaw:1998:M

[1666] Andrew Grimshaw, Adam Ferrari, Greg Lindahl, and Katherine Holcomb. Metasystems. Communications of the ACM, 41(11):46–55, November 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
11/p46-grimshas/.

Moore:1998:DIC

[1667] Reagan Moore, Thomas A. Prince, and Mark Ellisman. Data-intensive computing and digital libraries. Communications of the ACM, 41(11):56–62, November 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1998-41-11/p56-moore/.

Saltz:1998:PTE

[1668] Joel H. Saltz, Alan Sussman, Susan Graham, James Demmel, Scott Baden, and Jack Dongarra. Programming tools and environments. Communications of the ACM, 41(11):64–73, November 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
11/p64-sussman/.

Finley:1998:EOT

[1669] Amy Finley, Ann Redelfs, and Greg Moses. Education, outreach, and training. Communications of the ACM, 41(11):74–75, November 1998.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        11/p74-finley/.

Keil:1998:FIS

[1670] Mark Keil, Paul Cule, Kalle Lyytinen, and Roy Schmidt. A framework for identifying software projects risks. Communications of the ACM, 41(11):76–83, November 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        11/p76-keil/.

Kunz:1998:VDT

[1671] John C. Kunz, Tore R. Christiansen, Geoff P. Cohen, Yan Jin, and Raymond E. Levitt. The virtual design team. Communications of the ACM, 41(11):84–91, November 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        11/p84-kunz/.

Kuo:1998:NDC

[1672] Geng-Sheng Kuo and Jing-Pei Lin. New design concepts for an intelligent Internet. Communications of the ACM, 41(11):93–98, November 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        11/p93-kuo/.

Maisel:1998:ESR

[1673] Herbert Maisel and Catherine Gaddy.

Employment and salaries of recent CS doctorates. Communications of the ACM, 41(11):99–101, November 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
11/p99-maisel/.

Koczkodaj:1998:TOM

[1674] Waldemar W. Koczkodaj, Marian Orlowski, and Victor W. Marek. Technical opinion: myths about rough set theory. Communications of the ACM, 41(11):102–103, November 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
11/p102-koczkodaj/.

Schneider:1998:IRT

[1675] Fred B. Schneider. Inside risks: toward trustworthy networked information systems. Communications of the ACM, 41(11):144, November 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
11/p144-schneider/.

Crawford:1998:EPd

[1676] Diane Crawford. Editorial pointers.

Communications of the ACM, 41(12): 5, December 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1998-41-12/p5-crawford/.

Fox:1998:NTl

[1677] Robert Fox. News track. Communications of the ACM, 41(12):9–10, December 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1998-41-12/p9-fox/.

Glass:1998:PPH

[1678] Robert L. Glass. Practical programmer: how not to prepare for a consulting assignment, and other ugly consultancy truths. Communications of the ACM, 41(12):11–13, December 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1998-41-
                                        12/p11-glass/.

Aalberts:1998:LST

[1679] Robert J. Aalberts, Anthony M. Townsend, and Michael E. Whitman. Legally speaking: the threat of long-arm jurisdiction to electronic commerce. Communications of the ACM, 41(12):15–20, December 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org:80/pubs/citations/journals/
                                        cacm/1998-41-12/p15-aalberts/.

Kraut:1998:SSI

[1680] Robert Kraut, Sara Kiesler, Tridas Mukhopadhyay, William Scherlies, and Michael Patterson. On site: social impact of the Internet: what does it mean? Communications of the ACM, 41(12):21–22, December 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
12/p21-kraut/.

Crawford:1998:FCE

[1681] Diane Crawford. Forum: Computers in education: Observations and redress. Communications of the ACM, 41(12):23–26, December 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
12/p23-crawford/.

Cerf:1998:TRI

[1682] Vinton G. Cerf. Tribute: I remember IANA. Communications of the ACM, 41(12):27–28, December 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
12/p27-cerf/.

Karat:1998:VGR

[1683] Clare-Marie Karat. Viewpoint: guaranteeing rights for the user. Communications of the ACM, 41(12):29–31, December 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1998-41-12/p29-karat/.

Jarke:1998:RT

[1684] Matthias Jarke. Requirements tracing. Communications of the ACM, 41(12):32–36, December 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/

citations/journals/cacm/1998-41-
12/p32-jarke/.

Ramesh:1998:FIR

[1685] Balasubramaniam Ramesh. Factors influencing requirements traceability practice. Communications of the ACM, 41(12):37–44, December 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
12/p37-ramesh/.

Rose:1998:VAE

[1686] Thomas Rose. Visual assessment of engineering processes in virtual enterprises. Communications of the ACM, 41(12):45–52, December 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
12/p45-rose/.

Domges:1998:ATE

[1687] Ralf Dömges and Klaus Pohl. Adapting traceability environments to projectspecific needs. Communications of the ACM, 41(12):55–62, December 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
12/p55-domges/.

DeMichelis:1998:TFV

[1688] Giorgio De Michelis, Eric Dubois, Matthias Jarke, Florian Matthes, John Mylopoulos, Joachim W. Schmidt, Eric Yu, and Carson Woo. A three-faceted view of informations systems. Communications of the ACM, 41(12):64–70, December 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1998-41-12/p64-de_michelis/

.

Feijs:1998:VSA

[1689] Loe Feijs and Roel De Jong. 3D visualization of software architectures. Communications of the ACM, 41(12):73–78, December 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1998-41-12/p73-feijs/.

Noffsinger:1998:LOM

[1690] W. B. Noffsinger, Robert Niedbalski, Michael Blanks, and Niall Emmart. Legacy object modeling speeds software integration. Communications of the ACM, 41(12):80–89, December 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
12/p80-noffsinger/.

De:1998:ISI

[1691] Prabuddha De and Thomas W. Ferratt. An information system involving competing organizations. Communications of the ACM, 41(12):90–98, December 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1998-41-12/p90-de/.

Date:1998:TOT

[1692] C. J. Date and Hugh Darwen. Technical opinion: toward a model of type inheritance. Communications of the ACM, 41(12):99–101, December 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
12/p99-date/.

Ladkin:1998:IRR

[1693] Peter Ladkin. Inside risks: the risks of hubris. Communications of the ACM, 41(12):160, December 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
12/p160-ladkin/.

Fox:1999:NTa

[1694] Robert Fox. News track. Communications of the ACM, 42(1):9–10, January 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
1/p9-fox/.

Crawford:1999:FCP

[1695] Diane Crawford, Timothy D. Prime, Justus Pendleton, Greg Kirsch, Lyle P. Bickley, Markku Sakkinen, Jakob Engblom, Howard L. Funk, John Rieman, Joy Davis, and David Balok. Forum: Correlating patents and living standards. Communications of the ACM, 42(1):11–13, January 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
1/p11-crawford/.

Whitman:1999:LSC

[1696] Michael W. Whitman, Anthony M.

Townsend, and Robert J. Aalberts. Legally speaking: The Communications Decency Act is not as dead as you think. Communications of the ACM, 42(1):15–17, January 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
1/p15-townsend/.

Berghel:1999:DVV

[1697] Hal Berghel. Digital village: Valueadded publishing. Communications of the ACM, 42(1):19–23, January 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
1/p19-berghel/.

Hailperin:1999:VCB

[1698] Max Hailperin. Viewpoint: The COPA battle and the future of free speech. Communications of the ACM, 42(1):24–25, January 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
1/p24-hailperin/.

Simons:1999:PSW

[1699] Barbara Simons. From the President: Starr wars. Communications of the ACM, 42(1):26–27, January 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
1/p26-simons/.

Kaindl:1999:SMP

[1700] Hermann Kaindl and John M. Carroll. Symbolic modeling in practice. Communications of the ACM, 42(1):28–30, January 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org:80/pubs/citations/journals/
                                        cacm/1999-42-1/p26-simons/.

Mylopoulos:1999:OOG

[1701] John Mylopoulos, Lawrence Chung, and Eric Yu. From object-oriented to goal-oriented requirements analysis. Communications of the ACM, 42(1):31–37, January 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        journals/cacm/1999-42-1/p26-simons/

.

Butler:1999:CDS

[1702] Keith A. Butler, Chris Esposito, and Ron Hebron. Connecting the design of software to the design of work. Communications of the ACM, 42(1):38–46, January 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org:80/pubs/citations/journals/
                                        cacm/1999-42-1/p26-simons/.

Jarke:1999:SM

[1703] Matthias Jarke. Scenarios for modeling. Communications of the ACM, 42(1):47–48, January 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        1/p26-simons/.

Rosson:1999:IDT

[1704] Mary Beth Rosson. Integrating development of task and object models. Communications of the ACM, 42(1):49–56, January 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
1/p26-simons/.

Gaines:1999:EFK

[1705] Brian R. Gaines and Mildred L. G.

Shaw. Embedding formal knowledge models in active documents. Communications of the ACM, 42(1):57–63, January 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1999-42-1/p26-simons/.

Nielsen:1999:UID

[1706] Jakob Nielsen. User interface directions for the Web. Communications of the ACM, 42(1):65–72, January 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
1/p65-nielsen/.

Ballou:1999:EDQ

[1707] Donald P. Ballou and Giri Kumar Tayi. Enhancing data quality in data warehouse environments. Communications of the ACM, 42(1):73–78, January 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
1/p73-ballou/.

Nicol:1999:HIH

[1708] John R. Nicol, Yechezkal S. Gutfreund, Jim Paschetto, Kimberly S. Rush, and Christopher Martin. How the Internet helps build collaborative multimedia applications. Communications of the ACM, 42(1):79–85, January 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        1/p79-nicol/.

Robillard:1999:RKS

[1709] Pierre N. Robillard. The role of knowledge in software development. Communications of the ACM, 42(1):87–92, January 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org:80/pubs/citations/journals/
                                        cacm/1999-42-1/p87-robillard/.

Avison:1999:AR

[1710] David E. Avison, Francis Lau, Michael D. Myers, and Peter Axel Nielsen. Action research. Communications of the ACM, 42(1):94–97, January 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        1/p94-avison/.

Nambisan:1999:TOR

[1711] Satish Nambisan and Yu-Ming Wang. Technical opinion: Roadblocks to Web technology adoption? Communications of the ACM, 42(1):98–101, January 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:

80/pubs/citations/journals/cacm/
1999-42-1/p98-nambisan/.

Schneider:1999:IRE

[1712] Fred B. Schneider and Steven M.

Bellovin. Inside risks: Evolving telephone networks. Communications of the ACM, 42(1):160, January 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
1/p160-schneider/.

Fox:1999:NTb

[1713] Robert Fox. News track. Communications of the ACM, 42(2):9–11, February 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
2/p9-fox/.

Press:1999:PCN

[1714] Larry Press. Personal computing:

The next generation of business data processing. Communications of the ACM, 42(2):13–16, February 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
2/p13-press/.

Meeks:1999:EFP

[1715] Brock N. Meeks. Electronic frontier:

The privacy hoax. Communications of the ACM, 42(2):17–19, February 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
2/p17-meeks/.

Crawford:1999:FMU

[1716] Diane Crawford. Forum: More ugly truths about consulting assignments. Communications of the ACM, 42(2):21–24, February 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        2/p21-crawford/.

Jalloul:1999:SCL

[1717] Ghinwa Jalloul. On site: Computing in Lebanon. Communications of the ACM, 42(2):25–26, February 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        2/p25-jalloul/.

Cranor:1999:IP

[1718] Lorrie Faith Cranor. Internet privacy. Communications of the ACM, 42(2):28–38, February 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        2/p28-cranor/.

Reiter:1999:AWT

[1719] Michael K. Reiter and Aviel D. Rubin. Anonymous Web transactions with crowds. Communications of the ACM, 42(2):32–48, February 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        2/p32-reiter/.

Goldschlag:1999:OR

[1720] David Goldschlag, Michael Reed, and Paul Syverson. Onion routing. Communications of the ACM, 42(2):39–41, February 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1999-42-2/p39-goldschlag/.

Gabber:1999:CYA

[1721] Eran Gabber, Phillip B. Gibbons, David M. Kristol, Yossi Matias, and Alain Mayer. Consistent, yet anonymous, Web access with LPWA. Communications of the ACM, 42(2):42–47, February 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1999-42-2/p42-gabber/.

Reagle:1999:PPP

[1722] Joseph Reagle and Lorrie Faith Cranor. The platform for privacy preferences. Communications of the ACM, 42(2):48–55, February 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
2/p48-reagle/.

Benassi:1999:TOP

[1723] Paola Benassi. TRUSTe: An online privacy seal program. Communications of the ACM, 42(2):56–59, February 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
2/p56-benassi/.

Clarke:1999:IPC

[1724] Roger Clarke. Internet privacy concerns confirm the case for intervention. Communications of the ACM, 42(2):60–67, February 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        2/p60-clarke/.

Gowan:1999:YCD

[1725] J. Arthur Gowan, Chris Jesse, and Richard G. Mathieu. Y2K compliance and the distributed enterprise. Communications of the ACM, 42(2):68–73, February 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org:80/pubs/citations/journals/
                                        cacm/1999-42-2/p68-gowan/.

Glass:1999:RST

[1726] Robert L. Glass. The realities of software technology payoffs. Communications of the ACM, 42(2):74–79, February 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1999-42-2/p74-glass/.

Sabherwal:1999:RTO

[1727] Rajiv Sabherwal. The role of trust in outsourced is development projects. Communications of the ACM, 42(2):80–86, February 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        2/p80-sabherwal/.

Hasselbring:1999:TOD

[1728] Wilhelm Hasselbring. Technical opinion: On defining computer science terminology. Communications of the ACM, 42(2):88–91, February 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
2/p88-hasselbring/.

Neumann:1999:IRRa

[1729] Peter G. Neumann. Inside risks: Robust open-source software. Communications of the ACM, 42(2):128, February 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1999-42-2/p128-neumann/.

Fox:1999:NTc

[1730] Robert Fox. News track. Communications of the ACM, 42(3):11–12, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
3/p11-fox/.

Lyytinen:1999:FUS

[1731] Kalle Lyytinen and Seymour Goodman. Finland: the unknown soldier on the IT front. Communications of the ACM, 42(3):13–17, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
3/p13-lyytinen/.

Samuelson:1999:GNB

[1732] Pamela Samuelson. Good news and bad news on the intellectual property front. Communications of the ACM, 42(3):19–24, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
3/p19-samuelson/.

Crawford:1999:FUV

[1733] Diane Crawford, Clay Shirky, David Webber, Jakob Nielson, Joseph M. Newcomer, The HomeNet Group, W. M. Jaworski, Waldemar W. Koczkodaj, Marian Orlowski, and Victor W. Marek. Forum: Usability vs. the Web. Communications of the ACM, 42(3):27–32, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
3/p27-crawford/.

Rickert:1999:VFG

[1734] Donald E. Rickert and Thomas Nagy. Viewpoint: focus groups, theory or the kid in the garage? Communications of the ACM, 42(3):34–36, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
3/p34-rickert/.

Joshi:1999:MSN

[1735] Anupam Joshi and Munindar P. Singh. Multiagent systems on the net. Communications of the ACM, 42(3):38–40, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1999-42-3/p38-joshi/.

Gustavsson:1999:AP

[1736] Rune Gustavsson. Agents with power. Communications of the ACM, 42(3):41–47, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
3/p41-gustavsson/.

Drashansky:1999:NAS

[1737] Tzvetan Drashansky, Elias N. Houstis, Naren Ramakrishnan, and John R. Rice. Networked agents for scientific computing. Communications of the ACM, 42(3):48–52, 54, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
3/p48-drashansky/.

Dyer:1999:MSD

[1738] Douglas E. Dyer. Multiagent systems and DARPA. Communications of the ACM, 42(3):53, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
3/p53-dyer/.

Hattori:1999:SMS

[1739] Fumio Hattori, Takeshi Ohguro, Makoto Yokoo, Shigeo Matsubara, and Sen Yoshida. Socialware: multiagent systems for supporting network communities. Communications of the ACM, 42(3):55–61, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        3/p55-hattori/.

Chaturvedi:1999:SEM

[1740] Alok R. Chaturvedi and Shailendra R. Mehta. Simulations in economics and management. Communications of the ACM, 42(3):60–61, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        3/p60-chaturvedi/.

Jain:1999:APC

[1741] Anuj K. Jain, Manuel Aparico IV, and Munindar P. Singh. Agents for process coherence in virtual enterprises. Communications of the ACM, 42(3):62–69, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1999-42-3/p62-jain/.

Green:1999:MDC

[1742] Richard Green and Sangam Pant. Multiagent data collection in Lycos. Communications of the ACM, 42(3):70, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1999-42-3/p70-green/.

Sloman:1999:BCR

[1743] Aaron Sloman and Brian Logan. Building cognitively rich agents using the SIM Agent toolkit. Communications of the ACM, 42(3):71–73, 75, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
3/p71-sloman/.

Baxter:1999:ATB

[1744] Jeremy Baxter and Richard Helewhite. Agents in tank battle simulations. Communications of the ACM, 42(3):74–75, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
3/p74-baxter/.

Ma:1999:AC

[1745] Moses Ma. Agents in E-commerce.

Communications of the ACM, 42 (3):78–80, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//www.acm.org:80/pubs/citations/
journals/cacm/1999-42-3/p78-ma/.

Maes:1999:ABS

[1746] Pattie Maes, Robert H. Guttman, and Alexandros G. Moukas. Agents that buy and sell. Communications of the ACM, 42(3):81–87, 90–91, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
3/p81-maes/.

Sandholm:1999:AN

[1747] Tuomas Sandholm. Automated negotiation. Communications of the ACM, 42(3):84–85, March 1999.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
3/p84-sandholm/.

Lange:1999:SGR

[1748] Danny B. Lange and Mitsuru Oshima. Seven good reasons for mobile agents. Communications of the ACM, 42(3):88–89, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
3/p88-lange/.

Wong:1999:JBM

[1749] David Wong, Noemi Paciorek, and Dana Moore. Java-based mobile agents. Communications of the ACM, 42(3):92–95, 97–102, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
3/p92-wong/.

Koblick:1999:C

[1750] Reuven Koblick. Concordia. Communications of the ACM, 42(3):96–97, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1999-42-3/p96-koblick/.

Tai:1999:AP

[1751] Hideki Tai and Kazuya Kosaka. The Aglets project. Communications of the ACM, 42(3):100–101, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
3/p100-tai/.

Hayes-Roth:1999:SWI

[1752] Barbara Hayes-Roth, Vaughan Johnson, Robert van Gent, and Keith Wescourt. Staffing the Web with interactive characters. Communications of the ACM, 42(3):103–105, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
3/p103-hayes-roth/.

Glushko:1999:XFA

[1753] Robert J. Glushko, Jay M. Tenenbaum, and Bart Meltzer. An XML framework for agent-based E-commerce. Communications of the ACM, 42 (3):106–109, 111–114, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
3/p106-glushko/.

Smith:1999:SOX

[1754] Howard Smith and Kevin Poulter.

Share the ontology in XML-based trading architectures. Communications of the ACM, 42(3):110–111, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
3/p110-smith/.

Alderson:1999:TOV

[1755] Albert Alderson and Hanifa Shah.

Technical opinion: Viewpoints on legacy systems. Communications of the ACM, 42(3):115–116, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        3/p115-alderson/.

Weinstein:1999:IRB

[1756] Lauren Weinstein. Inside risks: Bitrot roulette. Communications of the ACM, 42(3):144, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        3/p144-weinstein/.

Yeandel:1999:DIL

[1757] J. Yeandel, D. Thulborn, and S. Jones. The design and implementation of an on-line testable UART. Communications of the ACM, 42(3):187–198, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1999-42-3/p187-yeandel/.

Abdulla:1999:OMS

[1758] M. F. Abdulla, C. P. Ravikumar, and Anshul Kumar. Optimization of mutual and signature testing schemes for highly concurrent systems. Communications of the ACM, 42(3):199–216, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1999-42-3/p199-abdulla/.

Sahraoui:1999:TRN

[1759] Zohair Sahraoui, Francky Catthoor, Paul Six, and Hugo de Man. Techniques for reducing the number of decisions and backtracks in combinational test generation. Communications of the ACM, 42(3):217–238, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1999-42-3/p217-sahraoui/.

Pappu:1999:SDF

[1760] Lakshminarayana Pappu, Michael L.

Bushnell, Vishwani D. Agrawal, and Srinivas Mandyam-Komar. Statistical delay fault coverage estimation for synchronous sequential circuits. Communications of the ACM, 42(3):239–254, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1999-42-3/p239-pappu/.

Wang:1999:CAC

[1761] De-Qiang Wang and Lian-Chang Zhao.

Combinatorial analysis of check set construction for algorithm-based fault tolerance systems. Communications of the ACM, 42(3):255–260, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
3/p255-wang/.

Crawford:1999:EPa

[1762] Diane Crawford. Editorial pointers.

Communications of the ACM, 42(4):

5, April 1999. URL http://www.acm.
org:80/pubs/citations/journals/
cacm/1999-42-4/p5-crawford/.

Fox:1999:NTd

[1763] Robert Fox. News track. Communications of the ACM, 42(4):11–12, April

                                        1999. URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1999-42-4/p11-fox/.

Crawford:1999:FRB

[1764] Diane Crawford, John Ehrt, Joe Saltiel, Robert Glass, John Wubbel, John Willmore, and Tom Moran. Forum: a righteous balance of Internet freedom; don’t compare business majors to hackers; “popular” press; empirical studies vs. armchair argument. Communications of the ACM, 42(4):13–16, April

                                        1999. URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1999-42-4/p13-crawford/.

Glass:1999:PPI

[1765] Robert L. Glass. Practical programmer: inspections—some surprising findings. Communications of the ACM, 42(4):17–19, April 1999.

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        4/p17-glass/.

Rada:1999:SSI

[1766] Roy Rada. Sharing standards: IT skills standards. Communications of the ACM, 42(4):21–26, April 1999.

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        4/p21-rada/.

Bagert:1999:VTL

[1767] Donald J. Bagert. Viewpoint: taking the lead in licensing software engineers. Communications of the ACM, 42(4):27–29, April 1999.

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
4/p27-bagert/.

White:1999:ADL

[1768] John White. ACM digital library enhancements. Communications of the ACM, 42(4):30–31, April 1999.

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
4/p30-white/.

OReilly:1999:LOS

[1769] Tim O’Reilly. Lessons from opensource software development. Communications of the ACM, 42(4):32–37,

April 1999. URL http://www.acm.
org:80/pubs/citations/journals/
cacm/1999-42-4/p32-o_reilly/.

Torvalds:1999:LE

[1770] Linus Torvalds. The Linux edge. Communications of the ACM, 42(4):38–39,

April 1999. URL http://www.acm.
org:80/pubs/citations/journals/
cacm/1999-42-4/p38-torvalds/.

Wall:1999:OCL

[1771] Larry Wall. The origin of the camel lot in the breakdown of the bilingual Unix. Communications of the ACM, 42(4):40–41, April 1999.

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
4/p40-wall/.

Fielding:1999:SLA

[1772] Roy T. Fielding. Shared leadership in the Apache project. Communications of the ACM, 42(4):42–43, April 1999.

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
4/p42-fielding/.

Ousterhout:1999:FSN

[1773] John Ousterhout. Free software needs profit. Communications of the ACM, 42(4):44–45, April 1999.

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        4/p44-ousterhout/.

Yao:1999:FPE

[1774] Xin Yao. Following the path of evolvable hardware. Communications of the ACM, 42(4):46–49, April 1999.

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        4/p46-yao/.

Sipper:1999:QVE

[1775] Moshe Sipper, Daniel Mange, and Eduardo Sanchez. Quo Vadis evolvable hardware? Communications of the ACM, 42(4):50–56, April 1999.

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        4/p50-sipper/.

Marchal:1999:FPG

[1776] Pierre Marchal. Field-programmable gate arrays. Communications of the ACM, 42(4):57–59, April 1999.

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        4/p57-marchal/.

Higuchi:1999:EHC

[1777] Tetsuya Higuchi and Nobuki Kajihara. Evolvable hardware chips for industrial applications. Communications of the ACM, 42(4):60–66, April 1999.

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        4/p60-higuchi/.

Hikage:1999:HES

[1778] Tomofumi Hikage, Hitoshi Hemmi, and Katsunori Shimohara. Hardware evolution system AdAM. Communications of the ACM, 42(4): 64–65, 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1999-42-4/p64-hikage/.

Lohn:1999:EES

[1779] Jason D. Lohn. Experiments on evolving software models of analog circuits. Communications of the ACM, 42 (4):67–69, 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/1999-42-4/p67-lohn/.

Thompson:1999:AUE

[1780] Adrian Thompson and Paul Layzell.

Analysis of unconventional evolved electronics. Communications of the ACM, 42(4):71–79, April 1999.

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
4/p71-thompson/.

Hoffman:1999:BCT

[1781] Donna L. Hoffman, Thomas P. Novak, and Marcos Peralta. Building consumer trust online. Communications of the ACM, 42(4):80–85, April 1999.

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
4/p80-hoffman/.

Imielinski:1999:GBG

[1782] Tomasz Imielinski and Julio C. Navas.

GPS-based geographic addressing, routing, and resource discovery. Communications of the ACM, 42(4):86–92,

April 1999. URL http://www.acm.
org:80/pubs/citations/journals/
cacm/1999-42-4/p86-imielinski/.

Ahituv:1999:TMT

[1783] Niv Ahituv, Moshe Zviran, and Chanan Glezer. Top management toolbox for managing Corporate IT. Communications of the ACM, 42(4):93–99,

April 1999. URL http://www.acm.
org:80/pubs/citations/journals/
cacm/1999-42-4/p93-ahituv/.

Guerraoui:1999:TOOa

[1784] Rachid Guerraoui and Mohamed E. Fayed. Thinking objectively: OO distributed programming is not distributed OO programming. Communications of the ACM, 42(4):101–104,

April 1999. URL http://www.acm.
org:80/pubs/citations/journals/
cacm/1999-42-4/p101-guerraoui/.

Weinstein:1999:IRJ

[1785] Lauren Weinstein. Inside risks: just a matter of bandwidth. Communications of the ACM, 42(4):120, April 1999.

URL http://www.acm.org:80/pubs/
articles/journals/cacm/1999-42-
4/p120-weinstein/p120-weinstein.
pdf;   http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
4/p120-weinstein/.

Crawford:1999:EPb

[1786] Diane Crawford. Editorial pointers. Communications of the ACM, 42(5):5, May 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:

80/pubs/citations/journals/cacm/
1999-42-5/p5-crawford/.

Fox:1999:NTe

[1787] Robert Fox. News track. Communications of the ACM, 42(5):9–10, May 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
5/p9-fox/.

Berghel:1999:DVH

[1788] Hal Berghel. Digital village: how Xday figures in the Y2K countdown. Communications of the ACM, 42(5):11–15, May 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1999-42-5/p11-berghel/.

Post:1999:PCH

[1789] Gerald V. Post. Personal computing:

how often should a firm buy new PCs? Communications of the ACM, 42(5): 17–21, May 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1999-42-5/p17-post/.

Kappelman:1999:VSO

[1790] Leon A. Kappelman. Viewpoint: saving our sacred honor. Communications of the ACM, 42(5):23–25, May 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
5/p23-kappelman/.

Fogg:1999:PT

[1791] B. J. Fogg. Persuasive technologies. Communications of the ACM, 42(5): 26–29, May 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org:80/pubs/citations/journals/
                                        cacm/1999-42-5/p26-fogg/.

King:1999:LPT

[1792] Phillip King and Jason Tester. The landscape of persuasive technologies. Communications of the ACM, 42(5): 31–38, May 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org:80/pubs/citations/journals/
                                        cacm/1999-42-5/p31-king/.

Tseng:1999:CCT

[1793] Shawn Tseng and B. J. Fogg. Credibility and computing technology. Communications of the ACM, 42(5):39–44, May 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1999-42-5/p39-tseng/.

Khaslavsky:1999:USE

[1794] Julie Khaslavsky and Nathan Shedroff. Understanding the seductive experience. Communications of the ACM, 42(5):45–49, May 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        5/p45-khaslavsky/.

Berdichevsky:1999:TEP

[1795] Daniel Berdichevsky and Erik Neunschwander. Toward an ethics of persuasive technology. Communications of the ACM, 42(5):51–58, May 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
5/p51-berdichevsky/.

Madsen:1999:DUP

[1796] Kim Halskov Madsen. The diversity of usability practices. Communications of the ACM, 42(5):60–62, May 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
5/p60-madsen/.

Buur:1999:RUT

[1797] Jacob Buur and Kirsten Bagger.

Replacing usability testing with user dialogue. Communications of the ACM, 42(5):63–66, May 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
5/p63-buur/.

Vredenburg:1999:IEU

[1798] Karel Vredenburg. Increasing ease of use. Communications of the ACM, 42(5):67–71, May 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//www.acm.org:80/pubs/citations/
journals/cacm/1999-42-5/p67-vredenburg/

.

Baerentsen:1999:CDP

[1799] Klaus B. Bærentsen and Henning Slavensky. A contribution to the design process. Communications of the ACM, 42(5):72–77, May 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        5/p72-baerentsen/.

Gardner:1999:SFU

[1800] Julia Gardner. Strengthening the focus on users’ working practices. Communications of the ACM, 42(5):79–82, May 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        5/p79-gardner/.

Dolan:1999:FAT

[1801] pWilliam Dolan and Joseph Dumas. A flexible approach to thirdparty usability. Communications of the ACM, 42(5):83–85, May 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        5/p83-dolan/.

Muller:1999:OUW

[1802] Michael Muller and Mary P. Czerwinski. Organizing usability work to fit the full product range. Communications of the ACM, 42(5):87–90, May 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        5/p87-muller/.

Borgholm:1999:CUP

[1803] Thea Borgholm and Kim Halskov Madsen. Cooperative usability practices. Communications of the ACM, 42(5):

91–97, May 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1999-42-5/p91-borgholm/.

Poulin:1999:TOR

[1804] Jeffrey S. Poulin. Technical opinion:

reuse: been there, done that. Communications of the ACM, 42(5):98–100, May 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1999-42-5/p98-poulin/.

Parnas:1999:IRT

[1805] David Lorge Parnas. Inside risks: ten myths about Y2K inspections. Communications of the ACM, 42(5):128, May 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1999-42-5/p128-parnas/.

Crawford:1999:EPc

[1806] Diane Crawford. Editorial pointers.

Communications of the ACM, 42(6):5, June 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1999-42-6/p5-crawford/.

Fox:1999:NTf

[1807] Robert Fox. News track. Communications of the ACM, 42(6):9–10, June 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
6/p9-fox/.

Glass:1999:PPC

[1808] Robert L. Glass. Practical programmer: is criticism of computing academe inevitably divisive? Communications of the ACM, 42(6):11–13, June 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        6/p11-glass/.

Heeks:1999:IPS

[1809] Richard B. Heeks. International perspectives: software strategies in developing countries. Communications of the ACM, 42(6):15–20, June 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-

6/p15-heeks/. See correction [1851].

Crawford:1999:FDF

[1810] Diane Crawford, Jason Steffler, Jane D. Bernstein, John Chapin, David M. Rockwell, and Kenneth Bell. Forum: Don’t forget the intangibles in improving software inspections; OO programming; licensing a unique field. Communications of the ACM, 42(6):21–22, June 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1999-42-6/p21-crawford/.

Wang:1999:VII

[1811] Wei-Lung Wang. Viewpoint: Information in the information age. Communications of the ACM, 42(6):23–24, June 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
6/p23-wang/.

Simons:1999:PMM

[1812] Barbara Simons. From the president: Melissa’s message. Communications of the ACM, 42(6):25–26, June 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
6/p25-simons/.

Mueller:1999:EII

[1813] Milton Mueller. Emerging Internet infrastructures worldwide. Communications of the ACM, 42(6):28–30, June 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
6/p28-mueller/.

Petrazzini:1999:IDC

[1814] Ben Petrazzini and Mugo Kibati. The Internet in developing countries. Communications of the ACM, 42(6):31–36, June 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1999-42-6/p31-petrazzini/.

Cohen:1999:MTN

[1815] Robert B. Cohen. Moving toward a non-U.S.-centric international Internet. Communications of the ACM, 42(6): 37–40, June 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1999-42-6/p37-cohen/.

Mueller:1999:CII

[1816] Milton Mueller. Commentary: ICANN and Internet regulation. Communications of the ACM, 42(6):41–43, June 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        6/p41-mueller/.

Tan:1999:CSC

[1817] Zixiang (Alex) Tan, William Foster, and Seymour Goodman. China’s state-coordinated Internet infrastructure. Communications of the ACM, 42(6):44–52, June 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        journals/cacm/1999-42-6/p44-tan/

.

Agarwal:1999:BIN

[1818] P. K. Agarwal. Building India’s national Internet backbone. Communications of the ACM, 42(6):53–58, June 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        6/p53-agarwal/.

Kibati:1999:WLL

[1819] Mugo Kibati and Donyaprueth Krairit. The wireless local loop in developing regions. Communications of the ACM, 42(6):60–66, June 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        6/p60-kibati/.

Peha:1999:LHI

[1820] Jon A. Peha. Lessons from Haiti’s Internet development. Communications of the ACM, 42(6):67–72, June 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
6/p67-peha/.

Chatterjee:1999:AON

[1821] Samir Chatterjee and Suzanne Pawlowski.

All-optical networks. Communications of the ACM, 42(6):74–83, June 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
6/p74-chatterjee/.

Beer:1999:UAR

[1822] Randall D. Beer, Hillel J. Chiel, and Richard F. Drushel. Using Autonomous Robotics to teach science and engineering. Communications of the ACM, 42(6):85–92, June 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
6/p85-beer/.

Tsichritzis:1999:RU

[1823] Dennis Tsichritzis. Reengineering the University. Communications of the ACM, 42(6):93–100, June 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
6/p93-tsichritzis/.

Whitman:1999:CET

[1824] Michael E. Whitman, Anthony M. Townsend, and Robert J. Aalberts. Considerations for an effective telecommunications-use policy. Communications of the ACM, 42(6):101– 108, June 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1999-42-6/p101-whitman/.

Sauter:1999:IDM

[1825] Vicki L. Sauter. Intuitive decisionmaking. Communications of the ACM, 42(6):109–115, June 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
6/p109-sauter/.

Wise:1999:CAR

[1826] David S. Wise. Changes in ACM’s revised copyright policy. Communications of the ACM, 42(6):116–119, June 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
6/p116-wise/.

Noll:1999:TOD

[1827] A. Michael Noll. Technical opinion: does data traffic exceed voice traffic? Communications of the ACM, 42(6):121–124, June 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
6/p121-noll/.

Neumann:1999:IRRb

[1828] Peter G. Neumann and Declan McCullagh. Inside risks: risks of Y2K. Communications of the ACM, 42(6):144, June 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1999-42-6/p144-mccullagh/.

Crawford:1999:EPd

[1829] Diane Crawford. Editorial pointers.

Communications of the ACM, 42(7):5, July 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1999-42-7/p5-crawford/.

Fox:1999:NTg

[1830] Robert Fox. News track. Communications of the ACM, 42(7):9–10, July 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
7/p9-fox/.

Crawford:1999:FCD

[1831] Diane Crawford, Tom DeMarco, Don Bagert, Joel C. Ewing, Hal Berghel, Ken Kashmarek, Hank Walker, and Don Unger. Forum: Certification or decertification?; rule as myth; earning trust; misleading numbers; correction. Communications of the ACM, 42(7): 11–13, July 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1999-42-7/p11-crawford/.

McGinity:1999:SCT

[1832] Meg McGinity. Staying connected: Telecom act, scene one. Communications of the ACM, 42(7):15–18, July 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        7/p15-mcginity/.

Munro:1999:WIA

[1833] Neil Munro. From washington: Infowar: AK-47s, lies, and videotape. Communications of the ACM, 42(7): 19–22, July 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org:80/pubs/citations/journals/
                                        cacm/1999-42-7/p19-munro/.

Tilley:1999:SNS

[1834] Scott Tilley. On site: the need for speed. Communications of the ACM, 42(7):23–26, July 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        7/p23-tilley/.

Gallaugher:1999:VCN

[1835] John Gallaugher. Viewpoint: challenging the new conventional wisdom of net commerce strategies. Communications of the ACM, 42(7):27–29, July 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        7/p27-gallaugher/.

Panda:1999:DIW

[1836] Brajendra Panda and Joseph Giordano. Defensive information warfare. Communications of the ACM, 42(7): 30–32, July 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1999-42-7/p30-panda/.

Chin:1999:HCD

[1837] Shiu-Kai Chin. High-confidence design for security: don’t trust — verify. Communications of the ACM, 42(7):33–37, July 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1999-42-7/p33-chin/.

Ghosh:1999:ISS

[1838] Anup K. Ghosh and Jeffrey M. Voas.

Inoculating software for survivability. Communications of the ACM, 42(7): 38–44, July 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1999-42-7/p38-ghosh/.

Goan:1999:CBC

[1839] Terrance Goan. A cop on the beat: collecting and appraising intrusion evidence. Communications of the ACM, 42(7):46–52, July 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
7/p46-goan/.

Durst:1999:TEC

[1840] Robert Durst, Terrence Champion, Brian Witten, Eric Miller, and Luigi Spagnuolo. Testing and evaluating computer intrusion detection systems. Communications of the ACM, 42(7): 53–61, July 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1999-42-7/p53-durst/. See ad-

dendum [1870].

Stillerman:1999:IDD

[1841] Matthew Stillerman, Carla Marceau, and Maureen Stillman. Intrusion detection for distributed applications. Communications of the ACM, 42(7):62–69, July 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1999-42-7/p62-stillerman/.

Jajodia:1999:TR

[1842] Sushil Jajodia, Catherine D. McCollum, and Paul Ammann. Trusted recovery. Communications of the ACM, 42(7):71–75, July 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//www.acm.org:80/pubs/citations/
journals/cacm/1999-42-7/p71-jajodia/

.

Waldo:1999:JAN

[1843] Jim Waldo. The Jini architecture for network-centric computing. Communications of the ACM, 42(7):76–82, July 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/

citations/journals/cacm/1999-42-
7/p76-waldo/.

Plummer:1999:RLD

[1844] David Plummer, Larry J. Dalton, and Frank Peter. The recodable locking device. Communications of the ACM, 42(7):83–87, July 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
7/p83-plummer/.

Sipior:1999:DSE

[1845] Janice C. Sipior and Burke T. Ward.

The dark side of employee email. Communications of the ACM, 42(7):88–95, July 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1999-42-7/p88-sipior/.

Kock:1999:CAP

[1846] Ned Kock. A case of academic plagiarism. Communications of the ACM, 42(7):96–104, July 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
7/p96-kock/.

Neumann:1999:IRI

[1847] Peter G. Neumann. Inside risks: information is a double-edged sword. Communications of the ACM, 42(7):120, July 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1999-42-7/p120-neumann/.

Crawford:1999:EPe

[1848] Diane Crawford. Editorial pointers. Communications of the ACM, 42(8):5, August 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1999-42-8/p5-crawford/.

Fox:1999:NTh

[1849] Robert Fox. News track. Communications of the ACM, 42(8):9–10, August 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        8/p9-fox/.

Crawford:1999:FFL

[1850] Diane Crawford, Leon Kappelman, Ken Mandefrot, Chuck Crawford, Dennis E. Hamilton, Orville E. Wheeler, Albert L. LeDuc, Max Hailperin, Michael E. Whitman, Anthony M. Townsend, Robert J. Aalberts, Charles Hixson, Daniel Berdichevsky, Erik Neunschwander, Bength Klenbergs, and Linux Torvalds. Forum: The folly of laws limiting Y2K liability; regarding Glass; duly noted; persuasive technologies; Linux goal. Communications of the ACM, 42(8):11–15, August 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        8/p11-crawford/.

Crawford:1999:CRH

[1851] Diane Crawford. Correction: Richard Heeks’ column. Communications of the ACM, 42(8):15, August 1999.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
8/p11-crawford/. See [1809].

Glass:1999:PPB

[1852] Robert L. Glass. Practical programmer: buzzwordism and the epic $150 million software debacle. Communications of the ACM, 42(8):17–19, August 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
8/p17-glass/.

Soloway:1999:LES

[1853] Elliot Soloway, Wayne Grant, Robert Tinger, Jeremy Roschelle, Mike Mills, Mitchell Resnick, Robbie Berg, and Mike Eisenberg. Log on education: science in the palms of their hands. Communications of the ACM, 42(8):21–26, August 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1999-42-8/p21-soloway/.

Rosenbloom:1999:TII

[1854] Andrew Rosenbloom. Toward an image indistinguishable from reality. Communications of the ACM, 42(8):28–30, August 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1999-42-8/p28-rosenbloom/.

Terzopoulos:1999:ALC

[1855] Demetri Terzopoulos. Artificial life for computer graphics. Communications of the ACM, 42(8):32–42, August 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        8/p32-terzopoulos/.

Greenberg:1999:FRI

[1856] Donald P. Greenberg. A framework for realistic image synthesis. Communications of the ACM, 42(8):44–53, August 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        8/p44-greenberg/.

Fitzmaurice:1999:SSI

[1857] George W. Fitzmaurice, Ravin Balakrisnan, and Gordon Kurtenbach. Sampling, synthesis, and input devices. Communications of the ACM, 42(8):54–63, August 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        8/p54-fitzmaurice/.

Badler:1999:ACR

[1858] Norman I. Badler, Martha S. Palmer, and Rama Bindiganavale. Animation control for real-time virtual humans. Communications of the ACM, 42(8):64–73, August 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        8/p64-badler/.

Salisbury:1999:MGP

[1859] J. Kenneth Salisbury, Jr. Making graphics physically tangible. Communications of the ACM, 42(8):74–81, August 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1999-42-8/p74-salisbury/.

Bryson:1999:VEG

[1860] Steve Bryson, David Kenwright, Michael Cox, David Ellsworth, and Robert Haimes. Visually exploring gigabyte data sets in real time. Communications of the ACM, 42(8):82–90, August 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1999-42-8/p82-bryson/.

Singh:1999:WHP

[1861] Surendra N. Singh and Nikunj P.

Dalal. Web home pages as advertisements. Communications of the ACM, 42(8):91–98, August 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
8/p91-singh/.

Ebert:1999:RVU

[1862] David Ebert, Edward Bedwell, Stephen Maher, Laura Smoliar, and Elizabeth Downing. Realizing 3D visualization using grossed-beam volumetric displays. Communications of the ACM, 42(8):100–107, August 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        8/p100-ebert/.

Vedder:1999:CCP

[1863] Richard G. Vedder, Michael T. Vanecek, C. Stephen Guynes, and James J. Cappel. CEO and CIO perspectives on competitive intelligence. Communications of the ACM, 42(8):108–116, August 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        8/p108-vedder/.

Truex:1999:GSE

[1864] Duane P. Truex, Richard Baskerville, and Heinz Klein. Growing systems in emergent organizations. Communications of the ACM, 42(8):117–123, August 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1999-42-8/p117-truex/.

Guerraoui:1999:TOOb

[1865] Rachid Guerraoui and Mohamed E. Fayad. Thinking objectively: objectoriented abstractions for distributed programming. Communications of the ACM, 42(8):125–127, August 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        8/p125-guerraoui/.

Schneier:1999:IRU

[1866] Bruce Schneier. Inside risks: the uses and abuses of biometrics. Communications of the ACM, 42(8):136, August 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
8/p136-schneier/.

Crawford:1999:EPf

[1867] Diane Crawford. Editorial pointers.

Communications of the ACM, 42(9):5, September 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1999-42-9/p5-crawford/.

Fox:1999:NTi

[1868] Robert Fox. News track. Communications of the ACM, 42(9):9–10, September 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1999-42-9/p9-fox/.

Crawford:1999:FVD

[1869] Diane Crawford, Bob Quinn, Daniel Dulitz, Andrew Odlyzko, Kerry Coffman, Michael Noll, James E. Harvey, and Misha Vaughan. Forum: Voice and data traffic jam; buying new PCs; ACM’s copyright policy. Communications of the ACM, 42(9):11–15, September 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1999-42-9/p11-crawford/.

Durst:1999:FAT

[1870] Robert Durst, Terrence Champion, Eric Miller, Luigi Spagnuolo, and Brian Witten. Forum: Addendum to “Testing and evaluating computer intrusion detection systems”. Communications of the ACM, 42(9):15, September 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
9/p11-crawford/. See [1840].

Samuelson:1999:LSW

[1871] Pamela Samuelson. Legally speaking: why the anticircumvention regulations need revision. Communications of the ACM, 42(9):17–21, September 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
9/p17-samuelson/.

McHenry:1999:IPY

[1872] William K. McHenry. International perspectives: Y2K international status and opportunities. Communications of the ACM, 42(9):23–27, September 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
9/p23-mchenry/.

Phoha:1999:VCC

[1873] Vir V. Phoha. Viewpoint: Can a course be taught entirely via email? Communications of the ACM, 42(9):29–30, September 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1999-42-9/p29-phoha/.

Palmer:1999:DNE

[1874] Jonathan W. Palmer and Lars Bo Eriksen. Digital newspapers explore marketing on the Internet. Communications of the ACM, 42(9):32–40, September 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1999-42-9/p32-palmer/.

Robertson:1999:EBY

[1875] Stewart Robertson and Philip Powell. Exploiting the benefits of Y2K preparation. Communications of the ACM, 42(9):42–48, September 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
9/p42-robertson/.

Dhamija:1999:TCM

[1876] Rachna Dhamija, Rachelle Heller, and Lance J. Hoffman. Teaching e-commerce to a multidisciplinary class. Communications of the ACM, 42(9):50–55, September 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
9/p50-dhamija/.

Gupta:1999:ENM

[1877] Alok Gupta, Dale O. Stahl, and Andrew B. Whinston. The economics of network management. Communications of the ACM, 42(9):57–63, September 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org:80/pubs/citations/journals/
                                        cacm/1999-42-9/p57-gupta/.

Ramesh:1999:DMT

[1878] R. Ramesh and Dee H. Andrews. Distributed mission training: teams, virtual reality, and real-time networking. Communications of the ACM, 42(9):64–67, September 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        9/p64-ramesh/.

Carroll:1999:MIT

[1879] Lynn A. Carroll. Multimodal integrated team training. Communications of the ACM, 42(9):68–71, September 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        9/p68-carroll/.

Bell:1999:EDM

[1880] Herbert H. Bell. The effectiveness of distributed mission training. Communications of the ACM, 42(9):72–78, September 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org:80/pubs/citations/journals/
                                        cacm/1999-42-9/p72-bell/.

Dahmann:1999:RAS

[1881] Judith S. Dahmann, James O. Calvin, and Richard M. Weatherly. A reusable architecture for simulations. Communications of the ACM, 42(9):79–84, September 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1999-42-9/p79-dahmann/.

Peppler:1999:IGS

[1882] Philipp W. Peppler and Steve Stephens.

Image generation systems in virtual training platforms. Communications of the ACM, 42(9):85–88, September 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
9/p85-peppler/.

Crane:1999:IDM

[1883] Peter Crane. Implementing distributed mission training. Communications of the ACM, 42(9):90–94, September 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
9/p90-crane/.

Tatham:1999:TOG

[1884] Eric W. Tatham. Technical opinion: getting the best of both real and virtual worlds. Communications of the ACM, 42(9):96–98, September 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
9/p96-tatham/.

Schneier:1999:IRT

[1885] Bruce Schneier. Inside risks: the Trojan horse race. Communications of the ACM, 42(9):128, September 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
9/p128-schneier/.

Crawford:1999:EPg

[1886] Diane Crawford. Editorial pointers. Communications of the ACM, 42(10): 5, October 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1999-42-10/p5-crawford/.

Fox:1999:NTj

[1887] Robert Fox. News track. Communications of the ACM, 42(10):9–10, October 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/journals/cacm/
1999-42-10/p9-fox/.

Crawford:1999:FSPa

[1888] Diane Crawford, Saul I. Gass, Rommert J. Casimir, Adrian Bowyer, Anonymous, Glen Henshaw, Scott M. Brown, and William Adams. Forum: Solving the plagiarism problem; the limits of LEGOs; what about information dominance warfare?; the licensing issue. Communications of the ACM, 42(10):11–14, October 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
10/p11-crawford/.

Munro:1999:WOT

[1889] Neil Munro. From washington: the ongoing tug-of-war for e-commerce control. Communications of the ACM, 42(10):17–20, October 1999.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
10/p17-munro/.

Press:1999:PCP

[1890] Larry Press. Personal computing:

the post-PC era. Communications of the ACM, 42(10):21–24, October 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
10/p21-press/.

Simons:1999:PHA

[1891] Barbara Simons. From the President:

The heroes among us. Communications of the ACM, 42(10):25, 112, October 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
10/p25-simons/.

Booch:1999:UA

[1892] Grady Booch. UML in action. Communications of the ACM, 42(10):26– 28, October 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1999-42-10/p26-booch/.

Kobryn:1999:USO

[1893] Cris Kobryn. UML 2001: a standardization odyssey. Communications of the ACM, 42(10):29–37, October 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/

                                        citations/journals/cacm/1999-42-
                                        10/p29-kobryn/.

Larsen:1999:DCB

[1894] Grant Larsen. Designing componentbased frameworks using patterns in the UML. Communications of the ACM, 42(10):38–45, October 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        10/p38-larsen/.

Selic:1999:TCU

[1895] Bran Selic. Turning clockwise: using UML in the real-time domain. Communications of the ACM, 42(10):46– 54, October 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org:80/pubs/citations/journals/
                                        cacm/1999-42-10/p46-selic/.

Bell:1999:UEA

[1896] Alex E. Bell and Ryan W. Schmidt. UMLoquent expression of AWACS software design. Communications of the ACM, 42(10):55–61, October 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        10/p55-bell/.

Conallen:1999:MWA

[1897] Jim Conallen. Modeling Web application architectures with UML. Communications of the ACM, 42(10):63– 70, October 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1999-42-10/p63-conallen/.

Cusumano:1999:WNL

[1898] Micheal A. Cusumano and David B.

Yoffie. What Netscape learned from cross-platform software development. Communications of the ACM, 42(10):72–78, October 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
10/p72-cusumano/.

Mattsson:1999:FIP

[1899] Michael Mattsson, Jan Bosch, and Mohamed E. Fayad. Framework integration problems, causes, solutions. Communications of the ACM, 42(10):80– 87, October 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org:80/pubs/citations/journals/
cacm/1999-42-10/p80-mattsson/.

Lau:1999:PII

[1900] Tessa Lau, Oren Etzioni, and Daniel S.

Weld. Privacy interfaces for information management. Communications of the ACM, 42(10):88–94, October 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1999-42-
10/p88-lau/.

Lie:1999:MWP

[1901] Håkon Wium Lie and Janne Saarela.

Multipurpose Web publishing using HTML, XML, and CSS. Communications of the ACM, 42(10):95–101, October 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1999-42-10/p95-lie/.

Gotterbarn:1999:SEC

[1902] Don Gotterbarn, Keith Miller, and Simon Rogerson. Software engineering code of ethics is approved. Communications of the ACM, 42(10):102–107, October 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org:
                                        80/pubs/citations/journals/cacm/
                                        1999-42-10/p102-gotterbarn/.

Prechelt:1999:TOC

[1903] Lutz Prechelt. Technical opinion: comparing Java vs. C/C++ efficiency differences to interpersonal differences. Communications of the ACM, 42(10):109–112, October 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        10/p109-prechelt/.

Schneier:1999:IRR

[1904] Bruce Schneier. Inside risks: Risks of relying on cryptography. Communications of the ACM, 42(10):144, October 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org:80/pubs/
                                        citations/journals/cacm/1999-42-
                                        10/p144-schneier/.

Crawford:1999:EPh

[1905] Diane Crawford. Editorial pointers. Communications of the ACM, 42(11): 5, November 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/1999-42-11/p5-crawford/p5-
crawford.pdf;   http://www.acm.
org/pubs/citations/journals/cacm/
1999-42-11/p5-crawford/.

Fox:1999:NTk

[1906] Robert Fox. News track. Communications of the ACM, 42(11):9–10, November 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
1999-42-11/p9-fox/p9-fox.pdf; http:
//www.acm.org/pubs/citations/journals/
cacm/1999-42-11/p9-fox/.

Berghel:1999:DVC

[1907] Hal Berghel. Digital village: the cost of having analog executives in a digital world. Communications of the ACM, 42(11):11–15, November 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/1999-42-11/p11-berghel/p11-
berghel.pdf;   http://www.acm.org/
pubs/citations/journals/cacm/1999-
42-11/p11-berghel/.

Glass:1999:ENT

[1908] Robert L. Glass. Evolving a new theory of project success. Communications of the ACM, 42(11):17–19, November 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
1999-42-11/p17-glass/p17-glass.
pdf;   http://www.acm.org/pubs/

citations/journals/cacm/1999-42-
11/p17-glass/.

Crawford:1999:FSPb

[1909] Diane Crawford, Dino Mandrioli, Paul Zeiger, Dan Chisham, Mark Wallace, Ralph E. Johnson, Robert Glass, Robert A. Ellis, Myles Losch, and Scott Tilley. Forum: Software project failure lessons learned; more powerful Internet connections. Communications of the ACM, 42(11):21–24, November 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
1999-42-11/p21-crawford/p21-crawford.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/1999-42-
11/p21-crawford/.

Munakata:1999:KD

[1910] Toshinori Munakata. Knowledge discovery. Communications of the ACM, 42(11):26–29, November 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/articles/journals/
cacm/1999-42-11/p26-munakata/p26-
munakata.pdf;   http://www.acm.
org/pubs/citations/journals/cacm/
1999-42-11/p26-munakata/.

Mitchell:1999:MLD

[1911] Tom M. Mitchell. Machine learning and data mining. Communications of the ACM, 42(11):30–36, November 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
1999-42-11/p30-mitchell/p30-mitchell.
pdf;   http://www.acm.org/pubs/

citations/journals/cacm/1999-42-
11/p30-mitchell/.

Valdes-Perez:1999:DTS

[1912] Raúl E. Valdés-Pérez. Discovery tools for science apps. Communications of the ACM, 42(11):37–41, November 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
1999-42-11/p37-valdes-perez/p37-
valdes-perez.pdf;   http://www.
acm.org/pubs/citations/journals/
cacm/1999-42-11/p37-valdes-perez/

.

Muggleton:1999:SKD

[1913] Stephen Muggleton. Scientific knowledge discovery using inductive logic programming. Communications of the ACM, 42(11):42–46, November 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/cacm/
1999-42-11/p42-muggleton/p42-muggleton.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/1999-42-
11/p42-muggleton/.

Fu:1999:KDB

[1914] LiMin Fu. Knowledge discovery based on neural networks. Communications of the ACM, 42(11):47–50, November 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
1999-42-11/p47-fu/p47-fu.pdf; http:
//www.acm.org/pubs/citations/journals/
cacm/1999-42-11/p47-fu/.

DeJong:1999:ECD

[1915] Kenneth A. De Jong. Evolutionary computation for discovery. Communications of the ACM, 42(11):51–53, November 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        1999-42-11/p51-de\_jong/p51-de\
                                        _jong.pdf;   http://www.acm.org/
                                        pubs/citations/journals/cacm/1999-
                                        42-11/p51-de\_jong/;   http://www.
                                        acm.org/pubs/citations/journals/
                                        cacm/1999-42-11/p51-de_jong/.

Ziarko:1999:DTR

[1916] Wojciech Ziarko. Discovery through rough set theory. Communications of the ACM, 42(11):54–57, November 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/cacm/
                                        1999-42-11/p54-zirko/p54-zirko.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/cacm/1999-42-
                                        11/p54-zirko/.

Knight:1999:MOT

[1917] Kevin Knight. Mining online text. Communications of the ACM, 42(11): 58–61, November 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        cacm/1999-42-11/p58-knight/p58-
                                        knight.pdf;   http://www.acm.org/
                                        pubs/citations/journals/cacm/1999-
                                        42-11/p58-knight/.

Schulze-Kremer:1999:DHG

[1918] Steffen Schulze-Kremer. Discovery in the human genome project. Communications of the ACM, 42(11):62–64, November 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
1999-42-11/p62-schulze-kremer/p62-
schulze-kremer.pdf;   http://www.
acm.org/pubs/citations/journals/
cacm/1999-42-11/p62-schulze-kremer/

.

Campbell:1999:KDD

[1919] Murray Campbell. Knowledge discovery in deep blue. Communications of the ACM, 42(11):65–67, November 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
1999-42-11/p65-campbell/p65-campbell.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/1999-42-
11/p65-campbell/.

Ganesan:1999:MAP

[1920] Ravi Ganesan. The messyware advantage: the precipitous rise of the middleman. Communications of the ACM, 42(11):68–73, November 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/1999-42-11/p68-ganesan/p68-
ganesan.pdf;   http://www.acm.org/
pubs/citations/journals/cacm/1999-
42-11/p68-ganesan/.

Oviatt:1999:TMM

[1921] Sharon Oviatt. Ten myths of multimodal interaction. Communications of the ACM, 42(11):74–81, November 1999. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        cacm/1999-42-11/p74-oviatt/p74-
                                        oviatt.pdf;   http://www.acm.org/
                                        pubs/citations/journals/cacm/1999-
                                        42-11/p74-oviatt/.

Nezlek:1999:IAE

[1922] George S. Nezlek, Hemant K. Jain, and Derek L. Nazareth. An integrated approach to enterprise computing architectures. Communications of the ACM, 42(11):82–90, November 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        cacm/1999-42-11/p82-nezlek/p82-
                                        nezlek.pdf;   http://www.acm.org/
                                        pubs/citations/journals/cacm/1999-
                                        42-11/p82-nezlek/.

Gregg:1999:POD

[1923] Dawn G. Gregg and Michael Goul. A proposal for an open DSS protocol. Communications of the ACM, 42 (11):91–96, November 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        cacm/1999-42-11/p91-gregg/p91-gregg.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/cacm/1999-42-
                                        11/p91-gregg/.

Felber:1999:POD

[1924] Pascal Felber, Rachid Guerraoui, and Mohamed E. Fayad. Putting OO distributed programming to work. Communications of the ACM, 42(11): 97–101, November 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/articles/journals/
cacm/1999-42-11/p97-felber/p97-
felber.pdf;   http://www.acm.org/
pubs/citations/journals/cacm/1999-
42-11/p97-felber/.

Neumann:1999:IRRc

[1925] Peter G. Neumann and Lauren Weinstein. Inside risks: risks of content filtering. Communications of the ACM, 42(11):152, November 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/cacm/
1999-42-11/p152-neumann/p152-neumann.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/1999-42-
11/p152-neumann/.

Crawford:1999:EPi

[1926] Diane Crawford. Editorial pointers.

Communications of the ACM, 42(12): 5, December 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/1999-42-12/p5-crawford/p5-
crawford.pdf;   http://www.acm.
org/pubs/citations/journals/cacm/
1999-42-12/p5-crawford/.

Fox:1999:NTl

[1927] Robert Fox. News track. Communications of the ACM, 42(12):9–10, December 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
1999-42-12/p9-fox/p9-fox.pdf; http:
//www.acm.org/pubs/citations/journals/
cacm/1999-42-12/p9-fox/.

Crawford:1999:FRU

[1928] Diane Crawford, Peter Revesz, Philip Gilbert, and S. W. Ng. Forum: Reengineering the university. Communications of the ACM, 42(12):11–13, December 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        1999-42-12/p11-crawford/p11-crawford.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/cacm/1999-42-
                                        12/p11-crawford/.

Godwin:1999:WNW

[1929] Mike Godwin. From Washington: net to worry. Communications of the ACM, 42(12):15–17, December 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        cacm/1999-42-12/p15-godwin/p15-
                                        godwin.pdf;   http://www.acm.org/
                                        pubs/citations/journals/cacm/1999-
                                        42-12/p15-godwin/.

McGinity:1999:SCF

[1930] Meg McGinity. Staying connected: flying wireless, with a net. Communications of the ACM, 42(12):19–21, December 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        1999-42-12/p19-mcginity/p19-mcginity.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/cacm/1999-42-
                                        12/p19-mcginity/.

Walsh:1999:SEA

[1931] Patrick Walsh and Adamantios Koumpis. On site: the Esprit ATLAS project.

Communications of the ACM, 42(12): 23–26, December 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/articles/journals/
cacm/1999-42-12/p23-walsh/p23-walsh.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/1999-42-
12/p23-walsh/.

El-Kadi:1999:VSD

[1932] Amr El-Kadi. Viewpoint: stop that divorce! Communications of the ACM, 42(12):27–28, December 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/1999-42-12/p27-el-kadi/p27-
el-kadi.pdf;   http://www.acm.org/
pubs/citations/journals/cacm/1999-
42-12/p27-el-kadi/.

Frailey:1999:LSE

[1933] Dennis J. Frailey. Licensing software engineers. Communications of the ACM, 42(12):29–30, December 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/1999-42-12/p29-frailey/p29-
frailey.pdf;   http://www.acm.org/
pubs/citations/journals/cacm/1999-
42-12/p29-frailey/.

Bellman:1999:POB

[1934] Steven Bellman, Gerald L. Lohse, and Eric J. Johnson. Predictors of online buying behavior. Communications of the ACM, 42(12):32–38, December 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        cacm/1999-42-12/p32-bellman/p32-
                                        bellman.pdf;   http://www.acm.org/
                                        pubs/citations/journals/cacm/1999-
                                        42-12/p32-bellman/.

Adams:1999:UE

[1935] Anne Adams and Martina Angela Sasse. Users are not the enemy. Communications of the ACM, 42(12):40–46, December 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        1999-42-12/p40-adams/p40-adams.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/cacm/1999-42-
                                        12/p40-adams/.

Dutton:1999:CSN

[1936] William H. Dutton, Anita Elberse, and Matthew Hale. A case study of a Netizen’s guide to elections. Communications of the ACM, 42(12):48–54, December 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        cacm/1999-42-12/p48-dutton/p48-
                                        dutton.pdf;   http://www.acm.org/
                                        pubs/citations/journals/cacm/1999-
                                        42-12/p48-dutton/.

Larsen:1999:VTI

[1937] Kai R. T. Larsen. Voting technology implementation. Communications of the ACM, 42(12):55–57, December 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        cacm/1999-42-12/p55-larsen/p55-

larsen.pdf;   http://www.acm.org/
pubs/citations/journals/cacm/1999-
42-12/p55-larsen/.

Watson:1999:TLG

[1938] Richard T. Watson, Sigmund Akselsen, Benete Evjemo, and Nils Aarsæther. Teledemocracy in local government. Communications of the ACM, 42(12):58–63, December 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/1999-42-12/p58-watson/p58-
watson.pdf;   http://www.acm.org/
pubs/citations/journals/cacm/1999-
42-12/p58-watson/.

Igbaria:1999:DFV

[1939] Magid Igbaria. The driving forces in the virtual society. Communications of the ACM, 42(12):64–70, December 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/1999-42-12/p64-igbaria/p64-
igbaria.pdf;   http://www.acm.org/
pubs/citations/journals/cacm/1999-
42-12/p64-igbaria/.

Tellioglu:1999:SC

[1940] Hilda Tellioğlu and Ina Wagner. Software cultures. Communications of the ACM, 42(12):71–77, December 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/cacm/
1999-42-12/p71-tellioglu/p71-tellioglu.
pdf;   http://www.acm.org/pubs/

citations/journals/cacm/1999-42-
12/p71-tellioglu/.

Brereton:1999:FS

[1941] Pearl Brereton, David Budgen, Keith Bennnett, Malcolm Munro, Paul Layzell, Linda MaCaulay, David Griffiths, and Charles Stannett. The future of software. Communications of the ACM, 42(12):78–84, December 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
1999-42-12/p78-brereton/p78-brereton.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/1999-42-
12/p78-brereton/.

Payton:1999:HCH

[1942] Fay Cobb Payton and Patricia Flatley Brennan. How a community health information network is really used. Communications of the ACM, 42(12):85–89, December 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/1999-42-12/p85-payton/p85-
payton.pdf;   http://www.acm.org/
pubs/citations/journals/cacm/1999-
42-12/p85-payton/.

Cuccias:1999:WRF

[1943] Frank Cuccias. Are we ready?: the FAA’s Y2K preparedness. Communications of the ACM, 42(12):91–92, December 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/1999-42-12/p91-cuccias/p91-
cuccias.pdf;   http://www.acm.org/

pubs/citations/journals/cacm/1999-
42-12/p91-cuccias/.

Kappelman:1999:WRR

[1944] Leon A. Kappelman. Are we ready?:

risk, reality, and readiness. Communications of the ACM, 42(12):93–94, December 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
1999-42-12/p93-kappelman/p93-kappelman.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/1999-42-
12/p93-kappelman/.

Neumann:1999:IRRd

[1945] Peter G. Neumann. Inside risks:

risks of insiders. Communications of the ACM, 42(12):160, December 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/cacm/
1999-42-12/p160-neumann/p160-neumann.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/1999-42-
12/p160-neumann/.

Karp:1981:ERPa

[1946] R. M. Karp and M. O. Rabin. Efficient randomized pattern-matching algorithms. Technical report TR-31-81, Harvard University, Cambridge, MA, USA, ?? 1981. ?? pp. An incremental hash function is described for application to the string search problem. See [381].

Blinn:1997:JBC

[1947] James F. Blinn. Jim Blinn’s corner:

Fugue for MMX. IEEE Computer Graphics and Applications, 17(2):88– 93, March/April 1997. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). Makes several cogent comments about deficiencies in the Intel MMX pixel-processing instruction set [1193] for use in image compositing.

Knuth:1977:FPM

[1948] Donald E. Knuth, J. H. Morris, and V. R. Pratt. Fast pattern matching in strings. SIAM Journal on Computing, 6(2):323–350, June 1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See also [1, 85, 381].

Dittmer:1976:IEP

[1949] Ingo Dittmer. Implementation eines Einschrittcompilers für die Programmiersprache PASCAL auf der Rechenanlage IBM/360 der Universität Münster. (English title: Implementation of a one-step compiler for the programming language PASCAL on the IBM/360 of the University of Muenster). Diplomearbeit, Universität Münster, Münster, Germany, ?? 1976. ?? pp. Diplomearbeit Münster 1976 und doert angegebene Literatur (English: Muenster diploma work 1976 and the literature cited therein). The hashing method was rediscovered fourteen years later by Pearson [63], and then commented on by several authors [244, 245, 246, 247].

Karp:1981:ERP

[1950] R. M. Karp and M. O. Rabin. Efficient randomized pattern-matching algorithms. Technical report TR-31-81, Harvard University, Cambridge, MA, USA, ?? 1981. ?? pp. An incremental hash function is described for application to the string search problem. See [381].

Knuth:1992:LP

[1951] Donald E. Knuth. Literate Programming. CSLI Lecture Notes Number 27. Stanford University Center for the Study of Language and Information, Stanford, CA, USA, 1992. ISBN 0-937073-80-6 (paperback), 0-937073-81-4 (hardcover). xv + 368 pp. LCCN QA76.6 .K644 1992. US$24.95 (paperback), US$59.95 (hardcover). Distributed by the University of Chicago Press.